A Bibliography of Papers in Lecture Notes in Computer Science (1999), Part 1 of 2

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: http://www.math.utah.edu/~beebe/

11 June 2016 Version 1.14

Title word cross-reference

(+, <) [90]. (n, m) [696]. (t, s) [391]. 1 [1023, 591]. 1/2 [590]. 2 [419, 75, 591, 1555]. 26 [697]. 3 [1529, 121, 691, 307, 417, 1061, 124, 703, 714, 1453, 123, 1454, 724, 818, 105, 125]. c [659]. d [1403]. e [205, 206]. H [1135, 865]. i [1012]. j [591]. j + 1 [591]. K [899, 1150, 390, 613, 600, 659]. L2 [355]. λ [1005]. m [1336, 577]. μ [666, 619]. N [445, 1182, 846, 1181, 847, 708, 390, 475]. N^{0.292} [1403]. O [728]. p [588, 1143]. π [778]. V [1132].

-Approximation [590]. -ary [390]. -body [445, 846, 847]. -Calculus [619, 666, 778, 1005]. -Cubes [390, 696]. -D [419, 417, 703, 123]. -Decorrelation [355]. -Dimensional [1529, 75]. -formulae [659]. -means [659, 899]. -Median [1143]. -Minimal [728]. -Nearest [1150]. -Permutation [1336]. -person [1182, 1181]. -Protocol [1012]. -Rule [865]. -Safe [1023]. -Section [588]. -Sequences [391]. -sets [1132]. -Stable [613]. -Subsumption [1135]. -Surfaces [697]. -th [475]. -versus- [591].

1.1 [1590]. 1995 [435].

2 [442]. 2000 [288]. 2D [701]. 2n0n [370].

3XOR [381].

’97 [371].

= [1117].

A-Teams [343]. ABACO [169]. ABC [1327]. Abound [788]. Abramsky [1080]. Absence [1345]. Abstract [483, 159, 341, 1080, 1373, 130, 1084, 163]. Abstraction [145, 595, 1019]. Abstractions [133, 742, 743]. Abuse [198]. Academic [204]. Accelerate [482]. Accelerating [1334]. Acceleration [721, 1344]. Acceptance [333]. Access [193, 192, 194, 269, 1560, 833, 265, 285, 1233, 1353, 1628, 272, 37]. Accesses [1290]. Accessibility [1507]. Accessible [37]. Accounting [1609, 634, 770]. Accuracy [885]. Accurate [359, 410, 869]. Achieve [1244]. Acme [1583]. Acme-Based [1583]. Acquire [554]. Acquired [1145]. Acquiring [1492]. Acquisition [756, 465, 126]. Across [1244]. Action [1037, 169, 114, 113, 318, 170, 681, 682, 234, 1356]. Actions [1171, 1172]. Active [1493, 121, 120, 647, 1584, 119, 116, 1513, 313, 1565, 108, 225, 1573, 272, 126, 1159]. Activities [1572]. Activity [1556, 1581, 279]. Adapt [1416]. Adaptation [1238, 1145]. Adaptive [1260, 1410, 1230, 1236, 1425, 258, 1189, 127, 1446, 1198, 1300, 1161, 1146, 1258, 1284, 1255, 430, 1329, 272]. Adding [1047, 71]. Addition [596]. Addressable [1292]. Addressing [1318]. Adequacy [1133]. Adiabatic [51]. ADIPS [1660]. Administration [637]. Admissible [1366]. ADORE [127]. ADSL [1628]. Advanced [1550, 1549, 1044]. Advances [248]. Advection [851, 1477]. Adversarial [462]. Adversary [1425]. Aerial [683]. after [90]. Against [1410, 521, 1425, 375, 1421, 460, 1336, 166]. Ageing [1440]. Agencies [518]. Agenda [853, 473, 474]. Agent [334, 1618, 1621, 768, 350, 769, 1658, 1589, 1632, 338, 771, 551, 341, 773, 351, 1670, 347, 1586, 1620, 348, 1669, 671, 1561, 1662, 1663, 915, 346, 340, 1654, 335, 1619, 343, 1666, 336, 1622, 344, 353, 349, 1648, 1657, 1661, 329, 1671, 1664, 1173, 1656, 216, 774, 1587]. Agent-Based [771, 1648]. Agent-Mediated [773, 1671]. Agent-Oriented [347]. Agent-Systems [344]. Agentis [348]. Agents [550, 1562, 549, 1560, 1171, 1172, 777, 630, 1667, 1000, 1668, 339, 1638, 1215, 1174, 1646, 1647, 1665, 1645, 776, 543, 1614, 564, 552, 335, 354, 913, 914]. Aggregate [91, 1211]. Agreement [380, 1335]. Ahead [725]. AI [649, 317, 648]. AI-STRATA [317]. Aided [766, 427, 673, 904, 99]. AIDI [1530]. AIDI-Heart [1530]. Air [727, 368, 660]. Air-Traffic [727]. Airborne [1236, 110]. Aircraft [763]. Airflow [818]. Airport [1443]. Ajdukiewicz [1092]. Alamos [52]. Album [297]. Album-Oriented [297]. Algebra [1025, 580, 1362, 1364, 151, 1456, 150]. Algebraic [1358, 137, 992, 154, 1537, 1070, 152, 601, 557, 558, 1115, 157, 166, 170, 978, 130, 153, 1378, 169]. algebras [1363]. ALGOL [1005]. Algorithm [1220, 899, 588, 820, 385, 839, 923, 337, 23, 1460, 1165, 997, 1536, 1143, 670, 389, 1150, 901, 256, 255, 1152, 1416, 1556, 1201, 907, 257, 713, 898, 1312, 590, 1447, 1454, 1009, 1124, 390, 1227, 828, 889, 1379, 873, 791, 924, 860, 1481, 897]. Algorithm-Dependant [839]. Algorithms [1225, 1157, 445, 814, 852, 446, 386, 1307, 1148, 1179, 1180, 246, 1189, 612, 419, 15, 805, 1196, 792, 1293, 20, 1170, 1470, 1186, 936, 1267, 370, 1154, 620, 82, 564, 1187, 1144, 468, 938, 455, 440, 103, 497, 1147]. Alignment [1157, 1634]. All-Optical [1289]. All-or-Nothing [526]. All-Pairs [90]. Allocation [1149, 237, 936, 1645, 1169, 1521]. Almost [814]. Alpha [457, 458, 1554]. Alpha-Beta [457, 458]. Alternatives [5, 360]. Ambients [999]. Among [392, 1171, 1172, 1576, 1053, 919]. Amount [884]. Amplitude [23]. Analog [743, 744]. Analogy [543]. Analyses [1449]. Analysis [1225, 173, 1582, 1033, 1016, 708, 1528, 993, 1035, 139, 667, 158, 986, 494, 736, 114, 397, 398, 478, 908, 285, 113, 953, 837, 673, 726, 1308, 727, 452, 1017, 931, 821, 1480, 951, 1187, 1429, 524, 1486, 172, 174, 108, 960, 745, 606, 1019, 1353, 668, 556, 1032, 1390, 830, 841, 686, 917, 749, 476, 1096]. Analytic [1543]. Analytical [1484, 76]. Analyzing [1010, 1487]. Ancient [478, 627]. Angiographic [1454]. Angiographical [1529]. Animating [1186]. Animation [217]. AnnoDomini [1064]. Announcement [266]. Announcer [301]. Anonymity [506]. Anonymous [509, 1414]. ANSI [947, 1504]. Answer [1434]. Answering [94]. Ant [1224]. Anytime [772]. Aperture [423]. Appendix [1400]. Applicability [10, 861]. Applicable [43]. Application [1021, 187, 550, 1248, 1632, 692, 186, 188, 557, 1150, 789, 1636, 1252, 373, 465, 1040, 1452, 273, 1270, 703, 1044, 1200, 284, 658, 1628, 234]. Application-Oriented [187, 186, 188]. Applications [1063, 725, 1305, 1328, 1547, 429, 1476, 1523, 1544, 1589, 1507, 1309, 1241, 307, 1433, 1210, 974, 1287, 754, 582, 1525, 276, 224, 1320, 1480, 82, 1247, 1548, 25, 1318, 1578, 39, 1587, 235, 1254, 406]. Applied [1228, 1190]. Applying [233, 1170, 294, 638]. Approach [559, 1545, 733, 29, 129, 1240, 1593, 1584, 414, 220, 653, 783, 352, 1064, 986, 48, 806, 867, 1176, 558, 515, 1482, 1031, 609, 760, 857, 1326, 978, 1373, 415, 1219, 153, 862, 1566, 462, 125, 1664]. Approaches [444, 1669, 1634, 931, 401, 68]. Approximable [592]. Approximate [1302]. Approximating [665, 589, 726, 680]. Approximation [588, 785, 438, 590, 745, 1183]. Approximations [734]. AQUA [225]. AQUA-Based [225]. Arbitrary [23, 702].

Arbitration [1355]. Architecting [625]. Architectural [155, 1292, 5, 401, 222, 221, 943]. Architecture [117, 1613, 1476, 1494, 1610, 414, 115, 1652, 1306, 1583, 1276, 1245, 1297, 1263, 116, 1213, 1442, 740, 1044, 340, 1271, 1272, 342, 415, 343, 1043, 440, 1329, 878, 1659, 1505]. Architectures [1571, 1259, 341, 1237, 1264, 1662, 374, 977, 402, 1262, 834, 1607]. Arising [828]. Arithmetic [1275, 507, 370]. Arithmetic-Based [1275]. ARMCI [1256]. Arnoldi [807]. Array [396, 1242, 1308, 1256, 387]. Arrays [53, 430, 1344]. Arrows [165]. Art [326]. Article [301]. Articulation [886]. Artificial [560, 1201, 335]. ary [390]. Ascender [124]. ASCI [1519]. ASF [940]. Aspect [507, 825]. Aspects [32, 1471, 1630, 417, 251, 1052, 1093]. Assembly [1470, 1217]. Assessing [1053]. Assessment [11, 1480]. Assessments [2]. Assets [1169]. Assignment [650, 762, 1009, 345]. Assistance [34]. Assisted [854, 765]. Associated [917]. Association [859, 867, 863, 904, 92, 864, 861, 873, 872, 862, 860, 866]. Associations [856]. Associative [1199, 1484]. Assumptions [1407]. Asymmetric [55, 514]. Asymptotic [791]. Asynchronous [1225, 814, 425, 996, 1156, 1024, 1318, 135]. AT&T [182, 181, 183]. ATM [442, 448, 268]. Attack [1410, 1336, 1337, 378, 363]. Attacks [517, 521, 1343, 1427, 1421, 369]. Attempt [518]. Attempto [479]. Attempts [495]. Attention [553]. Attitudes [339]. Attractors [1184]. Attribute [789, 244, 242, 870, 869]. Attribute-Efficient [789]. Attribute-Oriented [242]. Attributes [922, 859, 872, 1563]. Auction [777, 770]. Auction-Based [777]. Audio [1641]. Audio/Video [1641]. Audiovisual [317].

Authenticated [380]. Authentication [1589, 511, 1355]. Author [532, 563, 780, 1102, 1357, 1673]. Authorisation [1589]. Authorization [1352]. Auto [529]. Auto-Recoverable [529]. Automata [729, 730, 995, 665, 734, 1471, 986, 787, 675, 740, 1200, 1155, 680, 683, 138, 1551, 686]. Automata-Theoretic [986]. Automated [1016, 1014, 213, 1530, 775, 881, 1628, 21]. Automatic [829, 302, 113, 311, 1017, 883, 324, 1209]. Automatically [301, 215]. Automating [1149]. Automaton [1095]. Autonomous [333, 110, 735, 335, 912, 1273, 1580, 97, 99]. Average [572, 1556, 606]. Average-Case [606]. averaged [364]. Averaging [793]. Avionics [725]. Avoidance [1438]. Avoiding [1300, 795]. Award [849, 848]. Awareness [277, 1659]. Axial [453]. Axiomatization [1073].

Back [1603]. Backward [820]. Backward-Facing [820]. Backwards [1188]. BACO [170]. Bags [997]. Balance [1310]. Balanced [582]. Balancing [604, 1479, 838, 822, 437, 1463, 1478, 1477]. Bandwidth [589]. Barriers [145, 525]. Base [757, 1204, 1587]. Based [899, 559, 1317, 1175, 304, 1582, 164, 550, 753, 923, 1275, 1054, 1528, 625, 1621, 1562, 1588, 372, 296, 237, 295, 1476, 61, 350, 664, 1314, 246, 306, 1294, 1658, 1584, 689, 169, 1261, 1232, 147, 509, 751, 653, 1469, 217, 761, 771, 254, 895, 1433, 690, 280, 777, 1583, 1424, 1612, 752, 408, 275, 409, 1466, 412, 1609, 114, 351, 767, 1497, 1301, 289, 1483, 1625, 302, 367, 1586, 1221, 531, 1620, 762, 1160, 933, 1428, 1174, 672, 1201, 1234, 484, 844, 1442, 907, 760, 1354, 1223, 1222, 1475, 713, 321, 898, 170, 1341, 1342, 962, 1272]. Based [651, 314, 294, 1429, 1530, 1419, 108, 1414, 263, 960, 1420, 1155, 864, 317, 1447, 1578, 1574, 281, 292, 122, 1284, 299, 1107, 1104, 764, 97, 1257, 889, 1058, 225, 349, 1032, 1648, 1203, 1657, 1338, 921, 881, 890, 659, 924, 860, 1481, 884, 897, 530, 1226]. Bases [336]. Basing [1407]. Basis [1524, 361, 1215, 507, 570, 349]. Basket [908]. Batch [1350]. Bayesian [1175, 909, 112, 893, 111]. BDDs [1020]. BDI [328, 330]. Be [552, 600, 840]. Bearing [1446]. Become [1598]. Bed [1058]. Behaving [327]. Behavior [320, 1258, 791]. Behavioral [643]. Behaviors [661]. Behaviour [113, 595, 681, 746, 682, 747]. Behavioural [830]. Behaviours [97]. Belief [329]. Belnap [1127]. Below [717]. Benchmark [1605, 288]. Benchmarking [213]. Benefits [1026]. Best [849, 848, 459, 1232, 460]. Best-Effort [1232]. Best-First [459]. Beta [457, 458]. Better [1244]. Between [1359, 146, 204, 842, 525, 330, 995, 303, 448, 512, 1057, 19, 165, 1145, 476, 648]. Beyond [298, 218]. Bézier [719]. Bias [921]. Bibliography [1110]. Bid [775]. Bidding [777]. Binary [1267, 570, 1268]. Binding [970, 228, 1641]. Biological [1215, 636]. Biologically [1214]. Biology [1228]. Bioremediation [1551]. Bisimulation [733, 1026, 1001, 1027]. Bisimulations [674]. Bit [1407, 1266, 1430, 406]. Bit-Rate [406]. Black [25, 1603]. Black-Holes [1603]. Blasting [1467]. Blend [546]. Blending [545]. Blind [513, 1339]. Block [443, 806, 1264, 537, 407, 56, 1542, 406, 1541, 496]. Block-Jacobi [806]. Block-Matching [406]. Block-Processing [1264]. Blocking [394]. Blue [1519]. Blue-Pacific [1519]. body [445, 846, 847]. Boiler [971, 817]. Boltzmann [1468]. Boolean [1331, 1330, 1332, 376]. Boosting [781, 893, 869]. Border [708]. both [657]. Bottom [504]. Bottom-Up [504]. Bound [1123, 602, 1313]. Boundary [722, 393, 1539]. Bounded [1021, 72, 1663, 608, 135]. Bounds [799, 1406, 621, 601, 622, 1113, 801, 798]. Box [628, 356, 25]. boxes [1423]. Branch [1313]. Branch-and-Bound [1313]. Branching [1337, 1376, 570, 613]. Branchwidth [581]. Breaking [628]. Breeding [1162]. Bridging [328]. Brief [662]. Bringing [1383]. Broadband [1444, 1652]. Broadcast [309, 259, 1486, 263, 366, 1555, 1434]. Broadcasting [312]. Broker [1646, 1647]. Brokerage [1611, 1612]. Brothers [995]. Browsing [316]. BRRA [923, 924]. BSP [1284, 1327, 1498]. Buffer [1485]. Build [1386]. Building [547, 768, 1230, 769, 142, 1049, 742, 743, 1614, 143, 553, 125, 1659]. Bunched [1081]. Burrows [281]. Bus [1246, 1233]. Bused [1296]. Buses [1295]. BVIEW [1557]. Bytecode [1013].

C [947, 1504, 1491, 942, 294, 958, 864]. C-Prototypes [294]. CA [528]. Cache [1291, 1290, 258, 836, 428, 1457, 1483, 944, 1379]. Cached [259]. Caches [1484]. CAD [1445, 1177]. CAD/CFD [1445]. Calculate [1308]. Calculation [807]. Calculations [850]. Calculi [972]. Calculus [137, 1384, 129, 1072, 147, 1075, 1024, 1004, 1081, 619, 1376, 666, 994, 778, 1005]. Calibration [126]. Call [488, 1075, 1079, 132, 931]. Call-by-Push-Value [1079]. Call-by-Value [1075]. Call-Coverage [132]. Call/Post [488]. Camera [126]. Can [35, 1184, 552, 25, 600, 1246]. Cancer [767]. Candidate [860]. Capacities [38]. Capacity [1296]. Capturing [79, 313]. Car [1445]. Card [202]. Cardiac [1453]. Cards [201, 1430]. Carlo [1460, 393, 391]. Cartesian [654]. Case [820, 753, 1054, 1544, 1232, 653, 1212, 10, 1457, 605, 673, 760, 1655, 1318, 606, 1578, 1605, 1227]. Case-Based [753, 653, 760]. Cases [657, 759, 902]. Cash [1350, 515, 1349]. CASL [155, 1371, 1372]. Casting [723]. Casy [120]. Categorical [159, 998]. Categories [728]. Causal [906, 909]. Causally [1257]. Cavities [1552]. CCS [156]. Cell [1541, 689, 690, 1272]. Cellular [1152, 1200, 1155, 430, 1551, 270]. Center [1043]. Centered [317, 1661]. Centralised [270]. Centralized [1505]. Centred [1517]. Certificates [516]. Certification [193, 192, 194, 203]. Certified [1346]. CFD [1445]. Chain [503]. Chains [1437]. Challenge [382]. Challenges [625, 1620, 636]. Change [796]. Changing [1345, 1145]. Channel [1021, 1233]. Channels [37, 1345, 1434, 38]. Chaotic [1174, 1173]. Chapter [1399, 1396, 1394, 1395, 1398, 1397, 1393, 1392]. Character [327, 101]. Characterisation [578]. Characterising [1085]. Characteristic [1416, 1125]. Characteristics [1561, 1527]. Characterization [992, 1355, 892]. Characters [327, 1145]. Charts [1003]. Chase [70]. CHAT [220]. Cheating [1340]. Checkers [467]. Checking [1358, 337, 1020, 596, 146, 1022, 1000, 138, 60]. Chemical [1154, 1080]. Chervonenkis [798]. Chess [1158]. Chess-Endgame [1158]. Chimera [1538]. Chinese [877]. CHIP [158]. Choice [1119]. Cholesky [455, 1458]. Chop [616, 617]. Chosen [1410, 363]. CIM [1531]. Cipher [374]. Ciphers [1427, 537, 365, 355]. Ciphertext [1410]. Circles [693]. Circuit [569, 1555]. Circuit-Switched [1555]. Circuits [15, 1556, 21, 135]. Circular [1001]. CLAM [1580]. Class [290, 665, 738, 739, 388, 745, 1053]. Classes [1585, 1136, 63, 701, 794, 1419]. Classical [1088]. Classification [856, 922, 301, 895, 408, 409, 901, 903, 800, 302, 1151, 898, 651, 658, 893, 354]. Classifier [1193, 869]. Classifiers [802, 884]. Classify [1158]. Classifying [568, 902]. clause [1100]. Clean [1606]. Cleansing [767]. Client [954, 815, 227]. Client-Server [954, 815]. Client-Side [227]. Climate [1449]. Climbing [1330]. Clinical [766]. Clocks [971]. Cloning [33, 353]. Closed [794, 92]. Cloud [895, 926]. CLP [1022, 238, 239, 498]. Cluster [411, 435, 436, 1315, 412, 1252, 1492, 864, 1458, 1227, 1319, 1467]. Clustering [856, 899, 1175, 1311, 896, 903, 555, 294, 900, 897]. Clusters [441, 1226, 1524, 437, 1464, 1263, 1513, 1451, 1548, 1318, 1204, 900, 1462, 1254, 818]. CMOS [122]. Co [1192, 1193, 1574, 660, 1364]. co- [1364]. Co-evolution [1192]. Co-operative [1193]. Co-ordination [1574, 660]. Coarse [1312]. Code [947, 942, 817, 1070, 1490, 231, 230, 421, 935, 851, 964, 229]. Coded [864]. Codes [1461, 41, 789, 1427, 1355, 42]. Coding [269, 795, 106, 406]. Cog [547]. Cognition [644]. Cognitive [770]. Coherence [1483]. Coherent [38]. Coin [1414]. Coin-Based [1414]. Coinductive [1134, 1001]. Collaborative [280]. Collapse [591]. Collection [1525, 440]. Collections [1614]. Collective [1514, 1205, 1231, 1216]. Collector [1325]. Collision [1438]. Collusion [1413]. Collusion-Resistant [1413]. Colonies [1224]. Color [306]. Color-Based [306]. Com [326]. Combination [1208]. Combinational [1556]. Combinator [1599]. Combinatorial [885, 1224]. Combinators [218]. Combinatory [486, 487]. Combined [387]. Combiner [1346]. Combining [139, 876, 1416, 153]. Commerce [779, 773, 515, 1535]. Commercial [204]. Commitment [535, 1407]. Committed [509]. Committee [784]. Committees [869]. Communicating [1000, 595]. Communication [1550, 1511, 1514, 1499, 47, 1431, 1435, 350, 17, 1315, 1309, 448, 261, 1028, 632, 1014, 1334, 1513, 561, 1510, 262, 1318, 349, 1203, 1254, 300, 1434, 1672, 969]. Communication-Intensive [1254]. Communication-Storage [1435].

Communications [1251, 1243, 1231].

commutative [1065, 1083]. Compact [966]. Companies [762]. Companion [1347]. Comparative [1012]. Comparing [1571, 1417]. Comparison [1543, 283, 1646, 1647, 931, 909, 864, 842, 938, 1145, 720]. Comparisons [995]. Compatible [845]. Competence [757]. Competitive [575]. Compilation [1082, 927, 940]. Compiler [102, 1256, 938, 941, 1203, 169, 170]. Compilers [928]. Compiling [493, 171]. Complementarity [667, 668]. Complementation [610]. Complete [15, 1368, 1001, 606, 26]. Completeness [618]. Complex [887, 711, 742, 743, 1162, 1540]. Complexes [689, 690]. Complexity [27, 1066, 580, 569, 1136, 17, 471, 579, 460, 72, 796, 572, 609, 82, 1122, 1135, 1127, 608, 1113]. Compliant [1645]. Component [1582, 1624, 1396, 1398, 1397, 484, 1442, 1535]. Component-Based [1582, 484, 1442]. Components [1049, 1627, 978, 1540]. Composing [485, 1569]. Composite [77, 848, 1419]. Composition [187, 486, 154, 1322, 1309, 186, 188, 487, 1432, 1580]. Compositional [338, 1017]. Compositionality [1018]. Compression [408, 409, 420, 1271, 413, 281, 106]. Computability [1125]. Computable [579, 716]. Computation [49, 721, 51, 1511, 732, 574, 29, 534, 396, 692, 40, 815, 320, 420, 450, 1174, 543, 44, 387, 453, 1140, 1552, 1173, 16]. Computational [27, 816, 1471, 399, 199, 200, 484, 360, 1475, 68, 1122, 1255, 1133, 655]. Computationally [1431]. Computations [594, 162, 1425, 1456, 1546, 1369, 1473, 1304]. Compute [572]. Computer [559, 117, 447, 1314, 838, 52, 766, 1126, 1370, 1440, 28, 854, 473, 474, 1272, 54, 112, 25, 765, 101, 100, 221, 103, 1288]. Computer-Aided [766]. Computers [839, 812, 1456, 417, 388, 832]. Computing [441, 849, 650, 435, 1260, 436, 46, 1282, 1280, 665, 1285, 53, 1261, 48, 102, 382, 1520, 259, 1452, 276, 1027, 857, 264, 1465, 1455, 776, 56, 843, 1438, 1284, 1534, 1206, 748, 1467, 1522, 826]. concatenation [1099]. Concept [867, 794, 199, 200, 894]. Concepts [823, 824, 1292, 218, 759]. Conceptual [624, 1528, 625, 626, 627, 647, 646, 630, 631, 633, 645, 644, 556, 638, 251, 1661, 919, 1101]. CONCERT [1059]. Concise [332]. Conciseness [1168]. Conclusion [1109]. Concurrency [256, 255, 1234, 1375, 330, 277, 134]. Concurrent [973, 1210, 1497, 1596, 1586, 215, 960, 1377, 1432, 1005, 987, 1606, 1572]. Condensing [161, 160]. Condition [1167]. Conditional [1408, 1409, 974, 36]. Conditions [393]. Conference [1335, 1577]. Conferences [1653]. Configurable [1276, 1268]. Configuration [752, 1635, 235, 830, 1268]. Configurations [1582, 695]. Conflicts [833, 428, 264]. Conformance [1028]. Congruence [1027]. Congruences [156]. Conjunctive [70]. Connected [442, 844]. Connection [1095, 1604]. Connections [1585, 704, 544]. Connectors [1591]. Cons [1639]. Consequences [1087]. Conservative [974]. Considerations [1223, 1044]. Consistency [176, 175, 177, 433, 1379, 148]. Consistent [1257, 1087]. Constant [791]. Constrained [1148, 1141, 873]. Constraint [237, 352, 494, 236, 80, 95, 493, 1137, 960]. Constraint-Based [237]. Constraints [238, 239, 1117, 70]. Construct [1119]. Constructing [375, 282, 314, 597, 661]. Construction [480, 425, 125, 315, 937]. Constructions [1362, 582]. Constructor [956]. Constructs [965]. Consumer [1200]. Container [1058]. Container-Based [1058]. Contemporary [537]. Content [304, 1292, 260, 317, 349, 323]. Content-Addressable [1292]. Content-Based [304, 317, 349]. Contents [211, 322]. Context [1589, 1237, 1648]. Continuous [922, 1189, 318, 673, 1247, 745, 680, 681, 746, 682, 747, 969]. Contract [772]. Contracting [775]. Contracts [539]. Contributors [210]. Control [730, 6, 1239, 664, 1560, 351, 256, 1170, 1489, 255, 672, 726, 1234, 1146, 740, 728, 1238, 951, 679, 268, 101, 680, 1353, 1163, 1164, 50, 685, 687, 688]. Control-Synchronous [1489]. Controll [680]. Controllable [1423, 1575]. Controlled [51, 55, 479, 514, 673, 1203]. Controlled-NOT [51, 55]. Controller [735, 669, 684, 1190]. Controllers [748]. Controlling [219]. Convergence [799, 276]. Convergent [1378]. Converse [1252]. Conversion [1048, 942, 361, 1128]. Convex [717, 894]. Convolutional [41, 1427]. Cooperating [1540]. Cooperation [6, 1181]. Cooperative [1476, 1171, 1172, 313, 281, 1173]. Coordinated [1565]. Coordinating [1584, 774, 1662]. Coordination [1063, 1592, 1571, 1570, 1562, 1568, 1588, 1589, 1590, 1560, 1579, 152, 1586, 1663, 1569, 1567, 1576, 1581, 1666, 1578, 1197, 336, 1587, 1577, 1572, 1563, 1664]. Cope [528]. Copy [220, 1256, 266]. Copy-Hybrid [220]. Copying [32]. Copyright [519]. CORBA [1623, 1617, 228, 1616, 268, 1443, 1628]. CORBA/TC [1616]. Correction [41, 43, 1144]. Correlated [43, 1185]. Correlating [244]. Correlation [512, 1427, 19, 1331]. Correlations [392, 19]. Coscheduling [1316]. Cost [510, 95, 840, 821, 598, 599, 523, 597]. Costs [607, 770]. Counterexamples [787]. Countermeasures [1347]. Counting [1069]. Cover [622]. Coverage [132]. Covering [802]. CoW [439]. COWs [1497]. CPS [963]. CPU [1190]. Crashworthiness [822]. Cray [385]. Creation [1633, 1638, 1634, 322, 1639]. Creativity [558]. Crew [1436]. Criterion [1219]. Critical [725]. Cross [552].

Cross-Cultural [552]. Crossbar [1263]. Crowds [324]. Cryptanalysis [1404, 381, 1403, 379, 375, 1428, 365, 371, 377]. Cryptographic [1017, 540]. Cryptography [1417, 30, 536, 518, 31, 541, 191, 542, 189, 190]. Cryptoprocessor [523]. Cryptosystem [372, 1410, 371, 1346, 1356]. Cryptosystems [1424, 1354, 524, 1419, 1420, 1415, 369, 529]. CSP [1324, 1327]. CSP/FDR [1324]. Cube [991]. Cubes [252, 1536, 390, 696]. Cultural [552]. Current [1048, 647, 271]. Curry [1082]. Curvature [692]. Curve [717, 692, 702, 1417, 1416, 1354, 370, 369]. Curves [719, 1338]. Customizable [1282]. Customization [1625]. Customizing [1626]. Cut [1088, 457, 458]. Cut-Elimination [1088]. Cuts [456]. Cyclic [470]. Cycling [1343].

D [121, 708, 691, 419, 307, 417, 1061, 124, 703, 714, 1453, 123, 1454, 724, 818, 105, 1555, 125]. Dame [475]. Darwinian [1145]. Data [245, 856, 1550, 309, 1493, 290, 571, 1317, 483, 86, 85, 1388, 950, 247, 1384, 1045, 1549, 1547, 1494, 876, 246, 404, 1070, 252, 1536, 1507, 159, 283, 986, 896, 254, 831, 1524, 835, 136, 911, 96, 265, 905, 767, 261, 244, 908, 285, 1211, 1050, 671, 1061, 259, 844, 1452, 857, 257, 1554, 910, 213, 854, 1258, 287, 979, 662, 1492, 291, 68, 123, 855, 1510, 1375, 1084, 262, 1043, 253, 637, 944, 292, 248, 927, 888, 873, 639, 640, 251, 872, 277, 1177, 1522, 1217, 1038, 878, 272, 305, 1498, 319, 315, 1159, 884, 874, 913]. Data [914]. Data-Flow [986, 1510]. Data-Mining [913, 914]. Data-Parallel [1550]. Database [91, 24, 1292, 1047, 628, 1506, 1383, 1234, 635, 281, 292, 440, 925, 279, 270, 891, 310, 845, 272, 897, 444]. Databases [304, 316, 1352, 923, 887, 303, 1387, 629, 630, 1116, 243, 1382, 72, 80, 875, 863, 79, 907, 69, 505, 877, 242, 636, 76, 912, 293, 1032, 1056, 890, 78, 924, 865].

Davidson [808]. DBMS [224, 1508]. Dead [229]. Deadlock [1023, 1496]. Deal [1038]. Dealing [709, 1430]. Debugging [238, 239, 505, 928]. Decentralized [336]. Decidability [73, 584, 1127]. Decidable [950, 739]. Decision [110, 467, 527, 256, 255, 1531, 343, 872, 482]. Decision-Making [110]. Declarations [496]. Declarative [480, 226, 240, 235, 241]. Decoders [1188]. Decoherence [43, 54, 50]. Decompilation [962]. Decomposed [1219]. Decomposing [691]. Decomposition [816, 419, 424, 417, 1540, 345]. Decoration [1071]. Decorrelation [355]. Decrement [604]. DEDICA [525]. Dedicated [1254]. Deduction [243, 1370, 1083]. Deductive [243, 505, 225, 482]. Deep [655]. DEFACTO [1260]. Default [892]. Defence [460]. Definability [1074, 72, 1121]. Definite [979, 1100]. Definition [66, 1308, 1216, 866]. Definitions [657, 1073, 1053]. Deformable [709]. Degree [831, 1419, 1302]. Delay [1276, 135]. Delay-Intensive [1276]. Delegating [195]. Delegation [201, 202, 196, 197, 199, 200, 516, 203]. Deletions [90]. Delivery [260, 261, 263, 262]. Demand [444, 1601]. Demonstrating [10]. Demonstration [9, 1049]. Denotational [62]. Dense [455]. Density [897]. Density-Based [897]. Dependability [1239]. Dependant [839]. dependencies [1097]. Dependency [404, 953, 925]. Dependent [356, 1602, 1420, 798, 229]. Deploying [1573]. Deployment [1635, 1249]. Depth [477]. Derivational [1094]. Derivations [1364]. Derive [233]. Derived [397, 398]. Deriving [626]. Describe [977]. Describing [950, 58]. Description [69, 317]. Descriptions [1636]. Descriptive [579, 72, 609, 1113]. Descriptors [1041]. DESEO [119]. Desiderata [1662]. Design [624, 948, 1033, 249, 1054, 1528, 1228, 1260, 1491, 254, 670, 557, 1264, 672, 451, 273, 131, 1330, 1616, 1372, 743, 744, 770, 1009, 941, 240, 1661, 1628, 845, 21, 749, 132]. Design-Space [1264]. Designing [1191, 1057]. Desktop [298]. Detailed [1399]. Detect [264]. Detection [1070, 887, 829, 1150, 119, 122]. Determine [839]. Determinism [1192]. Deterministic [1313, 1002, 677, 958, 1124, 1369]. Developing [67, 1287, 235]. Development [725, 216, 1195, 763, 1636, 976, 484, 984, 1628, 891]. Developmental [553]. Devices [1545, 1269]. DFA [1031]. DFA-Frameworks [1031]. DFA-Generators [1031]. DHeartView [1529]. Diagnosis [751, 670, 754]. Diagnostic [1530]. Diagnostics [1011]. Diagram [718, 937]. Diakoptic [29]. Dictation [301]. Difference [463, 611, 791]. Differences [204, 363, 919]. Different [709, 701, 919]. Differential [364, 393, 1422, 375, 363, 291, 524, 454]. Differentials [1404]. Diffie [380]. Diffraction [721, 1455]. Diffusion [389, 1108]. Digital [1493, 697, 707, 691, 698, 695, 1277, 700, 1554, 1062, 321, 1451, 312]. Digitization [719]. Digitizer [703]. Dilemma [1178, 1181]. Dimension [7, 702, 80]. Dimensional [1529, 40, 75, 314, 305]. Dimensions [709, 82]. Dipole [56]. Dipole-Dipole [56]. Direct [1460, 812, 792]. directed [1093]. Directional [714]. Directions [753, 646, 8, 751, 752, 209, 1622, 750]. Directly [864]. Directory [1631]. DIS [1443]. Disassembly [93]. Discernibility [882]. disciplinary [1507]. Discipulus [1273]. Discovered [880, 888]. Discoveries [878]. Discovering [867, 875, 907, 92, 919]. Discovery [918, 631, 243, 915, 910, 854, 242, 912, 495, 921, 881, 890]. Discrete [721, 706, 709, 692, 719, 1424, 701, 743, 744, 679, 680, 723, 704, 724, 568, 693]. Discrete-Log [1424]. Discrete-Time [679].

Discretely [673]. Discretization

[922, 817, 720]. Discussion

[187, 185, 176, 173, 182, 193, 179, 194, 202, 197, 180, 183, 177, 188, 204, 208, 195, 200, 198, 174, 203, 191, 206]. Disjunctions [94, 792]. Disk [1290]. Display [217, 298]. Displays [298]. Dissection [1302]. Dissemination [309]. Distance [901, 424]. Distance-Hereditary [424]. Distances [657, 602]. Distributed [1225, 571, 411, 446, 1511, 1613, 1228, 1543, 1035, 1588, 1239, 1280, 429, 1476, 1240, 1248, 1494, 1303, 1236, 418, 1610, 1515, 1392, 1649, 410, 1424, 467, 911, 1313, 1553, 1211, 1057, 1660, 1556, 1234, 1442, 1452, 1475, 1062, 313, 224, 518, 821, 1207, 1247, 1426, 1238, 1256, 516, 1205, 1639, 268, 1557, 912, 1229, 1439, 383, 281, 1463, 1587, 1219, 1462, 1509, 308, 1329, 826]. Distributed-Memory [429]. Distributing [1244]. Distribution [23, 1456, 1635, 1258, 299, 801, 1209, 798, 1498]. Distribution-Dependent [798]. Distributions [608]. Diverse [919]. Divertible [1339]. Divisible [449]. DMKD [926]. DNA [1174, 1217]. do [340]. Doacross [1503]. Document [1175, 875, 218, 282, 1578]. Document-Based [1578]. Documents [630, 1391]. Domain [816, 1269, 494, 877, 651, 482]. Domain-Decomposition [816]. Domain-Specific [482]. Domains [832]. Domino [120]. Don’t [1188]. DOT [1652]. DOT/MAT [1652]. Dots [55]. Double [46, 1364, 523]. Double-Size [523]. Down [993, 892]. Downward [591]. Dreams [1598]. Drifts [1187]. Driven [505, 981, 1577]. DSM [1483, 1258, 1257]. DSM-Runtime [1258]. DSMs [1552]. DSP [1277]. Dual [507]. Duopolistic [1200]. Duplication [1497]. Duration [129]. During [107]. DVIZ [905]. DYANA [1033]. Dynamic [62, 447, 1248, 1309, 838, 601, 822, 975, 275, 261, 1245, 603, 83, 871, 1473, 1146, 967, 313, 1233, 742, 743, 1492, 1602, 262, 1463, 100, 1478, 1477, 345, 921, 232]. Dynamical [678]. Dynamically [1271, 1272]. Dynamics [816, 1490, 1506, 726, 56, 1463, 16]. Dynamite [1467]. Dynamite-Blasting [1467]. DyPERS [100].

E-Commerce [515]. E-mail [1016]. E/R [250]. EARL [1487]. Earth [1450, 1306, 1044, 850]. Easier [332]. Easily [1030]. Easy [1327]. Eavesdropping [31]. ECLiPSe [241]. Ecology [1152]. ECOPLA [167]. Eden [1598]. Edge [1061]. EDIFACT [525]. Editing [297, 765]. Editor [167]. Education [1149, 1046, 1531, 634, 764, 826]. Effect [54]. Effective [1311, 1501, 298, 1116, 336, 1477, 138]. Effectiveness [521]. Effects [269, 20, 1605]. Efficiency [817, 1657]. Efficient [899, 1253, 721, 1251, 1350, 1435, 1410, 1425, 1536, 1469, 1484, 789, 285, 361, 1599, 907, 507, 1354, 257, 879, 840, 1271, 1302, 1281, 455, 1257, 1552, 1203, 862, 860, 305, 937]. Effort [1232]. Egyptian [627]. Ehrenfeucht [1122]. Ehrenfeucht-Fraiesse [1122]. Eigenproblem [805]. Eigenproblems [809]. Eigenvalue [804, 810, 806, 811, 28, 808]. ELAN [1369]. Elastic [1451]. Election [1573]. Electromagnetic [1539]. Electron [46]. Electron-Nuclear [46]. Electronic [1350, 768, 769, 779, 773, 367, 1475, 1535, 1349, 913, 914]. Electronically [539]. Elegance [240]. Element [1478, 400]. Elementary [1066, 1133]. Elements [1056]. Eli [928]. Eli-Generated [928]. Eliminating [619]. Elimination [443, 1138, 1070, 934, 1088, 229]. Elite [1189]. Elliptic [1537, 1417, 1416, 1354, 370, 1338, 369]. Elliptical [53]. Embedded [989, 1033, 1328, 1184, 1567, 830]. Embedding [491]. Embodiment [549, 548]. Emergence [1526]. Emergent [320, 339, 561]. Emerging [536]. Empirical [560]. Enabled [1642]. Encoder [442, 412]. Encoders [1188]. Encrypted [511]. Encryption [1435, 358, 1408, 1409, 510, 1341, 1342, 366, 514]. End [1324, 1209]. End-to-End [1324]. Endgame [478, 1158]. Engine [947]. Engineering [625, 216, 1507, 254, 1154, 1036, 750, 969]. English [479, 556]. Enhance [510]. Enhanced [100]. Enlarged [388]. Ensemble [1015]. Ensembles [1433]. Entanglement [14, 17, 34]. Enterprise [216, 1494, 1658, 853, 633, 1654]. Entropy [36, 803]. Enumeration [1331]. Environment [1033, 1260, 1240, 1623, 1658, 1515, 1241, 307, 1287, 1609, 1625, 1553, 1637, 1061, 1561, 1645, 1475, 1455, 1062, 324, 888, 153, 548, 1551, 315]. Environments [1518, 418, 1474, 261, 1533, 1202, 1247, 262, 1578, 1145, 1086, 1235, 1659]. EO [1040]. EO/GIS [1040]. Epidemiological [767]. Epistemic [918, 338]. Equalities [1087]. Equational [137, 999, 145, 70]. Equations [389, 393, 854, 454, 912, 585, 1163, 1164, 1087]. Equilibria [605]. equilibrium [39]. Equipped [552]. Equivalence [1321, 1092, 1001]. Equivalences [1027, 156]. Equivalent [164, 527, 522]. ER’97 [643]. Erratic [1567]. Error [574, 41, 43, 800]. Errors [496]. Escrow [518, 1333, 529]. Establishment [1334, 1233]. Estimating [299, 475]. Estimation [410, 800, 407, 1556, 920, 28, 454, 1508, 406]. eta [1128]. eta-conversion [1128]. Ethernet [441, 442]. Ethernet-Connected [442]. ETI [1037]. Euclidean [657]. European [1641]. Evaluated [765]. Evaluating [524, 225, 1106]. Evaluation [441, 848, 464, 1239, 359, 1243, 936, 1246, 1645, 758, 981, 1231, 663, 775, 1657, 1206, 1319]. Evaluator [431, 432]. Event [990, 1018, 61, 1565, 916, 292, 1577, 1487]. Event-Based [61]. Event-Driven [1577]. Every [1071]. Evolution [1199, 1179, 1180, 1608, 1306, 1193, 1151, 1161, 1221, 1040, 1186, 978, 909, 1142, 1197, 1145, 279, 1351, 1192]. Evolution-Based [1221]. Evolutionary [1148, 1195, 1171, 1172, 1153, 1191, 1176, 1170, 582, 1160, 1174, 1146, 1154, 776, 1141, 1169, 1181, 1177, 1218, 878, 1194, 1173, 1190]. Evolvable [1273]. Evolving [1171, 1172, 1155, 1158]. Exact [788]. Example [1452, 1375, 344, 1522]. Examples [1322, 680]. Exceptions [879]. Executable [972, 214]. Execution [1561, 1208, 433, 1082]. Exhaustive [503]. Exhaustive-Search [503]. Existence [1128, 42]. Existential [1084]. Existing [1651]. Expanding [991]. Expansion [933]. Expansion-Based [933]. Experience [109, 688]. Experiences [1047, 761, 767, 766, 764]. Experimental [1514, 1292, 551, 669, 931, 1278]. Experimentation [1249]. Experiments [386, 1316, 1475]. Expert [763, 793]. Explicit [602, 998, 977, 388, 1085, 1086, 1606]. Explicitly [1075]. Exploitation [944]. Exploiting [1512]. Exploration [1034, 1264, 904]. Exploring [551, 878]. Exponent [521]. Exponential [594]. Exponentiation [1429, 523, 1279]. Expressing [965]. Expressions [790, 84, 958, 1030, 663, 1056]. Expressive [961]. Expressiveness [1570, 76]. Extended [667, 1440, 312, 668, 472]. Extendible [167]. Extending [945, 591, 861, 250]. Extensible [1488, 1030, 118, 1487]. Extension [974, 1575]. Extensions [1060]. External [176, 175, 177, 593]. Extract [915]. Extracted [301]. Extracting [1389, 675, 877, 325, 292]. Extraction [1317, 1183, 858].

Face [106]. Face-Tracking [106]. Facial [451, 325]. Facility [1063]. Facing [820]. Factories [974]. Factoring [522]. Factorisation [828]. Factorization [455, 1458, 1459]. Factorizing [164]. Factors [362]. Fail [1338]. Fail-Stop [1338]. Failed [495]. Failure [960]. Fair [513, 518, 1351]. Farey [696]. Fast [441, 594, 418, 307, 15, 243, 1427, 375, 1468, 373, 1160, 1416, 1014, 713, 371, 516, 453, 1058, 900, 897]. Fast-Track [1014]. Faster [929, 369, 529]. Fault [139, 1239, 670, 40, 1245, 1221, 1222, 524, 1574, 1284]. Fault-Tolerant [139, 40, 1245, 1221, 1222, 1284]. FDR [1324]. FEAST [823, 824]. Feature [304, 654, 1193, 871, 325, 1053, 1041, 1099]. Features [464, 1177, 132]. FEC [269]. Federated [630]. Federation [1646, 1647]. Feedback [306]. Feistel [355]. FEM [1450, 821]. FFT [450]. FGS [910]. Fiber [1231]. fibring [1099]. Fiedler [1301]. Field [361, 1160, 507, 828, 1344]. Fieldbus [1235]. Fiffo [138]. Fight [913, 914]. Fighting [1012]. Figure [104]. Figure-Ground [104]. Figures [554]. File [258, 836]. Fill [583]. Fill-In [583]. Filtering [704]. Finding [598, 599, 1391, 1500]. Fine [1318]. Fingerprinting [509, 1414, 191, 189, 190]. Finite [1461, 1364, 1073, 494, 361, 674, 1024, 1122, 1478, 400, 1114]. Finite-Domain [494]. Finite-Size-Particle [1461]. Firewalls [511]. First [849, 73, 463, 459, 1115, 79, 1139, 854, 71, 644]. First-Order [73, 1115, 79, 1139, 71]. Fish [778]. Fishtank [321]. Fixed [942, 1010, 1131]. Fixed-Point [1131]. Fixed-Priority [1010]. Fixpoint [1384, 616, 617]. FL [648]. Fleet [763]. Flexible [1652, 1474, 1660, 1208, 1205, 1204, 232, 406]. Flight [735]. Floating [942]. Flow [820, 1541, 734, 986, 951, 1510, 952, 1217, 1194, 660]. Flowfield [1445]. Flows [729, 1460]. Fluid [1541, 816, 16]. FLYSIG [1276]. FO [90]. Focal [1242]. Fock [397, 398]. Folding [1605]. For-Loops [71]. Forces [546]. Forecasting [1201]. Forecasts [876, 1449]. Forget [615, 614]. Forging [544]. Fork [77]. ForkLight [1489]. Formal [57, 970, 1323, 973, 1320, 910, 1429, 64, 1216]. Formatting [218]. formedness [1101]. Forms [1128, 1360]. Formula [74, 1445, 570]. formulae [659]. Formulas [786, 140]. Fortran [1502]. Forward [315]. Foundation [1396]. Foundations [502, 501, 560]. Founded [1068]. Four [1544, 450]. Four-Step [450]. FPGA [1263, 1155, 1262, 1279]. FPGAs [373, 1278]. FR [1417]. Fractal [408, 409, 1271, 413]. Fraiesse [1122]. Framework [11, 1358, 970, 1642, 91, 1323, 1076, 1212, 1198, 1643, 1466, 1670, 1029, 1660, 1663, 1308, 124, 314, 1255, 1206, 232, 105, 642, 272]. Frameworks [1031, 484]. Franklin [1434]. Fraud [207, 208, 1150]. Free [569, 694, 30, 927]. Free-Space [30]. Frequencies [1552]. Frequent [92]. Friendly [66]. FRIL [652]. Frobenius [1416]. Front [1445, 1209]. Front-End [1209]. Frontiers [705]. Fudgets [1567]. Fuel [1541]. Full [570]. Fully [713]. Function [594, 697, 17, 1433, 1330, 519, 1183]. Functional [162, 233, 1593, 217, 1595, 949, 1213, 213, 925, 1130, 1060, 1356, 931]. Functionality [1383]. Functionals [1068]. Functions [464, 91, 1161, 572, 375, 1331, 1426, 1405, 453, 1602, 540, 1332, 613, 1058, 526, 376]. Functorial [1363]. Fundamental [700]. Fundamentals [650, 1394]. Further [622]. Fusion [671, 658, 655]. Future [753, 627, 646, 8, 751, 298, 752, 276, 209, 750, 1166, 1607]. Fuzz [1167]. Fuzzy [1138, 654, 657, 653, 649, 1112, 242, 662, 294, 663, 889, 1168, 1664, 1481, 648, 660]. Fuzzy-Neural [660].

G [1603]. G-Machine [1603]. GA [1175, 1226, 1156]. GA-based [1226]. GAIA [551]. Galley [218]. Game [179, 1179, 1180, 1072, 178, 180, 469, 1182, 468, 1181, 600, 1664, 476]. Games [470, 1178, 1122]. Gamma [1566]. Gang [1519]. Gang-Scheduling [1519]. Gap [328]. Garbage [1325, 1525]. Gas [1471, 1469]. Gasdynamics [1460]. Gate [51, 1344]. Gates [55]. Gateway [1616]. Gateways [1628]. GCR [1542]. Gene [1197]. General [1528, 485, 607, 1395, 1265, 1277, 590, 1421]. Generalised [1343]. Generality [1168]. Generalization [490, 489, 36, 802, 801]. Generalized [929, 23, 1092, 808]. Generate [552]. Generated [215, 1002, 928]. Generating [1516, 1241, 1301, 1163, 1164]. Generation [392, 1424, 75, 324, 655, 528, 860]. Generators [1523, 360, 1031]. Generic [114, 1445]. Generous [1178]. Genetic [1220, 1225, 1157, 654, 852, 1179, 1180, 246, 1189, 1143, 1196, 1150, 1170, 1152, 1201, 1186, 1222, 1187, 1144, 1227, 1219, 1163, 1164, 1147, 1159]. Genetic-Based [246, 1222]. Genome [636]. Genomes [233]. Genomic [765]. Genus [566, 567, 565]. Geo [896]. Geo-referenced [896]. Geodesic [657]. GeoFEM [1450]. Geographic [656]. Geographical [275]. Geometric [1518, 733, 783, 728, 1052]. Geometry [1066, 1507, 719, 427]. Geometry-Aided [427]. Geoprocessing [1062]. GEOPROVE [1518]. Geospatial [1042, 1050, 1053, 1058]. GF [370]. GHI [459]. GHZ [45]. GIS [1049, 1040]. GIScience [1046]. Glasgow [1597]. Global [1042, 626, 1444, 88, 438, 632, 1465]. Glossary [12]. Go [474, 476, 471, 465, 473, 466, 472, 475, 462]. Goal [758, 344]. Goal-Question-Metric [758]. Graceful [694]. Gradient [1107]. Gradient-Based [1107]. Grail [541]. Grained [1312, 1318]. Grammar [57, 1090, 1094]. Grammars [990, 1098, 1092, 1097, 1100]. Grammatical [1195]. Grande [1523]. Granular [857]. Granules [654]. Graph [705, 990, 1364, 1269, 1469, 1366, 1367, 1368, 1301, 908, 426, 1120, 907, 1295, 910, 982]. Graph-Based [907]. Graph-Structures [982]. Graphic [973, 167, 1050]. Graphics [307, 101]. Graphs [583, 424, 603, 1482, 589, 90, 590, 659]. Grasping [128]. Greedy [1178]. Grid [717, 850]. Gridless [1461]. Grids [1300, 1255]. Gries [980]. Grothendieck [1074]. Ground [104]. Grounds [1055]. Group [1584, 244, 278, 1014, 871, 1495, 700, 1339, 1346, 279, 1349, 1356]. Group-Oriented [1346]. Groups [762, 1205, 475]. Groupware [279]. Grover [23]. Growth [1534]. Guided [1372, 685, 498].

Haiku [545]. Halo [1302]. Halos [1305]. Handles [1041]. Handling [1491]. Handy [1370]. Hanoi [600]. Hardness [785, 573]. Hardware [1274, 1028, 1029, 417, 1267, 1272, 153]. Hardware-Algorithms [1267]. Hardware-Based [1272]. Hardware/Software [153]. Hardwired [1263]. Hardwired-Clusters [1263]. Harmonic [1310]. Harnessing [1388]. Hartree [397, 398]. Hash [1412, 1405, 540, 526]. Hash-and-Sign [1412]. Haskell [1597, 219, 1604]. Haskell/Java [1604]. HaskellScript [227]. Hausdorff [720]. HealthLine [1532]. Heap [469]. Heart [1530]. Heian [478]. HEIR [342]. Helicopter [735]. Hellman [380]. Hennessy [988]. Hereditary [424].

Heretical [648]. Hermitian [809]. Heterogeneous [447, 1476, 1248, 1365, 1171, 1172, 260, 437, 1456, 1056, 1038, 865]. Heuristic [1311, 244, 678, 1278]. Heuristics [460, 880]. Hidden [1362, 612, 28]. Hierarchical [1018, 1365, 115, 267, 1263, 407, 979, 342, 1168, 982, 1572]. Hierarchically [1476]. Hierarchies [425, 831, 490, 489, 426]. Hierarchy [963, 529]. High [1317, 849, 1282, 1491, 1292, 1285, 1296, 1450, 836, 434, 1464, 1270, 727, 1455, 1451, 1202, 1548, 1373, 843, 1436, 1448, 1552, 440, 1441, 1522, 1286, 305]. High-Level [727, 1202, 1373, 1286]. High-Performance [1317, 1491, 1464]. High-Speed [1292]. Higher [1149, 146, 40, 157, 1416, 363, 1601, 964]. Higher-Dimensional [40]. Higher-Order [146, 157, 964]. Highly [1257]. Hill [1330]. Himalayan [1045]. Hindu [1045]. Histograms [1491]. Historical [285]. Histories [433]. History [318]. HOL [980, 1013]. Holes [1603]. Holy [541]. Homeomorphic [491]. Homogeneous [1332]. Homotopic [709, 715]. Honourable [848]. Hop [1294]. Hopfield [1184]. Horizon [679]. Horn [611, 790]. Horticultural [99]. Hot [878]. Hour [107]. Howard [1082]. HPC [1328, 1453]. HPCN [1442]. HPF [1305, 1490, 395]. HTMT [1306]. HTTP [193, 192, 194, 1590]. HTTP/1.1 [1590]. Hull [717]. Hulls [894]. Human [559, 7, 112, 891, 1173]. Human-Computer [559]. Humanoid [547, 553]. Hybrid [729, 731, 733, 1165, 1291, 1290, 664, 665, 734, 666, 667, 220, 1143, 671, 672, 726, 674, 739, 740, 1246, 742, 741, 676, 743, 677, 679, 342, 668, 683, 274, 748, 1218, 685, 125]. Hybridized [1196]. Hybridizing [1302]. Hydrodynamics [828]. Hypercube [852, 449, 407]. Hyperelliptic [1415]. Hypermedia [312]. Hyperplanes [695]. Hyperrobust [795]. Hypertext [1387, 1385]. Hysteresis [737].

I/O [1546, 1548, 860, 429]. I/O-Intensive [1546]. ID [1042]. ID3 [651]. IDEA [374]. Idealised [1005]. Idealized [961]. Identification [654, 1428, 274]. Identifying [1056]. Identities [1041]. Identity [1354]. IEEE1394 [435, 436]. IEEE1394-1995 [435, 436]. II [124]. Illicit [913, 914]. Illumination [299]. Illustrating [1322]. Im [1407]. Image [411, 708, 1226, 306, 404, 710, 303, 410, 1242, 408, 409, 102, 1553, 422, 116, 451, 1270, 313, 712, 403, 1271, 402, 843, 658, 1104, 109, 416, 305]. Image-Based [1104]. Imagers [122]. Images [423, 321, 1451, 1530, 1454]. Imaging [1455]. Imitation [553]. Immersion [326]. Immune [1215, 1331]. Impact [1439]. Impartial [470]. Impediments [634]. Imperative [500]. Implementation [1348, 1592, 55, 1282, 418, 1047, 1536, 1400, 1395, 669, 434, 1198, 807, 373, 1211, 273, 1271, 864, 1447, 1043, 939, 1603, 1284, 1575, 1227, 1319, 1279, 1542, 241, 1217]. Implementations [164, 47, 1523, 844]. Implemented [550, 1455]. Implementing [290, 226, 1598, 1213, 1587, 1502, 982, 288]. Implications [989, 478, 1081, 335]. Implicit [975, 389]. Implicitly [807]. Importance [870]. Impossible [1404]. Improve [944]. Improved [1265, 1427, 370, 622, 866]. Improvement [932, 2, 837]. Improvements [1433]. Improving [162, 121, 1457, 491, 1249, 872, 893]. IN-TINA [1625]. Inaccuracies [54]. Including [1348]. Inclusion [881]. Incomplete [863]. Increasing [76, 353]. Increment [604]. Incremental [245, 867, 90, 1385, 862, 1496, 874]. Incrementalization [967]. Indefinite [320]. Independent [737, 715]. Index [290, 532, 563, 780, 1102, 1357, 1673, 1402, 210, 75, 83, 84, 305]. Indexing [302, 318, 286, 319]. Indirect [792].

Individually [1207]. Induction [1324, 242, 651, 894, 889, 868]. Inductive [1134, 173, 486, 1073, 487, 172, 174]. Inductive-Recursive [1073]. Inductive/Coinductive [1134]. Industrial [849, 1545, 1239, 1544]. Inequality [39]. Inertial [819]. Inference [146, 1600]. Inferring [493]. Infinite [470, 1100]. Influence [817, 222, 221, 834]. Information [35, 37, 233, 1042, 1431, 626, 32, 628, 1007, 779, 88, 656, 1667, 1668, 886, 632, 1669, 871, 36, 645, 636, 555, 68, 658, 39, 952, 1532, 1650, 251, 266, 329, 969, 38, 1147, 870]. Information-Passing [329]. Information-Theoretic [32, 870]. Infrastructure [1045, 771, 1643, 1040, 1651]. Infrastructures [184, 525, 969]. Inhabitants [1069]. Inherent [393]. Inheritance [1145]. Inhomogeneous [1543]. Initial [381, 23]. Initialization [529]. Ink [506]. Inner [17]. Innovations [1527]. Input [1255]. Input/Output [1255]. Insertions [90]. Insider [207, 208]. Inspired [1214]. Instance [1223]. Instance-Based [1223]. Instant [1590]. Instantiation [496]. Institution [1371]. Institutions [165]. Instruction [835]. Instructions [64]. Integer [1196, 935, 1436, 362]. Integral [1539]. Integrate [630]. Integrated [1264, 1445, 1040, 760, 1532, 105]. Integrating [1007, 745, 1019, 97, 109, 300, 1052]. Integration [756, 86, 85, 110, 779, 1670, 1638, 1669, 743, 744, 391, 111]. Integrative [773]. Integrator [846, 847]. Integrity [1117]. Intelligence [1611, 560, 1649, 1216]. Intelligent [1444, 1658, 549, 755, 630, 1215, 1174, 1614, 342, 1651, 354, 661]. Intensive [1276, 1546, 1452, 1254]. Intention [331]. Inter [404, 833, 452, 1616, 866]. Inter-procedural [452]. Inter-processor [404]. Inter-transaction [866]. Inter-vector-Conflicts [833].

Inter-working [1616]. Interaction [559, 1066, 549, 1398, 551, 298, 448, 278, 348, 112, 638, 642]. Interaction-Oriented [638]. Interactions [1016, 954]. Interactive [1593, 1382, 815, 113, 321, 322, 100, 1441]. Interchange [1488, 1583]. Interconnect [1262]. Interconnection [838, 384]. Interconnects [1296]. Interdisciplinary [629]. Interesting [904, 878]. Interestingness [880]. Interface [1283, 1047, 557, 1513, 877, 1058, 655]. Interfaces [1523]. Interfacing [1596]. Interiors [125]. Intermediate [225]. Internal [1274]. International [774, 518]. Internet [173, 1380, 1653, 1444, 1516, 1473, 172, 174, 1578, 1651]. Internet-Based [1578]. Interoperability [1054, 1048, 1046, 1040, 1059, 1055, 525]. Interoperable [1047, 1049, 1210]. Interoperation [957]. Interpolation [141, 390, 661]. Interpretation [1136, 1586, 1530, 163]. Interpretations [1081]. Interpreter [505, 503]. Interpreting [166]. Interprocedural [986, 931, 930, 951]. Interprocessor [1510]. Intersection [698, 794]. Intersection-Closed [794]. Interval [589, 687]. Interworking [1617, 1615]. intonation [1096]. Intra [349]. Intra-Agent [349]. Intractability [512]. Intranet [1670]. Introducing [1529, 1585]. Introduction [534, 1652, 557, 1103]. Intuitionistic [1077, 1083]. Invalidation [263]. Invalidation-Based [263]. Invariance [664]. Invariant [734, 672, 1121, 495]. InvenTcl [307]. Inverse [291]. Inversion [1337]. Inverted [1170, 1163, 1164]. Investigation [91, 963, 1152]. Invisible [300]. Invited [309, 304, 804, 816, 327, 689, 690, 298, 701, 716, 313, 326, 322, 843, 704, 323, 845, 826]. Involving [1241]. Ion [53, 52]. IP [1240, 1643, 271]. IQ [1216]. Irregular [1305, 1307, 429, 404, 1457, 1301, 1482, 426, 844, 1548]. Irregularly [1309, 1299]. Irrelevant [506, 788]. ISA [842]. Isabelle [980, 1013]. Isabelle/HOL [980, 1013]. IsaWin [984]. Island [1156, 1187]. Islands [1225]. Isolation [670]. Isomorphism [1082]. Isotropic [713]. ISPW [9]. Issues [756, 67, 249, 647, 8, 1400, 1395, 399, 1626, 1619, 1318, 939, 1622, 1036, 968]. Itemsets [92]. Iterated [25, 1181, 1421]. Iteration [979]. Iterative [1468, 1510]. IWIM [1588]. IWIM-Based [1588].

J [1642]. Jacobi [806, 808]. Jacopini [1087]. Jam [1500]. Japanese [556]. Java [1642, 62, 57, 1283, 1307, 66, 1549, 1280, 61, 1523, 973, 58, 1399, 1524, 65, 1525, 452, 959, 1013, 64, 1604, 1284, 1575, 1209, 59, 60, 1522]. Java-Based [1284]. Java-Targeted [973]. JavaPorts [1287]. Java^{T M} [67]. JIAJIA [1483, 1552]. Join [77, 440, 89]. Joint [1153, 553]. Jordan [698]. JTN [973]. Jumps [738]. Just [479].

K2 [1592]. KD [910]. KD-FGS [910]. KDCs [1426]. KDD [853]. Kernel [453]. Kernels [803]. Key [380, 517, 184, 1403, 372, 1410, 8, 510, 1424, 1643, 356, 536, 373, 538, 1334, 1335, 518, 1430, 371, 1419, 1420, 525, 191, 357, 366, 1333, 189, 190]. Key-Dependent [356]. Key-Search [373]. Key-Share [517]. KeyNote [184]. KGK [478]. kit [768, 769]. Knowledge [756, 1317, 1411, 535, 751, 761, 895, 631, 243, 752, 911, 114, 880, 767, 762, 907, 757, 633, 915, 877, 910, 1655, 242, 651, 1339, 108, 1432, 764, 750, 892, 891]. Knowledge-Based [751, 761, 752, 114, 767, 108, 764]. Kolmogorov [608, 819]. Kos [472]. Kripke [1127]. Kush [1045]. Kutta [388].

Labeling [883]. Labs [183, 182, 181].

Lagrange [1141, 390]. Lamarckian [1145]. Lambda [1074, 807, 219, 1071, 1075, 1081]. lambda-Calculus [1075, 1081]. lambdaeta [1072]. lambdaeta-Calculus [1072]. Lambek [1092]. Laminar [820]. Landscape [1152]. Language [162, 1063, 948, 1592, 652, 1488, 1588, 214, 350, 627, 1210, 479, 226, 1595, 102, 554, 1489, 965, 1213, 1561, 1564, 877, 1567, 1580, 1587, 1577, 1575, 225, 235, 957, 1101, 1093]. Languages [1571, 996, 3, 1387, 351, 1002, 346, 1429, 1574, 943]. Large [245, 571, 809, 896, 243, 811, 815, 1506, 1525, 907, 1451, 864, 344, 416, 89, 897]. Large-Scale [811, 1506]. Late [970]. Latency [1315, 1318]. Lattice [372, 521, 1471, 1469, 867, 1468, 860]. Lattice-Based [372, 860]. Lattice-Boltzmann [1468]. Lattice-Gas [1471]. Law [27]. Layered [420, 339, 740, 503]. Layers [1015]. Layouts [589]. Lazy [898]. LC [1078]. Leakage [528]. Learnability [786]. Learner [906]. Learners [783]. Learning [1199, 922, 463, 607, 787, 1153, 1612, 788, 792, 789, 1161, 796, 113, 1167, 790, 794, 795, 606, 782, 1166, 791, 1183, 869]. Left [1459]. Left-Right [1459]. Legacy [968]. Legged [1273]. Lemmas [161, 160]. Lens [399]. Leonardo [1273]. Less [1403]. Level [1460, 1035, 835, 397, 398, 1237, 727, 1512, 1371, 1202, 1373, 939, 1286, 1484]. Leveraging [783]. LGen [860]. Libraries [1499, 395, 1256]. Library [1493, 290, 396, 1469, 1479, 1256, 1502]. Life [549, 754, 320]. Life-Like [549, 320]. Lifecycle [9]. Light [1029, 597]. Light-Weight [1029]. Lighting [697]. Lightwave [1294]. Like [549, 1178, 320, 365]. Limit [1364]. Limited [1333]. Limits [1413, 835]. Lindblad [39]. Lindstroem [1119]. Line [717, 1275, 898, 770, 944, 1349, 1659, 792]. Linear [49, 814, 784, 665, 142, 667, 813, 812, 1077, 738, 935, 1456, 375, 578, 1080, 143, 454, 1083, 590, 383, 668, 1113, 134, 749, 1097]. Lined [312]. Lines [1144]. Linguistic [1093, 888, 1183]. Linguistics [1095]. Link [932]. Link-Time [932]. Linked [950]. Linker [1501]. Links [657]. Linux [1252]. Lithography [1447]. Live [321]. Livelock [139, 1012]. Load [1479, 822, 449, 437, 1201, 1463, 1478, 1477]. Load-Balancing [1463]. Loaded [1552]. Local [971, 709, 695, 238, 239, 936, 453]. Locality [831, 1457, 1546, 426, 1223]. Localization [123]. Localizing [456]. Locate [1184]. Locating [1646, 1647]. Location [267]. Locking [955, 256, 255]. Log [1424]. LOGCFL [609]. Logic [1065, 988, 162, 989, 1138, 73, 233, 618, 137, 950, 486, 488, 492, 918, 142, 649, 338, 479, 1116, 236, 1112, 487, 1023, 149, 493, 796, 1635, 1277, 484, 215, 1117, 141, 1095, 616, 617, 71, 143, 130, 1374, 960, 959, 1083, 1084, 1376, 1158, 1377, 481, 144, 889, 222, 221, 496, 1056, 1088, 504, 232, 1481, 648, 1096, 1097]. Logic-Automaton [1095]. Logical [1134, 500, 1359, 129, 1074, 1076, 1077, 1098, 578, 1001, 1217]. Logics [1361, 1119, 1000, 1370, 69, 1139, 1004, 330, 1099]. Logopedics [662]. LogP [1311]. Long [392]. Long-Range [392]. Looking [1459]. Loop [939, 1605]. Loop-Level [939]. Loops [71, 496, 1503]. Lossy [1021, 596]. Lost [87]. LOTOS [1375]. Low [521, 1315, 840, 1513, 821, 523, 406]. Low-Cost [523]. Low-Exponent [521]. Low-Latency [1315]. Lower [799, 1406, 602, 601, 1113, 767]. Lower-Saxony [767]. LR [929]. LU [385]. Lubrication [1545]. Lying [698].

MAC [526]. Machine [373, 1080, 64, 1603, 60]. Machine-Checking [60]. Machines [1220, 571, 852, 839, 446, 1311, 429, 1241, 1313, 578, 1463]. Macintosh [1347]. Macro [1002]. MaDCoWS [1515]. Magic [506, 243]. Maglev [688]. MAGNET [775]. Magnetised [851]. mail [1016]. Maintain [259]. Maintaining [907]. Maintenance [95, 257, 104, 90, 1385, 126]. Majorant [1125]. Majorant-Computability [1125]. Making [110, 467, 332, 1182, 883]. Management [185, 182, 1656, 184, 1621, 1653, 1547, 1610, 1632, 181, 183, 836, 1643, 1609, 267, 763, 278, 632, 1464, 1485, 1627, 633, 727, 1655, 287, 1654, 291, 281, 1577, 1641, 1657, 1441, 940, 405, 270, 1607, 1329, 660, 1629]. Manager [1628]. Managing [290, 1057]. Manifold [1061]. Manifolds [797]. Manipulation [115, 1241]. Manipulations [356, 850]. Manipulator [128]. Manufacturing [237]. Many [1139, 600, 902]. Many-Valued [1139]. Map [708, 1416, 883]. Mapper [1054]. Mapping [1220, 1269, 428]. Mappings [556]. Maps [883, 98]. Marching [1536]. Margin [801]. MARINE [1632]. Marked [584]. Market [1039, 771, 1200, 778]. Marketplaces [768, 769]. MarketSpace [771]. MAS [1659]. Mass [650]. Masses [1285]. Massive [410]. Massively [1515]. MAT [1652]. Matching [423, 527, 407, 712, 1451, 1003, 590, 623, 659, 406]. Materialize [95]. Materialized [94]. Mathematical [502, 731, 501, 1101, 1030]. Matrices [828]. Matrix [1457, 387, 1302]. Matrix-Vector [1457]. Maurer [359]. Max [588]. Maximum [364, 590]. May [996]. MBASE [625]. MCM [433]. MCORBA [228]. Meaning [562, 555]. Meaningful [81]. Means [899, 659]. Measure [971, 903, 1216]. Measurement [31]. Measurements [1503]. Measures [917, 870]. Measuring [1240, 711]. Mechanical [1325]. Mechanics [546]. Mechanism [1246, 266]. Mechanisms [553]. Media [318, 1247, 322]. Median [1143]. Mediated [773, 1671]. Mediation [1038, 89]. Mediator [1057, 1058]. Medical [1536, 1150, 1452, 1455, 1526, 1527, 1480]. Meet [552]. Meeting [754]. Megamodules [1580]. Memories [1457]. Memory [756, 1199, 571, 446, 1547, 429, 418, 945, 1515, 831, 428, 1594, 1313, 1525, 417, 757, 433, 758, 1256, 1492, 413, 1463, 1459, 1204, 940, 1509]. Mental [339]. Mention [848]. MEPHISTO [118]. Mercury [228, 480]. Mere [1006]. Merging [945]. Mesh [1303, 1265, 1245, 1266, 1538, 384, 825]. Meshes [694, 427]. Message [1618, 511, 278, 513, 1003, 1606, 1235, 1487, 1496]. Message-Passing [1496]. Messages [1513, 913, 914]. Messaging [1590]. Meta [245, 1585, 4, 1544, 254, 167, 505, 503, 1534]. Meta-Applications [1544]. Meta-Computing [1534]. Meta-Editor [167]. Meta-Interpreter [505, 503]. Meta-Mining [245]. Meta-Object [254]. Meta-Process [4]. Meta-Programming [1585]. Metacomputing [1474, 1466, 1472, 1286]. Metaheuristic [1218]. Metalanguage [134]. MetaML [961]. Metaphor [559, 550, 298, 558, 561, 552, 548, 546]. Metaphor-Based [550]. Metaphorizing [548]. Metaphorizing-in [548]. Metaphors [1228, 543, 562, 546]. Metasystems [1521]. Method [839, 887, 1171, 1172, 393, 1150, 807, 1301, 244, 1184, 1539, 1160, 699, 1470, 1455, 242, 1141, 1271, 651, 291, 454, 1375, 503, 281, 299, 1462, 872, 892, 1194, 1542]. Methodologies [347]. Methodology [480, 13, 1326, 830]. Methods [139, 806, 243, 408, 409, 1320, 1330, 286, 388, 64, 937]. Metric [710, 901, 758]. Metrics [1243, 909]. MG [1605]. MHPCC [1449]. Microprocessors [1277]. Migration [1253, 1575]. Migratory [1247]. MILAN [1520]. Milieu [662]. Milky [545]. Milner [988]. MIMD [418, 1539, 417]. Mimics [1152]. Mind [796]. Mini [515]. Mini-Cash [515]. Minima [1312]. Minimal [1123, 906, 728]. minimalist [1094]. Minimalistic [515]. Minimum [583, 510, 1302]. Mining [245, 856, 923, 1494, 246, 859, 895, 1524, 911, 905, 244, 857, 1492, 68, 855, 916, 864, 253, 873, 925, 862, 1522, 891, 878, 924, 865, 860, 874, 913, 914]. Minors [1120]. Missing [920, 890, 902]. Mission [683]. Mix [367]. Mix-Based [367]. Mixed [903]. Mixing [589]. Mixins [63]. ML [992]. MMIPPS [422]. Mnesia [224]. Mobile [1621, 1623, 1658, 1589, 258, 1560, 1632, 1630, 260, 999, 265, 271, 128, 1160, 1638, 259, 1234, 1334, 276, 273, 1645, 264, 1564, 1576, 268, 1574, 1622, 1657, 266, 1671, 78, 272]. Mobiles [274]. Mobility [1632, 1648, 1644, 1629]. MobiS [1564]. Modal [666, 140, 1533, 141, 616, 617, 1004]. Mode [1230, 676]. Model [1199, 559, 337, 707, 1020, 1528, 625, 596, 1562, 918, 306, 709, 906, 146, 1022, 280, 1306, 1023, 114, 1000, 815, 1215, 348, 465, 595, 675, 1440, 898, 742, 676, 1480, 743, 1187, 778, 885, 317, 952, 250, 1448, 1534, 251, 138, 890, 1661, 1572, 126, 660, 1100]. Model-Based [625, 898]. Model-Checking [146]. Modelica [946]. Modeling [625, 627, 647, 646, 667, 630, 631, 412, 1457, 558, 1178, 1050, 643, 727, 634, 645, 635, 636, 112, 1453, 1581, 644, 293, 668, 638, 1168, 323, 641, 642, 400]. Modelling [1529, 560, 817, 3, 1, 1398, 1484, 1006, 1627, 633, 1062, 1655, 123, 1162, 284, 1551]. Models [304, 654, 1570, 1543, 142, 628, 283, 895, 74, 1191, 998, 926, 157, 1626, 1326, 433, 328, 1202, 909, 76, 143, 1353, 1204, 858, 125]. Modern [832]. Modes [807, 496]. Modification [1336]. Modular [66, 1035, 1195, 523, 1377, 118, 1279, 1628]. Modularization [1650]. Module [1062]. Modules [1065, 1367]. Modulo [945]. Molecular [1460, 1490, 1506, 1463]. Molecule [1566]. Moments [453, 693]. Momentum [1468]. Monadic [1596, 1120]. Monitoring [1557]. Monoids [586, 587]. Monomials [606]. Monotone [1131]. Montages [948]. Monte [1460, 393, 391].

Monte-Carlo [393]. Moral [334].

Morphisms [1132]. Morphological [844]. Morphology [657]. MorphoSys [1270]. Mortals [1006]. Motion [410, 407, 294, 325, 219, 556, 406, 315]. Motion-Based [294]. MOV [1417]. Movable [1163, 1164]. Movies [444]. Movies-on-Demand [444]. Moving [1359, 1062, 287, 286, 324, 122]. MPEG [442, 1642, 412]. MPEG-2 [442]. MPEG-4 [1642, 412]. MPEG-J [1642]. MPI [1283, 1514, 434, 439, 846, 1447, 847, 1319]. MPI-Based [1447]. MPI-parallelized [439]. mpiJava [1283]. MSO [1095]. Mu [1004]. Mu-Calculus [1004]. Multi [37, 1541, 334, 1656, 457, 458, 983, 216, 14, 61, 1292, 1363, 1294, 1507, 1471, 338, 805, 774, 1484, 1670, 1182, 1237, 1533, 1263, 1669, 1662, 1663, 1645, 1666, 336, 1587, 1262, 1661, 329, 1555, 1664]. Multi-access [37]. Multi-Agent [334, 338, 1670, 1669, 1662, 1663, 1666, 336, 1661, 329, 1664, 1656, 216, 774, 1587]. Multi-algebras [1363]. Multi-block [1541]. Multi-cut [457, 458]. Multi-disciplinary [1507]. Multi-Epistemic [338]. Multi-FPGA [1263, 1262]. Multi-Hop [1294]. Multi-level [1484]. Multi-Level-Context [1237]. Multi-OPS [1294]. Multi-paradigm [983]. Multi-particle [14]. Multi-port [1555]. Multi-retailer [1645]. Multi-species [1471]. Multi-stage [1182]. Multi-sweep [805]. Multi-Threaded [61]. Multi-Wavelength [1292]. Multiagent [337, 1655, 638, 1672, 1612]. Multicast [1435, 1643, 1243, 1495, 1257]. Multicast-Based [1257]. Multicomputer [1485]. Multicomputers [419, 408, 409, 1312]. Multidimensional [1491, 250, 866]. Multiform [1156]. Multigraph [566, 567, 565]. Multihop [1297, 1486]. Multilevel [1537, 111, 825]. Multimedia [304, 1653, 1230, 1640, 1397, 1392, 307, 1636, 374, 645, 1238, 1249, 401, 323, 310]. Multimodal [551]. Multiparadigm [1031]. Multiparty [1425]. Multiple [1157, 626, 876, 392, 231, 230, 1184, 1221, 671, 433, 1233, 1141, 496, 472, 869]. Multiplicity [787]. Multiplier [1141]. Multipliers [1268]. Multiprocessor [1594, 407, 1222, 430, 405]. Multiprocessors [1291, 1515, 419, 1510, 1459]. Multiresolution [691, 689, 690, 403]. Multispectral [422]. Multitemporal [422]. Multitheaded [1306]. Multithreaded [1251]. Multitude [298]. Multiuser [1261]. MuSE [1208]. Mutations [1189]. Mutual [1199, 303, 871]. MW [1292]. MW-OCAPP [1292]. Myrinet [434]. Myrtle [505].

NAG [1479]. Naive [893, 696]. Naming [1631]. Narratives [629]. Narrowing [1368]. NAS [1605]. National [1655, 1043]. Natural [877, 123, 171, 1083, 1140, 1093]. Navigation [275, 1160, 98, 99]. NAVL [308]. NChiql [877]. Near [1391]. Near-Replicas [1391]. Nearest [81, 1150, 901, 1151]. Necessary [1430]. Negative [974]. Negatively [1185]. negligible [1510]. Negotiating [776]. Negotiation [773, 1663]. Neighbor [81, 1151]. Neighborhood [1153]. Neighbour [1150, 901]. Neighbourhood [618]. Nested [252, 1302]. Nestor [1502]. NetCache [1291]. Nets [1023, 672, 1078, 1373, 696]. Network [785, 1656, 447, 1291, 1290, 1610, 303, 1196, 1609, 1184, 886, 1482, 1201, 760, 1513, 1655, 1654, 466, 1651, 1657, 1235, 1628, 270, 884]. Network/Cache [1291]. Network/Write [1290]. Networked [1441]. Networking [1559, 1650]. Networks [604, 1317, 47, 269, 358, 1314, 1195, 1294, 669, 838, 1649, 810, 448, 1293, 1297, 572, 1170, 1456, 1485, 1185, 1222, 19, 1154, 1233, 1486, 916, 268, 1197, 1532, 384, 1183, 1573, 1052, 111, 1286, 1289, 18, 1505]. Neural [785, 1195, 887, 669, 1153, 1196, 1193, 1170, 886, 1482, 1201, 760, 1185, 1154, 885, 466, 1183, 1188, 660, 884]. Neuro [662]. Neuro-Fuzzy [662]. Neuronal [561]. Neuroprocessor [1275]. Neville [443]. News [301, 311]. Newscasts [310]. Newsgroups [96]. Newspapers [310]. Next [655, 827]. NFS [1348]. Niching [1184]. NMR [45]. No [993, 1618, 738]. Non [1065, 364, 1340, 1595, 1422, 1237, 83, 1369, 1213, 1061, 342, 1510, 1083, 383, 39, 266, 1254, 1298, 1351]. Non-averaged [364]. Non-commutative [1065, 1083]. Non-Dedicated [1254]. Non-deterministic [1369]. Non-equilibrium [39]. Non-Manifold [1061]. Non-negligible [1510]. Non-pipelined [1298]. Non-preemptive [1237]. Non-replicating [83]. Non-repudiation [1340, 1351]. Non-standard [383]. Non-storage [266]. Non-Strict [1595, 1213]. Non-XOR [1422]. Nonadditive [42]. Noncommutative [612]. Nondeterministic [997, 1116, 431, 432, 578]. Nonempty [1130]. Nonhomomorphicity [376]. Nonlinear [1196, 679, 748, 688]. Nonlinearity [1423]. Nonmonotonic [1116]. Nonparametric [799]. Nonsymmetric [813]. Noosa [928]. NOPE [431, 432]. Norm [333]. Normal [1128, 1360, 724]. Normalisation [1088]. Normalization [1129, 1131, 1378]. Not-So [201, 202]. Notation [973, 240, 1286]. Note [1413]. Notebook [1475]. Nothing [526]. Notification [1565]. Notions [706]. Novel [233, 1191, 68]. NP [592, 214, 1120, 1137, 26]. NP-complete [26]. Nuclear [46, 807]. Number [839, 1411, 1523, 392, 698, 360, 475, 528]. Numbers [569, 1430, 802, 803]. Numeric [859, 854]. Numerical [446, 395, 394, 393, 1540, 251, 661]. Numerics [382]. Nurse [1165]. Nursing [1531]. Nvestigators [205, 206]. NWCache [1290].

O [429, 1548, 860]. O-Intensive [1546]. O.O [500]. OASIS [1666]. OBDDs [613]. OBJ3 [1371]. Object [444, 1321, 1352, 1283, 652, 1511, 1042, 1228, 306, 396, 127, 254, 722, 128, 1506, 1201, 712, 1205, 1281, 225, 1032, 1462, 440, 943, 105, 845]. Object-Oriented [444, 1321, 1283, 652, 1506, 1201, 225, 1032, 1462, 440, 943]. Objects [1623, 709, 1584, 1399, 283, 318, 1057, 1001, 1207, 1565, 287, 286, 1576, 64, 122]. Oblivious [1408, 1407, 1409, 1406, 509]. Observation [320, 1027, 1044, 118]. Observational [149, 1377]. Observations [381]. Observed [802]. Observers [742, 743]. Obstacles [1467]. Obtained [503]. OCAPP [1292]. Ocean [1448]. ODMG [845]. Off [1349]. Off-Line [1349]. Office [298]. Oil [1418]. OLAP [252]. Oligopoly [1179, 1180]. Omission [475]. On-Line [1275, 898, 770, 792]. On-Stack [1208]. Once [613]. Oncology [766]. One [386, 574, 602, 1433, 512, 1430]. One-Bit [1430]. One-Sided [386, 574]. One-Way [1433]. One-Wayness [512]. Online [1037, 1007, 490, 489, 491, 623]. Only [1257]. onto [1510]. OO [994]. OOD [484]. Open [1613, 1610, 782]. OpenGIS [1048, 1051, 1039, 1049]. OpenMap [1047]. OPERA [1442]. Operating [1261]. Operation [561]. Operational [61, 963, 1076, 1595, 65, 912, 1448, 1100]. Operations [604, 1514, 1205, 1231, 1217]. operative [1193]. Operator [1147]. Operators [705, 844, 803]. Ophthalmic [399]. OPS [1294]. Optical [1290, 1292, 1296, 48, 1293, 1297, 1295, 50, 1289]. Optics [49, 1288]. Optimal [730, 839, 1091, 33, 692, 460, 1293, 267, 575, 83, 116, 620, 1355, 1510, 1124, 384, 576, 684, 1194, 1500]. Optimising [901, 825]. Optimistic [265, 205, 206]. Optimization [1189, 935, 1223, 1141, 678, 343, 1436, 1219, 1224]. Optimizations [1550, 1323, 1501]. Optimize [1169, 1498]. Optimizer [1446]. Optimizing [1305, 1290, 429, 399, 102, 943, 89]. Or-Parallelism [223]. Oracle [1412]. Oracles [527, 166]. Order [73, 146, 833, 1115, 157, 79, 1139, 363, 71, 1601, 715, 964]. Ordered [920]. Ordering [1302]. Ordinary [454]. ordination [1574, 660]. Organisation [555]. Organization [571, 1612]. Organizational [756, 757, 758]. Organized [303]. Organizing [883]. Oriented [444, 187, 1321, 1352, 1283, 652, 533, 186, 188, 347, 1506, 1201, 505, 242, 297, 1346, 225, 638, 1032, 830, 1462, 440, 943]. Origin2000 [841]. Orthogonal [732, 749]. Orthogonally [67]. Orthographic [80]. ORYX [377]. Other [346, 720]. OTIS [1294]. OTIS-Based [1294]. Outdoor [99]. Outliers [900]. Output [1255]. Overhead [1513]. Overlap [814]. Overlapped [450]. Overlapping [1511, 1538]. Overview [182, 1393, 181, 183, 1212, 1520, 741]. Owicki [980]. Owicki/Gries [980].

p [68, 1121]. P/poly [1121]. Pacific [1519]. Package [823, 824, 1491, 422]. Packages [741]. Packet [418, 419]. Pairs [164, 244, 90]. Pan [1641]. Pan-European [1641]. Panel [209]. Pangaea [1209]. Panmictic [1225]. Paper [309, 304, 849, 192, 184, 848, 201, 196, 689, 178, 181, 690, 175, 298, 186, 207, 701, 199, 716, 313, 322, 172, 704, 323, 205, 189, 290, 285, 282, 294, 284]. PARADEIS [396]. Paradigm [10, 1079, 250, 983]. Paradigms [419, 346]. Paradox [1188]. Parallel [441, 1550, 624, 442, 1220, 1225, 1157, 445, 820, 1545, 814, 592, 385, 1541, 852, 839, 816, 425, 809, 446, 386, 1307, 435, 436, 1543, 1280, 1547, 1476, 1226, 1593, 1303, 1292, 817, 418, 414, 53, 396, 1515, 147, 1469, 1366, 392, 829, 1579, 1241, 1479, 838, 1446, 1287, 1450, 813, 836, 1524, 806, 423, 812, 389, 1595, 577, 811, 807, 1497, 1156, 382, 424, 815, 1553, 422, 1437, 1520, 1489, 1597, 837, 1470, 844, 451, 713, 1326, 1320, 403, 1312, 821, 1202, 1480, 1187, 1548, 808, 1492, 1250, 454, 846, 415, 402, 388, 864, 832, 1447]. Parallel [1231, 1229, 1438, 1454, 1538, 383, 455, 390, 1463, 850, 1459, 391, 1478, 1227, 1552, 847, 222, 221, 819, 830, 1224, 440, 1299, 1254, 1508, 1467, 1218, 818, 1542, 845, 826, 400, 406, 1481, 1505, 1214]. Parallel/Distributed [1476]. Parallelisation [848, 421]. Parallelism [1321, 831, 410, 835, 393, 1512, 223, 405]. Parallelization [1460, 1461, 1490, 397, 398, 452, 939, 1310, 1458, 1503]. parallelized [439]. Parallelizing [1504, 840, 1448]. Parameter [1543, 1354]. Parameterisation [1361]. Parameters [1091, 448, 1146, 325, 222, 221]. Parametric [686]. Parametrized [428]. ParaPART [1303]. Parse [552]. Parsers [1599]. Parsing [929, 966, 1093, 1100]. Parsons [1126]. Part [757]. Partial [393, 1263, 933, 1665, 716, 148, 362]. Partial-Crossbar [1263]. Particle [1461, 1462, 14]. Partitioned [443]. Partitioning [899, 1303, 1469, 800, 1264, 1482, 427, 832, 153, 1203, 825]. Partitionings [82]. Partitions [1360, 1132]. Party [193, 192, 194]. Passing [146, 1027, 1606, 329, 1487, 1496]. Passive [1293]. Past [1166, 827]. PastSet [1509]. Patches [719]. Path [735, 1034, 930, 84, 70, 118]. Path-Conjunctive [70]. Paths [698, 598, 599, 90]. Patient [901]. Pattern [799, 1504, 465, 1447, 659]. Patterns [1389, 887, 1309, 711, 1626, 888]. Pay [531, 530]. Pay-per-View [531, 530]. Paying [539]. Payment [1671]. Payments [367]. PC [441, 1317, 411, 1464, 1513, 1492, 864].

PC-Based [1317]. PC-Cluster [411]. PCF [1070]. PCP [584]. PCS [1645, 1204]. Pedagogical [551]. Pediatric [766]. Peer [265]. Peer-to-Peer [265]. Peg [600]. PEI [1304]. Pendulum [1170, 1163, 1164]. Penna [1440]. People [107, 919]. PEP [168]. Perception [114]. Perception-Action [114]. Perceptron [1428]. Performance [1317, 849, 1592, 971, 1282, 1491, 1240, 1285, 1315, 829, 1450, 836, 434, 397, 398, 1252, 1464, 1645, 1270, 1455, 1451, 1480, 1405, 76, 1548, 981, 1492, 1486, 864, 843, 1538, 1436, 802, 353, 222, 221, 1415, 1206, 1503, 440, 893, 1441, 841, 1522, 288, 685]. Performances [1490, 448]. Period [1009, 1218]. Periodic [1591, 1244]. Permanent [573]. Permutation [358, 1336, 522, 1289]. Permutations [1433]. Permuted [1428]. Persistent [67, 1042, 1381]. Person [562, 300, 1182, 1181]. Personal [1314, 1631, 100, 1648]. Personalized [1667, 1668]. Perspective [1608, 1590, 629, 165, 855, 497]. Perspectives [635]. Peter [1126]. Petri [1023, 672, 1373]. Phase [688, 1558]. Phase-Space [688]. Phenomena [741]. Phenomenon [1216]. philosophical [1093]. phrase [1097]. Physical [27, 47]. Pi [1024, 156]. Pi-Calculus [1024]. Pi-Congruences [156]. PIC [1470]. Piecewise [738, 749]. Ping [1016]. Ping-Pong [1016]. Pipe [734]. Pipelined [1295, 1298]. Pipelining [1566, 937]. Pivots [385]. PJama [67]. Place [649]. Plan [1665]. Planar [79]. Plane [1242]. Planes [694, 696]. Planning [735, 766, 499, 683, 118, 462]. Plans [332]. Plant [761, 1162]. Plasma [1461, 851]. Plastic [1272]. Platform [67, 1618, 1465, 1641]. Platforms [260]. Plausible [881]. Play [460, 101, 1059]. Playing [120]. Plug [1059]. Pluggable [1062]. Plumber [1566]. PM2 [1253]. PMD [287]. PMD-Tree [287]. Point [942, 1131]. Points [717, 325, 1650]. Polarized [1078]. Policies [187, 186, 188]. poly [1121]. Polygonal [734, 736]. Polyhedra [732]. Polyhedral [729, 734]. Polyhedral-Invariant [734]. Polyhedrization [722]. Polylogarithmic [1431]. Polymer [1437]. Polymorphism [1075]. Polynomial [1136, 997, 1339]. Polynomial-Time [997]. Polytypic [966]. Pong [1016]. port [1555]. Portable [1256, 1575]. Portrait [387]. Pos [164]. Posedness [738]. Position [192, 184, 201, 196, 178, 181, 175, 186, 207, 199, 172, 205, 189]. Positions [1158]. Positive [974]. Positive/Negative [974]. Possibilistic [659]. Possibility [658, 1407]. Possible [730, 632, 475]. Post [488]. Power [848, 1115, 1194]. Practical [1608, 30, 375, 1480, 758]. Practice [533, 1048, 1064, 1638, 634]. Practice-Oriented [533]. Pragmatic [546, 1093]. Pragmatics [481]. Pre [872, 868]. Pre-processing [872, 868]. Preassigned [366]. Preconditioned [809, 1542]. Preconditioners [813]. Preconditioning [816]. Predicate [664, 1056]. Predicates [1077]. Prediction [1232, 926, 1185, 1154, 921, 858]. Prediction-Based [1232]. Predictions [793]. Predictive [1665]. preemptive [1237]. Preface [212]. Prefetching [1554]. Prefix [572, 1267]. Preliminary [1292, 1492, 778]. PREMO [1399, 1394, 1393, 1392]. Presence [1590]. Present [1166]. Presentation [1533, 700]. Presentations [1050]. Presented [371]. Preservation [1153]. Preserve [1085, 888]. Preserving [516]. Pressure [934]. Pretty [218]. Pretty-Printing [218]. Preventing [496]. Prevention [1496]. Primality [538]. Primes [1411, 1343, 538]. Primitive [453]. Principal [1069, 797]. Principle [854]. Principled [957]. Principles [1653, 13].

Printed [1007]. Printing [966, 218]. Priority [1010]. Priors [909]. Prisoner [1178, 1181]. Privacy [516, 1349]. Privacy-Protecting [1349]. Private [1403, 1431]. Prize [849]. Probabilistic [574, 1139, 1004]. Probabilities [1422]. Probability [364]. Probes [1518]. Probing [911]. Problem [849, 1545, 995, 459, 1165, 667, 1064, 1143, 1196, 702, 806, 901, 586, 587, 1428, 673, 760, 28, 335, 668, 1227, 1218]. Problem-Solving [335]. Problems [804, 785, 580, 214, 1048, 1148, 1269, 352, 829, 601, 810, 22, 811, 1023, 1457, 1301, 815, 1456, 1295, 808, 1420, 623, 248, 1278, 1224, 1299, 26, 1214]. procedural [452]. Procedures [482]. Proceeding [1054]. Process [9, 7, 1592, 4, 1025, 972, 3, 8, 1, 13, 1309, 2, 280, 1497, 1186, 198, 19, 978, 680, 1005, 148, 900, 270, 1286]. Process-Algebraic [978]. Process-Based [280]. Processes [993, 918, 146, 136, 595, 1027, 153, 400, 18]. Processing [411, 1292, 1242, 421, 102, 1497, 1553, 422, 1264, 1277, 116, 451, 401, 916, 402, 263, 1438, 1439, 655, 109, 416, 406, 1505, 1214, 872, 868]. Processor [1292, 831, 65, 289, 572, 1270, 1512, 1235, 1481, 404]. Processor-and-Memory [831]. Processors [392, 810, 811, 1221, 842]. Product [1411, 17, 1457, 235, 1289, 1659]. Product-Line-Architecture [1659]. PROFIBUS [1235]. Profiler [1597]. Profiling [1593, 930]. Program [502, 480, 501, 486, 154, 1322, 1501, 829, 490, 489, 487, 397, 398, 1211, 431, 432, 215, 984, 1082, 499, 466, 570, 1203]. Programmable [947, 1344, 1487]. Programmer [66, 63]. Programming [989, 1585, 652, 654, 233, 731, 1549, 1381, 159, 419, 1171, 1172, 1196, 236, 949, 935, 1489, 1160, 1126, 1213, 1146, 967, 1117, 1326, 1320, 346, 1202, 1250, 959, 1169, 1204, 1219, 889, 1163, 1164, 222, 221, 638, 1194, 1159].

Programming-Based [1160]. Programs [1550, 483, 500, 1504, 486, 488, 492, 485, 1568, 1593, 932, 147, 1212, 477, 487, 1023, 493, 796, 1221, 452, 1510, 960, 1162, 1158, 481, 503, 952, 613, 496, 163, 109, 504, 232, 1487, 1496, 1498, 498]. Progress [634]. Project [1493, 182, 547, 181, 183, 1520, 1530, 525]. Projections [736]. Projects [326]. Prolog [163]. PROMIS [941]. Promoting [1565]. Proof [1321, 145, 1078, 620, 1339, 499, 495, 1378, 1096]. Proof-Nets [1078]. proof-theoretic [1096]. Proofs [1015, 238, 239, 1432]. Propagation [1601, 1278]. Properties [1322, 1153, 999, 136, 1125, 672, 965, 79, 1154, 1019, 1011, 568]. Property [426, 1332, 526]. PROPLANT [761]. Proposal [286]. Proposition [1579, 274]. Pros [1639]. Protecting [1349]. Protection [761, 1336]. Protocol [1021, 173, 517, 1618, 1315, 1012, 1483, 1028, 1234, 172, 174]. Protocols [985, 176, 380, 1324, 535, 175, 177, 1586, 199, 200, 1334, 1017, 1565, 476]. Prototype [1049, 114, 1316, 765]. Prototypes [294]. Prototyping [948, 307, 1637, 215]. Provable [533, 358]. Prover [1032]. Prover-Based [1032]. Providing [1574]. Proving [1411, 58, 984, 1013, 1019, 59, 60]. Provision [1652, 1532]. Provisioning [368]. Proxy [414, 1340, 514, 415, 1379]. Pruning [457, 458]. PSEE [6, 10]. PSEEs [11]. Pseudo [306, 1426]. Pseudo-random [1426]. Pseudorandom [1433]. Public [184, 372, 1410, 510, 536, 538, 371, 1419, 1420, 525, 191, 189, 190]. Public-Key [184, 372, 510, 536, 1419, 191, 189, 190]. Publicly [1333]. Puglia [1493]. Pure [1128, 161, 160, 1089]. Purpose [607, 1277]. Push [1079, 263]. Putting [1603]. PVM [448]. PVS [1025]. Python [946].

QoS [1238, 1329]. Quadratic [585].

Quality [1244, 1520]. Quantified [786, 663]. Quantifier [1138]. Quantifiers [1118, 1115]. Quantifying [1153]. Quantitative [1067, 1176, 917]. Quantum [49, 35, 37, 51, 55, 46, 47, 33, 29, 32, 41, 24, 17, 53, 43, 612, 15, 48, 40, 22, 30, 52, 20, 36, 19, 56, 28, 54, 44, 25, 31, 42, 39, 541, 38, 21, 16, 18, 26]. Quasi [391]. Quasi-Monte [391]. Qubits [33, 50]. Queries [94, 592, 73, 91, 591, 263, 70, 89]. Query [784, 24, 1382, 226, 256, 289, 255, 440, 1505]. Querying [1384, 88]. Question [758, 1434]. Questions [782]. Queue [1558, 438]. Quorum [514]. Quotienting [953].

R [1049, 167, 264, 250]. R-W [264]. Radar [423, 1438]. Radial [1148]. Radiance [439]. Railway [107]. Raised [67]. Random [784, 1523, 508, 392, 1412, 360, 528, 1426]. Randomized [571, 582, 613]. Range [392, 83, 790, 1312, 123]. Ranking [880]. Rasiowa [1370]. Rasiowa-Sikorski [1370]. Raster [1060]. Rate [799, 406]. Rates [1145]. Ratio [825]. Ray [723]. Ray-Casting [723]. Rays [577]. RBAC [1348]. RC4 [365]. RC4-like [365]. RC5 [378]. Re [754, 1276, 514, 1565, 1106]. Re-Configurable [1276]. Re-encryption [514]. Re-evaluating [1106]. Re-usability [1565]. Re-use [754]. Reachability [736, 1023, 745, 1011]. Reaching [1308]. Reaction [113]. Reactive [117, 973, 352, 949, 131, 978, 101, 138, 497]. Reactor [807]. Read [993, 613]. Read-Once [613]. Read-Up [993]. Real [411, 1582, 971, 137, 129, 1232, 1010, 1242, 754, 901, 1636, 420, 1237, 1234, 1442, 1246, 1233, 1247, 104, 1231, 1229, 1438, 1009, 292, 299, 345, 830, 103, 300, 1441, 105, 78]. Real-Life [754]. Real-Time [411, 1582, 971, 137, 129, 1232, 1010, 1242, 1636, 420, 1237, 1234, 1442, 1246, 1233, 1247, 104, 1231, 1229, 1438, 1009, 345, 830, 103, 1441, 105]. Real-World [901]. Realisation [1470].

Realistic [205, 206]. Reality [35, 1516, 853, 1598, 326, 100, 308]. Reallocation [772]. Reals [1125]. Reapplication [240]. Reason [332]. Reasoning [550, 753, 653, 656, 987]. Receding [679]. Receding-Horizon [679]. Recognition [799, 1504, 127, 119, 128, 1665, 712, 1177, 659]. Recognizability [995, 138]. Recommendation [311]. Recomposition [311]. Reconfigurable [117, 1275, 1269, 1261, 1259, 1265, 102, 1277, 1295, 1270, 1266, 1271, 1272, 402, 1262, 827]. Reconfiguration [1014]. Reconsideration [331]. Reconsidered [326, 331]. Reconstruction [702, 701, 124, 1454]. Record [1130]. Recoverable [529]. Recovery [513]. Recruitment [546]. Rectangle [923, 924]. Rectangular [82, 745, 686]. Rectilinear [427]. Recurrent [1482, 18]. Recursion [988, 350, 619]. Recursive [425, 1073, 1207, 453, 944]. Reduce [833]. Reduced [1404, 566, 1167, 567, 565]. Reducing [43, 428]. Reduction [63]. Reductions [1406, 1417]. Redundancies [933]. Redundancy [934, 266]. Refer [639, 640]. Reference [945, 1151, 1416, 675, 1650, 1329]. referenced [896]. References [12, 1401, 1308]. Refinement [147, 746, 747]. Refinements [1367]. Reflection [1144, 1374]. Reflections [648]. Refutations [1123]. Region [1045]. Regions [388]. Register [934, 936]. Registers [33, 834]. Registration [1226]. Registry [767]. Regular [832]. Regularity [1223]. Regularized [797]. Reinforcement [1167, 782, 1166]. Related [859, 762, 586, 587]. Relating [570]. Relational [1292, 505, 242, 925]. Relations [1134, 1389, 1360, 1074, 1339, 144, 1060, 476]. Relationship [512, 346, 330, 648]. Relationships [923, 1057, 69, 924].

Relaxation [1468]. Release [1408, 1409]. Relevance [306, 456, 1651, 1349]. Relevant [923, 918, 1127, 924]. Reliability [848]. Reliable [1315, 879]. Remainder [900]. Remote [1493, 1241, 1497, 1256, 1492]. Removal [933]. Removing [525]. Renaming [485]. Rendering [1536, 1398, 1104]. Renovation [974, 968]. Replacement [1379]. Replay [1496]. Replicas [1391]. Replicated [76]. replicating [83]. Replication [265, 267]. Report [778, 644]. Repository [1549]. Representation [708, 732, 296, 295, 689, 690, 289, 699, 1061, 712, 1060]. representational [1094]. Representations [283]. Representatives [1495]. Representing [341]. Repudiation [1340, 1351]. Requirements [754, 215]. Requires [1055]. Research [309, 647, 52, 276, 634, 473, 474, 248, 968]. Resemblance [711]. Residuosity [1419]. Resistance [1421]. Resistant [1413]. Resolution [1123, 621, 1448]. Resolving [819]. Resonance [26]. Resonances [46]. Resonant [1552]. Resource [1656, 237, 1464, 1654, 1081, 1521, 353, 430, 1329]. Resources [1581, 266]. Respect [488, 831]. Response [1162, 1508]. Responsibility [196, 197]. Restarted [807]. Restricted [790]. Restructuring [226, 1546]. Results [1592, 785, 463, 1514, 1315, 1118, 669, 243, 586, 587, 915, 166, 1492, 1439, 1508]. Resumable [1603]. Retail [773]. Retailer [1648, 1645]. Retransmission [1021]. Retrieval [756, 304, 1611, 1431, 306, 303, 1667, 1668, 1174, 451, 68, 317]. Retrieving [1667, 1668]. Retrograde [478]. Return [240]. Reusability [1018]. Reusable [1274, 755, 1469]. Reuse [1624, 1627]. Reveal [883]. Revision [1117, 329]. Revisited [1067, 1090]. Revocation [506]. Rewrite [1136]. Rewriting [1366, 974, 1374, 144]. Rheumatology [764]. Right [340, 598, 599, 253, 1459]. Rigorous [670, 976]. Rings [1298]. Ripple [892]. Risk [762]. RMI [1282]. Road [1054, 1052]. Roaming [1630]. ROBDD [164]. ROBDD-Based [164]. Robot [547, 1160, 123, 1273, 553, 97, 1173]. Robotic [115]. Robots [219, 342]. Robust [896, 224, 1142, 105, 685]. Robustness [1192, 749]. Role [1586, 644, 345, 1661]. Role-Based [1586]. Role-Centered [1661]. Roles [1565]. Rolling [1446, 1170]. Rostering [1165]. Rough [545, 882, 855, 881, 890]. Round [375]. Rounding [718]. Rounds [1404]. Routed [1485]. Routine [1479, 766]. Routines [1477]. Routing [1444, 838, 1263, 1222, 349, 1218, 1289, 597]. RPO [1137]. RSA [1403, 520, 521, 1343, 1341, 1342, 1420, 1344]. RTR [1264]. Rule [351, 800, 864, 873, 235, 921, 1183, 925, 862, 858, 890, 865, 860, 866]. Rule-Based [351]. Rules [922, 859, 867, 1167, 863, 904, 92, 861, 1183, 872, 881, 892, 917]. Rumor [265]. Run [1256, 1250, 939]. Run-Time [1256, 1250, 939]. Runge [388]. Runtime [1253, 1251, 1257, 1258]. Rush [107]. Rush-Hour [107].

S [1423, 356]. S-Box [356]. S-boxes [1423]. SAC [1594, 371, 1605]. SACRES [725]. Safe [1411, 955, 1023, 957]. Safety [678, 1019, 60]. Salesman [702, 1227]. Sampling [243]. Sandglass [1503]. Sandglass-Type [1503]. SAREX [763]. Sassy [102]. SAT [1124]. Satellite [1444, 421, 658]. Satisfaction [350, 352, 344]. Satisfiability [1278]. Saxony [767]. SCALA [1206]. Scalable [1515, 671, 1267, 1472, 263, 1310, 1459, 1206]. Scale [811, 737, 815, 1506, 864, 637, 344, 416]. Scale-Independent [737]. Scale-Up [637]. Scaleable [241]. Scaling [1265]. Scattering [1539]. Scenario [9, 1441]. Scene [313, 108, 299, 126]. Scenes [123]. SCF [397, 398]. Schedule [357]. Scheduling [447, 237, 945, 755, 735, 669, 1446, 151, 412, 1293, 1008, 603, 1221, 1237, 1316, 837, 1246, 1519, 1510, 1581, 623, 1436, 1521, 1219, 103, 1235, 150, 1298, 1147]. Schema [87, 354, 874, 498]. Schema-Guided [498]. Schemas [86, 85, 88, 499]. Scheme [1253, 817, 932, 1336, 531, 1428, 1341, 1342, 679, 1463, 405, 1672, 530]. Schemes [571, 486, 535, 1340, 487, 1418, 1548, 366, 1338, 720]. SCI [1204]. SCI-VM [1204]. Science [1370, 850, 1288]. Scientific [1488, 1549, 918, 815]. SCREEN [1634]. Screening [520]. Script [291]. Scripting [1561, 227]. SDF [940]. SDL [1639]. Sea [545]. Seamless [1507, 1638]. Search [1613, 23, 457, 458, 459, 1165, 24, 1382, 477, 22, 87, 456, 373, 879, 1162, 503, 281, 541]. Searching [577, 83, 576]. Second [562]. Secret [615, 614, 1335, 1430, 528]. Secrets [883]. Section [588]. Secure [1348, 1562, 1410, 1425, 534, 1424, 1412, 65, 531, 1345, 952, 191, 1671, 1434, 1672, 189, 190, 530]. Security [985, 187, 176, 533, 1528, 1240, 358, 508, 520, 1407, 1640, 1474, 175, 177, 510, 186, 188, 204, 1643, 271, 375, 1341, 1342, 1622, 542]. Seeking [645]. Segmentation [403, 294, 104]. Selected [1400]. Selection [95, 871, 1354, 1495, 593, 1648, 775, 921, 1481, 1159, 869]. Selective [647]. Self [303, 1612, 1300, 610, 555, 883, 1346, 1573, 1217, 1268]. Self-Assembly [1217]. Self-Avoiding [1300]. Self-Certified [1346]. Self-Complementation [610]. Self-Deploying [1573]. Self-Organisation [555]. Self-Organization [1612]. Self-Organized [303]. Self-Organizing [883]. Selfish [564]. Semantic [1054, 1048, 296, 295, 1362, 933, 634, 1567, 981, 1053, 1052, 319]. Semantic-Driven [981]. Semantically [1056]. Semantics [62, 480, 990, 1067, 137, 66, 1588, 61, 350, 1363, 1072, 58, 152, 63, 1116, 1076, 1595, 65, 1023, 1139, 676, 171, 912, 1055, 1005, 292]. Semi [86, 85, 630, 873, 874]. Semi-structured [86, 85, 630, 873, 874]. Semiconductor [55]. Semiotics [557]. Semistructured [1384, 1382]. Sensing [1493]. Sensitive [934]. Sensor [671]. Sensors [298]. Sentiments [334]. Separate [1358]. Separating [140]. Separation [784]. Sequence [1157, 1003]. Sequences [579, 410, 765, 1374, 317, 391]. Sequencing [772, 1268]. Sequential [1000, 959, 952]. Sequoia [288]. Serialization [1281]. Series [926, 1185, 354]. Server [1047, 954, 815, 1386, 845]. Servers [414, 1316, 415, 1379]. Service [193, 192, 194, 1037, 1633, 1652, 1630, 1609, 1670, 1625, 1244, 531, 1635, 1520, 1637, 1638, 1634, 1631, 1530, 368, 1472, 1641, 274, 1648, 1573, 530]. Services [1642, 1016, 1380, 1621, 1640, 1397, 1649, 1612, 1620, 1626, 1639]. Session [185, 204, 1636, 209]. Set [1175, 1484, 157, 505, 855, 960, 284, 704, 890, 860]. Set-Associative [1484]. Set-Based [960]. Set-Oriented [505]. Set-Theoretic [157]. Set-Theory [284]. Sets [245, 592, 997, 896, 140, 1151, 701, 716, 882, 881, 1132]. Setting [136]. SGI [439, 841]. Shadows [299]. Shape [703]. Shape-from-Silhouette [703]. Shape-from-Silhouette/Stereo [703]. Shapes [691, 689, 690]. Share [517, 449, 953]. Shared [571, 1515, 147, 1594, 332, 417, 413, 912, 1459, 1204, 277, 1509, 528]. Shared-Variable [147]. Sharing [1042, 1050, 1335, 1430, 362]. Shell [179, 178, 180]. Shipping [774]. Shogi [463, 477, 478]. Short [290, 285, 282, 294, 284]. Shortest [90]. Shot [302]. Shusse [288]. Shusse-Uo [288]. Shuttle [234]. SI3CO [1058]. SICS [771]. Side [227]. Sided [386, 574]. Sign [1412]. Signal [421, 1277, 1438]. Signalling [1618]. Signature [152, 1340, 1418, 513, 1339, 1338].

Signatures [506, 1350, 1412, 1349]. Significance [1507]. Signing [539]. Sikorski [1370]. Silhouette/Stereo [703]. SIMD [1242, 420]. Similar [1056]. Similarities [1053]. Similarity [903, 1234]. Similarity-Based [1234]. Simple [464, 1308, 1463, 1575]. Simpler [961]. Simplex [1273]. Simplicial [289]. Simplified [1534]. Simplifying [1636]. Simulate [1179, 1180]. Simulating [437, 1182, 1437, 851, 54, 1200, 18]. Simulation [1541, 1460, 1461, 1543, 1265, 1446, 822, 1198, 1440, 635, 741, 676, 1443, 1447, 1310, 1439, 1463, 1577, 1534, 819, 1551, 103, 818]. Simulations [1199, 445, 1468, 1506, 1442, 1478, 1462, 1441, 26]. Single [24, 1669, 165, 1235]. Single-Agent [1669]. Situated [886]. Situation [275, 1376]. Six [1273]. Six-Legged [1273]. Size [1461, 523, 570, 613, 834]. Skeletal [1212]. Skeleton [713, 1498]. Skill [476]. Skills [553]. Skipjack [1404, 381]. Skipjack-3XOR [381]. Slab [1296]. Slicing [232]. Sliding [676]. Slope [686]. Slope-Parametric [686]. Slow [269]. Small [1039, 597]. Smallest [787]. Smart [201, 202]. SMARTFED [1441]. SMP [439, 1513, 1458, 1319]. SMT [1512]. SNMP [1628]. SNMP-CORBA [1628]. Social [335, 553]. Socially [549]. Societies [335]. Soft [1174, 299]. Software [7, 823, 824, 625, 1571, 216, 8, 1, 13, 755, 2, 974, 394, 1583, 1483, 422, 417, 484, 740, 1405, 1540, 240, 1257, 1552, 153, 1036, 937]. Software-Only [1257]. Soil [1551]. Solid [1450, 56]. Solidification [400]. Solids [716]. Solution [571, 459, 468, 383, 277]. Solutions [271, 1214]. Solve [1023]. Solved [600]. Solver [1579]. Solvers [809, 1537, 812, 384]. Solving [1545, 214, 1148, 404, 1269, 352, 1143, 494, 813, 1196, 810, 389, 811, 393, 1456, 760, 1295, 1137, 808, 335, 1278, 1299, 26]. Some [804, 823, 824, 580, 1118, 586, 587, 634, 1036, 1508, 648]. Soon [730]. Sophisticated [464]. Soundness [58, 1076, 1013, 59]. Soup [1566]. Source [1502]. Sources [626, 876, 508, 88, 1038]. Space [899, 1149, 621, 30, 1264, 314, 1574, 300, 234, 688]. Space-Partitioning [899]. Spaces [1562, 710, 277]. Span [1333]. Spanish [778]. Spanning [597]. Sparse [592, 809, 396, 1457, 1302, 383, 1458, 1459, 1304]. Spatial [316, 721, 1045, 1047, 831, 79, 555, 286, 556, 1056]. Spatial-Temporal [286]. Spatially [43]. Spatio [290, 283, 287]. Spatio-Temporal [290, 283, 287]. Spatiotemporal [289, 56, 291, 284]. SPEC [214]. Special [550, 144]. Specialisation [504]. Specialising [488]. Specialization [231, 230, 490, 489]. species [1471]. Specific [1269, 1614, 482]. Specification [1321, 971, 137, 983, 1359, 214, 479, 1115, 215, 1564, 1375, 1013, 64, 1377, 60, 241]. Specifications [985, 155, 488, 972, 1365, 1399, 1015, 157, 166, 1569, 1003, 148, 1378]. Specify [1387, 1374]. Specifying [1323]. Spectral [1480]. Spectroscopy [1176]. Spectrum [819]. Speculation [835]. Speculative [461, 1512, 1258, 939]. Speech [303, 554, 111]. Speed [1292, 1296, 408, 409, 25, 454, 379]. Speed-Up [408, 409]. Spin [31]. Spine [1603]. Spineless [1603]. Splicing [964]. Split [414, 438, 415]. Split-Proxy [414, 415]. Splitting [582]. SPMD [1497]. Sports [915, 1055]. SportsFinder [915]. Spots [878]. Spotting [303]. Sprite [217]. Sprite-Based [217]. SQL [864]. Square [569]. Square-Free [569]. SS.7 [1617]. Stability [388]. Stabilization [749]. Stable [1023, 675, 613, 869]. Stack [77, 1208]. stage [1182]. Standard [1392, 330, 383]. Standards [2, 536]. Star [1293, 586, 587]. Start [269]. State [1018, 1035, 975, 168, 1182, 977, 1024, 540, 1376, 1114, 937]. State-Diagram [937]. State/Event [1018].

Static [989, 1585, 1582, 993, 967, 1019, 232]. Statically [971]. Station [517, 107]. Station-to-Station [517]. Stations [274]. Statistics [193, 192, 194]. Stealing [1208]. Steam [971]. Steel [237]. Step [820, 480, 725, 450, 465, 388]. Steps [600]. Stepsize [791]. Stereo [703, 1107, 1106, 1104, 1108, 1111]. STL [1587]. Stochastic [1010, 684, 18, 869]. Stop [1338]. Storage [1435, 285, 361, 1472, 1255, 266]. Storage-Efficient [361]. Storyboard [297]. STRATA [317]. Strategic [1597, 345]. Strategies [1068, 1490, 777, 421, 1368, 1178, 1646, 1647, 259, 1142, 1162, 383, 685]. Strategy [848, 1446, 461, 397, 398, 575, 375, 1222, 1645, 1459, 576]. Stream [1427, 1641]. Streams [503]. Streets [575, 576]. Strength [641]. Strengthened [39]. Stretch [523]. Strict [1595, 1213]. Strictness [494]. String [226, 1002, 931]. Strong [989, 697, 918, 1343, 1092, 1027, 1131, 1088, 659]. Strong/Weak [1027]. Strongly [389]. Structural [1228, 61, 965, 821]. Structure [990, 691, 875, 1061, 909, 335, 927, 553, 305, 1097]. Structured [1225, 86, 85, 1309, 22, 908, 278, 282, 1326, 293, 1299, 1509, 1378, 630, 873, 874]. Structures [950, 1359, 254, 351, 83, 1211, 844, 728, 84, 213, 979, 1122, 944, 982, 919, 319]. Structuring [1248]. STS [517]. Student [849, 848, 764]. Studies [1544]. Study [1175, 820, 1054, 120, 1012, 1212, 10, 1457, 903, 1553, 320, 673, 1080, 1318, 1578, 1605, 1227, 1181]. Sub [9, 440]. Sub-collection [440]. Subgroup [28]. Subgroups [612]. Subiterations [714]. Submodels [1168]. Subpopulations [1156]. Subrange [819]. Subroutines [64]. Subscript [1308]. Subset [1193]. Substitution [358]. Substitution-Permutation [358]. Substitutions [998, 1600, 1085].

Subsuming [1079]. Subsumption [1135]. Subtypes [954]. Subtyping [956, 992, 1130, 1089]. Success [350]. Successes [492]. Sufficient [1430]. Suitability [246]. Suite [1388, 983]. Summarization [296, 295]. Summary [647, 643]. Summation [1266]. Sums [1074, 1267]. Supercomputing [1559]. Superdevelopments [1129]. Superimposed [628]. Superscalar [842]. Supervisory [672]. Support [948, 1549, 1035, 1658, 1210, 1630, 1594, 256, 1559, 255, 273, 741, 1247, 1654, 1249, 343, 300, 270, 803]. Supporting [604, 1623, 280, 631, 1645]. Supportive [1202]. Surface [698, 699, 1162]. Surfaces [697, 700]. Surprising [1118]. Surveillance [110, 1169, 122]. Survey [753, 751, 477, 752, 347, 1104, 750]. Sustainability [891]. Suva [762]. SVD [386]. sweep [805]. Swiss [1499]. Swiss-Tx [1499]. Switched [675, 745, 1555]. Switching [737, 1556, 684]. Symbolic [1021, 1020, 655]. Symmetric [805, 806, 1332]. Synchronized [1316, 1486]. Synchronous [1225, 814, 1568, 1489, 743, 744]. Synergy [146]. Synthesis [483, 500, 486, 664, 487, 113, 740, 843, 481, 1105, 1111, 1262, 482, 1190, 498]. Synthesizing [497]. Synthetic [327, 423, 1437, 837, 1061]. System [1253, 411, 1033, 654, 550, 117, 1656, 1228, 120, 1618, 447, 139, 1230, 1292, 1236, 169, 1261, 1118, 761, 1474, 1558, 774, 275, 905, 119, 1008, 1483, 763, 311, 762, 1215, 1506, 726, 259, 1208, 451, 1001, 273, 757, 882, 727, 131, 910, 1519, 821, 170, 1247, 678, 1249, 314, 112, 1438, 1009, 122, 764, 100, 1257, 118, 775, 1657, 266, 300, 1509, 310, 308, 1329, 406, 126]. Systematic [132]. Systematising [131]. Systems [1021, 624, 77, 814, 1582, 971, 1128, 1511, 137, 334, 1018, 731, 337, 1613, 1328, 1136, 848, 1359, 596, 1251, 216, 1543, 733, 1035, 1239, 1051, 129, 1248, 664, 1303, 665, 973, 258, 666, 667, 1232, 1397, 338, 813, 836, 1010, 975, 490, 489, 40, 1367, 102, 114, 267, 738, 1539, 422, 161, 632, 1263, 1660, 671, 1370, 672, 673, 1014, 674, 739, 1234, 538, 1662, 675, 1564, 56, 620, 1044, 1655, 328, 742, 676, 743, 677, 1238, 1256, 1250, 679, 454, 413, 745, 680, 1557, 1229, 1377, 681, 682, 160, 336, 383, 987]. Systems [1284, 668, 344, 222, 221, 638, 684, 251, 830, 748, 1349, 1661, 482, 968, 329, 405, 827, 685, 1672, 497, 1114, 89, 749, 687, 1190, 1089]. Systolic [387, 828].

T3E [385]. Table [1591, 1416]. Tableau [88]. Tabling [220, 223]. Tabu [1165]. Tabular [1030]. Tagless [1603]. Tales [1039]. Talk [804, 816, 327, 326, 843, 1188, 845, 826]. Tangent [692]. Tanks [673]. Targeted [973, 1270]. TAS [984]. Task [1220, 772, 1311, 128, 837, 353, 345, 1235]. Tasks [404, 115]. TC [1616]. TCP [269]. Teaching [794]. Teams [343]. Teamwork [345]. Technique [721, 1469, 871, 758, 1538, 944, 1508]. Techniques [856, 1035, 394, 88, 1026, 1017]. Technologies [1534]. Technology [1260, 1048, 1, 13, 1632, 763, 1470, 1453, 1619, 913, 914]. Telecollaboration [298]. Telecommunication [1626]. Telecommunications [224, 1607]. Telemedical [1533]. Telemedicine [1535, 1532]. Tell [35]. Templates [1274]. Temporal [245, 290, 463, 338, 283, 629, 1116, 136, 285, 1264, 133, 287, 1004, 555, 286, 1376, 284, 791, 568, 687]. Temporal-Difference [791]. Teraflops [382]. Term [1071, 974, 1368, 144]. Termination [493, 491, 1085]. Terms [1129, 802, 476, 1100]. Terrain [1062]. Territories [581]. Test [820, 359, 166, 1058]. Test-Bed [1058]. Testing [988, 569, 996, 1028, 1114]. Tests [538]. Text [545, 281]. Texts [759]. Textual [876]. th [475]. Theater [101]. Their [187, 283, 186, 188, 1646, 1647, 69, 346, 741, 284, 1349, 1628]. Theorem [698, 1112, 984, 1019, 1032, 1087]. Theorems [1126]. Theoretic [32, 1064, 986, 157, 1664, 870, 1096]. Theoretical [781, 782]. Theories [1364, 611, 1077]. Theory [35, 996, 1295, 1320, 1567, 1200, 284, 658, 39, 234]. Therapy [766]. Thermodynamic [1154]. Thermodynamics [39]. Thermography [472]. Thin [694]. Thing [340]. Things [253]. Thinning [714, 715]. Third [193, 192, 194, 518]. Thoughts [269]. Thread [1253, 1512]. Thread-Level [1512]. Threaded [61, 1213]. Threads [1237]. Threats [1347]. Three [67, 1667, 1668, 314]. Three-Dimensional [314]. Threshold [518, 1338, 1410]. Thresholds [1345]. Thrill [710]. Tight [606]. Tile [972]. Tiling [938]. Time [411, 730, 1582, 971, 137, 129, 1323, 997, 932, 1232, 692, 1010, 1242, 438, 1636, 1244, 572, 420, 926, 1237, 1234, 578, 1442, 1663, 1185, 1246, 19, 1233, 312, 677, 1247, 1256, 104, 1250, 679, 590, 1231, 939, 1229, 1438, 1009, 608, 1113, 556, 345, 830, 1508, 103, 1441, 105, 1333, 78]. Time-Bounded [1663, 608]. Time-Deterministic [677]. Time-Lined [312]. Time-Series [926]. TimeCube [285]. Timed [730, 1408, 1409, 672, 675, 1011, 135]. Timed-Release [1408, 1409]. Timestamp [528]. Timestamps [264]. Timetabling [236]. Timing [378]. TINA [1653, 1609, 1625, 1646, 1647, 1645, 1650, 1644]. TINA-Based [1609]. TLS [173, 172, 174]. TM [1047, 64]. TMN [1608]. Tolerance [1574]. Tolerant [139, 1239, 40, 1245, 1221, 1222, 1284]. Tomography [1455]. Tool [948, 1388, 983, 1035, 768, 1303, 769, 1034, 1028, 853, 1531, 1370, 673, 1557, 1502, 1032, 1390, 1441].

Tool-Aided [673]. Tool-kit [768, 769]. Toolbox [416, 688]. Toolkit [1487]. Tools [1012, 1387, 421, 1559, 984, 1475, 1036]. Toolset [1030]. Topic [1614]. Topic-Specific [1614]. Topics [647]. Topological [705, 708, 426, 699, 79, 44, 98, 1060]. Topology [707, 706, 1148, 1589]. Topos [1133]. Tori [1555]. Torus [1485]. TOSCA [1634]. Total [1068, 1002]. Tour [1372]. Tournaments [777]. Towers [600]. Town [1039]. Trace [147, 586, 587, 505]. Trace-Based [147]. Traceability [366]. Tracer [1477]. Traces [1487, 135]. Track [1014]. Tracking [121, 1236, 119, 325, 107, 122, 106, 105, 78]. Tradeoffs [1435]. Trading [777]. Traffic [110, 727, 1298, 660]. Trakhtenbrot [1112]. Transaction [280, 1374, 1439, 866]. Transactional [1380]. Transactions [77]. Transcript [506, 194, 202, 197, 180, 183, 177, 188, 208, 195, 200, 198, 174, 203, 191, 206]. Transcript-Irrelevant [506]. Transducers [1002]. Transfer [1408, 1407, 1409, 1406, 509, 1058]. Transform [418]. Transformation [947, 1367, 1211, 403, 281]. Transformational [1212, 984, 1094]. Transformations [502, 501, 1502, 943]. Transforming [1504]. Transforms [15]. Transition [677]. Translating [1371]. Translation [86, 85]. Transmitting [266]. Transparent [1253, 1204, 1575]. Transport [1618]. Transportation [1054, 344]. Trapdoor [522]. Trapped [52]. TRAPping [946]. Traps [53]. Traveling [702, 1227]. Treasure [883]. Treasure-Map [883]. treatment [1097]. Tree [445, 425, 995, 72, 582, 1002, 712, 287, 468, 872]. Tree-Width [72]. Trees [997, 1531, 468, 597]. Treewidth [583]. Trends [637]. Triangulated [583, 125]. Tribe [1433]. Tribology [849]. Tricks [795]. Tridiagonal [384]. Tropology [284].

TRUCE [1586]. Trust [185, 182, 179, 184, 178, 180, 181, 183, 195, 205, 206]. Trust-Management [182, 181, 183]. Trusted [518]. TSP [602]. Tsume [471, 477, 465, 466]. Tsume-Go [471, 465, 466]. Tsume-Shogi [477]. TTP [1351]. Tube [721]. Tumor [1534]. tune [1096]. Tuning [885]. Tunnel [694]. Tunnel-Free [694]. Tuple [1574]. Turbulence [819]. Turing [578]. Tutoring [764]. TV [311, 310]. Two [444, 386, 1460, 14, 1543, 574, 1411, 602, 256, 255, 465, 673, 82, 388, 680]. Two-Level [1460]. Two-Particle [14]. Two-Sided [386, 574]. Two-Step [465, 388]. Twofish [357]. Tx [1499]. Typability [992]. Typable [1129]. Type [1358, 506, 1128, 1069, 146, 158, 58, 1064, 1077, 1098, 161, 1090, 1600, 962, 1602, 160, 59, 266, 1503, 60, 1089, 1093]. Type-Based [962]. Type-Checking [1358]. type-directed [1093]. Type-Logical [1098]. Type-Safety [60]. Type-Theoretic [1064]. Typed [162, 1075, 1076, 1367]. Types [1134, 772, 483, 159, 955, 954, 157, 1131, 1602, 1375, 1084, 987, 1130, 229, 1070].

Ubiquitous [1630]. UKMO [1477]. Ultra [713]. Ultra-Fast [713]. UML [1528, 976, 1627]. UMTS [1631]. Unbalanced [1418]. unbounded [1097]. Uncertain [651]. Uncertainty [652]. Unconditional [542]. Undecidability [584, 586, 587]. Underspecification [1098]. Understanding [404, 632, 696, 111]. Undirected [90]. Unfolding [990]. Uni [1160]. Uni-vector [1160]. Unidirectional [1298]. Unification [1381, 997, 1600]. Unified [806, 1379]. Uniform [1373]. Uniformly [1244]. Unifying [1031, 281]. Unique [1042]. Uniqueness [1128]. Unit [1278]. Units [78]. Universal [580, 33, 1322, 359, 373, 1405, 608]. Universes [161, 160]. University [236].

Unknown [517, 524, 576]. Unmanned [683]. Unraveling [710]. Unroll [1500]. Unroll-and-Jam [1500]. Unseen [902]. Unsharp [31]. Unsolvable [1071]. Unsteady [1460]. Unstructured [1300, 427]. Unsupervised [1153, 294]. Untyped [1072]. Unusual [804]. Uo [288]. Update [266]. Upper [1123, 622]. Urn [74]. Urquhart [1127]. Usability [1620, 1565]. Use [485, 1653, 1322, 318, 538, 1308, 651, 546, 1522, 476, 754]. Used [662]. Useless [1070]. Useless-Code [1070]. User [1517, 631, 557, 275, 1646, 1647, 317, 655]. User-Centered [317]. User-Centred [1517]. Users [1631, 1527]. Using [94, 1157, 820, 301, 55, 654, 463, 1018, 1305, 1404, 1504, 691, 46, 1350, 1230, 1179, 1180, 734, 1364, 1324, 1189, 303, 1296, 669, 1287, 630, 236, 736, 477, 412, 1023, 807, 1026, 1156, 1636, 420, 1028, 1252, 837, 766, 672, 1061, 238, 239, 1665, 871, 1627, 1295, 1185, 977, 1475, 1330, 910, 363, 742, 1565, 743, 287, 758, 291, 1639, 714, 453, 846, 1144, 268, 101, 1374, 1169, 1443, 1438, 503, 987, 122, 191, 1108, 1111, 391, 1577, 683, 526, 847, 1386, 496, 1534, 118, 1056, 1278, 872, 482, 111, 1194, 1173, 1356, 189, 126, 190]. Using [1498, 719]. Utility [1230, 817]. Utilization [353, 1390, 430]. Utilizing [1453]. Utterance [301].

Valbum [297]. Valid [888]. Validation [258, 1637]. Value [146, 1075, 835, 1079, 1027]. Value-Passing [146, 1027]. Valued [1139]. Values [920, 890, 902]. Vapnik [798]. Variable [164, 147, 477]. Variable-Depth [477]. Variables [788]. Various [1145]. Vast [884]. Vector [596, 812, 811, 1457, 1301, 842, 834, 803, 833, 1160]. Vectors [724]. Vehicle [110, 99, 1218]. Vehicles [683]. Verifiable [199, 200, 1333, 362]. Verification [1021, 176, 1018, 731, 733, 734, 146, 1012, 338, 175, 177, 136, 1325, 1026, 1008, 1029, 133, 1017, 1024, 678, 979, 1036, 135]. Verifier [1013]. Verifying [1323, 1324]. vernacular [1101]. Versatile [851]. Version [1348, 256, 255, 600]. Versus [574, 1298, 591, 1376]. Vertex [622]. Very [1451]. via [731, 14, 1290, 1590, 1074, 298, 1427, 671, 967, 1258, 1004, 495, 1217, 919]. Video [316, 411, 296, 295, 414, 303, 412, 302, 420, 531, 321, 312, 297, 314, 415, 413, 293, 292, 1641, 106, 406, 319, 315, 530]. Video-Based [314]. Videoplex} [314]. View [989, 154, 93, 531, 257, 1105, 1111, 648, 530]. Viewing [297]. Views [94, 73, 626, 1387, 5, 95, 318, 282, 70, 781, 1385, 1094]. Vinegar [1418]. Virtual [839, 1518, 1547, 1516, 1658, 1507, 438, 1533, 1526, 326, 64, 944, 324, 300, 308, 315]. Viruses [1347]. Vision [117, 120, 110, 115, 119, 116, 313, 112, 101, 1111, 100, 97, 99, 103, 111, 126]. Vision-Aided [99]. Visual [121, 113, 133, 124, 123, 98, 323, 546, 433]. Visual-MCM [433]. VisualExpresso [1516]. Visualisation [1445]. Visualise [1534]. Visualising [433]. Visualization [247, 1517, 1473, 313]. Visualizing [905]. Visually [904]. Vitro [1534]. VIVRE [1517]. VLSI [406]. VM [1204]. Volatility [888]. Volumes [354]. Volunteer [1284]. Voronoi [718]. Voxel [722]. Voxels [716]. vs [410]. VSF [886].

W [264]. Walking [575, 1273]. Walks [784, 1300]. WAM [231, 230]. Warehouse [247, 254, 96]. Warehousing [249, 257, 253, 248]. Warp [1323, 438]. Watermarks [1413]. Watershed [403]. WATM [269]. Waveguide [1552]. Waveguides [1296]. Wavelength [1292]. Wavelet [418, 419, 15, 417, 416]. Way [1433, 545, 554, 840]. Wayness [512]. WDM [1293, 1298]. Weak [783, 1027]. Weakened [1407]. Weakly [583].

Weakness [610]. Weather [1449]. Web [1625, 1388, 247, 249, 1389, 1476, 1381, 1261, 260, 1466, 1667, 531, 1668, 915, 1465, 1475, 227, 1379, 1391, 1386, 1390, 1534, 277, 530]. Web-Based [1625, 1476, 1261, 531, 1475, 530]. WebFlow [1466]. WebSuite [1388]. Weight [1029]. Weighted [590]. WekI [762]. Welfare [1173]. Well [1068, 738, 1101]. well-formedness [1101]. Well-Founded [1068]. Well-Posedness [738]. Wheel [1445]. Wheeler [281]. which [1085]. While [1456]. Whole [1501]. Whole-Program [1501]. Wide [1389, 1381]. Wider [10]. Width [72]. Winged [1061]. Winged-Edge [1061]. Wireless [261, 267, 1486, 262]. WITH-Loop-Folding [1605]. within [1558, 223]. Without [1412, 1346, 1020]. WLAN [274]. Word [585]. Words [650, 656]. Work [280, 1208]. Work-Stealing [1208]. Workflow [635]. Working [253, 1616]. Workload [1244, 1439]. Workloads [75, 837]. Works [1173]. Workshop [1328, 1280, 1314, 1259, 643, 1320, 1202, 1250, 644, 1229, 1299, 1288, 1214]. Workstation [1524, 437, 818]. Workstations [442, 1317, 1314, 412, 1548, 1227, 1254, 1505]. World [901, 292, 1389, 1381]. Worlds [1526]. Wormhole [1485]. Wormhole-Routed [1485]. Worst [605]. Worst-Case [605]. WP [994]. WP-calculus [994]. Wright [1434]. Write [993, 1290]. Write-Down [993]. WUM [1390]. WWW [1669, 1383, 1530].

X.509 [525]. XOR [1422]. XSIL [1488].

Yao [74]. YAT [1386]. Year [1064]. Years [67, 827].

Zero [1411, 535, 1339, 1432]. Zero-Knowledge [1411, 535, 1339, 1432]. ZEUS [768, 769]. Zhang [1340].

References

Derniame:1999:SPM

[1] Jean-Claude Derniame, Badara Ali Kaba, and Brian Warboys. The software process: Modelling and technology. Lecture Notes in Computer Science, 1500:1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1500/15000001.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1500/15000001.pdf. Editor: Carlo

Montangero.

Finkelstein:1999:SPS

[2] Anthony Finkelstein, Alfonso Fuggetta, Carlo Montangero, and Jean-Claude Derniame. Software process — standards, assessments and improvement. Lecture Notes in Computer Science, 1500:15–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1500/15000015.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1500/15000015.

pdf. Editor: Wolfgang Emmerich.

Conradi:1999:PML

[3] Reidar Conradi and M. Letizia Jaccheri. Process modelling languages. Lecture Notes in Computer Science, 1500:27–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/

0558/bibs/1500/15000027.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1500/15000027.
pdf.

Avrillionis:1999:MP

[4] Denis Avrillionis, Reidar Conradi, Pierre-Yves Cunin, Minh Nguyen, and Ian Robertson. Meta-process. Lecture Notes in Computer Science, 1500:53– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1500/15000053.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1500/15000053.

pdf. Editor: Brian Warboys.

Fuggetta:1999:AVA

[5] Alfonso Fuggetta, Claude Godart, and Jens Jahnke. Architectural views and alternatives. Lecture Notes in Computer Science, 1500:95–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1500/15000095.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1500/15000095.

pdf. Editor: Wilhelm Schaefer.

Belkhatir:1999:CCP

[6] Noureddine Belkhatir, Antonio Carzaniga, Jacky Estublier, Elisabetta Di Nitto, Jens Jahnke, Patricia Lago, Wilhelm Schaefer, and Hala Skaf. Cooperation control in PSEE. Lecture Notes in Computer Science, 1500:117–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1500/15000117.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1500/15000117.

pdf. Editor: Claude Godart.

Arbaoui:1999:HDS

[7] Selma Arbaoui, Jacques Lonchamp, and Carlo Montangero. The human dimension of the software process. Lecture Notes in Computer Science, 1500: 165–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1500/15000165.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1500/15000165.

pdf. Editor: David Wastell.

Derniame:1999:SPK

[8] Jean-Claude Derniame, Mark Greenwood, Peter Kawalek, and Ian Roberston. Software process: Key issues and future directions. Lecture Notes in Computer Science, 1500:201– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1500/15000201.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1500/15000201.

pdf. Editors: David Wastell, Luuk Groenewegen.

Anonymous:1999:LSP

[9] Anonymous. Lifecycle (sub) process demonstration scenario (ISPW 9). Lecture Notes in Computer Science, 1500: 217–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1500/15000217.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1500/15000217.
                                        pdf.

Groenewegen:1999:CSD

[10] Luuk Groenewegen and Vincenzo Ambriola. Case study demonstrating the wider applicability of the PSEE paradigm. Lecture Notes in Computer Science, 1500:227–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1500/15000227.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1500/15000227.

pdf. Editor: Jean-Claude Derniame.

Alloui:1999:AFP

[11] Ilham Alloui, Selma Arbaoui, Denis Avrilionis, Christer Fernstrom, Samir Dami, Jacky Estublier, and Flavio Oquendo. Assessment framework for PSEEs. Lecture Notes in Computer Science, 1500:245–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1500/15000245.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1500/15000245.

pdf. Editor: Pierre-Yves Cunin.

Anonymous:1999:GR

[12] Anonymous. Glossary and references. Lecture Notes in Computer Science, 1500:277–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1500/15000277.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1500/15000277.
pdf.

Derniame:1999:SPP

[13] J. C. (Jean Claude) Derniame, Badara Ali Kaba, and David Wastell. Software process: Principles, methodology, and technology. Lecture Notes in Computer Science, 1500:i–xii, 1–307, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brassard:1999:MPE

[14] Gilles Brassard and Tal Mor. Multiparticle entanglement via two-particle entanglement. Lecture Notes in Computer Science, 1509:1–9, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1509/15090001.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1509/15090001.pdf.

Fijany:1999:QWT

[15] Amir Fijany and Colin P. Williams. Quantum wavelet transforms: Fast algorithms and complete circuits. Lecture Notes in Computer Science, 1509: 10–33, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1509/15090010.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1509/15090010.
pdf.

Yepez:1999:QCF

[16] Jeffrey Yepez. Quantum computation of fluid dynamics. Lecture Notes in Computer Science, 1509:34– 60, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1509/15090034.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1509/15090034.
pdf.

Cleve:1999:QEC

[17] Richard Cleve, Wim van Dam, Michael Nielsen, and Alain Tapp. Quantum entanglement and the communication complexity of the inner product function. Lecture Notes in Computer Science, 1509:61–74, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1509/15090061.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1509/15090061.
pdf.

Zak:1999:QRN

[18] Michail Zak and Colin P. Williams. Quantum recurrent networks for simulating stochastic processes. Lecture Notes in Computer Science, 1509:75– 88, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1509/15090075.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1509/15090075.
pdf.

Mahler:1999:CBC

[19] Günter Mahler and Ilki Kim. Correlation between correlations: Process and time in quantum networks. Lecture Notes in Computer Science, 1509:89– 102, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1509/15090089.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1509/15090089.
pdf.

Jozsa:1999:QEA

[20] Richard Jozsa. Quantum effects in algorithms. Lecture Notes in Computer Science, 1509:103–112, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1509/15090103.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1509/15090103.
pdf.

Williams:1999:ADQ

[21] Colin P. Williams and Alexander G. Gray. Automated design of quantum circuits. Lecture Notes in Computer Science, 1509:113–125, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1509/15090113.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1509/15090113.
pdf.

Grover:1999:QSS

[22] Lov K. Grover. Quantum search on structured problems. Lecture Notes in Computer Science, 1509:126–139, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1509/15090126.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1509/15090126.
pdf.

Biron:1999:GGS

[23] David Biron, Ofer Biham, Eli Biham, Markus Grassl, and Daniel A. Lidar. Generalized grover search algorithm for arbitrary initial amplitude distribution. Lecture Notes in Computer Science, 1509:140–147, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1509/15090140.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1509/15090140.
pdf.

Chi:1999:QDS

[24] Dong Pyo Chi and Jinsoo Kim. Quantum database search by a single query. Lecture Notes in Computer Science, 1509:148–151, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1509/15090148.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1509/15090148.
pdf.

Ozhigov:1999:QCC

[25] Y. Ozhigov. Quantum computer can not speed up iterated applications of a black box. Lecture Notes in Computer Science, 1509:152–159, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1509/15090152.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1509/15090152.
pdf.

Zak:1999:QRS

[26] Michail Zak. Quantum resonance for solving NP-complete problems by simulations. Lecture Notes in Computer Science, 1509:160–166, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1509/15090160.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1509/15090160.
pdf.

Abrams:1999:CCP

[27] Daniel S. Abrams and Seth Lloyd. Computational complexity and physical law. Lecture Notes in Computer Science, 1509:167–173, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1509/15090167.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1509/15090167.pdf.

Mosca:1999:HSP

[28] Michele Mosca and Artur Ekert. The hidden subgroup problem and eigenvalue estimation on a quantum computer. Lecture Notes in Computer Science, 1509:174–188, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1509/15090174.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1509/15090174.
pdf.

Castagnoli:1999:DAQ

[29] Giuseppe Castagnoli and Dalida Monti. A diakoptic approach to quantum computation. Lecture Notes in Computer Science, 1509:189–199, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1509/15090189.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1509/15090189.
pdf.

Hughes:1999:PFS

[30] R. J. Hughes, W. T. Buttler, P. G. Kwiat, S. K. Lamoreaux, G. G. Luther, G. L. Morgan, J. E. Nordholt, C. G. Peterson, and C. M. Simmons. Practical free-space quantum cryptography. Lecture Notes in Computer Science, 1509:200–213, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1509/15090200.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/1509/15090200.
pdf.

Roy:1999:QCE

[31] S. Roy and G. Kar. Quantum cryptography, eavesdropping, and unsharp spin measurement. Lecture Notes in Computer Science, 1509:214–217, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1509/15090214.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1509/15090214.
pdf.

Cerf:1999:ITA

[32] Nicolas J. Cerf. Information-theoretic aspects of quantum copying. Lecture Notes in Computer Science, 1509:218– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1509/15090218.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1509/15090218.
pdf.

Buzek:1999:UOC

[33] V. Buzek and M. Hillery. Universal optimal cloning of qubits and quantum registers. Lecture Notes in Computer Science, 1509:235–246, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1509/15090235.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1509/15090235.
pdf.

DiVincenzo:1999:EA

[34] David P. DiVincenzo, Christopher A. Fuchs, Hideo Mabuchi, John A. Smolin, Ashish Thapliyal, and Armin Uhlmann. Entanglement of assistance. Lecture Notes in Computer Science, 1509:247–257, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1509/15090247.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1509/15090247.pdf.

Adami:1999:WIT

[35] C. Adami and N. J. Cerf. What information theory can tell us about quantum reality. Lecture Notes in Computer Science, 1509:258–268, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1509/15090258.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1509/15090258.
                                        pdf.

Levitin:1999:QGC

[36] L. B. Levitin. Quantum generalization of conditional entropy and information. Lecture Notes in Computer Science, 1509:269–275, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1509/15090269.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1509/15090269.
                                        pdf.

Allahverdyan:1999:AIM

[37] A. E. Allahverdyan and D. B. Saakian. Accessible information in multi-access quantum channels. Lecture Notes in Computer Science, 1509:276–284, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1509/15090276.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1509/15090276.
pdf.

Westmoreland:1999:CQC

[38] Michael Westmoreland and Benjamin Schumacher. Capacities of quantum channels and quantum coherent information. Lecture Notes in Computer Science, 1509:285–295, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1509/15090285.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1509/15090285.
pdf.

Saakian:1999:SLI

[39] D. B. Saakian and A. E. Allahverdyan. Strengthened Lindblad inequality: Applications in non-equilibrium thermodynamics and quantum information theory. Lecture Notes in Computer Science, 1509:296– 301, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1509/15090296.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1509/15090296.
                                        pdf.

Gottesman:1999:FTQ

[40] Daniel Gottesman. Fault-tolerant quantum computation with higherdimensional systems. Lecture Notes in Computer Science, 1509:302–313, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1509/15090302.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1509/15090302.
                                        pdf.

Chau:1999:QCE

[41] H. F. Chau. Quantum convolutional error correction codes. Lecture Notes in Computer Science, 1509:314–324, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1509/15090314.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1509/15090314.
                                        pdf.

Roychowdhury:1999:ENQ

[42] Vwani P. Roychowdhury and Farrokh Vatan. On the existence of nonadditive quantum codes. Lecture Notes in Computer Science, 1509:325–336, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1509/15090325.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1509/15090325.
                                        pdf.

Duan:1999:QEC

[43] Lu-Ming Duan and Guang-Can Guo. Quantum error correction is applicable for reducing spatially correlated decoherence. Lecture Notes in Computer Science, 1509:337–340, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1509/15090337.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1509/15090337.
pdf.

Ogburn:1999:TQC

[44] R. Walter Ogburn and John Preskill. Topological quantum computation. Lecture Notes in Computer Science, 1509:341–356, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1509/15090341.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1509/15090341.pdf.

Laflamme:1999:NG

[45] R. Laflamme, E. Knill, W. H. Zurek, P. Catasti, and S. V. S. Mariappan. NMR GHZ. Lecture Notes in Computer Science, 1509:357–363, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1509/15090357.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1509/15090357.
pdf.

Bowden:1999:QCU

[46] Charles M. Bowden, Jonathan P. Dowling, and Steven P. Hotaling. Quantum computing using electron-nuclear double resonances. Lecture Notes in Computer Science, 1509:364–372, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1509/15090364.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1509/15090364.
pdf.

Briegel:1999:PIQ

[47] H.-J. Briegel, J. I. Cirac, W. Dür, S. J. van Enk, H. J. Kimble, H. Mabuchi, and P. Zoller. Physical implementations for quantum communication in quantum networks. Lecture Notes in Computer Science, 1509:373–382, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1509/15090373.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1509/15090373.
pdf.

Franson:1999:OAQ

[48] J. D. Franson and T. B. Pittman. An optical approach to quantum computing. Lecture Notes in Computer Science, 1509:383–390, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1509/15090383.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1509/15090383.pdf.

Adami:1999:QCL

[49] C. Adami and N. J. Cerf. Quantum computation with linear optics. Lecture Notes in Computer Science, 1509:391– 401, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1509/15090391.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1509/15090391.
pdf.

Vitali:1999:DCO

[50] D. Vitali and P. Tombesi. Decoherence control for optical qubits. Lecture Notes in Computer Science, 1509:402– 412, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1509/15090402.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1509/15090402.
pdf.

Averin:1999:ACG

[51] D. V. Averin. Adiabatic controlled-NOT gate for quantum computation. Lecture Notes in Computer Science, 1509:413–425, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1509/15090413.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1509/15090413.pdf.

James:1999:TIQ

[52] D. F. V. James, M. S. Gulley, M. H. Holzscheiter, R. J. Hughes, P. G.

Kwiat, S. K. Lamoreaux, C. G. Peterson, V. D. Sandberg, M. M. Schauer, C. M. Simmons, D. Tupa, P. Z. Wang, and A. G. White. Trapped ion quantum computer research at Los Alamos. Lecture Notes in Computer Science, 1509:426–437, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1509/15090426.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1509/15090426.
pdf.

DeVoe:1999:AEI

[53] Ralph G. DeVoe. Arrays of elliptical ion traps for parallel quantum computing. Lecture Notes in Computer Science, 1509:438–446, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1509/15090438.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1509/15090438.
pdf.

Obenland:1999:SED

[54] Kevin M. Obenland and Alvin M. Despain. Simulating the effect of decoherence and inaccuracies on a quantum computer. Lecture Notes in Computer Science, 1509:447–459, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1509/15090447.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/1509/15090447.
pdf.

Balandin:1999:IQC

[55] Alexander A. Balandin and Kang L. Wang. Implementation of quantum controlled-NOT gates using asymmetric semiconductor quantum dots. Lecture Notes in Computer Science, 1509: 460–467, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1509/15090460.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1509/15090460.
pdf.

Matsueda:1999:SDQ

[56] Hideaki Matsueda. Spatiotemporal dynamics of quantum computing solid dipole-dipole block systems. Lecture Notes in Computer Science, 1509:468– 477, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1509/15090468.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1509/15090468.
pdf.

Alves-Foss:1999:FGJ

[57] J. Alves-Foss and D. Frincke. Formal grammar for Java. Lecture Notes in Computer Science, 1523:1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Drossopoulou:1999:DSJ

[58] S. Drossopoulou and S. Eisenbach. Describing the semantics of Java and proving type soundness. Lecture Notes in Computer Science, 1523:41– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Syme:1999:PJT

[59] D. Syme. Proving Java type soundness. Lecture Notes in Computer Science, 1523:83–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VonOheimb:1999:MCJ

[60] D. Von Oheimb and T. Nipkow. Machine-checking the Java specification: Proving type-safety. Lecture Notes in Computer Science, 1523:119– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cenciarelli:1999:EBS

[61] P. Cenciarelli, A. Knapp, B. Reus, and M. Wirsing. An event-based structural operational semantics of multithreaded Java. Lecture Notes in Computer Science, 1523:157–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alves-Foss:1999:DDS

[62] J. Alves-Foss and F. S. Lam. Dynamic denotational semantics of Java. Lecture Notes in Computer Science, 1523: 201–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Flatt:1999:PRS

[63] M. Flatt, S. Krishnamurthi, and M. Felleisen. A programmer’s reduction semantics for classes and mixins. Lecture Notes in Computer Science, 1523:241–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Qian:1999:FSJ

[64] Z. Qian. A formal specification of Java[TM] virtual machine instructions for objects, methods and subroutines. Lecture Notes in Computer Science, 1523:271–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hartel:1999:OSJ

[65] P. H. Hartel, M. J. Butler, and M. Levy. The operational semantics of a Java secure processor. Lecture Notes in Computer Science, 1523:313–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boerger:1999:PFM

[66] E. Boerger and W. Schulte. A programmer friendly modular definition of the semantics of Java. Lecture Notes in Computer Science, 1523:353–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Atkinson:1999:IRT

[67] Malcolm Atkinson and Mick Jordan. Issues raised by three years of developing PJama: An orthogonally persistent platform for Java^{T M}. Lecture Notes in Computer Science, 1540:1– 30, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1540/15400001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1540/15400001.
pdf.

Papadimitriou:1999:NCA

[68] Christos H. Papadimitriou. Novel computational approaches to information retrieval and data mining p. 31. Lecture Notes in Computer Science, 1540: 31–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1540/15400031.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1540/15400031.
pdf.

Lenzerini:1999:DLT

[69] Maurizio Lenzerini. Description logics and their relationships with databases. Lecture Notes in Computer Science, 1540:32–38, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1540/15400032.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1540/15400032.
pdf.

Popa:1999:ECP

[70] Lucian Popa and Val Tannen. An equational chase for path-conjunctive queries, constraints, and views. Lecture Notes in Computer Science, 1540: 39–57, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1540/15400039.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1540/15400039.
pdf.

Neven:1999:ALF

[71] Frank Neven, Martin Otto, Jurek Tyszkiewicz, and Jan Van den Bussche. Adding for-loops to first-order logic. Lecture Notes in Computer Science, 1540:58–69, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1540/15400058.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1540/15400058.
pdf.

Grohe:1999:DDC

[72] Martin Grohe and Julian Mariño. Definability and descriptive complexity on databases of bounded tree-width. Lecture Notes in Computer Science, 1540: 70–82, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1540/15400070.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1540/15400070.
pdf.

Bailey:1999:DFO

[73] James Bailey and Guozhu Dong. Decidability of first-order logic queries over views. Lecture Notes in Computer Science, 1540:83–99, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1540/15400083.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1540/15400083.pdf.

Gardy:1999:UMY

[74] Danièle Gardy and Laurent Némirovski. Urn models and Yao’s formula. Lecture Notes in Computer Science, 1540:100– 112, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1540/15400100.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1540/15400100.
pdf.

Hellerstein:1999:GDI

[75] Joseph M. Hellerstein, Lisa Hellerstein, and George Kollios. On the generation of 2-dimensional index workloads. Lecture Notes in Computer Science, 1540:113–130, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1540/15400113.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1540/15400113.
pdf.

Nicola:1999:IEA

[76] Matthias Nicola and Matthias Jarke. Increasing the expressiveness of analytical performance models for replicated databases. Lecture Notes in Computer Science, 1540:131–149, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1540/15400131.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1540/15400131.
pdf.

Alonso:1999:TSF

[77] Gustavo Alonso, Armin Feßler, Guy Pardon, and Hans-Jörg Schek. Transactions in stack, fork, and join composite systems. Lecture Notes in Computer Science, 1540:150–168, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1540/15400150.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1540/15400150.
pdf.

Wolfson:1999:DTM

[78] Ouri Wolfson, Liqin Jiang, A. Prasad Sistla, Sam Chamberlain, Naphtali Rishe, and Minglin Deng. Databases for tracking mobile units in real time. Lecture Notes in Computer Science, 1540:169–186, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1540/15400169.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1540/15400169.pdf.

Kuijpers:1999:CFO

[79] Bart Kuijpers and Jan Van den Bussche. On capturing first-order topological properties of planar spatial databases. Lecture Notes in Computer Science, 1540:187–198, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1540/15400187.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1540/15400187.
                                        pdf.

Grumbach:1999:ODC

[80] Stéphane Grumbach, Philippe Rigaux, and Luc Segoufin. On the orthographic dimension of constraint databases. Lecture Notes in Computer Science, 1540: 199–216, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1540/15400199.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1540/15400199.
                                        pdf.

Beyer:1999:WNN

[81] Kevin Beyer, Jonathan Goldstein, Raghu Ramakrishnan, and Uri Shaft. When is “nearest neighbor” meaningful? Lecture Notes in Computer Science, 1540:217–235, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1540/15400217.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1540/15400217.pdf.

Muthukrishnan:1999:RPT

[82] S. Muthukrishnan, Viswanath Poosala, and Torsten Suel. On rectangular partitionings in two dimensions: Algorithms, complexity, and applications. Lecture Notes in Computer Science, 1540:236–256, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1540/15400236.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1540/15400236.
pdf.

Kanth:1999:ODR

[83] K. V. Ravi Kanth and Ambuj Singh. Optimal dynamic range searching in non-replicating index structures. Lecture Notes in Computer Science, 1540: 257–276, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1540/15400257.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1540/15400257.
pdf.

Milo:1999:ISP

[84] Tova Milo and Dan Suciu. Index structures for path expressions. Lecture Notes in Computer Science, 1540:277– 295, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1540/15400277.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1540/15400277.
pdf.

Beeri:1999:SIT

[85] Catriel Beeri and Tova Milo. Schemas for integration and translation of structured and semi-structured data. Lecture Notes in Computer Science, 1540: 296–313, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1540/15400296.htm;   http:
//link.springer-ny.com/link/service/

                                        series/0558/papers/1540/15400296.
                                        pdf.

Been:1999:SIT

[86] C. Been and T. Milo. Schemas for integration and translation of structured and semi-structured data. Lecture Notes in Computer Science, 1540: 296–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grumbach:1999:SLS

[87] Stéphane Grumbach and Giansalvatore Mecca. In search of the lost schema. Lecture Notes in Computer Science, 1540:314–331, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1540/15400314.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1540/15400314.
                                        pdf.

Grahne:1999:TTQ

[88] Gösta Grahne and Alberto O. Mendelzon. Tableau techniques for querying information sources through global schemas. Lecture Notes in Computer Science, 1540:332–347, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1540/15400332.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1540/15400332.
                                        pdf.

Yerneni:1999:OLJ

[89] Ramana Yerneni, Chen Li, Jeffrey Ullman, and Hector Garcia-Molina. Optimizing large join queries in mediation systems. Lecture Notes in Computer Science, 1540:348–364, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1540/15400348.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1540/15400348.
pdf.

Pang:1999:IFM

[90] Chaoyi Pang, Ramamohanarao Kotagiri, and Guozhu Dong. Incremental FO(+, <) maintenance of all-pairs shortest paths for undirected graphs after insertions and deletions. Lecture Notes in Computer Science, 1540:365– 382, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1540/15400365.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1540/15400365.
pdf.

Cabibbo:1999:FIA

[91] Luca Cabibbo and Riccardo Torlone. A framework for the investigation of aggregate functions in database queries. Lecture Notes in Computer Science, 1540:383–397, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1540/15400383.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1540/15400383.
                                        pdf.

Pasquier:1999:DFC

[92] Nicolas Pasquier, Yves Bastide, Rafik Taouil, and Lotfi Lakhal. Discovering frequent closed itemsets for association rules. Lecture Notes in Computer Science, 1540:398–416, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1540/15400398.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1540/15400398.
                                        pdf.

Godfrey:1999:VD

[93] Parke Godfrey and Jarek Gryz. View disassembly. Lecture Notes in Computer Science, 1540:417–434, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1540/15400417.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1540/15400417.
                                        pdf.

Afrati:1999:AQU

[94] Foto N. Afrati, Manolis Gergatsoulis, and Theodoros Kavalieros. Answering queries using materialized views with disjunctions. Lecture Notes in Computer Science, 1540:435–452, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1540/15400435.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1540/15400435.
pdf.

Gupta:1999:SVM

[95] Himanshu Gupta and Inderpal Singh Mumick. Selection of views to materialize under a maintenance cost constraint. Lecture Notes in Computer Science, 1540:453–470, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1540/15400453.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1540/15400453.
pdf.

Gupta:1999:DWN

[96] Himanshu Gupta and Divesh Srivastava. The data warehouse of newsgroups. Lecture Notes in Computer Science, 1540:471–488, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1540/15400471.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1540/15400471.
pdf.

Schlegel:1999:IVB

[97] C. Schlegel, J. Illman, H. Jaberg, and M. Schuster. Integrating vision based behaviours with an autonomous robot. Lecture Notes in Computer Science, 1542:1–20, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Santos-Victor:1999:TMV

[98] J. Santos-Victor, R. Vassallo, and H. Schneebeli. Topological maps for visual navigation. Lecture Notes in Computer Science, 1542:21–36, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Southall:1999:VAO

[99] B. Southall, T. Hague, J. A. Marchant, and B. F. Buxton. Vision-aided outdoor navigation of an autonomous horticultural vehicle. Lecture Notes in Computer Science, 1542:37–50, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schiele:1999:ICV

[100] B. Schiele, N. Oliver, T. Jebara, and A. Pentland. An interactive computer vision system DyPERS: Dynamic personal enhanced reality system. Lecture Notes in Computer Science, 1542:51– 65, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pinhanez:1999:UCV

[101] C. Pinhanez and A. Bobick. Using computer vision to control a reactive computer graphics character in a theater play. Lecture Notes in Computer Science, 1542:66–82, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hammes:1999:SLO

[102] J. P. Hammes, B. A. Draper, and A. P. Willem Boehm. Sassy: A language and optimizing compiler for image processing on reconfigurable computing systems. Lecture Notes in Computer Science, 1542:83–97, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Torres:1999:SSR

[103] F. Torres, F. A. Candelas, S. T. Puente, and L. M. Jimenez. Simulation and scheduling of real-time computer vision algorithms. Lecture Notes in Computer Science, 1542:98–114, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nordlund:1999:RTM

[104] P. Nordlund and J.-O. Eklundh. Realtime maintenance of figure-ground segmentation. Lecture Notes in Computer Science, 1542:115–134, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vincze:1999:IFR

[105] M. Vincze, M. Ayromlou, and W. Kubinger. An integrated framework for robust real-time 3D object tracking. Lecture Notes in Computer Science, 1542: 135–150, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vieux:1999:FTC

[106] W. E. Vieux, K. Schwerdt, and J. L. Crowley. Face-tracking and coding for video compression. Lecture Notes in Computer Science, 1542:151–161, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prassler:1999:TPR

[107] E. Prassler, J. Scholz, and A. Elfes. Tracking people in a railway station during rush-hour. Lecture Notes in Computer Science, 1542:162–179, 1999.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paulus:1999:AKB

[108] D. Paulus, U. Ahlrichs, B. Heigl, and J. Denzler. Active knowledgebased scene analysis. Lecture Notes in Computer Science, 1542:180–199, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thonnat:1999:EII

[109] M. Thonnat, S. Moisan, and M. Crubezy. Experience in integrating image processing programs. Lecture Notes in Computer Science, 1542:200–215, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coradeschi:1999:IVD

[110] S. Coradeschi, L. Karlsson, and K. Nordberg. Integration of vision and decision-making in an autonomous airborne vehicle for traffic surveillance. Lecture Notes in Computer Science, 1542:216–230, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wachsmuth:1999:MIV

[111] S. Wachsmuth, H. Brandt-Pook, G. Socher, and F. Kummert. Multilevel integration of vision and speech understanding using Bayesian networks. Lecture Notes in Computer Science, 1542: 231–254, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oliver:1999:BCV

[112] N. Oliver, B. Rosario, and A. Pentland. A Bayesian computer vision system for modeling human interaction. Lecture Notes in Computer Science, 1542:255–272, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jebara:1999:ARL

[113] T. Jebara and A. Pentland. Action reaction learning: Automatic visual analysis and synthesis of interactive behaviour. Lecture Notes in Computer Science, 1542:273–292, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hernandez-Sosa:1999:GMP

[114] D. Hernandez-Sosa, J. Lorenzo-Navarro, M. Hernandez-Tejera, and J. Cabrera-Gamez. A generic model for perception-action systems analysis of a knowledge-based prototype. Lecture Notes in Computer Science, 1542:293– 311, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Doods:1999:HVA

[115] Z. Doods, M. Jaegersand, G. Hager, and K. Toyama. A hierarchical vision architecture for robotic manipulation tasks. Lecture Notes in Computer Science, 1542:312–330, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krautgartner:1999:OIP

[116] P. Krautgartner and M. Vincze. Optimal image processing architecture for active vision. Lecture Notes in Computer Science, 1542:331–347, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benitez:1999:RCV

[117] D. Benitez and J. Cabrera. Reactive computer vision system with reconfigurable architecture. Lecture Notes in Computer Science, 1542:348–360, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steinhaus:1999:MME

[118] P. Steinhaus, M. Ehrenmann, and R. Dillmann. MEPHISTO: A modular and extensible path planning system using observation. Lecture Notes in Computer Science, 1542:361–375, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hernandez:1999:DAV

[119] M. Hernandez, J. Cabrera, A. Dominguez, and M. Castrillon. DESEO: An active vision system for detection, tracking and recognition. Lecture Notes in Computer Science, 1542:376–391, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bollmann:1999:PDC

[120] M. Bollmann, R. Hoischen, M. Jesikiewicz, and C. Justkowski. Playing domino: A casy study for an active vision system. Lecture Notes in Computer Science, 1542:392–411, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barreto:1999:IAV

[121] J. Barreto, P. Peixoto, J. Batista, and H. Araujo. Improving 3D active visual tracking. Lecture Notes in Computer Science, 1542:412–431, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

SantosConde:1999:SSB

[122] J. E. Santos Conde, A. Teuner, S.-B. Park, and B. J. Hosticka. Surveillance system based on detection and tracking of moving objects using CMOS imagers. Lecture Notes in Computer Science, 1542:432–449, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parra:1999:DMR

[123] C. Parra, R. Murrieta-Cid, M. Devy, and M. Briot. 3-D modelling and robot localization from visual and range data in natural scenes. Lecture Notes in Computer Science, 1542:450–468, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marengoni:1999:AIV

[124] M. Marengoni, C. Jaynes, A. Hanson, and E. Riseman. Ascender II, a visual framework for 3D reconstruction. Lecture Notes in Computer Science, 1542:469–488, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolfart:1999:HAC

[125] E. Wolfart, V. Sequeira, K. Ng, and S. Butterfield. Hybrid approach to the construction of triangulated 3D models of building interiors. Lecture Notes in Computer Science, 1542:489–508, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Young:1999:CCS

[126] R. Young, J. Matas, and J. Kittler. On camera calibration for scene model acquisition and maintenance using an active vision system. Lecture Notes in Computer Science, 1542:509–521, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Draper:1999:AAO

[127] B. A. Draper, J. Bins, and K. Baek. ADORE: Adaptive object recognition. Lecture Notes in Computer Science, 1542:522–537, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jonquieres:1999:ORG

[128] S. Jonquieres, M. Devy, F. Huynh, and M. Khatib. Object recognition for a grasping task by a mobile manipulator. Lecture Notes in Computer Science, 1542:538–551, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chaochen:1999:DCL

[129] Zhou Chaochen. Duration calculus, a logical approach to real-time systems. Lecture Notes in Computer Science, 1548:1–7, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1548/15480001.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1548/15480001.
pdf.

Pigozzi:1999:AAL

[130] Don Pigozzi. Abstract algebraic logic. Lecture Notes in Computer Science, 1548:8–16, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1548/15480008.htm;   http:
//link.springer-ny.com/link/service/

series/0558/papers/1548/15480008.
pdf.

Maibaum:1999:SRS

[131] Thomas S. E. Maibaum, Pauline Kan, and Kevin Lano. Systematising reactive system design. Lecture Notes in Computer Science, 1548:17– 22, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1548/15480017.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1548/15480017.
pdf.

Zave:1999:SDC

[132] Pamela Zave. Systematic design of call-coverage features. Lecture Notes in Computer Science, 1548:23– 27, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1548/15480023.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1548/15480023.
pdf.

Manna:1999:VAT

[133] Zohar Manna, Anca Browne, Henny B. Sipma, and Tomás E. Uribe. Visual abstractions for temporal verification. Lecture Notes in Computer Science, 1548:28–41, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1548/15480028.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1548/15480028.
pdf.

Winskel:1999:LMC

[134] Glynn Winskel. A linear metalanguage for concurrency. Lecture Notes in Computer Science, 1548:42– 58, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1548/15480042.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1548/15480042.
                                        pdf.

Yoneda:1999:VBD

[135] Tomohiro Yoneda, Bin Zhou, and Bernd-Holger Schlingloff. Verification of bounded delay asynchronous circuits with timed traces. Lecture Notes in Computer Science, 1548:59– 73, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1548/15480059.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1548/15480059.
                                        pdf.

Groote:1999:VTP

[136] Jan Friso Groote and Radu Mateescu. Verification of temporal properties of processes in a setting with data. Lecture Notes in Computer Science, 1548: 74–90, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1548/15480074.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1548/15480074.
                                        pdf.

Baum:1999:LRT

[137] Gabriel A. Baum, Marcelo F. Frias, and Thomas S. E. Maibaum. A logic for real-time systems specification, its algebraic semantics, and equational calculus. Lecture Notes in Computer Science, 1548:91–105, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1548/15480091.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1548/15480091.
pdf.

Sutre:1999:ERM

[138] Gregoire Sutre, Alain Finkel, Olivier Roux, and Franck Cassez. Effective recognizability and model checking of reactive fiffo automata. Lecture Notes in Computer Science, 1548:106– 123, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1548/15480106.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1548/15480106.
pdf.

Buth:1999:CML

[139] Bettina Buth, Jan Peleska, and Hui Shi. Combining methods for the livelock analysis of a faulttolerant system. Lecture Notes in Computer Science, 1548:124–139, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1548/15480124.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1548/15480124.
                                        pdf.

Heinemann:1999:SSM

[140] Bernhard Heinemann. Separating sets by modal formulas. Lecture Notes in Computer Science, 1548:140–153, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1548/15480140.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1548/15480140.
                                        pdf.

Marx:1999:IML

[141] Maarten Marx. Interpolation in modal logic. Lecture Notes in Computer Science, 1548:154–163, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1548/15480154.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1548/15480154.
                                        pdf.

DePaiva:1999:BML

[142] V. De Paiva and A. Schalk. Building models of linear logic. Lecture Notes in Computer Science, 1548:164–177, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paiva:1999:BML

[143] Valeria de Paiva and Andrea Schalk. Building models of linear logic. Lecture Notes in Computer Science, 1548: 164–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1548/15480164.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1548/15480164.
pdf.

Schorlemmer:1999:TRL

[144] W. Marco Schorlemmer. Term rewriting in a logic of special relations. Lecture Notes in Computer Science, 1548: 178–195, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1548/15480178.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1548/15480178.
pdf.

Hannay:1999:ABE

[145] Jo Erskine Hannay. Abstraction barriers in equational proof. Lecture Notes in Computer Science, 1548:196– 213, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1548/15480196.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1548/15480196.
pdf.

Debbabi:1999:SBM

[146] Mourad Debbabi, Abdelkader Benzakour, and Béchir Ktari. A synergy between model-checking and type inference for the verification of valuepassing higher-order processes. Lecture Notes in Computer Science, 1548:214– 230, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/1548/15480214.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1548/15480214.
                                        pdf.

Dingel:1999:TBR

[147] Jürgen Dingel. A trace-based refinement calculus for shared-variable parallel programs. Lecture Notes in Computer Science, 1548:231–247, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1548/15480231.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1548/15480231.
                                        pdf.

Steen:1999:CPP

[148] Maarten Steen, John Derrick, Eerke Boiten, and Howard Bowman. Consistency of partial process specifications. Lecture Notes in Computer Science, 1548:248–262, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1548/15480248.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1548/15480248.pdf.

Hennicker:1999:OL

[149] Rolf Hennicker and Michel Bidoit. Observational logic. Lecture Notes in Computer Science, 1548:263–277, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1548/15480263.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1548/15480263.
pdf.

VanGlabbeek:1999:SA

[150] R. Van Glabbeek and P. Rittgen. Scheduling algebra. Lecture Notes in Computer Science, 1548:278–292, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glabbeek:1999:SA

[151] Rob van Glabbeek and Peter Rittgen. Scheduling algebra. Lecture Notes in Computer Science, 1548:278–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1548/15480278.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1548/15480278.
pdf.

Fiadeiro:1999:ASC

[152] José Luiz Fiadeiro and Antónia Lopes. Algebraic semantics of coordination or what is in a signature? Lecture Notes in Computer Science, 1548:293– 307, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1548/15480293.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1548/15480293.
pdf.

Silva:1999:AAC

[153] Leila Silva, Augusto Sampaio, Edna Barros, and Juliano Iyoda. An algebraic approach to combining processes in a hardware/software partitioning environment. Lecture Notes in Computer Science, 1548:308–324, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1548/15480308.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1548/15480308.
pdf.

Cenciarelli:1999:AVP

[154] Pietro Cenciarelli. An algebraic view of program composition. Lecture Notes in Computer Science, 1548:325–340, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1548/15480325.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1548/15480325.
pdf.

Bidoit:1999:ASC

[155] Michel Bidoit, Donald Sannella, and Andrzej Tarlecki. Architectural specifications in CASL. Lecture Notes in Computer Science, 1548:341–357, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1548/15480341.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1548/15480341.
pdf.

Quaglia:1999:PCC

[156] Paola Quaglia. Pi-congruences as CCS equivalences. Lecture Notes in Computer Science, 1548:358–372, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1548/15480358.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1548/15480358.
pdf.

Kirchner:1999:ASH

[157] Hélène Kirchner and Peter D. Mosses. Algebraic specifications, higher-order types, and set-theoretic models. Lecture Notes in Computer Science, 1548: 373–388, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1548/15480373.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1548/15480373.
pdf.

Drabent:1999:TAC

[158] Włodzimierz Drabent and Paweł Pietrzak. Type analysis for CHIP. Lecture Notes in Computer Science, 1548: 389–405, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1548/15480389.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1548/15480389.
pdf.

Erwig:1999:CPA

[159] Martin Erwig. Categorical programming with abstract data types. Lecture Notes in Computer Science, 1548:406– 421, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1548/15480406.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1548/15480406.
                                        pdf.

RuizJimenez:1999:CLP

[160] B. C. Ruiz Jimenez. Condensing lemmas for pure type systems with universes. Lecture Notes in Computer Science, 1548:422–437, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jimenez:1999:CLP

[161] Blas C. Ruiz Jiménez. Condensing lemmas for pure type systems with universes. Lecture Notes in Computer Science, 1548:422–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1548/15480422.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1548/15480422.
                                        pdf.

Almendros-Jimenez:1999:ICT

[162] Jesús Manuel Almendros-Jiménez. Improving computations in a typed functional logic language. Lecture Notes in Computer Science, 1548:438–454, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1548/15480438.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1548/15480438.
                                        pdf.

Spoto:1999:AIP

[163] Fausto Spoto and Giorgio Levi. Abstract interpretation of Prolog programs. Lecture Notes in Computer Science, 1548:455–470, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1548/15480455.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1548/15480455.pdf.

Bagnara:1999:FEV

[164] Roberto Bagnara and Peter Schachte. Factorizing equivalent variable pairs in ROBDD-based implementations of pos. Lecture Notes in Computer Science, 1548:471–485, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1548/15480471.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1548/15480471.pdf.

Martini:1999:SPA

[165] Alfio Martini and Uwe Wolter. A single perspective on arrows between institutions. Lecture Notes in Computer Science, 1548:486–501, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1548/15480486.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1548/15480486.
pdf.

Machado:1999:OIT

[166] Patrícia D. L. Machado. On oracles for interpreting test results against algebraic specifications. Lecture Notes in Computer Science, 1548:502–518, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1548/15480502.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1548/15480502.
                                        pdf.

Gassmann:1999:REE

[167] Zoltan Gassmann, Luis Mandel, and Roshan Sembacuttiaratchy. R ECOPLA : An extendible graphic meta-editor. Lecture Notes in Computer Science, 1548:519–521, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1548/15480519.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1548/15480519.
                                        pdf.

Grahlmann:1999:SP

[168] Bernd Grahlmann. The state of PEP. Lecture Notes in Computer Science, 1548:522–526, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1548/15480522.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1548/15480522.pdf.

DeMoura:1999:ASA

[169] H. P. De Moura and L. C. De Sousa Menezes. The ABACO system — an Algebraic Based Action COmpiler. Lecture Notes in Computer Science, 1548:527–529, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moura:1999:BSA

[170] Hermano Perrelli de Moura and Luis Carlos de Sousa Menezes. The A BACO system — an algebraic based action COmpiler. Lecture Notes in Computer Science, 1548:527–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1548/15480527.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1548/15480527.
pdf.

Pettersson:1999:CNS

[171] M. Pettersson. Compiling natural semantics. Lecture Notes in Computer Science, 1549:i–xvi, 1–240, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paulson:1999:IAIa

[172] Lawrence C. Paulson. Inductive analysis of the Internet Protocol TLS (position paper). Lecture Notes in Computer Science, 1550:1–12, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1550/15500001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1550/15500001.
pdf.

Anonymous:1999:IAI

[173] Anonymous. Inductive analysis of the Internet Protocol TLS: Discussion. Lecture Notes in Computer Science, 1550:13–23, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paulson:1999:IAIb

[174] Lawrence C. Paulson. Inductive analysis of the Internet Protocol TLS (transcript of discussion). Lecture Notes in Computer Science, 1550:13– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1550/15500013.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1550/15500013.
                                        pdf.

Foley:1999:ECVa

[175] Simon N. Foley. External consistency and the verification of security protocols (position paper). Lecture Notes in Computer Science, 1550:24– 27, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1550/15500024.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1550/15500024.
                                        pdf.

Anonymous:1999:ECV

[176] Anonymous. External consistency and the verification of security protocols: Discussion. Lecture Notes in Computer Science, 1550:28–35, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Foley:1999:ECVb

[177] Simon N. Foley. External consistency and the verification of security protocols (transcript of discussion). Lecture Notes in Computer Science, 1550:28– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1550/15500028.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1550/15500028.
pdf.

Ellison:1999:TSGa

[178] Carl Ellison. The trust shell game (position paper). Lecture Notes in Computer Science, 1550:36–40, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1550/15500036.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1550/15500036.
pdf.

Anonymous:1999:TSG

[179] Anonymous. The trust shell game: Discussion. Lecture Notes in Computer Science, 1550:41–44, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ellison:1999:TSGb

[180] Carl Ellison. The trust shell game (transcript of discussion). Lecture Notes in Computer Science, 1550:41– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1550/15500041.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1550/15500041.
pdf.

Feigenbaum:1999:OALa

[181] Joan Feigenbaum. Overview of the AT&T Labs trust-management project (position paper). Lecture Notes in Computer Science, 1550:45– 50, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1550/15500045.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1550/15500045.
                                        pdf.

Anonymous:1999:OAL

[182] Anonymous. Overview of the AT&T Labs trust-management project: Discussion. Lecture Notes in Computer Science, 1550:51–58, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feigenbaum:1999:OALb

[183] Joan Feigenbaum. Overview of the AT&T labs trust-management project (transcript of discussion). Lecture Notes in Computer Science, 1550:51– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1550/15500051.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1550/15500051.
                                        pdf.

Blaze:1999:KTM

[184] Matt Blaze, Joan Feigenbaum, and Angelos D. Keromytis. KeyNote: Trust management for public-key infrastructures (position paper). Lecture Notes in Computer Science, 1550:59– 63, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1550/15500059.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1550/15500059.
pdf.

Anonymous:1999:DST

[185] Anonymous. Discussion session — trust management. Lecture Notes in Computer Science, 1550:64–66, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1550/15500064.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1550/15500064.
pdf.

Gligor:1999:AOSa

[186] Virgil D. Gligor and Serban I. Gavrila. Application-oriented security policies and their composition (position paper). Lecture Notes in Computer Science, 1550:67–74, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1550/15500067.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1550/15500067.
pdf.

Anonymous:1999:AOS

[187] Anonymous. Application-oriented security policies and their composition: Discussion. Lecture Notes in Computer Science, 1550:75–82, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gligor:1999:AOSb

[188] Virgil D. Gligor. Application-oriented security policies and their composition (transcript of discussion). Lecture Notes in Computer Science, 1550:75– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1550/15500075.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1550/15500075.
                                        pdf.

Yoshiura:1999:SFU

[189] Hiroshi Yoshiura, Ryoichi Sasaki, and Kazuo Takaragi. Secure fingerprinting using public-key cryptography (position paper). Lecture Notes in Computer Science, 1550:83– 89, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1550/15500083.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1550/15500083.
                                        pdf.

Yshiura:1999:SFU

[190] H. Yshiura, R. Sasaki, and K. Takaragi. Secure fingerprinting using public-key cryptography. Lecture Notes in Computer Science, 1550:83–??, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sasaki:1999:SFU

[191] Ryoichi Sasaki. Secure fingerprinting using public-key cryptography (transcript of discussion). Lecture Notes in Computer Science, 1550:90– 94, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1550/15500090.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1550/15500090.
pdf.

Bergadano:1999:TPCa

[192] F. Bergadano and P. De Mauro. Third party certification of HTTP service access statistics (position paper). Lecture Notes in Computer Science, 1550:95– 99, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1550/15500095.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1550/15500095.
pdf.

Anonymous:1999:TPC

[193] Anonymous. Third party certification of HTTP service access statistics: Discussion. Lecture Notes in Computer Science, 1550:100–107, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bergadano:1999:TPCb

[194] F. Bergadano. Third party certification of HTTP service access statistics (transcript of discussion). Lecture Notes in Computer Science, 1550:100– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1550/15500100.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1550/15500100.
pdf.

Harbison:1999:DTT

[195] William S. Harbison. Delegating trust (transcript of discussion). Lecture Notes in Computer Science, 1550:108– 117, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1550/15500108.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1550/15500108.
                                        pdf.

Crispo:1999:DRP

[196] Bruno Crispo. Delegation of responsibility (position paper). Lecture Notes in Computer Science, 1550:118–124, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1550/15500118.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1550/15500118.
                                        pdf.

Crispo:1999:DRT

[197] Bruno Crispo. Delegation of responsibility (transcript of discussion). Lecture Notes in Computer Science, 1550:125– 130, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1550/15500125.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1550/15500125.
                                        pdf.

Lomas:1999:APT

[198] Mark Lomas. Abuse of process (transcript of discussion). Lecture Notes in Computer Science, 1550:131–136, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1550/15500131.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1550/15500131.
pdf.

Landrock:1999:NCPa

[199] Peter Landrock. A new concept in protocols: Verifiable computational delegation (position paper). Lecture Notes in Computer Science, 1550:137– 145, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1550/15500137.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1550/15500137.
pdf.

Landrock:1999:NCPb

[200] Peter Landrock. A new concept in protocols: Verifiable computational delegation (transcript of discussion). Lecture Notes in Computer Science, 1550:146–153, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1550/15500146.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1550/15500146.
pdf.

Christianson:1999:DSCa

[201] Bruce Christianson and James A. Malcolm. Delegation and not-so smart cards (position paper). Lecture Notes in Computer Science, 1550:154–157, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1550/15500154.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1550/15500154.
                                        pdf.

Christianson:1999:DSCb

[202] Bruce Christianson. Delegation and not-so smart card (transcript of discussion). Lecture Notes in Computer Science, 1550:158–167, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1550/15500158.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1550/15500158.
                                        pdf.

Roe:1999:CDT

[203] Michael Roe. Certification and delegation (transcript of discussion). Lecture Notes in Computer Science, 1550:168– 176, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1550/15500168.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1550/15500168.
                                        pdf.

Gligor:1999:DSD

[204] Virgil D. Gligor, Peter Landrock, Mark Lomas, Raphael Yahalom, and John Warne. Discussion session: Differences between academic and commercial security. Lecture Notes in Computer Science, 1550:177–192, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1550/15500177.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1550/15500177.
pdf.

Yahalom:1999:OTRa

[205] Raphael Yahalom. Optimistic trust with realistic eNvestigators (position paper). Lecture Notes in Computer Science, 1550:193–202, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1550/15500193.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1550/15500193.
pdf.

Yahalom:1999:OTRb

[206] Raphael Yahalom. Optimistic trust with realistic eNvestigators (transcript of discussion). Lecture Notes in Computer Science, 1550:203–212, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1550/15500203.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1550/15500203.
pdf.

Gollmann:1999:IFP

[207] Dieter Gollmann. Insider fraud (position paper). Lecture Notes in Computer Science, 1550:213–219, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1550/15500213.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1550/15500213.
                                        pdf.

Gollmann:1999:IFT

[208] Dieter Gollmann. Insider fraud (transcript of discussion). Lecture Notes in Computer Science, 1550:220–226, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1550/15500220.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1550/15500220.
                                        pdf.

Lee:1999:PSF

[209] Stewart Lee, Joan Feigenbaum, Virgil D. Gligor, and Bruce Christianson. Panel session — future directions. Lecture Notes in Computer Science, 1550:227–240, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1550/15500227.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1550/15500227.
                                        pdf.

Anonymous:1999:IC

[210] Anonymous. Index of contributors. Lecture Notes in Computer Science, 1550:241–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/papers/1550/15500241.pdf.

Anonymous:1999:C

[211] Anonymous.  Contents.  Lecture

Notes in Computer Science, 1550: vii, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/1550/15500VII.pdf.

Anonymous:1999:P

[212] Anonymous.  Preface.  Lecture

Notes in Computer Science, 1550: v, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/1550/1550000V.pdf.

Moss:1999:ABF

[213] Graeme E. Moss and Colin Runciman. Automated benchmarking of functional data structures. Lecture Notes in Computer Science, 1551:1– 15, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1551/15510001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1551/15510001.
pdf.

Cadoli:1999:NSE

[214] Marco Cadoli, Luigi Palopoli, Andrea Schaerf, and Domenico Vasile. NP — SPEC: An executable specification language for solving all problems in NP. Lecture Notes in Computer Science, 1551:16–30, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1551/15510016.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1551/15510016.
pdf.

Letelier:1999:PRS

[215] Patricio Letelier, Pedro Sánchez, and Isidro Ramos. Prototyping a requirements specification through an automatically generated concurrent logic program. Lecture Notes in Computer Science, 1551:31–45, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1551/15510031.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1551/15510031.
pdf.

Bozzano:1999:MAS

[216] Marco Bozzano, Giorgio Delzanno, Maurizio Martelli, Viviana Mascardi, and Floriano Zini. Multi-agent systems development as a software engineering enterprise. Lecture Notes in Computer Science, 1551:46–60, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1551/15510046.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1551/15510046.
pdf.

Elliott:1999:FAS

[217] Conal Elliott. From functional animation to sprite-based display. Lecture Notes in Computer Science, 1551:61– 75, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1551/15510061.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1551/15510061.
pdf.

Kahl:1999:BPP

[218] Wolfram Kahl. Beyond pretty-printing: Galley concepts in document formatting combinators. Lecture Notes in Computer Science, 1551:76–90, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1551/15510076.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1551/15510076.
pdf.

Peterson:1999:LMC

[219] John Peterson, Paul Hudak, and Conal Elliott. Lambda in motion: Controlling robots with Haskell. Lecture Notes in Computer Science, 1551:91– 105, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1551/15510091.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1551/15510091.
pdf.

Demoen:1999:CCH

[220] Bart Demoen and Konstantinos Sagonas. CHAT: The copy-hybrid approach to tabling. Lecture Notes in Computer Science, 1551:106–121, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1551/15510106.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1551/15510106.
                                        pdf.

Silva:1999:IPC

[221] M. G. Silva, I. C. Dutra, R. Bianchini, and V. Santos Costa. The influence of parallel computer architectural parameters on the performance of parallel logic programming systems. Lecture Notes in Computer Science, 1551:122–136, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Silva:1999:IAP

[222] Marcio G. Silva, Inês C. Dutra, Ricardo Bianchini, and Vítor Santos Costa. The influence of architectural parameters on the performance of parallel logic programming systems. Lecture Notes in Computer Science, 1551:122– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1551/15510122.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1551/15510122.
                                        pdf.

Rocha:1999:PWT

[223] Ricardo Rocha, Fernando Silva, and Vítor Santos Costa. Or-parallelism within tabling. Lecture Notes in Computer Science, 1551:137–151, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1551/15510137.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1551/15510137.
pdf.

Mattsson:1999:MDR

[224] Håkan Mattsson, Hans Nilsson, and Claes Wikström. Mnesia — A distributed robust DBMS for telecommunications applications. Lecture Notes in Computer Science, 1551:152–163, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1551/15510152.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1551/15510152.
pdf.

Siddabathuni:1999:ABI

[225] Babu Siddabathuni, Suzanne W. Dietrich, and Susan D. Urban. An AQUAbased intermediate language for evaluating an active deductive objectoriented language. Lecture Notes in Computer Science, 1551:164–178, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1551/15510164.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1551/15510164.
pdf.

Hakli:1999:IDS

[226] Raul Hakli, Matti Nykänen, Hellis Tamm, and Esko Ukkonen. Implementing a declarative string query language with string restructuring. Lecture Notes in Computer Science, 1551:179– 195, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1551/15510179.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1551/15510179.
                                        pdf.

Meijer:1999:CSW

[227] Erik Meijer, Daan Leijen, and James Hook. Client-side Web scripting with HaskellScript. Lecture Notes in Computer Science, 1551:196–210, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1551/15510196.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1551/15510196.
                                        pdf.

Jeffery:1999:MCB

[228] David Jeffery, Tyson Dowd, and Zoltan Somogyi. MCORBA: A CORBA binding for mercury. Lecture Notes in Computer Science, 1551:211–227, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1551/15510211.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1551/15510211.
                                        pdf.

Xi:1999:DCE

[229] Hongwei Xi. Dead code elimination through dependent types. Lecture Notes in Computer Science, 1551:228– 242, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1551/15510228.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1551/15510228.
pdf.

Ferriera:1999:MSW

[230] M. Ferriera and L. Damas. Multiple specialization of WAM code. Lecture Notes in Computer Science, 1551:243– 258, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferreira:1999:MSW

[231] Michel Ferreira and Luís Damas. Multiple specialization of WAM code. Lecture Notes in Computer Science, 1551: 243–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1551/15510243.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1551/15510243.
pdf.

Vasconcelos:1999:FFD

[232] Wamberto Weber Vasconcelos. A flexible framework for dynamic and static slicing of logic programs. Lecture Notes in Computer Science, 1551:259– 274, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1551/15510259.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1551/15510259.
pdf.

Bansal:1999:ALP

[233] Arvind K. Bansal and Peer Bork. Applying logic programming to derive novel functional information of genomes. Lecture Notes in Computer Science, 1551:275–289, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1551/15510275.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1551/15510275.
pdf.

Watson:1999:AAT

[234] Richard Watson. An application of action theory to the space shuttle. Lecture Notes in Computer Science, 1551:290– 304, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1551/15510290.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1551/15510290.
pdf.

Soininen:1999:DDR

[235] Timo Soininen and Ilkka Niemelä. Developing a declarative rule language for applications in product configuration. Lecture Notes in Computer Science, 1551:305–319, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1551/15510305.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1551/15510305.
pdf.

Goltz:1999:UTU

[236] Hans-Joachim Goltz and Dirk Matzke. University timetabling using constraint logic programming. Lecture Notes in Computer Science, 1551:320–334, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1551/15510320.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1551/15510320.
pdf.

Carlsson:1999:CBR

[237] Mats Carlsson, Per Kreuger, and Emil Åström. Constraint-based resource allocation and scheduling in steel manufacturing. Lecture Notes in Computer Science, 1551:335–349, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1551/15510335.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1551/15510335.
pdf.

Lai:1999:UCL

[238] Claude Laï. Using constraints in local proofs for CLP debugging. Lecture Notes in Computer Science, 1551:350– 359, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1551/15510350.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1551/15510350.
pdf.

Laie:1999:UCL

[239] C. Laie. Using constraints in local proofs for CLP debugging. Lecture Notes in Computer Science, 1551:350– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmidt:1999:RER

[240] David A. Schmidt. A return to elegance: The reapplication of declarative notation to software design. Lecture Notes in Computer Science, 1551:360– 364, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1551/15510360.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1551/15510360.
                                        pdf.

Wallace:1999:EDS

[241] Mark Wallace and Joachim Schimpf. ECLiPSe: Declarative specification and scaleable implementation. Lecture Notes in Computer Science, 1551:365– 366, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1551/15510365.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1551/15510365.
                                        pdf.

Mouaddib:1999:FAO

[242] N. Mouaddib and G. Raschia. A fuzzy attribute-oriented induction method for knowledge discovery in relational databases. Lecture Notes in Computer Science, 1552:1–13, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goh:1999:FMM

[243] C.-L. Goh, M. Tsukamoto, and S. Nishio. Fast methods with magic sampling for knowledge discovery in deductive databases with large deduction results. Lecture Notes in Computer Science, 1552:14–28, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:1999:HMC

[244] Ceil Chua Eng Huang, R. H. L. Chiang, and E.-P. Lim. A heuristic method for correlating attribute group pairs in data mining. Lecture Notes in Computer Science, 1552:29–40, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abraham:1999:IMM

[245] T. Abraham and J. F. Roddick. Incremental meta-mining from large temporal data sets. Lecture Notes in Computer Science, 1552:41–54, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Choenni:1999:SGB

[246] S. Choenni. On the suitability of genetic-based algorithms for data mining. Lecture Notes in Computer Science, 1552:55–67, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhowmick:1999:DVW

[247] S. S. Bhowmick, S. K. Madria, W.-K. Ng, and E.-P. Lim. Data visualization in a Web warehouse. Lecture Notes in Computer Science, 1552:68–80, 1999.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Samtani:1999:RAR

[248] S. Samtani, M. Mohania, V. Kumar, and Y. Kambayashi. Recent advances and research problems in data warehousing. Lecture Notes in Computer Science, 1552:81–92, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhowmick:1999:WWD

[249] S. S. Bhowmick, S. K. Madria, W.-K. Ng, and E.-P. Lim. Web warehousing: Design and issues. Lecture Notes in Computer Science, 1552:93–104, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sapia:1999:ERM

[250] C. Sapia, M. Blaschka, G. Hoefling, and B. Dinter. Extending the E/ R model for the multidimensional paradigm. Lecture Notes in Computer Science, 1552:105–116, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stumme:1999:NAD

[251] G. Stumme and K. E. Wolff. Numerical aspects in the data model of conceptual information systems. Lecture Notes in Computer Science, 1552:117–128, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dekeyser:1999:NDC

[252] S. Dekeyser, B. Kuijpers, J. Paredaens, and J. Wijsen. Nested data cubes for OLAP. Lecture Notes in Computer Science, 1552:129–140, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roddick:1999:DWD

[253] J. F. Roddick. Data warehousing and data mining — are we working on the right things? Lecture Notes in Computer Science, 1552:141–144, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Estrella:1999:DED

[254] F. Estrella, Z. Kovacs, J.-M. Le Goff, and R. McClatchey. The design of an engineering data warehouse based on meta-object structures. Lecture Notes in Computer Science, 1552:145–156, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1999:TVC

[255] H. Kim and S. Park. Two version concurrency control algorithm with query locking for decision support. Lecture Notes in Computer Science, 1552:157– 168, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoewon:1999:TVC

[256] K. Hoewon and P. Seog. Two version concurrency control algorithm with query locking for decision support. Lecture Notes in Computer Science, 1552: 157–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ling:1999:EVM

[257] Tok Wang Ling and Y. Liu. An efficient view maintenance algorithm for data warehousing. Lecture Notes in Computer Science, 1552:169–180, 1999.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cuce:1999:ACV

[258] S. Cuce and A. Zaslavsky. Adaptive cache validation for mobile file systems. Lecture Notes in Computer Science, 1552:181–192, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lam:1999:BSM

[259] K.-Y. Lam, E. Chan, and J. Chun-Hung Yuen. Broadcast strategies to maintain cached data for mobile computing system. Lecture Notes in Computer Science, 1552:193–204, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gaedke:1999:WCD

[260] M. Gaedke, M. Beigl, H.-W. Gellersen, and C. Segor. Web content delivery to heterogeneous mobile platforms. Lecture Notes in Computer Science, 1552:205–217, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hu:1999:DDD

[261] Q. Hu, Dik Lun Lee, and W.-C. Lee. Dynamic data delivery in wireless communication environments. Lecture Notes in Computer Science, 1552:218– 229, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Qinglong:1999:DDD

[262] H. Qinglong, Dik Lun Lee, and W.-C. Lee. Dynamic data delivery in wireless communication environments. Lecture Notes in Computer Science, 1552:

218–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pitoura:1999:SIB

[263] E. Pitoura. Scalable invalidation-based processing of queries in broadcast push delivery. Lecture Notes in Computer Science, 1552:230–241, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Madria:1999:TDR

[264] S. K. Madria. Timestamps to detect R-W conflicts in mobile computing. Lecture Notes in Computer Science, 1552:242–253, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guy:1999:RMD

[265] R. Guy, P. Reiher, D. Rather, and M. Gunter. Rumor: Mobile data access through optimistic peer-to-peer replication. Lecture Notes in Computer Science, 1552:254–265, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tagashira:1999:CUM

[266] S. Tagashira, K. Saisho, F. Inada, and A. Fukuda. A copy update mechanism for a mobile information announcement system — transmitting non-storage type resources with redundancy. Lecture Notes in Computer Science, 1552:266–277, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hwang:1999:TOR

[267] S.-Y. Hwang and J.-K. Chiu. Towards optimal replication for hierarchical location management in wireless systems. Lecture Notes in Computer Science, 1552:278–289, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pillai:1999:DCM

[268] R. Radhakrishna Pillai, M. Rangnath, R. Agrawal, and W. Wang. Distributed control of mobile ATM networks using CORBA. Lecture Notes in Computer Science, 1552:290–300, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cameron:1999:NTE

[269] F. Cameron, M. Zukerman, and M. Gitlits. New thoughts on effects of TCP slow start and FEC coding in WATM access networks. Lecture Notes in Computer Science, 1552:301–310, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Verroca:1999:CCD

[270] F. Verroca, C. Eynard, G. Ghinamo, and G. Gentile. A centralised cellular database to support network management process. Lecture Notes in Computer Science, 1552:311–322, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacobs:1999:SCM

[271] S. Jacobs. Security of current mobile IP solutions. Lecture Notes in Computer Science, 1552:323–334, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wu:1999:ADF

[272] S.-Y. Wu and C.-S. Chang. An active database framework for adaptive mobile data access. Lecture Notes in Computer Science, 1552:335–346, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leung:1999:DIM

[273] C.-H. Leung, K.-M. Cheung, and T.- F. Ngai. Design and implementation of a mobile application support system. Lecture Notes in Computer Science, 1552:347–357, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Soutou:1999:PIS

[274] C. Soutou, T. Val, F. Peyrard, and J.-J. Mercier. Proposition of an identification service on a hybrid WLAN with mobiles stations. Lecture Notes in Computer Science, 1552:358–367, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hamano:1999:DNS

[275] T. Hamano, H. Takakura, and Y. Kambayashi. A dynamic navigation system based on user’s geographical situation. Lecture Notes in Computer Science, 1552:368–379, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1999:FMC

[276] W.-C. Lee. Future of mobile computing: Convergence of research and applications. Lecture Notes in Computer Science, 1552:380–381, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thamm:1999:WSC

[277] J. Thamm, S. Wilke, and L. Wegner. A Web solution to concurrency awareness in shared data spaces. Lecture Notes in Computer Science, 1552:382–395, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Inoue:1999:SMM

[278] S. Inoue and M. Iwaihara. Structured message management for group interaction. Lecture Notes in Computer Science, 1552:396–407, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tarumi:1999:GAD

[279] H. Tarumi, T. Matsuyama, and Y. Kambayashi. Group activity database for groupware evolution. Lecture Notes in Computer Science, 1552: 408–420, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Furukawa:1999:SCW

[280] T. Furukawa, H. Xu, and Y. Shi. Supporting collaborative work by processbased transaction model. Lecture Notes in Computer Science, 1552:421–433, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sadakane:1999:CDT

[281] K. Sadakane and H. Imai. A cooperative distributed text database management method unifying search and compression based on the Burrows– Wheeler transformation. Lecture Notes in Computer Science, 1552:434–445, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kato:1999:CSD

[282] H. Kato and M. Yoshikawa. Constructing structured document views (short paper). Lecture Notes in Computer Science, 1552:446–453, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Erwig:1999:TOS

[283] M. Erwig, M. Schneider, and R. H. Gueting. Temporal objects for spatiotemporal data models and a comparison of their representations. Lecture Notes in Computer Science, 1552:454– 465, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Renolen:1999:TST

[284] A. Renolen. The temporal set-theory, tropology, and their application in spatiotemporal modelling (short paper). Lecture Notes in Computer Science, 1552:466–473, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ishii:1999:TES

[285] Y. Ishii, T. Ishizaka, N. Mohan, and J. Feng. TimeCube: Efficient storage, access and analysis of temporal (historical) data (short paper). Lecture Notes in Computer Science, 1552:474–483, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nishida:1999:PST

[286] S. Nishida, H. Nozawa, and N. Saiwaki. Proposal of spatial-temporal indexing methods for moving objects. Lecture Notes in Computer Science, 1552:484– 495, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakamura:1999:STD

[287] Y. Nakamura, H. Dekihara, and R. Furukawa. Spatio-temporal data management for moving objects using the PMD-tree. Lecture Notes in Computer Science, 1552:496–507, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1999:ISB

[288] B. Wang, H. Horinokuchi, S. Kuroki, and K. Kaneko. Implementing Sequoia 2000 benchmark on Shusse-Uo and its performance. Lecture Notes in Computer Science, 1552:508–519, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Horinokuchi:1999:SQP

[289] H. Horinokuchi, B. Wang, S. Kuroki, and A. Makinouchi. A spatiotemporal query processor based on simplicial representation. Lecture Notes in Computer Science, 1552:520–531, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amagasa:1999:ICL

[290] T. Amagasa, M. Aritsugi, T. Tanaka, and Y. Kanamori. Implementing class library and index for managing spatiotemporal data (short paper). Lecture Notes in Computer Science, 1552:532– 541, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohsawa:1999:SDM

[291] Y. Ohsawa and K. Kim. A spatiotemporal data management method using inverse differential script. Lecture Notes in Computer Science, 1552:542–553, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salev:1999:EES

[292] K. Salev, T. Tomii, and H. Arisawa. Extracting event semantics from video data based on real world database. Lecture Notes in Computer Science, 1552:554–567, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ryu:1999:SMV

[293] E. Ryu, Y. Kim, M.-Y. Lee, and K. Lee. Structured modeling for video databases. Lecture Notes in Computer Science, 1552:568–579, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nitsuwat:1999:AUF

[294] S. Nitsuwat and J. S. Jin. Applying unsupervised fuzzy C-prototypes clustering in motion-based segmentation (short paper). Lecture Notes in Computer Science, 1552:580–590, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carolos:1999:VSB

[295] R. P. Carolos and K. Uehara. Video summarization based on semantic representation. Lecture Notes in Computer Science, 1554:1–16, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carlos:1999:VSB

[296] Rafael Paulin Carlos and Kuniaki Uehara. Video summarization based on semantic representation. Lecture Notes in Computer Science, 1554:1– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1554/15540001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1554/15540001.
pdf.

Niikura:1999:VAO

[297] Yasuhiro Niikura, Yukinobu Taniguchi, Akihito Akutsu, and Yoshinobu Tonomura. Valbum: Album-oriented video storyboard for editing and viewing video. Lecture Notes in Computer Science, 1554:17–29, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1554/15540017.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1554/15540017.
pdf.

Fuchs:1999:BDM

[298] Henry Fuchs. Beyond the desktop metaphor: Toward more effective display, interaction, and telecollaboration in the office of the future via a multitude of sensors and displays (invited paper). Lecture Notes in Computer Science, 1554:30– 43, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1554/15540030.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1554/15540030.
pdf.

Sato:1999:MEI

[299] Imari Sato, Yoichi Sato, and Katsushi Ikeuchi. A method for estimating illumination distribution of a real scene based on soft shadows. Lecture Notes in Computer Science, 1554:44– 58, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1554/15540044.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1554/15540044.
pdf.

Tsukamoto:1999:IRS

[300] Masahiko Tsukamoto. Integrating real space and virtual space in the ‘invisible person’ communication support system. Lecture Notes in Computer Science, 1554:59–74, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1554/15540059.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1554/15540059.
pdf.

Ariki:1999:NDA

[301] Yasuo Ariki, Jun Ogata, and Masafumi Nishida. News dictation and article classification using automatically extracted announcer utterance. Lecture Notes in Computer Science, 1554:75– 86, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1554/15540075.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1554/15540075.
pdf.

Ide:1999:AVI

[302] Ichiro Ide, Koji Yamamoto, and Hidehiko Tanaka. Automatic video indexing based on shot classification. Lecture Notes in Computer Science, 1554:87– 102, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1554/15540087.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1554/15540087.
                                        pdf.

Endo:1999:MSR

[303] Takashi Endo, JianXin Zhang, Masakyuki Nakazawa, and Ryuichi Oka. Mutual spotting retrieval between speech and video image using self-organized network databases. Lecture Notes in Computer Science, 1554:103–118, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1554/15540103.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1554/15540103.
                                        pdf.

Apers:1999:CBR

[304] Peter Apers and Martin Kersten. Content-based retrieval in multimedia databases based on feature models (invited paper). Lecture Notes in Computer Science, 1554:119–130, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1554/15540119.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1554/15540119.
pdf.

Yoo:1999:EIS

[305] Jae Soo Yoo, Myung Keun Shin, Seok Hee Lee, Kil Seong Choi, Ki Hyung Cho, and Dae Young Hur. An efficient index structure for high dimensional image data. Lecture Notes in Computer Science, 1554:131–144, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1554/15540131.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1554/15540131.
pdf.

Chua:1999:CBP

[306] Tat-Seng Chua and Chun-Xin Chu. Color-based pseudo object model for image retrieval with relevance feedback. Lecture Notes in Computer Science, 1554:145–160, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1554/15540145.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1554/15540145.pdf.

Fels:1999:IFP

[307] Sidney Fels and Kenji Mase. InvenTcl: A fast prototyping environment for 3D graphics and multimedia applications. Lecture Notes in Computer Science, 1554:161–176, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/1554/15540161.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1554/15540161.
                                        pdf.

Wedlake:1999:NDV

[308] Martine Wedlake, Kin F. Li, and Fayez El Guibaly. The NAVL distributed virtual reality system. Lecture Notes in Computer Science, 1554:177– 193, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1554/15540177.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1554/15540177.
                                        pdf.

Aksoy:1999:RDB

[309] Demet Aksoy, Mehmet Altinel, Rahul Bose, Ugur Cetintemel, Michael Franklin, Jane Wang, and Stan Zdonik. Research in data broadcast and dissemination (invited paper). Lecture Notes in Computer Science, 1554:194– 207, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1554/15540194.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1554/15540194.
                                        pdf.

Watanabe:1999:MDS

[310] Yasuhiko Watanabe, Yoshihiro Okada, Kengo Kaneji, and Yoshitaka Sakamoto. Multimedia database system for TV newscasts and newspapers. Lecture Notes in Computer Science, 1554:208– 220, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1554/15540208.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1554/15540208.
pdf;   http://link.springer-ny.
com/link/service/series/0558/papers/
1554/15540208s.pdf.

Kamahara:1999:TNR

[311] Junzo Kamahara, Yuji Nomura, Kazunori Ueda, Keishi Kandori, Shinji Shimojo, and Hideo Miyahara. A TV news recommendation system with automatic recomposition. Lecture Notes in Computer Science, 1554:221–235, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1554/15540221.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1554/15540221.
pdf.

Nabeshima:1999:EDV

[312] Shinji Nabeshima, Kazuo Okamura, Takashi Kakiuchi, Kazutoshi Sumiya, Naoya Takao, and Yoshiyuki Miyabe. Extended digital video broadcasting with time-lined hypermedia. Lecture Notes in Computer Science, 1554:236– 251, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1554/15540236.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1554/15540236.
pdf.

Matsuyama:1999:AIC

[313] Takashi Matsuyama, Toshikazu Wada, and Shogo Tokai. Active image capturing and dynamic scene visualization by cooperative distributed vision (invited paper). Lecture Notes in Computer Science, 1554:252–288, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1554/15540252.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1554/15540252.
pdf;   http://link.springer-ny.
com/link/service/series/0558/papers/
1554/15540252s.pdf.

Nishimura:1999:VNS

[314] Go Nishimura, Tamio Kihara, and Ryoji Kataoka. Videoplex: A new system framework for constructing videobased three-dimensional space. Lecture Notes in Computer Science, 1554:289– 300, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1554/15540289.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1554/15540289.
pdf.

Zhang:1999:CVE

[315] Xiaohua Zhang, Hiroki Takahashi, and Masayuki Nakajima. Construction of virtual environment from video data with forward motion. Lecture Notes in Computer Science, 1554:301–312, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1554/15540301.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/1554/15540301.
pdf.

Arikawa:1999:SBV

[316] Masatoshi Arikawa and Tetsu Kamiyama. Spatial browsing for video databases. Lecture Notes in Computer Science, 1554:313–327, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1554/15540313.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1554/15540313.pdf.

Prie:1999:ASU

[317] Yannick Prié, Alain Mille, and Jean-Marie Pinon. AI-STRATA: A usercentered model for content-based description and retrieval of audiovisual sequences. Lecture Notes in Computer Science, 1554:328–343, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1554/15540328.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1554/15540328.
pdf.

Katayama:1999:UAH

[318] Kaoru Katayama, Osami Kagawa, Yasuhiro Kamiya, Hideki Tsushima, Takuya Yoshihiro, and Yahiko Kambayashi. Use of action history views for indexing continuous media objects. Lecture Notes in Computer Science, 1554:344–355, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1554/15540344.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1554/15540344.
                                        pdf.

Zettsu:1999:SSV

[319] Koji Zettsu, Kuniaki Uehara, and Katsumi Tanaka. Semantic structures for video data indexing. Lecture Notes in Computer Science, 1554:356– 369, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1554/15540356.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1554/15540356.
                                        pdf.

Imai:1999:SEC

[320] Michita Imai and Tsutomu Miyasato. A study of emergent computation of life-like behavior by indefinite observation. Lecture Notes in Computer Science, 1554:370–385, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1554/15540370.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1554/15540370.
                                        pdf.

Masaki:1999:IDF

[321] Toshihiro Masaki, Tetsuya Yamaguchi, and Yoshifumi Kitamura. An interactive digital fishtank based on live video images. Lecture Notes in Computer Science, 1554:386–396, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1554/15540386.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1554/15540386.
pdf.

Nakatsu:1999:CCI

[322] Ryohei Nakatsu. Contents creation for interactive media (invited paper). Lecture Notes in Computer Science, 1554:397–405, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1554/15540397.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1554/15540397.
pdf.

Terzopoulos:1999:VMM

[323] Demetri Terzopoulos. Visual modeling for multimedia content (invited paper). Lecture Notes in Computer Science, 1554:406–421, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1554/15540406.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1554/15540406.
pdf.

Saiwaki:1999:AGM

[324] Naoki Saiwaki, Toshiaki Komatsu, and Shogo Nishida. Automatic generation of moving crowds in the virtual environment. Lecture Notes in Computer Science, 1554:422–432, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1554/15540422.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1554/15540422.
                                        pdf.

Otsuka:1999:EFM

[325] Takahiro Otsuka and Jun Ohya. Extracting facial motion parameters by tracking feature points. Lecture Notes in Computer Science, 1554:433–444, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1554/15540433.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1554/15540433.
                                        pdf;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1554/15540433s.pdf.

Meschkat:1999:IRV

[326] Steffen Meschkat. Immersion reconsidered: Virtual reality projects of art+com (invited talk). Lecture Notes in Computer Science, 1554:445–450, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1554/15540445.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1554/15540445.
                                        pdf;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1554/15540445s.pdf.

Blumberg:1999:SCB

[327] Bruce M. Blumberg. Synthetic characters: Behaving in character (invited talk). Lecture Notes in Computer Science, 1554:451–452, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1554/15540451.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1554/15540451.
pdf.

Mora:1999:BMS

[328] M. Mora, J. Lopes, R. Viccari, and H. Coelho. BDI models and systems: Bridging the gap. Lecture Notes in Computer Science, 1555:11–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanEijk:1999:IPB

[329] R. Van Eijk, F. De Boer, W. Van Der Hoek, and J.-J. C. Meyer. Informationpassing and belief revision in multiagent systems. Lecture Notes in Computer Science, 1555:29–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schild:1999:RBB

[330] K. Schild. On the relationship between BDI logics and standard logics of concurrency. Lecture Notes in Computer Science, 1555:47–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wooldridge:1999:IRR

[331] M. Wooldridge and S. Parsons. Intention reconsideration reconsidered. Lecture Notes in Computer Science, 1555: 63–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hunsberger:1999:MSP

[332] L. Hunsberger. Making shared plans more concise and easier to reason about. Lecture Notes in Computer Science, 1555:81–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Conte:1999:ANA

[333] R. Conte, C. Castelfranchi, and F. Dignum. Autonomous norm acceptance. Lecture Notes in Computer Science, 1555:99–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bazzan:1999:MSM

[334] A. Bazzan, R. Bordini, and J. Campbell. Moral sentiments in multi-agent systems. Lecture Notes in Computer Science, 1555:113–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ossowski:1999:SSA

[335] S. Ossowski and A. Garcia-Serrano. Social structure in artificial agent societies: Implications for autonomous problem-solving agents. Lecture Notes in Computer Science, 1555:133–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rustogi:1999:BEC

[336] S. Rustogi and M. Singh. The bases of effective coordination in decentralized multi-agent systems. Lecture Notes in Computer Science, 1555:149–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benerecetti:1999:MCA

[337] M. Benerecetti, F. Giunchiglia, and L. Serafini. A model checking algorithm for multiagent systems. Lecture Notes in Computer Science, 1555:163– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Engelfriet:1999:CVM

[338] J. Engelfriet, C. Jonker, and J. Treur. Compositional verification of multiagent systems in temporal multiepistemic logic. Lecture Notes in Computer Science, 1555:177–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jung:1999:EMA

[339] C. Jung. Emergent mental attitudes in layered agents. Lecture Notes in Computer Science, 1555:195–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mueller:1999:RAA

[340] J. P. Mueller. The right agent (architecture) to do the right thing. Lecture Notes in Computer Science, 1555:211– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fisher:1999:RAA

[341] M. Fisher. Representing abstract agent architectures. Lecture Notes in Computer Science, 1555:227–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Piaggio:1999:HNH

[342] M. Piaggio. HEIR — A non hierarchical hybrid architecture for intelligent robots. Lecture Notes in Computer Science, 1555:243–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rachlin:1999:TAA

[343] J. Rachlin, R. Goodwin, S. Murthy, and R. Akkiraju. A-Teams: An agent architecture for optimization and decision support. Lecture Notes in Computer Science, 1555:261–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shehory:1999:GSL

[344] O. Shehory, S. Kraus, and O. Yadgar. Goal satisfaction in large scale agentsystems: A transportation example. Lecture Notes in Computer Science, 1555:277–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stone:1999:TDD

[345] P. Stone and M. Veloso. Task decomposition and dynamic role assignment for real-time strategic teamwork. Lecture Notes in Computer Science, 1555: 293–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meyer:1999:ALT

[346] J.-J. C. Meyer. Agent languages and their relationship to other programming paradigms. Lecture Notes in Computer Science, 1555:309–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iglesias:1999:SAO

[347] C. Iglesias, M. Garrijo, and J. Gonzalez. A survey of agent-oriented methodologies. Lecture Notes in Computer Science, 1555:317–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kinny:1999:AAI

[348] D. Kinny. The Agentis agent interaction model. Lecture Notes in Computer Science, 1555:331–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Skarmeas:1999:CBR

[349] N. Skarmeas and K. Clark. Contentbased routing as the basis for intraagent communication. Lecture Notes in Computer Science, 1555:345–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chaib-Draa:1999:ACL

[350] B. Chaib-Draa and D. Vanderveken. Agent communication language: Toward a semantics based on success, satisfaction, and recursion. Lecture Notes in Computer Science, 1555:363– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hindriks:1999:CSR

[351] K. Hindriks, F. De Boer, W. Van Der Hoek, and J.-J. C. Meyer. Control structures of rule-based agent languages. Lecture Notes in Computer Science, 1555:381–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dury:1999:RAS

[352] A. Dury, F. Le Ber, and V. Chevrier. A reactive approach for solving constraint satisfaction problems. Lecture Notes in Computer Science, 1555:397–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shehory:1999:IRU

[353] O. Shehory, K. Sycara, P. Chalasani, and S. Jha. Increasing resource utilization and task performance by agent cloning. Lecture Notes in Computer Science, 1555:413–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wooldridge:1999:CSV

[354] M. Wooldridge and J. P. Mueller. A classification schema to volumes I to 5 of the intelligent agents series. Lecture Notes in Computer Science, 1555: 427–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vaudenay:1999:FCD

[355] S. Vaudenay. Feistel ciphers with L2-decorrelation. Lecture Notes in Computer Science, 1556:1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harris:1999:KDB

[356] S. Harris and C. Adams. Keydependent S-box manipulations. Lecture Notes in Computer Science, 1556: 15–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schneier:1999:TKS

[357] B. Schneier, J. Kelsey, D. Whiting, and D. Wagner. On the Twofish key schedule. Lecture Notes in Computer Science, 1556:27–42, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1999:TPS

[358] Z.-G. Chen and S. E. Tavares. Toward provable security of substitutionpermutation encryption networks. Lecture Notes in Computer Science, 1556: 43–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coron:1999:AEM

[359] J.-S. Coron and D. Naccache. An accurate evaluation of Maurer’s universal test. Lecture Notes in Computer Science, 1556:57–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MRaihi:1999:CAR

[360] D. M’Raihi, D. Naccache, D. Pointcheval, and S. Vaudenay. Computational alternatives to random number generators. Lecture Notes in Computer Science, 1556:72–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaliski:1999:SEF

[361] B. S. Kaliski and Yiqun Lisa Yin. Storage-efficient finite field basis conversion. Lecture Notes in Computer Science, 1556:81–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wenbo:1999:VPS

[362] M. Wenbo. Verifiable partial sharing of integer factors. Lecture Notes in Computer Science, 1556:94–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moriai:1999:HOD

[363] S. Moriai, T. Shimoyama, and T. Kaneko. Higher order differential attack using chosen higher order differences. Lecture Notes in Computer Science, 1556:106–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aoki:1999:MNA

[364] K. Aoki. On maximum non-averaged differential probability. Lecture Notes in Computer Science, 1556:118–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mister:1999:CRL

[365] S. Mister and S. E. Tavares. Cryptanalysis of RC4-like ciphers. Lecture Notes in Computer Science, 1556:131– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stinson:1999:KPT

[366] D. R. Stinson and R. Wei. Key preassigned traceability schemes for broadcast encryption. Lecture Notes in Computer Science, 1556:144–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jakobsson:1999:MBE

[367] M. Jakobsson and D. M’Raihi. Mixbased electronic payments. Lecture Notes in Computer Science, 1556:157– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Patel:1999:ASP

[368] S. Patel. Over the air service provisioning. Lecture Notes in Computer Science, 1556:174–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wiener:1999:FAE

[369] M. J. Wiener and R. J. Zuccherato. Faster attacks on elliptic curve cryptosystems. Lecture Notes in Computer Science, 1556:190–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lopez:1999:IAE

[370] J. Lopez and R. Dahab. Improved algorithms for elliptic curve arithmetic in GF(2n0n). Lecture Notes in Computer Science, 1556:201–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nguyen:1999:CFP

[371] P. Nguyen and J. Stern. Cryptanalysis of a fast public key cryptosystem presented at SAC ’97. Lecture Notes in Computer Science, 1556:213–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cai:1999:LBP

[372] J.-Y. Cai and T. W. Cusick. A latticebased public-key cryptosystem. Lecture Notes in Computer Science, 1556:219– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaps:1999:FIF

[373] J.-P. Kaps and C. Paar. Fast DES implementation for FPGAs and its application to a universal key-search machine. Lecture Notes in Computer Science, 1556:234–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lipmaa:1999:ICM

[374] H. Lipmaa. IDEA: A cipher for multimedia architectures? Lecture Notes in Computer Science, 1556:248–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kanda:1999:SCF

[375] M. Kanda, Y. Takashima, T. Matsumoto, and K. Aoki. A strategy for constructing fast round functions with practical security against differential and linear cryptanalysis. Lecture Notes in Computer Science, 1556:264– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1999:NBF

[376] X.-M. Zhang and Y. Zheng. The nonhomomorphicity of Boolean functions. Lecture Notes in Computer Science, 1556:280–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wagner:1999:CO

[377] D. Wagner, L. Simpson, E. Dawson, and J. Kelsey. Cryptanalysis of ORYX. Lecture Notes in Computer Science, 1556:296–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Handschuh:1999:TAR

[378] H. Handschuh and H. M. Heys. A timing attack on RC5. Lecture Notes in Computer Science, 1556:306–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hall:1999:CS

[379] C. Hall, J. Kelsey, V. Rijmen, and B. Schneier. Cryptanalysis of SPEED. Lecture Notes in Computer Science, 1556:319–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blake-Wilson:1999:ADH

[380] S. Blake-Wilson and A. Menezes. Authenticated Diffie–Hellman key agreement protocols. Lecture Notes in Computer Science, 1556:339–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biham:1999:IOS

[381] E. Biham, A. Biryukov, O. Dunkelman, and E. Richardson. Initial observations on Skipjack: Cryptanalysis of Skipjack-3XOR. Lecture Notes in Computer Science, 1556:362–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hossfeld:1999:TCC

[382] Friedel Hossfeld. Teraflops computing: A challenge to parallel numerics? Lecture Notes in Computer Science, 1557: 1–12, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1557/15570001.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/1557/15570001.
pdf.

Saad:1999:NSP

[383] Yousef Saad and Maria Sosonkina. Non-standard parallel solution strategies for distributed sparse linear systems. Lecture Notes in Computer Science, 1557:13–27, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1557/15570013.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1557/15570013.
pdf.

Santos:1999:OTS

[384] Eunice E. Santos. Optimal tridiagonal solvers on mesh interconnection networks. Lecture Notes in Computer Science, 1557:28–37, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1557/15570028.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1557/15570028.
pdf.

Asenjo:1999:PPL

[385] Rafael Asenjo and Emilio L. Zapata. Parallel pivots LU algorithm on the Cray T3E. Lecture Notes in Computer Science, 1557:38–47, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1557/15570038.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/1557/15570038.
pdf.

Becka:1999:EPO

[386] Martin Becka, Sophie Robert, and Marian Vajtersic. Experiments with parallel one-sided and two-sided algorithms for SVD. Lecture Notes in Computer Science, 1557:48–57, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1557/15570048.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1557/15570048.
pdf.

Oksa:1999:CSA

[387] Gabriel Oksa. Combined systolic array for matrix portrait computation. Lecture Notes in Computer Science, 1557: 58–67, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1557/15570058.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1557/15570058.
pdf.

Podhaisky:1999:CET

[388] Helmut Podhaisky and Rüdiger Weiner. A class of explicit twostep Runge–Kutta methods with enlarged stability regions for parallel computers. Lecture Notes in Computer Science, 1557:68–77, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1557/15570068.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1557/15570068.
                                        pdf.

Halada:1999:PSI

[389] Ladislav Halada and Mária Lucká. A parallel strongly implicit algorithm for solving of diffusion equations. Lecture Notes in Computer Science, 1557:78– 84, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1557/15570078.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1557/15570078.
                                        pdf.

Sarbazi-Azad:1999:PAL

[390] Hamid Sarbazi-Azad, Lewis M. Mackenzie, and Mohamed Ould-Khaoua. A parallel algorithm for Lagrange interpolation on k-ary n-cubes. Lecture Notes in Computer Science, 1557:85–95, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1557/15570085.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1557/15570085.
                                        pdf.

Schmid:1999:PQM

[391] Wolfgang Ch. Schmid and Andreas Uhl. Parallel quasi-Monte Carlo integration using (t, s)-sequences. Lecture Notes in Computer Science, 1557:96– 106, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1557/15570096.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1557/15570096.
pdf.

Entacher:1999:PRN

[392] Karl Entacher, Andreas Uhl, and Stefan Wegenkittl. Parallel random number generation: Longrange correlations among multiple processors. Lecture Notes in Computer Science, 1557:107–116, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1557/15570107.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1557/15570107.
pdf.

Hausenblas:1999:MCM

[393] Erika Hausenblas. A Monte-Carlo method with inherent parallelism for numerical solving partial differential equations with boundary conditions. Lecture Notes in Computer Science, 1557:117–126, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1557/15570117.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1557/15570117.pdf.

Gansterer:1999:BTN

[394] Wilfried N. Gansterer, Dieter F. Kvasnicka, and Christoph W. Ueberhuber. Blocking techniques in numerical software. Lecture Notes in Computer Science, 1557:127–139, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1557/15570127.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1557/15570127.
                                        pdf.

Ehold:1999:HNL

[395] Harald J. Ehold, Wilfried N. Gansterer, Dieter F. Kvasnicka, and Christoph W. Ueberhuber. HPF and numerical libraries. Lecture Notes in Computer Science, 1557:140–152, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1557/15570140.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1557/15570140.
                                        pdf.

Delaplace:1999:POL

[396] Franck Delaplace and Didier Remy. PARADEIS: An object library for parallel sparse array computation. Lecture Notes in Computer Science, 1557:153– 162, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1557/15570153.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1557/15570153.
                                        pdf.

Hoefinger:1999:PAD

[397] S. Hoefinger, O. Steinhauser, and P. Zinterhof. Performance analysis and derived parallelization strategy for a SCF program at the Hartree Fock level. Lecture Notes in Computer Science, 1557:163–172, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hofinger:1999:PAD

[398] Siegfried Höfinger, Othmar Steinhauser, and Peter Zinterhof. Performance analysis and derived parallelization strategy for a SCF program at the Hartree Fock level. Lecture Notes in Computer Science, 1557:163– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1557/15570163.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1557/15570163.
pdf.

Fontdecaba:1999:CIO

[399] Enric Fontdecaba, José M. Cela, and Juan C. Dürsteler. Computational issues in optimizing ophthalmic lens. Lecture Notes in Computer Science, 1557:173–182, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1557/15570173.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1557/15570173.
pdf.

Wyrzykowski:1999:PFE

[400] Roman Wyrzykowski, Norbert Sczygiol, Tomasz Olas, and Juri Kanevski. Parallel finite element modeling of solidification processes. Lecture Notes in Computer Science, 1557:183–195, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1557/15570183.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1557/15570183.
                                        pdf.

Panchanathan:1999:AAM

[401] Sethuraman Panchanathan. Architectural approaches for multimedia processing. Lecture Notes in Computer Science, 1557:196–210, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1557/15570196.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1557/15570196.
                                        pdf.

Pissaloux:1999:PRA

[402] Edwige Pissaloux. On parallel reconfigurable architectures for image processing. Lecture Notes in Computer Science, 1557:211–225, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1557/15570211.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1557/15570211.
                                        pdf.

Moga:1999:PMI

[403] Alina N. Moga. Parallel multiresolution image segmentation with watershed transformation. Lecture Notes in Computer Science, 1557:226–235, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/1557/15570226.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1557/15570226.
pdf.

Chung:1999:SII

[404] Yongwha Chung and Jin-Won Park. Solving irregular inter-processor data dependency in image understanding tasks. Lecture Notes in Computer Science, 1557:236–245, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1557/15570236.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1557/15570236.pdf.

Verians:1999:NPM

[405] Xavier Verians, Jean-Didier Legat, Jean-Jacques Quisquater, and Benoit Macq. A new parallelism management scheme for multiprocessor systems. Lecture Notes in Computer Science, 1557:246–256, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1557/15570246.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1557/15570246.
pdf.

Xu:1999:FVP

[406] Donglai Xu and Reza Sotudeh. A flexible VLSI parallel processing system for block-matching motion estimation in low bit-rate video coding applications. Lecture Notes in Computer Science, 1557:257–264, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1557/15570257.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1557/15570257.
                                        pdf.

Konstantopoulos:1999:HBM

[407] Charalampos Konstantopoulos, Andreas Svolos, and Christos Kaklamanis. Hierarchical block matching motion estimation on a hypercube multiprocessor. Lecture Notes in Computer Science, 1557:265–275, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1557/15570265.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1557/15570265.
                                        pdf.

Haemmerle:1999:CBS

[408] J. Haemmerle and A. Uhl. Classification based speed-up methods for fractal image compression on multicomputers. Lecture Notes in Computer Science, 1557:276–285, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hammerle:1999:CBS

[409] Jutta Hämmerle and Andreas Uhl. Classification based speed-up methods for fractal image compression on multicomputers. Lecture Notes in Computer Science, 1557:276–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/1557/15570276.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1557/15570276.
pdf.

Gatineau:1999:AME

[410] Laurent Gatineau and François Meunier. Accurate motion estimation in image sequences: Massive vs. distributed parallelism. Lecture Notes in Computer Science, 1557:286–295, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1557/15570286.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1557/15570286.
pdf.

Arita:1999:RTD

[411] Daisaku Arita, Yoshio Hamada, and Rin ichiro Taniguchi. A real-time distributed video image processing system on PC-cluster. Lecture Notes in Computer Science, 1557:296–305, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1557/15570296.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1557/15570296.
pdf.

He:1999:MSM

[412] Yong He, Ishfaq Ahmad, and Ming L. Liou. Modeling and scheduling for MPEG-4 based video encoder using a cluster of workstations. Lecture Notes in Computer Science, 1557:306– 316, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1557/15570306.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1557/15570306.
                                        pdf.

Pommer:1999:FVC

[413] Andreas Pommer. Fractal video compression on shared memory systems. Lecture Notes in Computer Science, 1557:317–326, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1557/15570317.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1557/15570317.pdf.

DePietro:1999:SPA

[414] G. De Pietro and M. Lerro. The splitproxy approach: A new architecture for parallel video servers. Lecture Notes in Computer Science, 1557:327–336, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pietro:1999:SPA

[415] Giuseppe De Pietro and Marco Lerro. The split-proxy approach: A new architecture for parallel video servers. Lecture Notes in Computer Science, 1557: 327–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1557/15570327.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1557/15570327.
                                        pdf.

Uytterhoeven:1999:WTL

[416] Geert Uytterhoeven, Dirk Roose, and Adhemar Bultheel. A wavelet toolbox for large scale image processing. Lecture Notes in Computer Science, 1557: 337–346, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1557/15570337.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1557/15570337.
pdf.

Kutil:1999:HSA

[417] Rade Kutil and Andreas Uhl. Hardware and software aspects for 3-D wavelet decomposition on shared memory MIMD computers. Lecture Notes in Computer Science, 1557:347–356, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1557/15570347.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1557/15570347.
pdf.

Corsaro:1999:PIF

[418] Stefania Corsaro, Luisa D’Amore, and Almerico Murli. On the parallel implementation of the fast wavelet packet transform on MIMD distributed memory environments. Lecture Notes in Computer Science, 1557:357–366, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1557/15570357.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1557/15570357.
                                        pdf.

Feil:1999:APP

[419] Manfred Feil and Andreas Uhl. Algorithms and programming paradigms for 2-D wavelet packet decomposition on multicomputers and multiprocessors. Lecture Notes in Computer Science, 1557:367–376, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1557/15570367.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1557/15570367.
                                        pdf.

Jensen:1999:RTL

[420] Morten Vadskær Jensen and Brian Nielsen. Real-time layered video compression using SIMD computation. Lecture Notes in Computer Science, 1557: 377–387, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1557/15570377.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1557/15570377.
                                        pdf.

Glendinning:1999:PSS

[421] Ian Glendinning. Parallelisation of a satellite signal processing code — strategies and tools. Lecture Notes in Computer Science, 1557:388–397, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1557/15570388.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1557/15570388.
pdf.

Janoth:1999:MSP

[422] Jürgen Janoth, Markus M. Eisl, Erwin M. Bakker, Rogier van Sterkenburg, Roberta Borgia, Salvo Sabina, and Fabio Volpe. MMIPPS — A software package for multitemporal and multispectral image processing on parallel systems. Lecture Notes in Computer Science, 1557:398–407, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1557/15570398.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1557/15570398.
pdf.

Goller:1999:PMS

[423] Alois Goller. Parallel matching of synthetic aperture radar images. Lecture Notes in Computer Science, 1557:408– 416, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1557/15570408.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1557/15570408.
pdf.

Hsieh:1999:PDD

[424] Sun yuan Hsieh. Parallel decomposition of distance-hereditary graphs. Lecture Notes in Computer Science, 1557: 417–426, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/

                                        0558/bibs/1557/15570417.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1557/15570417.
                                        pdf.

Bartz:1999:APC

[425] Dirk Bartz and Wolfgang Straßer. Asynchronous parallel construction of recursive tree hierarchies. Lecture Notes in Computer Science, 1557:427– 436, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1557/15570427.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1557/15570427.
                                        pdf.

Kofler:1999:LPT

[426] Helmut Kofler, Ernst J. Haunschmid, Wilfried N. Gansterer, and Christoph W. Ueberhuber. The locality property in topological irregular graph hierarchies. Lecture Notes in Computer Science, 1557:437–449, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1557/15570437.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1557/15570437.pdf.

Koppler:1999:GAR

[427] Rainer Koppler. Geometry-aided rectilinear partitioning of unstructured meshes. Lecture Notes in Computer Science, 1557:450–459, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/1557/15570450.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1557/15570450.
pdf.

Genius:1999:RCC

[428] Daniela Genius and Jörn Eisenbiegler. Reducing cache conflicts by a parametrized memory mapping. Lecture Notes in Computer Science, 1557: 460–469, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1557/15570460.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1557/15570460.
pdf.

Carretero:1999:OIA

[429] Jesús Carretero, Jaechun No, and Alok Choudhary. Optimizing I/O for irregular applications on distributedmemory machines. Lecture Notes in Computer Science, 1557:470–479, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1557/15570470.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1557/15570470.
pdf.

Stolberg:1999:CMA

[430] Hans-Joachim Stolberg, Martin Ohmacht, and Peter Pirsch. Cellular multiprocessor arrays with adaptive resource utilization. Lecture Notes in Computer Science, 1557:480–489, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1557/15570480.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1557/15570480.
                                        pdf.

Kranzlmueller:1999:NNP

[431] D. Kranzlmueller and J. Volkert. NOPE: A nondeterministic program evaluator. Lecture Notes in Computer Science, 1557:490–499, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kranzlmuller:1999:NNP

[432] Dieter Kranzlmüller and Jens Volkert. NOPE: A nondeterministic program evaluator. Lecture Notes in Computer Science, 1557:490–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1557/15570490.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1557/15570490.
                                        pdf.

Melo:1999:VMV

[433] Alba Cristina Melo and Simone Cintra Chagas. Visual-MCM: Visualising execution histories on multiple memory consistency models. Lecture Notes in Computer Science, 1557:500– 509, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1557/15570500.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1557/15570500.
                                        pdf.

Golebiewski:1999:HPI

[434] Maciej Golebiewski, Markus Baum, and Rolf Hempel. High performance implementation of MPI for Myrinet. Lecture Notes in Computer Science, 1557:510–521, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1557/15570510.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1557/15570510.
pdf.

Boeszoermenyi:1999:PCC

[435] L. Boeszoermenyi, G. Hoelzl, and E. Pirker. Parallel cluster computing with IEEE1394-1995. Lecture Notes in Computer Science, 1557:522–532, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boszormenyi:1999:PCC

[436] László Böszörményi, Günther Hölzl, and Emanuel Pirker. Parallel cluster computing with IEEE1394-1995. Lecture Notes in Computer Science, 1557: 522–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1557/15570522.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1557/15570522.
pdf.

Hlavacs:1999:SLB

[437] Helmut Hlavacs and Christoph W. Ueberhuber. Simulating load balancing on heterogeneous workstation clusters. Lecture Notes in Computer Science, 1557:533–540, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1557/15570533.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1557/15570533.
pdf.

Hagenauer:1999:GVT

[438] Helge Hagenauer. Global virtual time approximation for split queue time warp. Lecture Notes in Computer Science, 1557:541–548, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1557/15570541.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1557/15570541.
pdf.

Koholka:1999:MPR

[439] Roland Koholka, Heinz Mayer, and Alois Goller. MPI-parallelized radiance on SGI CoW and SMP. Lecture Notes in Computer Science, 1557:549– 558, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1557/15570549.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1557/15570549.
pdf.

Taniar:1999:PSC

[440] David Taniar and J. Wenny Rahayu. Parallel sub-collection join query algorithms for a high performance objectoriented database architecture. Lecture Notes in Computer Science, 1557:559– 569, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1557/15570559.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1557/15570559.
pdf.

Acacio:1999:EPC

[441] Manuel Acacio, Oscar Cánovas, José M. García, and Pedro E. López de Teruel. An evaluation of parallel computing in PC clusters with fast ethernet. Lecture Notes in Computer Science, 1557:570–571, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1557/15570570.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1557/15570570.
pdf.

Akramullah:1999:PME

[442] Shahriar M. Akramullah, Ishfaq Ahmad, and Ming L. Liou. Parallel MPEG-2 encoder on ATM and Ethernet-connected workstations. Lecture Notes in Computer Science, 1557: 572–574, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1557/15570572.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1557/15570572.
pdf.

Alonso:1999:BPN

[443] Pedro Alonso and Juan M. Peña.

Block and partitioned Neville elimination. Lecture Notes in Computer Science, 1557:575–576, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1557/15570575.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1557/15570575.pdf.

Amalfitano:1999:OOD

[444] Francesco Amalfitano, Antonio d’Acierno, Ivana Marra, and Lucio Sansone. An object-oriented DataBase for movieson-demand: Two approaches. Lecture Notes in Computer Science, 1557:577– 578, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1557/15570577.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1557/15570577.
                                        pdf.

Antonuccio-Delogu:1999:PTA

[445] Vincenzo Antonuccio-Delogu, Ugo Becciani, Mario Gambera, and Antonio Pagliaro. Parallel tree algorithms for N-body simulations. Lecture Notes in Computer Science, 1557:579–580, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1557/15570579.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1557/15570579.
                                        pdf.

Bassomo:1999:PNA

[446] Pierre Bassomo, Ibrahima Sakho, and Annie Corbel. Parallel numerical algorithms for distributed memory machines. Lecture Notes in Computer Science, 1557:581–583, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1557/15570581.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1557/15570581.
pdf.

Brest:1999:DSN

[447] Janez Brest, Viljem Zumer, and Milan Ojstersek. Dynamic scheduling on a network heterogeneous computer system. Lecture Notes in Computer Science, 1557:584–585, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1557/15570584.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1557/15570584.
pdf.

Giordano:1999:IBP

[448] Maurizio Giordano, Mario Mango Furnari, and Francesco Vitobello. Interaction between PVM parameters and communication performances on ATM networks. Lecture Notes in Computer Science, 1557:586–587, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1557/15570586.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1557/15570586.
pdf.

Glazek:1999:HSD

[449] Włodzimierz Głazek. How to share a divisible load in a hypercube. Lecture Notes in Computer Science, 1557:588– 589, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1557/15570588.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1557/15570588.
                                        pdf.

Karner:1999:OFS

[450] Herbert Karner and Christoph W. Ueberhuber. Overlapped four-step FFT computation. Lecture Notes in Computer Science, 1557:590–591, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1557/15570590.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1557/15570590.
                                        pdf.

Lee:1999:DPP

[451] Hyung Lee, Kyung-Ae Moon, and Jong-Won Park. Design of parallel processing system for facial image retrieval. Lecture Notes in Computer Science, 1557:592–593, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1557/15570592.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1557/15570592.
                                        pdf.

Magnaghi:1999:IPA

[452] Antonio Magnaghi, Shuichi Sakai, and Hidehiko Tanaka. Inter-procedural analysis for parallelization of Java programs. Lecture Notes in Computer Science, 1557:594–595, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1557/15570594.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1557/15570594.pdf.

Palenichka:1999:FRC

[453] Roman M. Palenichka. Fast recursive computation of local axial moments by using primitive kernel functions. Lecture Notes in Computer Science, 1557:596–597, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1557/15570596.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1557/15570596.
pdf.

Pavlus:1999:SEP

[454] Miron Pavlus. Speed up estimation for a parallel method for systems of linear ordinary differential equations. Lecture Notes in Computer Science, 1557:598– 599, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1557/15570598.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1557/15570598.
pdf.

Santos:1999:EPA

[455] Eunice E. Santos and Pei-Yue Pauline Chu. Efficient parallel algorithms for dense Cholesky factorization. Lecture Notes in Computer Science, 1557:600– 602, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1557/15570600.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1557/15570600.
                                        pdf.

Junghanns:1999:RCL

[456] Andreas Junghanns and Jonathan Schaeffer. Relevance cuts: Localizing the search. Lecture Notes in Computer Science, 1558:1–14, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1558/15580001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1558/15580001.
                                        pdf.

Bjoernsson:1999:MCP

[457] Y. Bjoernsson and T. Marsland. Multicut pruning in alpha-beta search. Lecture Notes in Computer Science, 1558: 15–24, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bjornsson:1999:MCP

[458] Yngvi Björnsson and Tony Marsland. Multi-cut pruning in alpha-beta search. Lecture Notes in Computer Science, 1558:15–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1558/15580015.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1558/15580015.
pdf.

Breuker:1999:SGP

[459] Dennis M. Breuker, H. Jaap van den Herik, Jos W. H. M. Uiterwijk, and L. Victor Allis. A solution to the GHI problem for best-first search. Lecture Notes in Computer Science, 1558:25– 49, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1558/15580025.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1558/15580025.
pdf.

Frank:1999:OPA

[460] Ian Frank and David Basin. Optimal play against best defence: Complexity and heuristics. Lecture Notes in Computer Science, 1558:50–73, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1558/15580050.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1558/15580050.
pdf.

Gao:1999:SS

[461] Xinbo Gao, Hiroyuki Iida, Jos W. H. M. Uiterwijk, and H. Jaap van den Herik. A speculative strategy. Lecture Notes in Computer Science, 1558:74– 92, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1558/15580074.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1558/15580074.
                                        pdf.

Willmott:1999:APA

[462] Steven Willmott, Julian Richardson, Alan Bundy, and John Levine. An adversarial planning approach to Go. Lecture Notes in Computer Science, 1558: 93–112, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1558/15580093.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1558/15580093.
                                        pdf.

Beal:1999:FRU

[463] Donald F. Beal and Martin C. Smith. First results from using temporal difference learning in Shogi. Lecture Notes in Computer Science, 1558:113– 125, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1558/15580113.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1558/15580113.
                                        pdf.

Buro:1999:SFS

[464] Michael Buro. From simple features to sophisticated evaluation functions. Lecture Notes in Computer Science, 1558: 126–145, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/

0558/bibs/1558/15580126.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1558/15580126.
pdf.

Kojima:1999:TSM

[465] Takuya Kojima and Atsushi Yoshikawa. A two-step model of pattern acquisition: Application to Tsume-Go. Lecture Notes in Computer Science, 1558: 146–166, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1558/15580146.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1558/15580146.
pdf.

Sasaki:1999:NNP

[466] Nobusuke Sasaki, Yasuji Sawada, and Jin Yoshimura. A neural network program of Tsume-Go. Lecture Notes in Computer Science, 1558:167–182, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1558/15580167.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1558/15580167.
pdf.

Giraldez:1999:DDM

[467] J. Ignacio Giráldez and Daniel Borrajo. Distributed decision making in checkers. Lecture Notes in Computer Science, 1558:183–194, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1558/15580183.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1558/15580183.
                                        pdf.

Pijls:1999:GTA

[468] Wim Pijls and Arie de Bruin. Game tree algorithms and solution trees. Lecture Notes in Computer Science, 1558:195–204, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1558/15580195.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1558/15580195.
                                        pdf.

Fraenkel:1999:NHG

[469] Aviezri S. Fraenkel and Dmitri Zusman. A new heap game. Lecture Notes in Computer Science, 1558:205– 211, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1558/15580205.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1558/15580205.
                                        pdf.

Fraenkel:1999:ICI

[470] Aviezri S. Fraenkel and Ofer Rahat. Infinite cyclic impartial games. Lecture Notes in Computer Science, 1558:212– 221, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1558/15580212.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1558/15580212.
                                        pdf.

Crasmaru:1999:CTG

[471] Marcel Crâsmaru. On the complexity of Tsume-Go. Lecture Notes in Computer Science, 1558:222–231, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1558/15580222.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1558/15580222.
pdf.

Spight:1999:ETM

[472] William L. Spight. Extended thermography for multiple kos in Go. Lecture Notes in Computer Science, 1558:232– 251, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1558/15580232.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1558/15580232.
pdf.

Mueller:1999:CGR

[473] M. Mueller. Computer Go: A research agenda. Lecture Notes in Computer Science, 1558:252–264, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muller:1999:CGR

[474] Martin Müller. Computer go: A research agenda. Lecture Notes in Computer Science, 1558:252–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1558/15580252.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1558/15580252.
pdf.

Tajima:1999:EPO

[475] Morihiko Tajima and Noriaki Sanechika. Estimating the possible omission number for groups in Go by the number of n-th Dame. Lecture Notes in Computer Science, 1558:265–281, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1558/15580265.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1558/15580265.
pdf.

Yoshikawa:1999:RBS

[476] Atsushi Yoshikawa, Takuya Kojima, and Yasuki Saito. Relations between skill and the use of terms — an analysis of protocols of the game of go. Lecture Notes in Computer Science, 1558:282– 299, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1558/15580282.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1558/15580282.
pdf.

Grimbergen:1999:STS

[477] Reijer Grimbergen. A survey of Tsume-Shogi programs using variabledepth search. Lecture Notes in Computer Science, 1558:300–317, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/1558/15580300.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1558/15580300.
pdf.

Iida:1999:RAK

[478] Hiroyuki Iida, Jin Yoshimura, Kazuro Morita, and Jos W. H. M. Uiterwijk. Retrograde analysis of the KGK endgame in Shogi: Its implications for ancient Heian Shogi. Lecture Notes in Computer Science, 1558:318– 335, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1558/15580318.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1558/15580318.
pdf.

Fuchs:1999:ACE

[479] Norbert E. Fuchs, Uta Schwertel, and Rolf Schwitter. Attempto Controlled English — not just another logic specification language. Lecture Notes in Computer Science, 1559:1– 20, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1559/15590001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1559/15590001.
pdf.

Baldan:1999:STM

[480] Dante Baldan, Baudouin Le Charlier, Christophe Leclère, and Isabelle Pollet. A step towards a methodology for Mercury program construction: A declarative semantics for Mercury. Lecture Notes in Computer Science, 1559:21– 40, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1559/15590021.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1559/15590021.
                                        pdf.

Robertson:1999:PSL

[481] David Robertson and Jaume Agustí. Pragmatics in the synthesis of logic programs. Lecture Notes in Computer Science, 1559:41–60, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1559/15590041.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1559/15590041.
                                        pdf.

VanBaalen:1999:UDP

[482] Jeffrey Van Baalen and Steven Roach. Using decision procedures to accelerate domain-specific deductive synthesis systems. Lecture Notes in Computer Science, 1559:61–80, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1559/15590061.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1559/15590061.
                                        pdf.

Avellone:1999:SPA

[483] Alessandro Avellone, Mauro Ferrari, and Pierangelo Miglioli. Synthesis of programs in abstract data types. Lecture Notes in Computer Science, 1559:

81–100, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1559/15590081.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1559/15590081.
pdf.

Lau:1999:OFC

[484] Kung-Kiu Lau and Mario Ornaghi. OOD frameworks in component-based software — development in computational logic. Lecture Notes in Computer Science, 1559:101–123, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1559/15590101.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1559/15590101.
pdf.

Brogi:1999:URC

[485] Antonio Brogi, Simone Contiero, and Franco Turini. The use of renaming in composing general programs. Lecture Notes in Computer Science, 1559:124– 142, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1559/15590124.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1559/15590124.
pdf.

Bossi:1999:ISL

[486] A. Bossi and S. Rossi. Inductive synthesis of logic programs by composition of combinatory program schemes. Lecture Notes in Computer Science, 1559:

143–158, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hamfelt:1999:ISL

[487] Andreas Hamfelt and Jørgen Fischer Nilsson. Inductive synthesis of logic programs by composition of combinatory program schemes. Lecture Notes in Computer Science, 1559:143– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1559/15590143.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1559/15590143.
                                        pdf.

Bossi:1999:SLPa

[488] Annalisa Bossi and Sabina Rossi. Specialising logic programs with respect to call/post specifications. Lecture Notes in Computer Science, 1559:159– 178, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1559/15590159.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1559/15590159.
                                        pdf.

Glueck:1999:GHO

[489] R. Glueck, J. Hatcliff, and J. Joergensen. Generalization in hierarchies of online program specialization systems. Lecture Notes in Computer Science, 1559:179–198, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gluck:1999:GHO

[490] Robert Glück, John Hatcliff, and Jesper Jørgensen. Generalization in hierarchies of online program specialization systems. Lecture Notes in Computer Science, 1559:179–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1559/15590179.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1559/15590179.
pdf.

Leuschel:1999:IHE

[491] Michael Leuschel. Improving homeomorphic embedding for online termination. Lecture Notes in Computer Science, 1559:199–218, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1559/15590199.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1559/15590199.pdf.

Bossi:1999:SLPb

[492] Annalisa Bossi and Nicoletta Cocco. Successes in logic programs. Lecture Notes in Computer Science, 1559:219– 239, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1559/15590219.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1559/15590219.
pdf.

Hoarau:1999:ICT

[493] Sébastien Hoarau and Fred Mesnard. Inferring and compiling termination for constraint logic programs. Lecture Notes in Computer Science, 1559:240– 254, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1559/15590240.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1559/15590240.
                                        pdf.

Gabric:1999:SAF

[494] Tihomir Gabric, Kevin Glynn, and Harald Søndergaard. Strictness analysis as finite-domain constraint solving. Lecture Notes in Computer Science, 1559:255–270, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1559/15590255.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1559/15590255.pdf.

Stark:1999:IDF

[495] Jamie Stark and Andrew Ireland. Invariant discovery via failed proof attempts. Lecture Notes in Computer Science, 1559:271–288, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1559/15590271.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1559/15590271.pdf.

Smaus:1999:PIE

[496] Jan-Georg Smaus, Pat Hill, and Andy King. Preventing instantiation errors and loops for logic programs with multiple modes using block declarations. Lecture Notes in Computer Science, 1559:289–307, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1559/15590289.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1559/15590289.
pdf.

Wolper:1999:ASR

[497] Pierre Wolper. Algorithms for synthesizing reactive systems: A perspective. Lecture Notes in Computer Science, 1559:308–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1559/15590308.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1559/15590308.
pdf.

Zidoum:1999:SGS

[498] Hamza Zidoum, Pierre Flener, and Brahim Hnich. Schema-guided synthesis of CLP programs. Lecture Notes in Computer Science, 1559:309– 312, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1559/15590309.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1559/15590309.
pdf.

Richardson:1999:PPP

[499] Julian Richardson. Proof planning with program schemas. Lecture Notes in Computer Science, 1559:313–315, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1559/15590313.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1559/15590313.
                                        pdf.

Bellot:1999:LSI

[500] Patrick Bellot and Bernard Robinet. Logical synthesis of imperative O.O. programs. Lecture Notes in Computer Science, 1559:316–318, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1559/15590316.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1559/15590316.
                                        pdf.

BenAyed:1999:MFP

[501] R. Ben Ayed, J. Desharnais, M. Frappier, and A. Mill. Mathematical foundations for program transformations. Lecture Notes in Computer Science, 1559:319–321, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ayed:1999:MFP

[502] R. Ben Ayed, J. Desharnais, M. Frappier, and A. Mili. Mathematical foundations for program transformations. Lecture Notes in Computer Science, 1559:319–??, 1999. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1559/15590319.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1559/15590319.
pdf.

Rosenblueth:1999:ESM

[503] David A. Rosenblueth. An exhaustivesearch method using layered streams obtained through a meta-interpreter for chain programs. Lecture Notes in Computer Science, 1559:322–324, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1559/15590322.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1559/15590322.
pdf.

Vanhoof:1999:BSL

[504] Wim Vanhoof, Danny De Schreye, and Bern Martens. Bottom-up specialisation of logic programs. Lecture Notes in Computer Science, 1559:325– 327, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1559/15590325.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1559/15590325.
pdf.

Mallet:1999:MSO

[505] Sarah Mallet and Mireille Ducassé. Myrtle: A set-oriented meta-interpreter driven by a “relational” trace for deductive databases debugging. Lecture Notes in Computer Science, 1559:328– 330, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1559/15590328.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1559/15590328.
                                        pdf.

Bao:1999:NTM

[506] Feng Bao and Robert H. Deng. A new type of “magic ink” signatures — towards transcript-irrelevant anonymity revocation. Lecture Notes in Computer Science, 1560:1–11, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1560/15600001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1560/15600001.
                                        pdf.

Lee:1999:NAD

[507] Chang-Hyi Lee and Jong-In Lim. A new aspect of dual basis for efficient field arithmetic. Lecture Notes in Computer Science, 1560:12–28, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1560/15600012.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1560/15600012.
                                        pdf.

Coron:1999:SRSa

[508] Jean-Sébastien Coron. On the security of random sources. Lecture Notes in Computer Science, 1560:29– 42, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1560/15600029.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1560/15600029.
pdf.

Domingo-Ferrer:1999:AFB

[509] Josep Domingo-Ferrer. Anonymous fingerprinting based on committed oblivious transfer. Lecture Notes in Computer Science, 1560:43–52, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1560/15600043.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1560/15600043.
pdf.

Fujisaki:1999:HES

[510] Eiichiro Fujisaki and Tatsuaki Okamoto. How to enhance the security of publickey encryption at minimum cost. Lecture Notes in Computer Science, 1560: 53–68, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1560/15600053.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1560/15600053.
pdf.

Gamage:1999:EMA

[511] Chandana Gamage, Jussipekka Leiwo, and Yuliang Zheng. Encrypted message authentication by firewalls. Lecture Notes in Computer Science, 1560: 69–81, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1560/15600069.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1560/15600069.
                                        pdf.

Hada:1999:RBO

[512] Satoshi Hada and Toshiaki Tanaka. A relationship between one-wayness and correlation intractability. Lecture Notes in Computer Science, 1560:82– 96, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1560/15600082.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1560/15600082.
                                        pdf.

Lee:1999:MRF

[513] Hyung-Woo Lee and Tai-Yun Kim. Message recovery fair blind signature. Lecture Notes in Computer Science, 1560:97–111, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1560/15600097.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1560/15600097.
                                        pdf.

Jakobsson:1999:QCA

[514] Markus Jakobsson. On quorum controlled asymmetric proxy reencryption. Lecture Notes in Computer Science, 1560:112–121, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1560/15600112.

htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1560/15600112.pdf.

Jakobsson:1999:MCM

[515] Markus Jakobsson. Mini-cash: A minimalistic approach to E-commerce. Lecture Notes in Computer Science, 1560: 122–135, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1560/15600122.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1560/15600122.
pdf.

Nikander:1999:PPD

[516] Pekka Nikander, Yki Kortesniemi, and Jonna Partanen. Preserving privacy in distributed delegation with fast certificates. Lecture Notes in Computer Science, 1560:136–153, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1560/15600136.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1560/15600136.
pdf.

Blake-Wilson:1999:UKS

[517] Simon Blake-Wilson and Alfred Menezes. Unknown key-share attacks on the station-to-station (STS) protocol. Lecture Notes in Computer Science, 1560:154–170, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1560/15600154.
htm;   http://link.springer-ny.

                                        com/link/service/series/0558/papers/
                                        1560/15600154.pdf.

Miyazaki:1999:TFI

[518] Shingo Miyazaki, Ikuko Kuroda, and Kouichi Sakurai. Toward fair international key escrow: An attempt by distributed trusted third agencies with threshold cryptography. Lecture Notes in Computer Science, 1560:171– 187, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1560/15600171.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1560/15600171.
                                        pdf.

Naccache:1999:HCF

[519] David Naccache, Adi Shamir, and Julien P. Stern. How to copyright a function ? Lecture Notes in Computer Science, 1560:188–196, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1560/15600188.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1560/15600188.
                                        pdf.

Coron:1999:SRSb

[520] Jean-Sébastien Coron and David Naccache. On the security of RSA screening. Lecture Notes in Computer Science, 1560:197–203, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1560/15600197.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1560/15600197.
pdf.

Coupe:1999:ELA

[521] Christophe Coupé, Phong Nguyen, and Jacques Stern. The effectiveness of lattice attacks against low-exponent RSA. Lecture Notes in Computer Science, 1560:204–218, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1560/15600204.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1560/15600204.
pdf.

Paillier:1999:TPE

[522] Pascal Paillier. A trapdoor permutation equivalent to factoring. Lecture Notes in Computer Science, 1560:219– 222, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1560/15600219.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1560/15600219.
pdf.

Paillier:1999:LCD

[523] Pascal Paillier. Low-cost double-size modular exponentiation or how to stretch your cryptoprocessor. Lecture Notes in Computer Science, 1560:223– 234, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1560/15600223.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1560/15600223.
                                        pdf.

Paillier:1999:EDF

[524] Pascal Paillier. Evaluating differential fault analysis of unknown cryptosystems. Lecture Notes in Computer Science, 1560:235–244, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1560/15600235.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1560/15600235.
                                        pdf.

Rubia:1999:RIB

[525] Montse Rubia, Juan Carlos Cruellas, and Manel Medina. Removing interoperability barriers between the X.509 and EDIFACT public key infrastructures: The DEDICA project. Lecture Notes in Computer Science, 1560:245– 262, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1560/15600245.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1560/15600245.
                                        pdf.

Shin:1999:HFM

[526] Sang Uk Shin, Kyung Hyune Rhee, and Jae Woo Yoon. Hash functions and the MAC using all-or-nothing property. Lecture Notes in Computer Science, 1560:263–275, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/1560/15600263.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1560/15600263.
pdf.

Handschuh:1999:DOE

[527] Helena Handschuh, Yiannis Tsiounis, and Moti Yung. Decision oracles are equivalent to matching oracles. Lecture Notes in Computer Science, 1560:276– 289, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1560/15600276.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1560/15600276.
pdf.

Watanabe:1999:SGR

[528] Yuji Watanabe and Hideki Imai. Shared generation of random number with timestamp: How to cope with the leakage of the CA’s secret. Lecture Notes in Computer Science, 1560:290– 305, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1560/15600290.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1560/15600290.
pdf.

Young:1999:ARC

[529] Adam Young and Moti Yung. Autorecoverable cryptosystems with faster initialization and the escrow hierarchy. Lecture Notes in Computer Science, 1560:306–314, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/

                                        series/0558/bibs/1560/15600306.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1560/15600306.pdf.

Zhou:1999:SPP

[530] Jianying Zhou and Kwok-Yan Lam. A secure pay-per-view scheme for Webbased video service. Lecture Notes in Computer Science, 1560:315–326, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1560/15600315.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1560/15600315.
                                        pdf.

Jianying:1999:SPP

[531] Z. Jianying and K.-Y. Lam. A secure pay-per-view scheme for Webbased video service. Lecture Notes in Computer Science, 1560:315–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1999:AIa

[532] Anonymous. Author index. Lecture Notes in Computer Science, 1560:327– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/1560/15600327.pdf.

Bellare:1999:POP

[533] Mihir Bellare. Practice-oriented provable security. Lecture Notes in Computer Science, 1561:1–15, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1561/15610001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1561/15610001.
pdf.

Cramer:1999:ISC

[534] Ronald Cramer. Introduction to secure computation. Lecture Notes in Computer Science, 1561:16–62, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1561/15610016.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1561/15610016.
pdf.

Damgaard:1999:CSZ

[535] Ivan Damgård. Commitment schemes and zero-knowledge protocols. Lecture Notes in Computer Science, 1561:63– 86, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1561/15610063.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1561/15610063.
pdf.

Kaliski:1999:ESP

[536] Burton S. Kaliski Jr. Emerging standards for public-key cryptography. Lecture Notes in Computer Science, 1561: 87–104, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1561/15610087.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1561/15610087.
                                        pdf.

Knudsen:1999:CBC

[537] Lars R. Knudsen. Contemporary block ciphers. Lecture Notes in Computer Science, 1561:105–126, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1561/15610105.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1561/15610105.
                                        pdf.

Landrock:1999:PTU

[538] Peter Landrock. Primality tests and use of primes in public key systems. Lecture Notes in Computer Science, 1561:127–133, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1561/15610127.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1561/15610127.pdf.

Pedersen:1999:SCP

[539] Torben P. Pedersen. Signing contracts and paying electronically. Lecture Notes in Computer Science, 1561:134– 157, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1561/15610134.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1561/15610134.
                                        pdf.

Preneel:1999:SCH

[540] Bart Preneel. The state of cryptographic hash functions. Lecture Notes in Computer Science, 1561:158– 182, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1561/15610158.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1561/15610158.
pdf.

Salvail:1999:SHG

[541] Louis Salvail. The search for the Holy Grail in quantum cryptography. Lecture Notes in Computer Science, 1561: 183–216, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1561/15610183.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1561/15610183.
pdf.

Wolf:1999:USC

[542] Stefan Wolf. Unconditional security in cryptography. Lecture Notes in Computer Science, 1561:217–250, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1561/15610217.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1561/15610217.
pdf.

Nehaniv:1999:CMA

[543] Chrystopher L. Nehaniv. Computation for metaphors, analogy and agents.

Lecture Notes in Computer Science, 1562:1–10, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1562/15620001.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1562/15620001.
pdf.

Turner:1999:FC

[544] Mark Turner. Forging connections. Lecture Notes in Computer Science, 1562:11–26, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1562/15620011.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1562/15620011.
pdf.

Hiraga:1999:RSM

[545] Masako K. Hiraga. Rough sea and the milky way: ‘blending’ in a haiku text. Lecture Notes in Computer Science, 1562:27–36, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1562/15620027.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1562/15620027.
pdf.

Veale:1999:PFM

[546] Tony Veale. Pragmatic forces in metaphor use: The mechanics of blend recruitment in visual metaphors. Lecture Notes in Computer Science, 1562: 37–51, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1562/15620037.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1562/15620037.
pdf.

Brooks:1999:CPB

[547] Rodney A. Brooks, Cynthia Breazeal, Matthew Marjanovic, Brian Scassellati, and Matthew M. Williamson. The cog project: Building a humanoid robot. Lecture Notes in Computer Science, 1562:52–87, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1562/15620052.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1562/15620052.
pdf.

Stojanov:1999:EMM

[548] Georgi Stojanov. Embodiment as metaphor: Metaphorizing-in the environment. Lecture Notes in Computer Science, 1562:88–101, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1562/15620088.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1562/15620088.pdf.

Dautenhahn:1999:EIS

[549] Kerstin Dautenhahn. Embodiment and interaction in socially intelligent life-like agents. Lecture Notes in Computer Science, 1562:102–142, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1562/15620102.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1562/15620102.
                                        pdf.

Barnden:1999:ISM

[550] John A. Barnden. An implemented system for metaphor-based reasoning — with special application to reasoning about agents. Lecture Notes in Computer Science, 1562:143–153, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1562/15620143.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1562/15620143.
                                        pdf.

Fenton-Kerr:1999:GEP

[551] Tom Fenton-Kerr. GAIA: An experimental pedagogical agent for exploring multimodal interaction. Lecture Notes in Computer Science, 1562:154– 164, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1562/15620154.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1562/15620154.
                                        pdf.

ONeill-Brown:1999:WAM

[552] Patricia O’Neill-Brown. When agents meet cross-cultural metaphor: Can they be equipped to parse and generate it? Lecture Notes in Computer Science, 1562:165–175, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1562/15620165.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1562/15620165.
pdf.

Scassellati:1999:IMJ

[553] Brian Scassellati. Imitation and mechanisms of joint attention: A developmental structure for building social skills on a humanoid robot. Lecture Notes in Computer Science, 1562:176– 195, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1562/15620176.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1562/15620176.
pdf.

Kauppinen:1999:FSW

[554] Anneli Kauppinen. Figures of speech, a way to acquire language. Lecture Notes in Computer Science, 1562:196– 208, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1562/15620196.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1562/15620196.
pdf.

Nehmzow:1999:MTC

[555] Ulrich Nehmzow. “meaning” through clustering by self-organisation of spatial and temporal information. Lecture Notes in Computer Science, 1562:209– 229, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1562/15620209.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1562/15620209.
                                        pdf.

Shinohara:1999:CMS

[556] Kazuko Shinohara. Conceptual mappings from spatial motion to time: Analysis of English and Japanese. Lecture Notes in Computer Science, 1562: 230–241, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1562/15620230.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1562/15620230.
                                        pdf.

Goguen:1999:IAS

[557] Joseph Goguen. An introduction to algebraic semiotics, with application to user interface design. Lecture Notes in Computer Science, 1562:242– 291, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1562/15620242.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1562/15620242.
                                        pdf.

Indurkhya:1999:AAM

[558] Bipin Indurkhya. An algebraic approach to modeling creativity of metaphor. Lecture Notes in Computer Science, 1562:292–306, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1562/15620292.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1562/15620292.
pdf.

Alty:1999:MHC

[559] J. L. Alty and R. P. Knott. Metaphor and human-computer interaction: A model based approach. Lecture Notes in Computer Science, 1562:307–321, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1562/15620307.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1562/15620307.
pdf.

Beynon:1999:EMF

[560] Meurig Beynon. Empirical modelling and the foundations of artificial intelligence. Lecture Notes in Computer Science, 1562:322–364, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1562/15620322.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1562/15620322.
pdf.

Nasuto:1999:CEM

[561] Slawomir J. Nasuto, Kerstin Dautenhahn, and Mark Bishop. Communication as an emergent metaphor for neuronal operation. Lecture Notes in Computer Science, 1562:365–379, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1562/15620365.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1562/15620365.
                                        pdf.

Nehaniv:1999:SPM

[562] Chrystopher L. Nehaniv. The second person — meaning and metaphors. Lecture Notes in Computer Science, 1562: 380–388, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1562/15620380.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1562/15620380.
                                        pdf.

Anonymous:1999:AIb

[563] Anonymous. Author index. Lecture Notes in Computer Science, 1562:389– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/1562/15620389.pdf.

Nisan:1999:ASA

[564] Noam Nisan. Algorithms for selfish agents. Lecture Notes in Computer Science, 1563:1–15, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1563/15630001.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1563/15630001.pdf.

OssonadeMendez:1999:RGM

[565] P. Ossona de Mendez. The reduced genus of a multigraph. Lecture Notes in Computer Science, 1563:16–31, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeMendez:1999:RGM

[566] P. O. De Mendez. The reduced genus of a multigraph. Lecture Notes in Computer Science, 1563:16–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mendez:1999:RGM

[567] Patrice Ossona de Mendez. The reduced genus of a multigraph. Lecture Notes in Computer Science, 1563:16– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1563/15630016.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1563/15630016.
pdf.

Wilke:1999:CDT

[568] Thomas Wilke. Classifying discrete temporal properties. Lecture Notes in Computer Science, 1563:32–46, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1563/15630032.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1563/15630032.
pdf.

Bernasconi:1999:CCT

[569] Anna Bernasconi and Igor Shparlinski. Circuit complexity of testing square-free numbers. Lecture Notes in Computer Science, 1563:47– 56, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1563/15630047.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1563/15630047.
pdf.

Sauerhoff:1999:RBP

[570] Martin Sauerhoff, Ingo Wegener, and Ralph Werchner. Relating branching program size and formula size over the full binary basis. Lecture Notes in Computer Science, 1563:57– 67, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1563/15630057.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1563/15630057.
pdf.

Andreev:1999:MOS

[571] Alexander E. Andreev, Andrea E. F. Clementi, Paolo Penna, and José D. P. Rolim. Memory organization schemes for large shared data: A randomized solution for distributed memory machines. Lecture Notes in Computer Science, 1563:68–77, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1563/15630068.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1563/15630068.
pdf.

Jakoby:1999:ATC

[572] Andreas Jakoby. The average time complexity to compute prefix functions in processor networks. Lecture Notes in Computer Science, 1563:78– 89, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1563/15630078.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1563/15630078.
pdf.

Cai:1999:HP

[573] Jin-Yi Cai, A. Pavan, and D. Sivakumar. On the hardness of permanent. Lecture Notes in Computer Science, 1563:90–99, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1563/15630090.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1563/15630090.
pdf.

Buhrman:1999:OSV

[574] Harry Buhrman and Lance Fortnow. One-sided versus two-sided error in probabilistic computation. Lecture Notes in Computer Science, 1563:100– 109, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1563/15630100.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1563/15630100.
pdf.

Icking:1999:OCS

[575] Christian Icking, Rolf Klein, and Elmar Langetepe. An optimal competitive strategy for walking in streets. Lecture Notes in Computer Science, 1563:110– 120, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1563/15630110.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1563/15630110.
                                        pdf.

Schuierer:1999:OSS

[576] Sven Schuierer and Ines Semrau. An optimal strategy for searching in unknown streets. Lecture Notes in Computer Science, 1563:121–131, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1563/15630121.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1563/15630121.
                                        pdf.

Hammar:1999:PSR

[577] Mikael Hammar, Bengt J. Nilsson, and Sven Schuierer. Parallel searching on m rays. Lecture Notes in Computer Science, 1563:132–142, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1563/15630132.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1563/15630132.
                                        pdf.

Lautemann:1999:LCL

[578] Clemens Lautemann, Nicole Schweikardt, and Thomas Schwentick. A logical characterisation of linear time on nondeterministic Turing machines. Lecture Notes in Computer Science, 1563:143– 152, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1563/15630143.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1563/15630143.
pdf.

Durand:1999:DCC

[579] Bruno Durand, Alexander Shen, and Nikolai Vereshagin. Descriptive complexity of computable sequences. Lecture Notes in Computer Science, 1563: 153–162, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1563/15630153.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1563/15630153.
pdf.

Bergman:1999:CSP

[580] Clifford Bergman and Giora Slutzki. Complexity of some problems in universal algebra. Lecture Notes in Computer Science, 1563:163–172, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1563/15630163.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1563/15630163.
pdf.

Kloks:1999:NBT

[581] Ton Kloks, Jan Kratochvíl, and Haiko Müller. New branchwidth territories. Lecture Notes in Computer Science, 1563:173–183, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1563/15630173.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1563/15630173.pdf.

Kao:1999:BRT

[582] Ming-Yang Kao, Andrzej Lingas, and Anna Östlin. Balanced randomized tree splitting with applications to evolutionary tree constructions. Lecture Notes in Computer Science, 1563:184– 196, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1563/15630184.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1563/15630184.
                                        pdf.

Bouchitte:1999:TMF

[583] Vincent Bouchitté and Ioan Todinca. Treewidth and minimum fill-in of weakly triangulated graphs. Lecture Notes in Computer Science, 1563:197– 206, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1563/15630197.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1563/15630197.
                                        pdf.

Halava:1999:DUM

[584] Vesa Halava, Mika Hirvensalo, and Ronald de Wolf. Decidability and undecidability of marked PCP. Lecture Notes in Computer Science, 1563:207– 216, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1563/15630207.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1563/15630207.
pdf.

Robson:1999:QWE

[585] John Michael Robson and Volker Diekert. On quadratic word equations. Lecture Notes in Computer Science, 1563: 217–226, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1563/15630217.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1563/15630217.
pdf.

Kirsten:1999:SUR

[586] Daniel Kirsten. Some undecidability results related to the star problem in trace monoids. Lecture Notes in Computer Science, 1563:227–236, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1563/15630227.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1563/15630227.
pdf.

Kitsten:1999:SUR

[587] D. Kitsten. Some undecidability results related to the star problem in trace monoids. Lecture Notes in Computer Science, 1563:227–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andersson:1999:AAM

[588] Gunnar Andersson. An approximation algorithm for Max p-Section. Lecture Notes in Computer Science, 1563:237– 247, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1563/15630237.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1563/15630237.
pdf.

Kratsch:1999:ABM

[589] Dieter Kratsch and Lorna Stewart. Approximating bandwidth by mixing layouts of interval graphs. Lecture Notes in Computer Science, 1563:248– 258, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1563/15630248.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1563/15630248.
pdf.

Preis:1999:LTA

[590] Robert Preis. Linear time 1/2approximation algorithm for maximum weighted matching in general graphs. Lecture Notes in Computer Science, 1563:259–269, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1563/15630259.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1563/15630259.
pdf.

Hemaspaandra:1999:EDC

[591] Edith Hemaspaandra, Lane A. Hemaspaandra, and Harald Hempel. Extending downward collapse from 1-versus-2 queries to j-versus-j+1 queries. Lecture Notes in Computer Science, 1563:270– 280, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1563/15630270.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1563/15630270.
pdf.

Arvind:1999:SSA

[592] Vikraman Arvind and Jacobo Torán. Sparse sets, approximable sets, and parallel queries to NP. Lecture Notes in Computer Science, 1563:281–290, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1563/15630281.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1563/15630281.
pdf.

Sibeyn:1999:ES

[593] Jop F. Sibeyn. External selection. Lecture Notes in Computer Science, 1563: 291–301, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1563/15630291.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1563/15630291.
pdf.

Ahrendt:1999:FCE

[594] Timm Ahrendt. Fast computations of the exponential function. Lecture Notes in Computer Science, 1563:302– 312, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1563/15630302.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1563/15630302.
                                        pdf.

Koutny:1999:MBA

[595] Maciej Koutny and Giuseppe Pappalardo. A model of behaviour abstraction for communicating processes. Lecture Notes in Computer Science, 1563: 313–322, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1563/15630313.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1563/15630313.
                                        pdf.

Bouajjani:1999:MCL

[596] Ahmed Bouajjani and Richard Mayr. Model checking lossy vector addition systems. Lecture Notes in Computer Science, 1563:323–333, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1563/15630323.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1563/15630323.
                                        pdf.

Wu:1999:CLS

[597] Bang Ye Wu, Kun-Mao Chao, and Chuan Yi Tang. Constructing light spanning trees with small routing cost. Lecture Notes in Computer Science, 1563:334–344, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1563/15630334.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1563/15630334.
pdf.

Nykaenen:1999:FPR

[598] M. Nykaenen and E. Ukkonen. Finding paths with the right cost. Lecture Notes in Computer Science, 1563:345–355, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nykanen:1999:FPR

[599] Matti Nykänen and Esko Ukkonen. Finding paths with the right cost. Lecture Notes in Computer Science, 1563: 345–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1563/15630345.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1563/15630345.
pdf.

Szegedy:1999:HMS

[600] Mario Szegedy. In how many steps the k peg version of the Towers of Hanoi game can be solved? Lecture Notes in Computer Science, 1563:356– 361, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1563/15630356.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1563/15630356.
pdf.

Frandsen:1999:LBD

[601] Gudmund Skovbjerg Frandsen, Johan P. Hansen, and Peter Bro Miltersen. Lower bounds for dynamic algebraic problems. Lecture Notes in Computer Science, 1563:362–372, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1563/15630362.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1563/15630362.
pdf.

Engebretsen:1999:ELB

[602] Lars Engebretsen. An explicit lower bound for TSP with distances one and two. Lecture Notes in Computer Science, 1563:373–382, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1563/15630373.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1563/15630373.
pdf.

Jakoby:1999:SDG

[603] Andreas Jakoby, Maciej Liskiewicz, and Rüdiger Reischuk. Scheduling dynamic graphs. Lecture Notes in Computer Science, 1563:383–392, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1563/15630383.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1563/15630383.
pdf.

Aiello:1999:SID

[604] William Aiello, Costas Busch, Maurice Herlihy, Marios Mavronicolas, Nir Shavit, and Dan Touitou. Supporting increment and decrement operations in balancing networks. Lecture Notes in Computer Science, 1563:393– 403, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1563/15630393.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1563/15630393.
pdf.

Koutsoupias:1999:WCE

[605] Elias Koutsoupias and Christos Papadimitriou. Worst-case equilibria. Lecture Notes in Computer Science, 1563:404–413, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1563/15630404.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1563/15630404.pdf.

Reischuk:1999:CTA

[606] Rüdiger Reischuk and Thomas Zeugmann. A complete and tight averagecase analysis of learning monomials. Lecture Notes in Computer Science, 1563:414–423, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1563/15630414.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1563/15630414.pdf.

Case:1999:CGP

[607] John Case, Keh-Jiann Chen, and Sanjay Jain. Costs of general purpose learning. Lecture Notes in Computer Science, 1563:424–433, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1563/15630424.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1563/15630424.
                                        pdf.

Schuler:1999:UDT

[608] Rainer Schuler. Universal distributions and time-bounded Kolmogorov complexity. Lecture Notes in Computer Science, 1563:434–443, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1563/15630434.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1563/15630434.
                                        pdf.

Lautemann:1999:DCA

[609] Clemens Lautemann, Pierre McKenzie, Thomas Schwentick, and Heribert Vollmer. The descriptive complexity approach to LOGCFL. Lecture Notes in Computer Science, 1563:444– 454, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1563/15630444.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1563/15630444.
pdf.

Kupferman:1999:WSC

[610] Orna Kupferman and Moshe Y. Vardi. The weakness of selfcomplementation. Lecture Notes in Computer Science, 1563:455–466, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1563/15630455.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1563/15630455.
pdf.

Eiter:1999:DHT

[611] Thomas Eiter, Toshihide Ibaraki, and Kazuhisa Makino. On the difference of horn theories. Lecture Notes in Computer Science, 1563:467–477, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1563/15630467.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1563/15630467.
pdf.

Ettinger:1999:QAN

[612] Mark Ettinger and Peter Høyer. On quantum algorithms for noncommutative hidden subgroups. Lecture Notes in Computer Science, 1563:478– 487, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1563/15630478.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1563/15630478.
                                        pdf.

Sauerhoff:1999:SRO

[613] Martin Sauerhoff. On the size of randomized OBDDs and readonce branching programs for kstable functions. Lecture Notes in Computer Science, 1563:488–499, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1563/15630488.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1563/15630488.
                                        pdf.

DiCrescenzo:1999:HFS

[614] G. Di Crescenzo, N. Ferguson, R. Impagliazzo, and M. Jakobsson. How to forget a secret. Lecture Notes in Computer Science, 1563:500–509, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crescenzo:1999:HFS

[615] Giovanni Di Crescenzo, Niels Ferguson, Russell Impagliazzo, and Markus Jakobsson. How to forget a secret. Lecture Notes in Computer Science, 1563: 500–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1563/15630500.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1563/15630500.
                                        pdf.

Mueller-Olm:1999:MFL

[616] M. Mueller-Olm. A modal fixpoint logic with chop. Lecture Notes in Computer Science, 1563:510–520, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muller-Olm:1999:MFL

[617] Markus Müller-Olm. A modal fixpoint logic with chop. Lecture Notes in Computer Science, 1563:510–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1563/15630510.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1563/15630510.
pdf.

Barua:1999:CNL

[618] Rana Barua, Suman Roy, and Zhou Chaochen. Completeness of neighbourhood logic. Lecture Notes in Computer Science, 1563:521–530, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1563/15630521.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1563/15630521.
pdf.

Otto:1999:ERS

[619] Martin Otto. Eliminating recursion in the μ-calculus. Lecture Notes in Computer Science, 1563:531–540, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/1563/15630531.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1563/15630531.
pdf.

Messner:1999:OAO

[620] Jochen Messner. On optimal algorithms and optimal proof systems. Lecture Notes in Computer Science, 1563: 541–550, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1563/15630541.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1563/15630541.
pdf.

Esteban:1999:SBR

[621] Juan Luis Esteban and Jacobo Torán. Space bounds for resolution. Lecture Notes in Computer Science, 1563:551– 560, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1563/15630551.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1563/15630551.
pdf.

Niedermeier:1999:UBV

[622] Rolf Niedermeier and Peter Rossmanith. Upper bounds for vertex cover further improved. Lecture Notes in Computer Science, 1563:561–570, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1563/15630561.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/1563/15630561.
pdf.

Riedel:1999:OMS

[623] Marco Riedel. Online matching for scheduling problems. Lecture Notes in Computer Science, 1563:571–580, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1563/15630571.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1563/15630571.
pdf.

Akoka:1999:CDP

[624] Jacky Akoka. Conceptual design of parallel systems. Lecture Notes in Computer Science, 1565:1–23, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1565/15650001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1565/15650001.
pdf.

Boehm:1999:CMC

[625] Barry Boehm and Dan Port. Conceptual modeling challenges for modelbased architecting and software engineering (MBASE). Lecture Notes in Computer Science, 1565:24–43, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1565/15650024.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1565/15650024.
pdf.

Castano:1999:DGC

[626] Silvana Castano and Valeria De Antonellis. Deriving global conceptual views from multiple information sources. Lecture Notes in Computer Science, 1565:44–55, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1565/15650044.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1565/15650044.
                                        pdf.

Chen:1999:AEL

[627] Peter P. Chen. From Ancient Egyptian language to future conceptual modeling. Lecture Notes in Computer Science, 1565:56–64, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1565/15650056.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1565/15650056.
                                        pdf.

Delcambre:1999:BDB

[628] Lois M. L. Delcambre. Breaking the database out of the box: Models for superimposed information. Lecture Notes in Computer Science, 1565:65– 72, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1565/15650065.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1565/15650065.
                                        pdf.

Furtado:1999:NTD

[629] Antonio L. Furtado. Narratives and temporal databases: An interdisciplinary perspective. Lecture Notes in Computer Science, 1565:73–86, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1565/15650073.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1565/15650073.
pdf.

Gardarin:1999:UCM

[630] Georges Gardarin and Fei Sha. Using conceptual modeling and intelligent agents to integrate semi-structured documents in federated databases. Lecture Notes in Computer Science, 1565: 87–99, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1565/15650087.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1565/15650087.
pdf.

Goan:1999:SUC

[631] Terrance Goan. Supporting the user: Conceptual modeling & knowledge discovery. Lecture Notes in Computer Science, 1565:100–104, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1565/15650100.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1565/15650100.
pdf.

Kangassalo:1999:GUC

[632] Hannu Kangassalo. Are global understanding, communication, and information management in information systems possible? Lecture Notes in Computer Science, 1565:105–122, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1565/15650105.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1565/15650105.
pdf.

Loucopoulos:1999:EKM

[633] Pericles Loucopoulos and Vagelio Kavakli. Enterprise knowledge management and conceptual modelling. Lecture Notes in Computer Science, 1565:123–143, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1565/15650123.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1565/15650123.pdf.

McCarthy:1999:SMA

[634] William E. McCarthy. Semantic modeling in accounting education, practice, and research: Some progress and impediments. Lecture Notes in Computer Science, 1565:144–153, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1565/15650144.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1565/15650144.
pdf.

Miller:1999:PMS

[635] John A. Miller, Amit P. Sheth, and Krys J. Kochut. Perspectives in modeling: Simulation, database, and workflow. Lecture Notes in Computer Science, 1565:154–167, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1565/15650154.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1565/15650154.
pdf.

Navathe:1999:CMB

[636] Shamkant B. Navathe and Andreas M. Kogelnik. The challenges of modeling biological information for genome databases. Lecture Notes in Computer Science, 1565:168–182, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1565/15650168.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1565/15650168.
pdf.

Rosenthal:1999:TSD

[637] Arnon Rosenthal and Len Seligman. Trends and scale-up for data administration. Lecture Notes in Computer Science, 1565:183–194, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1565/15650183.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1565/15650183.
pdf.

Singh:1999:CMM

[638] Munindar P. Singh. Conceptual modeling for multiagent systems: Applying interaction-oriented programming. Lecture Notes in Computer Science, 1565:195–210, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1565/15650195.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1565/15650195.
                                        pdf.

Soelvberg:1999:DWT

[639] A. Soelvberg. Data and what they refer to. Lecture Notes in Computer Science, 1565:211–226, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Solvberg:1999:DWT

[640] Arne Sølvberg. Data and what they refer to. Lecture Notes in Computer Science, 1565:211–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1565/15650211.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1565/15650211.
                                        pdf.

Thalheim:1999:SM

[641] Bernhard Thalheim. The strength of ER modeling. Lecture Notes in Computer Science, 1565:227–242, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1565/15650227.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1565/15650227.
pdf.

Wegner:1999:IFM

[642] Peter Wegner and Dina Goldin. Interaction as a framework for modeling. Lecture Notes in Computer Science, 1565:243–257, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1565/15650243.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1565/15650243.pdf.

Liddle:1999:SEW

[643] Stephen W. Liddle, Stephen W. Clyde, and Scott N. Woodfield. A summary of the ER’97 workshop on behavioral modeling. Lecture Notes in Computer Science, 1565:258–271, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1565/15650258.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1565/15650258.
pdf.

Ramesh:1999:WRC

[644] Venkataraman Ramesh, Jeffrey Parsons, and Glenn J. Browne. What is the role of cognition in conceptual modeling? A report on the first workshop on cognition and conceptual modeling. Lecture Notes in Computer Science, 1565:272–280, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1565/15650272.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1565/15650272.
                                        pdf.

Meghini:1999:CMM

[645] Carlo Meghini and Fabrizio Sebastiani. Conceptual modeling in multimedia information seeking. Lecture Notes in Computer Science, 1565:281– 286, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1565/15650281.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1565/15650281.
                                        pdf.

Chen:1999:FDC

[646] Peter P. Chen, Bernhard Thalheim, and Leah Y. Wong. Future directions of conceptual modeling. Lecture Notes in Computer Science, 1565:287– 302, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1565/15650287.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1565/15650287.
                                        pdf.

Chen:1999:CIC

[647] Peter P. Chen. Current issues of conceptual modeling: A summary of selective active research topics. Lecture Notes in Computer Science, 1565:IX– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1565/156500IX.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1565/156500IX.
pdf.

Zadeh:1999:SRR

[648] L. A. Zadeh. Some reflections on the relationship between AI and fuzzy logic (FL) — A heretical view. Lecture Notes in Computer Science, 1566:1–8, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dubois:1999:PFL

[649] D. Dubois and H. Prade. The place of fuzzy logic in AI. Lecture Notes in Computer Science, 1566:9–21, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baldwin:1999:MAF

[650] J. F. Baldwin. Mass assignment fundamentals for computing with words. Lecture Notes in Computer Science, 1566: 22–44, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Narazaki:1999:MUU

[651] H. Narazaki and L. Shigaki. A method to use uncertain domain knowledge in the induction of classification knowledge based on ID3. Lecture Notes in Computer Science, 1566:45–61, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baldwin:1999:FLO

[652] J. F. Baldwin, T. P. Martin, and M. Vargas-Vera. FRIL++ a language for object-oriented programming with uncertainty. Lecture Notes in Computer Science, 1566:62–78, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dubois:1999:CBR

[653] D. Dubois, F. Esteva, P. Garcia, and L. Godo. Case-based reasoning: A fuzzy approach. Lecture Notes in Computer Science, 1566:79–90, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baldwin:1999:SIF

[654] J. F. Baldwin, T. P. Martin, and J. G. Shanahan. System identification of fuzzy Cartesian granules feature models using genetic programming. Lecture Notes in Computer Science, 1566:91–116, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tano:1999:DFS

[655] S. Tano. Deep fusion of symbolic and computational processing for next generation user interface. Lecture Notes in Computer Science, 1566:117–132, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guesgen:1999:RWA

[656] H. W. Guesgen. Reasoning with words about geographic information. Lecture Notes in Computer Science, 1566:133– 148, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bloch:1999:FMF

[657] I. Bloch. Fuzzy morphology and fuzzy distances: New definitions and links in both Euclidean and geodesic cases. Lecture Notes in Computer Science, 1566:149–165, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roux:1999:APT

[658] L. Roux. An application of possibility theory information fusion to satellite image classification. Lecture Notes in Computer Science, 1566:166–179, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wendling:1999:PRS

[659] L. Wendling and J. Desachy. Pattern recognition of strong graphs based on possibilistic c-means and k-formulae matching. Lecture Notes in Computer Science, 1566:180–189, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zerrouki:1999:FNM

[660] L. Zerrouki, B. Bouchon-Meunier, and R. Fondacci. A fuzzy-neural model for co-ordination in air traffic flow management. Lecture Notes in Computer Science, 1566:190–204, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1999:NIC

[661] J. Zhang and A. Knoll. From numerical interpolation to constructing intelligent behaviors. Lecture Notes in Computer Science, 1566:205–221, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Niskanen:1999:BLD

[662] V. A. Niskanen. A brief logopedics for the data used in a neuro-fuzzy milieu. Lecture Notes in Computer Science, 1566:222–233, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ralescu:1999:EFQ

[663] A. L. Ralescu, D. A. Ralescu, and K. Hirota. Evaluation of fuzzy quantified expressions. Lecture Notes in Computer Science, 1566:234–244, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1999:CSH

[664] Haoxun Chen and Hans-Michael Hanisch. Control synthesis of hybrid systems based on predicate invariance. Lecture Notes in Computer Science, 1567:1–15, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1567/15670001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1567/15670001.
pdf.

Chutinan:1999:CAA

[665] Alongkrit Chutinan and Bruce H. Krogh. Computing approximating automata for a class of linear hybrid systems. Lecture Notes in Computer Science, 1567:16–37, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1567/15670016.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1567/15670016.
pdf.

Davoren:1999:HSM

[666] J. M. Davoren. On hybrid systems and the modal μ-calculus. Lecture Notes in Computer Science, 1567:38– 69, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1567/15670038.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1567/15670038.
pdf.

DeSchutter:1999:ELC

[667] B. De Schutter and B. De Moor. The extended linear complementarity problem and the modeling and analysis of hybrid systems. Lecture Notes in Computer Science, 1567:70–85, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schutter:1999:ELC

[668] Bart De Schutter and Bart De Moor. The extended linear complementarity problem and the modeling and analysis of hybrid systems. Lecture Notes in Computer Science, 1567:70– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1567/15670070.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1567/15670070.
pdf.

Ferreira:1999:CSU

[669] Enrique D. Ferreira and Bruce H. Krogh. Controller scheduling using neural networks: Implementation and experimental results. Lecture Notes in Computer Science, 1567:86– 99, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1567/15670086.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1567/15670086.
pdf.

Gao:1999:RDF

[670] Jianping Gao and Qiwen Xu. Rigorous design of a fault diagnosis and isolation algorithm. Lecture Notes in Computer Science, 1567:100–121, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1567/15670100.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1567/15670100.
pdf.

Kohn:1999:SDS

[671] Wolf Kohn, Anil Nerode, and Jeffrey B. Remmel. Scalable data and sensor fusion via multiple agent hybrid systems. Lecture Notes in Computer Science, 1567:122–141, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1567/15670122.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1567/15670122.
pdf.

Koutsoukos:1999:HCS

[672] Xenofon D. Koutsoukos and Panos J. Antsaklis. Hybrid control systems using timed Petri nets: Supervisory control design based on invariant properties. Lecture Notes in Computer Science, 1567:142–162, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1567/15670142.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1567/15670142.
pdf.

Kowalewski:1999:CST

[673] S. Kowalewski, O. Stursberg, M. Fritz, H. Graf, I. Hoffmann, J. Preußig, M. Remelhe, S. Simon, and H. Treseler. A case study in tool-aided analysis of discretely controlled continuous systems: The two tanks problem. Lecture Notes in Computer Science, 1567:163– 185, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1567/15670163.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1567/15670163.
pdf.

Lafferriere:1999:HSF

[674] Gerardo Lafferriere, George J. Pappas, and Shankar Sastry. Hybrid systems with finite bisimulations. Lecture Notes in Computer Science, 1567:186– 203, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1567/15670186.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1567/15670186.
pdf.

Lemmon:1999:EST

[675] Michael Lemmon and Christopher J. Bett. Extracting stable timed automata from switched model reference systems. Lecture Notes in Computer Science, 1567:204–217, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1567/15670204.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1567/15670204.
                                        pdf.

Mosterman:1999:SMM

[676] Pieter J. Mosterman, Feng Zhao, and Gautam Biswas. Sliding mode model semantics and simulation for hybrid systems. Lecture Notes in Computer Science, 1567:218–237, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1567/15670218.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1567/15670218.
                                        pdf.

Nadjm-Tehrani:1999:TDH

[677] Simin Nadjm-Tehrani. Time-deterministic hybrid transition systems. Lecture Notes in Computer Science, 1567:238– 250, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1567/15670238.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1567/15670238.
                                        pdf.

Neller:1999:HOD

[678] Todd W. Neller. Heuristic optimization and dynamical system safety verification. Lecture Notes in Computer Science, 1567:251–261, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1567/15670251.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1567/15670251.
pdf.

Parisini:1999:HRH

[679] T. Parisini and S. Sacone. A hybrid receding-horizon control scheme for nonlinear discrete-time systems. Lecture Notes in Computer Science, 1567: 262–278, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1567/15670262.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1567/15670262.
pdf.

Raisch:1999:AAD

[680] Jörg Raisch, Eberhard Klein, Siu O’Young, Christian Meder, and Alexander Itigin. Approximating automata and discrete control for continuous systems — two examples from process controll. Lecture Notes in Computer Science, 1567:279–303, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1567/15670279.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1567/15670279.
                                        pdf.

Roenkkoe:1999:ASC

[681] M. Roenkkoe and A. P. Ravn. Action systems with continuous behaviour. Lecture Notes in Computer Science, 1567:304–323, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ronkko:1999:ASC

[682] Mauno Rönkkö and Anders P. Ravn. Action systems with continuous behaviour. Lecture Notes in Computer Science, 1567:304–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1567/15670304.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1567/15670304.
                                        pdf.

Seibel:1999:TUH

[683] Conrado W. Seibel, Jean-Marie Farines, and José E. R. Cury. Towards using hybrid automata for the mission planning of unmanned aerial vehicles. Lecture Notes in Computer Science, 1567:324–340, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1567/15670324.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1567/15670324.
                                        pdf.

Skafidas:1999:OCS

[684] E. Skafidas, R. J. Evans, I. M. Y. Mareels, and A. Nerode. Optimal controller switching for stochastic systems. Lecture Notes in Computer Science, 1567:341–355, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1567/15670341.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1567/15670341.
pdf.

Wang:1999:RCH

[685] Le Yi Wang, Pramod P. Khargonekar, and Ali Beydoun. Robust control of hybrid systems: Performance guided strategies. Lecture Notes in Computer Science, 1567:356–389, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1567/15670356.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1567/15670356.
pdf.

Wong-Toi:1999:ASP

[686] Howard Wong-Toi. Analysis of slopeparametric rectangular automata. Lecture Notes in Computer Science, 1567: 390–413, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1567/15670390.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1567/15670390.
pdf.

Yu:1999:CIT

[687] Xinyao Yu and Zongji Chen. Control of interval temporal systems. Lecture Notes in Computer Science, 1567:414– 428, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1567/15670414.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1567/15670414.
                                        pdf.

Zhao:1999:PSN

[688] Feng Zhao, Shiou C. Loh, and Jeff A. May. Phase-space nonlinear control toolbox: The maglev experience. Lecture Notes in Computer Science, 1567:429–444, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1567/15670429.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1567/15670429.
                                        pdf.

DeFloriani:1999:IPM

[689] L. De Floriani, P. Magillo, and E. Puppo. Invited paper: Multiresolution representation of shapes based on cell complexes. Lecture Notes in Computer Science, 1568:3–18, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Floriani:1999:IPM

[690] Leila De Floriani, Paola Magillo, and Enrico Puppo. Invited paper: Multiresolution representation of shapes based on cell complexes. Lecture Notes in Computer Science, 1568:3– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1568/15680003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1568/15680003.
pdf.

Borgefors:1999:DDS

[691] Gunilla Borgefors, Gabriella Sanniti di Baja, and Stina Svensson. Decomposing digital 3D shapes using a multiresolution structure. Lecture Notes in Computer Science, 1568:19– 30, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1568/15680019.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1568/15680019.
pdf.

Feschet:1999:OTC

[692] Fabien Feschet and Laure Tougne. Optimal time computation of the tangent of a discrete curve: Application to the curvature. Lecture Notes in Computer Science, 1568:31– 40, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1568/15680031.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1568/15680031.
pdf.

Zunic:1999:DMC

[693] Jovisa Zunic. The discrete moments of the circles. Lecture Notes in Computer Science, 1568:41–49, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1568/15680041.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1568/15680041.
                                        pdf.

Brimkov:1999:GPT

[694] Valentin E. Brimkov and Reneta P. Barneva. Graceful planes and thin tunnel-free meshes. Lecture Notes in Computer Science, 1568:53–64, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1568/15680053.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1568/15680053.
                                        pdf.

Gerard:1999:LCD

[695] Yan Gérard. Local configurations of digital hyperplanes. Lecture Notes in Computer Science, 1568:65–75, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1568/15680065.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1568/15680065.
                                        pdf.

Vittone:1999:CFN

[696] J. Vittone and J. M. Chassery. (n, m)-cubes and Farey nets for naive planes understanding. Lecture Notes in Computer Science, 1568:76–87, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1568/15680076.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1568/15680076.
pdf.

Ayala:1999:DLF

[697] R. Ayala, E. Domínguez, A. R. Francés, and A. Quintero. A digital lighting function for strong 26-surfaces. Lecture Notes in Computer Science, 1568:91– 103, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1568/15680091.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1568/15680091.
pdf.

Fourey:1999:INP

[698] Sébastien Fourey and Rémy Malgouyres. Intersection number of paths lying on a digital surface and a new Jordan theorem. Lecture Notes in Computer Science, 1568:104–117, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1568/15680104.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1568/15680104.
pdf.

Kovalevsky:1999:TMS

[699] Vladimir Kovalevsky. A topological method of surface representation. Lecture Notes in Computer Science, 1568: 118–135, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1568/15680118.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1568/15680118.
                                        pdf.

Malgouyres:1999:PFG

[700] Rémy Malgouyres. Presentation of the fundamental group in digital surfaces. Lecture Notes in Computer Science, 1568:136–150, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1568/15680136.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1568/15680136.
                                        pdf.

Kuba:1999:IPR

[701] Attila Kuba. Invited paper: Reconstruction in different classes of 2D discrete sets. Lecture Notes in Computer Science, 1568:153–163, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1568/15680153.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1568/15680153.
                                        pdf.

Giesen:1999:CRA

[702] Joachim Giesen. Curve reconstruction in arbitrary dimension and the traveling salesman problem. Lecture Notes in Computer Science, 1568:164– 176, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1568/15680164.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1568/15680164.
pdf.

Matsumoto:1999:SSS

[703] Y. Matsumoto, K. Fujimura, and T. Kitamura. Shape-from-silhouette/ stereo and its application to 3-D digitizer. Lecture Notes in Computer Science, 1568:177–188, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1568/15680177.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1568/15680177.
pdf.

Serra:1999:IPS

[704] Jean Serra. Invited paper: Set connections and discrete filtering. Lecture Notes in Computer Science, 1568:191– 206, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1568/15680191.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1568/15680191.
pdf.

Ahronovitz:1999:TOT

[705] Ehoud Ahronovitz, Christophe Fiorio, and Sylvain Glaize. Topological operators on the topological graph of frontiers. Lecture Notes in Computer Science, 1568:207–217, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1568/15680207.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1568/15680207.
                                        pdf.

Bertrand:1999:NND

[706] Gilles Bertrand. New notions for discrete topology. Lecture Notes in Computer Science, 1568:218–228, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1568/15680218.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1568/15680218.
                                        pdf.

Bertrand:1999:MDT

[707] Gilles Bertrand and Michel Couprie. A model for digital topology. Lecture Notes in Computer Science, 1568:229– 241, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1568/15680229.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1568/15680229.
                                        pdf.

Bertrand:1999:BMT

[708] Yves Bertrand, Christophe Fiorio, and Yann Pennaneach. Border map: A topological representation for n D image analysis. Lecture Notes in Computer Science, 1568:242–257, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1568/15680242.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1568/15680242.
pdf.

Cointepas:1999:DHD

[709] Yann Cointepas, Isabelle Bloch, and Line Garnero. A discrete homotopic deformable model dealing with objects with different local dimensions. Lecture Notes in Computer Science, 1568:258– 271, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1568/15680258.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1568/15680258.
pdf.

Eklund:1999:UTM

[710] Patrik Eklund and Fredrik Georgsson. Unraveling the thrill of metric image spaces. Lecture Notes in Computer Science, 1568:275–285, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1568/15680275.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1568/15680275.
pdf.

Hagedoorn:1999:MRC

[711] Michiel Hagedoorn and Remco Veltkamp. Measuring resemblance of complex patterns. Lecture Notes in Computer Science, 1568:286–298, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1568/15680286.
htm;   http://link.springer-ny.

                                        com/link/service/series/0558/papers/
                                        1568/15680286.pdf.

Mattes:1999:TRI

[712] Julian Mattes, Mathieu Richard, and Jacques Demongeot. Tree representation for image matching and object recognition. Lecture Notes in Computer Science, 1568:299–309, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1568/15680298.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1568/15680298.
                                        pdf.

Manzanera:1999:UFS

[713] Antoine Manzanera, Thierry M. Bernard, Françoise Prêteux, and Bernard Longuet. Ultra-fast skeleton based on an isotropic fully parallel algorithm. Lecture Notes in Computer Science, 1568:313–324, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1568/15680313.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1568/15680313.
                                        pdf.

Palagyi:1999:DTU

[714] Kálmán Palágyi and Attila Kuba. Directional 3D thinning using 8 subiterations. Lecture Notes in Computer Science, 1568:325–336, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/1568/15680325.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1568/15680325.
pdf.

Ranwez:1999:OIH

[715] Vincent Ranwez and Pierre Soille. Order independent homotopic thinning. Lecture Notes in Computer Science, 1568:337–346, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1568/15680337.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1568/15680337.pdf.

Lieutier:1999:IPC

[716] André Lieutier. Invited paper: Computable partial solids and voxels sets. Lecture Notes in Computer Science, 1568:349–360, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1568/15680349.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1568/15680349.pdf.

Balza-Gomez:1999:CHG

[717] Hélymar Balza-Gomez, Jean-Michel Moreau, and Dominique Michelucci. Convex hull of grid points below a line or a convex curve. Lecture Notes in Computer Science, 1568:361–374, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1568/15680361.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1568/15680361.
                                        pdf.

Devillers:1999:RVD

[718] Olivier Devillers and Pierre-Marie M. Gandoin. Rounding Voronoi diagram. Lecture Notes in Computer Science, 1568:375–387, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1568/15680375.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1568/15680375.pdf.

Figueiredo:1999:DBC

[719] Oscar Figueiredo, Jean-Pierre Reveillès, and Roger D. Hersch. Digitization of Bézier curves and patches using discrete geometry. Lecture Notes in Computer Science, 1568:388–398, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1568/15680388.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1568/15680388.
                                        pdf.

Tajine:1999:HDC

[720] Mohamed Tajine, Daniel Wagner, and Christian Ronse. Hausdorff discretization and its comparison to other discretization schemes. Lecture Notes in Computer Science, 1568:399–410, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1568/15680399.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1568/15680399.
pdf.

Aveneau:1999:DTS

[721] Lilian Aveneau, Eric Andres, and Michel Mériaux. The discrete tube: A spatial acceleration technique for efficient diffraction computation. Lecture Notes in Computer Science, 1568:413– 424, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1568/15680413.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1568/15680413.
pdf.

Francon:1999:PBV

[722] Jean Françon and Laurent Papier. Polyhedrization of the boundary of a voxel object. Lecture Notes in Computer Science, 1568:425–434, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1568/15680425.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1568/15680425.
pdf.

Salam:1999:DRC

[723] Ivan Salam, Philippe Nehlig, and Eric Andres. Discrete ray-casting. Lecture Notes in Computer Science, 1568:435– 446, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1568/15680435.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/1568/15680435.
pdf.

Tellier:1999:DNV

[724] Pierre Tellier and Isabelle Debled-Rennesson. 3D discrete normal vectors. Lecture Notes in Computer Science, 1568:447–458, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1568/15680447.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1568/15680447.pdf.

Baufreton:1999:SSA

[725] Philippe Baufreton. SACRES: A step ahead in the development of critical avionics applications. Lecture Notes in Computer Science, 1569:1– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1569/15690001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1569/15690001.
pdf.

Krogh:1999:AHS

[726] Bruce H. Krogh. Approximating hybrid system dynamics for analysis and control. Lecture Notes in Computer Science, 1569:2–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1569/15690002.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1569/15690002.
pdf.

Lynch:1999:HLM

[727] Nancy Lynch. High-level modeling and analysis of an air-traffic management system. Lecture Notes in Computer Science, 1569:3–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1569/15690003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1569/15690003.
pdf.

Miller:1999:GCM

[728] Chris Miller. Geometric categories, Ominimal structures and control. Lecture Notes in Computer Science, 1569: 4–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1569/15690004.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1569/15690004.
pdf.

Alur:1999:PFH

[729] Rajeev Alur, Sampath Kannan, and Salvatore La Torre. Polyhedral flows in hybrid automata. Lecture Notes in Computer Science, 1569:5– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1569/15690005.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1569/15690005.
pdf.

Asarin:1999:SPT

[730] Eugene Asarin and Oded Maler. As soon as possible: Time optimal control for timed automata. Lecture Notes in Computer Science, 1569:19– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1569/15690019.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1569/15690019.
pdf.

Bemporad:1999:VHS

[731] Alberto Bemporad and Manfred Morari. Verification of hybrid systems via mathematical programming. Lecture Notes in Computer Science, 1569: 31–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1569/15690031.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1569/15690031.
pdf.

Bournez:1999:OPR

[732] Olivier Bournez, Oded Maler, and Amir Pnueli. Orthogonal polyhedra: Representation and computation. Lecture Notes in Computer Science, 1569: 46–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1569/15690046.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1569/15690046.
pdf.

Broucke:1999:GAB

[733] Mireille Broucke. A geometric approach to bisimulation and verification of hybrid systems. Lecture Notes in Computer Science, 1569:61– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1569/15690061.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1569/15690061.
pdf.

Chutinan:1999:VPI

[734] Alongkrit Chutinan and Bruce H. Krogh. Verification of polyhedralinvariant hybrid automata using polygonal flow pipe approximations. Lecture Notes in Computer Science, 1569:76– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1569/15690076.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1569/15690076.
pdf.

Egerstedt:1999:PPF

[735] M. Egerstedt, T. J. Koo, F. Hoffmann, and S. Sastry. Path planning and flight controller scheduling for an autonomous helicopter. Lecture Notes in Computer Science, 1569:91– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1569/15690091.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1569/15690091.
pdf.

Greenstreet:1999:RAU

[736] Mark R. Greenstreet and Ian Mitchell. Reachability analysis using polygonal projections. Lecture Notes in Computer Science, 1569:103–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1569/15690103.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1569/15690103.
                                        pdf.

Hespanha:1999:SIH

[737] J. P. Hespanha and A. Stephen Morse. Scale-independent hysteresis switching. Lecture Notes in Computer Science, 1569:117–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1569/15690117.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1569/15690117.
                                        pdf.

Imura:1999:WPC

[738] Jun ichi Imura and Arjan van der Schaft. Well-posedness of a class of piecewise linear systems with no jumps. Lecture Notes in Computer Science, 1569:123–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1569/15690123.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1569/15690123.
                                        pdf.

Lafferriere:1999:NCD

[739] Gerardo Lafferriere, George J. Pappas, and Sergio Yovine. A new class of decidable hybrid systems. Lecture Notes in Computer Science, 1569:137– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1569/15690137.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1569/15690137.
pdf.

Lin:1999:SCS

[740] Man Lin. Synthesis of control software in a layered architecture from hybrid automata. Lecture Notes in Computer Science, 1569:152–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1569/15690152.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1569/15690152.
pdf.

Mosterman:1999:OHS

[741] Pieter J. Mosterman. An overview of hybrid simulation phenomena and their support by simulation packages. Lecture Notes in Computer Science, 1569: 165–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1569/15690165.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1569/15690165.
pdf.

Mosterman:1999:BHO

[742] Pieter J. Mosterman and Gautam Biswas. Building hybrid observers for complex dynamic systems using model abstractions. Lecture Notes in Computer Science, 1569:178–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1569/15690178.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1569/15690178.
                                        pdf.

Nadjm-Tehrani:1999:BHO

[743] S. Nadjm-Tehrani. Building hybrid observers for complex dynamic systems using model abstractions integration of analog and discrete synchronous design. Lecture Notes in Computer Science, 1569:193–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nadjm-Tehrani:1999:IAD

[744] Simin Nadjm-Tehrani. Integration of analog and discrete synchronous design. Lecture Notes in Computer Science, 1569:193–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1569/15690193.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1569/15690193.
                                        pdf.

Preussig:1999:RAC

[745] J. Preußig, O. Stursberg, and S. Kowalewski. Reachability analysis of a class of switched continuous systems by integrating rectangular approximation and rectangular analysis. Lecture Notes in Computer Science, 1569:209– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1569/15690209.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1569/15690209.
pdf.

Roenkkoe:1999:RCB

[746] M. Roenkkoe and K. Sere. Refinement and continuous behaviour. Lecture Notes in Computer Science, 1569: 223–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ronkko:1999:RCB

[747] Mauno Rönkkö and Kaisa Sere. Refinement and continuous behaviour. Lecture Notes in Computer Science, 1569: 223–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1569/15690223.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1569/15690223.
pdf.

Tomlin:1999:CCN

[748] Claire Tomlin, John Lygeros, and Shankar Sastry. Computing controllers for nonlinear hybrid systems. Lecture Notes in Computer Science, 1569:238– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/1569/15690238.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1569/15690238.
                                        pdf.

Yfoulis:1999:SOP

[749] C. A. Yfoulis, A. Muir, P. E. Wellstead, and N. B. O. L. Pettit. Stabilization of orthogonal piecewise linear systems: Robustness analysis and design. Lecture Notes in Computer Science, 1569: 256–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1569/15690256.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1569/15690256.
                                        pdf.

Studer:1999:KES

[750] R. Studer, D. Fensel, S. Decker, and V. R. Benjamins. Knowledge engineering — survey and future directions. Lecture Notes in Computer Science, 1570:1–23, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dressler:1999:KBD

[751] O. Dressler and F. Puppe. Knowledgebased diagnosis — survey and future directions. Lecture Notes in Computer Science, 1570:24–46, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guenter:1999:KBC

[752] A. Guenter and C. Kuehn. Knowledgebased configuration — survey and future directions. Lecture Notes in Computer Science, 1570:47–66, 1999.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bartsch-Spoerl:1999:CBR

[753] B. Bartsch-Spoerl, M. Lenz, and A. Huebner. Case-based reasoning — survey and future directions. Lecture Notes in Computer Science, 1570:67– 89, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guckenbiehl:1999:MRU

[754] T. Guckenbiehl, H. Milde, B. Neumann, and P. Struss. Meeting re-use requirements of real-life diagnosis applications. Lecture Notes in Computer Science, 1570:90–100, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dorn:1999:TRI

[755] J. Dorn. Towards reusable intelligent scheduling software. Lecture Notes in Computer Science, 1570:101–112, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abecker:1999:OMK

[756] A. Abecker and S. Decker. Organizational memory: Knowledge acquisition, integration, and retrieval issues. Lecture Notes in Computer Science, 1570:113–124, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liao:1999:CKB

[757] M. Liao, K. Hinkelmann, A. Abecker, and M. Sintek. A competence knowledge base system as part of the organizational memory. Lecture Notes in Computer Science, 1570:125–137, 1999.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nick:1999:PEO

[758] M. Nick and C. Tautz. Practical evaluation of an organizational memory using the goal-question-metric technique. Lecture Notes in Computer Science, 1570:138–147, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lenz:1999:TCC

[759] M. Lenz and A. Glintschert. On texts, cases, and concepts. Lecture Notes in Computer Science, 1570:148–156, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lees:1999:ICB

[760] B. Lees and J. Corchado. Integrated case-based neural network approach to problem solving. Lecture Notes in Computer Science, 1570:157–166, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Epke:1999:PEK

[761] K. Epke. PROPLANT — experiences with a knowledge-based system for plant protection. Lecture Notes in Computer Science, 1570:167–175, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaufmann:1999:WKB

[762] G. Kaufmann, F. Braun, K. Egli, and M. Hodel. WekI — A knowledge based system of suva for the assignment of companies to risk related groups. Lecture Notes in Computer Science, 1570:176–185, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Imhof:1999:EST

[763] H. Imhof. Expert system technology in aircraft fleet management — the development of SAREX. Lecture Notes in Computer Science, 1570:186–192, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schewe:1999:EKB

[764] S. Schewe, B. Reinhardt, and C. Betz. Experiences with a knowledge-based tutoring system for student education in rheumatology. Lecture Notes in Computer Science, 1570:193–200, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pfisterer:1999:CAE

[765] T. Pfisterer and T. Wetter. Computer assisted editing of genomic sequences — why and how we evaluated a prototype. Lecture Notes in Computer Science, 1570:201–209, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knaup:1999:EUC

[766] P. Knaup, T. Wiedemann, A. Bachert, and U. Creutzig. Experiences of using a computer-aided therapy planning for pediatric oncology in clinical routine. Lecture Notes in Computer Science, 1570:210–217, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hinrichs:1999:EKB

[767] H. Hinrichs and K. Panienski. Experiences with knowledge-based data cleansing at the Epidemiological Cancer Registry of Lower-Saxony. Lecture Notes in Computer Science, 1570:218– 226, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Callis:1999:BEM

[768] J. C. Callis and L. C. Lee. Building electronic marketplaces with the ZEUS agent tool-kit. Lecture Notes in Computer Science, 1571:1–24, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Collis:1999:BEM

[769] Jaron C. Collis and Lyndon C. Lee. Building electronic marketplaces with the ZEUS agent tool-kit. Lecture Notes in Computer Science, 1571:1– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1571/15710001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1571/15710001.
pdf.

Parkes:1999:ACC

[770] David C. Parkes, Lyle H. Ungar, and Dean P. Foster. Accounting for cognitive costs in on-line auction design. Lecture Notes in Computer Science, 1571: 25–40, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1571/15710025.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1571/15710025.
pdf.

Eriksson:1999:SMA

[771] Joakim Eriksson, Niclas Finne, and Sverker Janson. SICS MarketSpace — an agent-based market infrastructure. Lecture Notes in Computer Science, 1571:41–53, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1571/15710041.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1571/15710041.
pdf.

Andersson:1999:SCT

[772] Martin R. Andersson and Tuomas W. Sandholm. Sequencing of contract types for anytime task reallocation. Lecture Notes in Computer Science, 1571:54–69, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1571/15710054.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1571/15710054.
pdf.

Guttman:1999:AMI

[773] Robert H. Guttman and Pattie Maes. Agent-mediated integrative negotiation for retail electronic commerce. Lecture Notes in Computer Science, 1571:70– 90, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1571/15710070.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1571/15710070.
pdf.

Goldsmith:1999:MAS

[774] Steven Y. Goldsmith, Laurence R. Phillips, and Shannon V. Spires. A multi-agent system for coordinating international shipping. Lecture Notes in Computer Science, 1571:91–104, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1571/15710091.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1571/15710091.
                                        pdf.

Steinmetz:1999:BES

[775] Erik Steinmetz, John Collins, Scott Jamison, Rashmi Sundareswara, Bamshad Mobasher, and Maria Gini. Bid evaluation and selection in the MAGNET automated contracting system. Lecture Notes in Computer Science, 1571:105– 125, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1571/15710105.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1571/15710105.
                                        pdf.

Matos:1999:ECN

[776] Noyda Matos and Carles Sierra. Evolutionary computing and negotiating agents. Lecture Notes in Computer Science, 1571:126–150, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1571/15710126.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1571/15710126.
pdf.

Garcia:1999:BST

[777] Pere Garcia, Eduard Giménez, Lluís Godo, and Juan A. Rodríguez-Aguilar. Bidding strategies for trading agents in auction-based tournaments. Lecture Notes in Computer Science, 1571:151– 165, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1571/15710151.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1571/15710151.
pdf.

Padget:1999:CMS

[778] Julian Padget and Russell Bradford. A π-calculus model of a Spanish fish market — preliminary report. Lecture Notes in Computer Science, 1571:166– 188, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1571/15710166.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1571/15710166.
pdf.

Ghidini:1999:IIE

[779] Chiara Ghidini and Luciano Serafini. Information integration for electronic commerce. Lecture Notes in Computer Science, 1571:189–206, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1571/15710189.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1571/15710189.
                                        pdf.

Anonymous:1999:AIc

[780] Anonymous. Author index. Lecture Notes in Computer Science, 1571:207– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/1571/1571auth.pdf.

Schapire:1999:TVB

[781] Robert E. Schapire. Theoretical views of boosting. Lecture Notes in Computer Science, 1572:1–10, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1572/15720001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1572/15720001.
                                        pdf.

Sutton:1999:OTQ

[782] Richard S. Sutton. Open theoretical questions in reinforcement learning. Lecture Notes in Computer Science, 1572:11–17, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1572/15720011.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1572/15720011.
                                        pdf.

Duffy:1999:GAL

[783] Nigel Duffy and David Helmbold. A geometric approach to leveraging weak learners. Lecture Notes in Computer Science, 1572:18–33, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1572/15720018.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1572/15720018.
pdf.

Bachrach:1999:QCL

[784] Ran Bachrach, Shai Fine, and Eli Shamir. Query by committee, linear separation and random walks. Lecture Notes in Computer Science, 1572:34– 49, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1572/15720034.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1572/15720034.
pdf.

Bartlett:1999:HRN

[785] Peter Bartlett and Shai Ben-David. Hardness results for neural network approximation problems. Lecture Notes in Computer Science, 1572:50– 62, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1572/15720050.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1572/15720050.
pdf.

Dalmau:1999:LQF

[786] Víctor Dalmau and Peter Jeavons. Learnability of quantified formulas. Lecture Notes in Computer Science, 1572:63–78, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1572/15720063.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1572/15720063.
                                        pdf.

Forster:1999:LMA

[787] Jürgen Forster. Learning multiplicity automata from smallest counterexamples. Lecture Notes in Computer Science, 1572:79–90, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1572/15720079.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1572/15720079.
                                        pdf.

Guijarro:1999:ELW

[788] David Guijarro, Víctor Lavín, and Vijay Raghavan. Exact learning when irrelevant variables abound. Lecture Notes in Computer Science, 1572:91– 100, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1572/15720091.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1572/15720091.
                                        pdf.

Hofmeister:1999:ACA

[789] Thomas Hofmeister. An application of codes to attribute-efficient learning. Lecture Notes in Computer Science, 1572:101–110, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/

series/0558/bibs/1572/15720101.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1572/15720101.pdf.

Khardon:1999:LRR

[790] Roni Khardon. Learning range restricted horn expressions. Lecture Notes in Computer Science, 1572:111– 125, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1572/15720111.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1572/15720111.
pdf.

Tadic:1999:ABC

[791] Vladislav Tadic. On the asymptotic behavior of a constant stepsize temporaldifference learning algorithm. Lecture Notes in Computer Science, 1572:126– 137, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1572/15720126.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1572/15720126.
pdf.

Helmbold:1999:DIA

[792] D. P. Helmbold, S. Panizza, and M. K. Warmuth. Direct and indirect algorithms for on-line learning of disjunctions. Lecture Notes in Computer Science, 1572:138–152, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1572/15720138.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1572/15720138.
                                        pdf.

Kivinen:1999:AEP

[793] Jyrki Kivinen and Manfred K. Warmuth. Averaging expert predictions. Lecture Notes in Computer Science, 1572:153–167, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1572/15720153.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1572/15720153.pdf.

Kuhlmann:1999:TLI

[794] Christian Kuhlmann. On teaching and learning intersection-closed concept classes. Lecture Notes in Computer Science, 1572:168–182, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1572/15720168.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1572/15720168.
                                        pdf.

Ott:1999:ACT

[795] Matthias Ott and Frank Stephan. Avoiding coding tricks by hyperrobust learning. Lecture Notes in Computer Science, 1572:183–197, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1572/15720183.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1572/15720183.
pdf.

Jain:1999:MCC

[796] Sanjay Jain and Arun Sharma. Mind change complexity of learning logic programs. Lecture Notes in Computer Science, 1572:198–213, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1572/15720198.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1572/15720198.
pdf.

Smola:1999:RPM

[797] Alex J. Smola, Robert C. Williamson, Sebastian Mika, and Bernhard Schölkopf. Regularized principal manifolds. Lecture Notes in Computer Science, 1572: 214–229, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1572/15720214.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1572/15720214.
pdf.

Vayatis:1999:DDV

[798] Nicolas Vayatis and Robert Azencott. Distribution-dependent Vapnik– Chervonenkis bounds. Lecture Notes in Computer Science, 1572:230–240, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1572/15720230.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1572/15720230.
                                        pdf.

Antos:1999:LBR

[799] András Antos. Lower bounds on the rate of convergence of nonparametric pattern recognition. Lecture Notes in Computer Science, 1572:241– 252, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1572/15720241.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1572/15720241.
                                        pdf.

Horvath:1999:EEP

[800] Márta Horváth. On error estimation for the partitioning classification rule. Lecture Notes in Computer Science, 1572:253–262, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1572/15720253.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1572/15720253.
                                        pdf.

Shawe-Taylor:1999:MDB

[801] John Shawe-Taylor and Nello Cristianini. Margin distribution bounds on generalization. Lecture Notes in Computer Science, 1572:263–273, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1572/15720263.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/1572/15720263.
pdf.

Shawe-Taylor:1999:GPC

[802] John Shawe-Taylor and Robert C. Williamson. Generalization performance of classifiers in terms of observed covering numbers. Lecture Notes in Computer Science, 1572:274– 284, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1572/15720274.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1572/15720274.
pdf.

Williamson:1999:ENO

[803] Robert C. Williamson, Alex J. Smola, and Bernhard Schölkopf. Entropy numbers, operators and support vector kernels. Lecture Notes in Computer Science, 1572:285–300, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1572/15720285.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1572/15720285.
pdf.

Bai:1999:SUE

[804] Z. Bai and G. H. Golub. Some unusual eigenvalue problems (invited talk). Lecture Notes in Computer Science, 1573:4–19, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gansterer:1999:MSA

[805] W. N. Gansterer, D. F. Kvasnicka, and C. W. Ueberhuber. Multi-sweep algorithms for the symmetric eigenproblem. Lecture Notes in Computer Science, 1573:20–28, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gimenez:1999:UAP

[806] D. Gimenez, V. Hernandez, and A. M. Vidal. A unified approach to parallel block-Jacobi methods for the symmetric eigenvalue problem. Lecture Notes in Computer Science, 1573:29– 42, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hernandez:1999:CLM

[807] V. Hernandez, J. E. Roman, A. M. Vidal, and V. Vidal. Calculation of lambda modes of a nuclear reactor: A parallel implementation using the implicitly restarted Arnoldi method. Lecture Notes in Computer Science, 1573:43–57, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nool:1999:PJD

[808] M. Nool and A. Van der Ploeg. Parallel Jacobi–Davidson for solving generalized eigenvalue problems. Lecture Notes in Computer Science, 1573:58– 70, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Basermann:1999:PPS

[809] A. Basermann. Parallel preconditioned solvers for large sparse Hermitian eigenproblems. Lecture Notes in Computer Science, 1573:71–84, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gimenez:1999:SEP

[810] D. Gimenez, C. Jimenez, M. J. Majado, and N. Marin. Solving eigenvalue problems on networks of processors. Lecture Notes in Computer Science, 1573:85–99, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harrar:1999:SLS

[811] D. L. Harrar and M. R. Osborne. Solving large-scale eigenvalue problems on vector parallel processors. Lecture Notes in Computer Science, 1573:100– 113, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grund:1999:DLS

[812] F. Grund. Direct linear solvers for vector and parallel computers. Lecture Notes in Computer Science, 1573:114– 127, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garcia-Loureiro:1999:PPS

[813] A. J. Garcia-Loureiro, T. F. Pena, J. M. Lopez-Gonzalez, and L. P. Viñas. Parallel preconditioners for solving nonsymmetric linear systems. Lecture Notes in Computer Science, 1573: 128–141, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arnal:1999:SAP

[814] J. Arnal, V. Migallon, and J. Penades. Synchronous and asynchronous parallel algorithms with overlap for almost linear systems. Lecture Notes in Computer Science, 1573:142–155, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Husbands:1999:PPS

[815] P. Husbands and C. Isbell. The parallel problems server: A client-server model for interactive large scale scientific computation. Lecture Notes in Computer Science, 1573:156–170, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barth:1999:PDD

[816] T. J. Barth, T. F. Chan, and W.-P. Tang. Parallel domain-decomposition preconditioning for computational fluid dynamics (invited talk). Lecture Notes in Computer Science, 1573:176–202, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coelho:1999:IDS

[817] P. J. Coelho. Influence of the discretization scheme on the parallel efficiency of a code for the modelling of a utility boiler. Lecture Notes in Computer Science, 1573:203–214, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vicaire:1999:PAS

[818] J.-B. Vicaire, L. Prylli, G. Perrot, and B. Tourancheau. Parallel 3D airflow simulation on workstation clusters. Lecture Notes in Computer Science, 1573:215–226, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Strietzel:1999:PTS

[819] M. Strietzel and T. Gerz. Parallel turbulence simulation: Resolving the inertial subrange of the Kolmogorov spectrum. Lecture Notes in Computer Science, 1573:227–237, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Areal:1999:SPA

[820] P. M. Areal and J. M. L. M. Palma. The study of a parallel algorithm using the laminar backward-facing step flow as a test case. Lecture Notes in Computer Science, 1573:238–249, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moretti:1999:LCD

[821] C. O. Moretti, T. N. Bittencourt, and L. F. Martha. A low cost distributed system for FEM parallel structural analysis. Lecture Notes in Computer Science, 1573:250–262, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Galbas:1999:DLB

[822] H. G. Galbas and O. Kolp. Dynamic load balancing in crashworthiness simulation. Lecture Notes in Computer Science, 1573:263–270, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Becker:1999:SCS

[823] C. Becker, S. Kilian, and S. Turek. Some concepts of the software package FEAST. Lecture Notes in Computer Science, 1573:271–284, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Becket:1999:SCS

[824] C. Becket, S. Kilian, and S. Turek. Some concepts of the software package FEAST. Lecture Notes in Computer Science, 1573:271–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Walshaw:1999:MMP

[825] C. Walshaw, M. Cross, R. Diekmann, and F. Schlimbach. Multilevel mesh partitioning for optimising aspect ratio. Lecture Notes in Computer Science, 1573:285–300, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Welch:1999:PDC

[826] P. H. Welch. Parallel and distributed computing in education (invited talk). Lecture Notes in Computer Science, 1573:301–330, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vuillemin:1999:RSP

[827] J. Vuillemin. Reconfigurable systems: Past and next 10 years. Lecture Notes in Computer Science, 1573:334–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seo:1999:SAF

[828] S.-G. Seo, M. J. Downie, G. E. Hearn, and C. Phillips. A systolic algorithm for the factorisation of matrices arising in the field of hydrodynamics. Lecture Notes in Computer Science, 1573:355–364, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Espinosa:1999:ADP

[829] A. Espinosa, T. Margalef, and E. Luque. Automatic detection of parallel program performance problems. Lecture Notes in Computer Science, 1573:365–377, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Suarez:1999:BAM

[830] F. J. Suarez and D. F. Garcia. Behavioural analysis methodology oriented to configuration of parallel, realtime and embedded systems. Lecture Notes in Computer Science, 1573:378– 395, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Figueiredo:1999:SDL

[831] R. J. O. Figueiredo, J. A. B. Fortes, and Z. B. Miled. Spatial data locality with respect to degree of parallelism in processor-and-memory hierarchies. Lecture Notes in Computer Science, 1573:396–410, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prieto-Matias:1999:PRD

[832] M. Prieto-Matias, I. Martin-Llorente, and F. Tirado. Partitioning regular domains on modern parallel computers. Lecture Notes in Computer Science, 1573:411–424, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DelCorral:1999:NAO

[833] A. M. Del Corral and J. M. Llaberia. New access order to reduce intervector-conflicts. Lecture Notes in Computer Science, 1573:425–438, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Villa:1999:RSI

[834] L. Villa, R. Espasa, and M. Valero. Registers size influence on vector architectures. Lecture Notes in Computer Science, 1573:439–451, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gonzalez:1999:LIL

[835] J. Gonzalez and A. Gonzalez. Limits of instruction level parallelism with data value speculation. Lecture Notes in Computer Science, 1573:452–465, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garcia:1999:HPC

[836] F. Garcia, J. Carretero, F. Perez, and P. De Miguel. High performance cache management for parallel file systems. Lecture Notes in Computer Science, 1573:466–479, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kitajima:1999:USW

[837] J. P. Kitajima and S. Porto. Using synthetic workloads for parallel task scheduling improvement analysis. Lecture Notes in Computer Science, 1573:480–493, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Franco:1999:DRB

[838] D. Franco, I. Garces, and E. Luque. Dynamic routing balancing in parallel computer interconnection networks. Lecture Notes in Computer Science, 1573:494–507, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barbosa:1999:ADM

[839] J. Barbosa and A. Padilha. Algorithmdependant method to determine the optimal number of computers in parallel virtual machines. Lecture Notes in Computer Science, 1573:508–521, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martin:1999:LCP

[840] M. Martin and B. Chopard. Low cost parallelizing: A way to be efficient. Lecture Notes in Computer Science, 1573:522–533, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanderSteen:1999:PAS

[841] A. J. Van der Steen and R. Van der Pas. A performance analysis of the SGI Origin2000. Lecture Notes in Computer Science, 1573:534–547, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Quintana:1999:ICB

[842] F. Quintana, R. Espasa, and M. Valero. An ISA comparison between superscalar and vector processors. Lecture Notes in Computer Science, 1573:548– 560, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Priol:1999:HPC

[843] T. Priol. High performance computing for image synthesis (invited talk). Lecture Notes in Computer Science, 1573:563–578, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laurent:1999:PIM

[844] C. Laurent and J. Roman. Parallel implementations of morphological connected operators based on irregular data structures. Lecture Notes in Computer Science, 1573:579–592, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Watson:1999:DOC

[845] P. Watson. The design of an ODMG compatible parallel object database server (invited talk). Lecture Notes in Computer Science, 1573:593–622, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pereira:1999:PBI

[846] N. S. A. Pereira. A parallel N-body integrator using MPI. Lecture Notes in Computer Science, 1573:627–639, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sidonio:1999:PBI

[847] N. Sidonio and A. Pereira. A parallel N-body integrator using MPI. Lecture Notes in Computer Science, 1573: 627–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borges:1999:PSP

[848] C. L. T. Borges and D. M. Falcao. A parallelisation strategy for power systems composite reliability evaluation (best student paper award: Honourable mention). Lecture Notes in Computer Science, 1573:640–651, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arenaz:1999:HPC

[849] M. Arenaz, R. Doallo, G. Garcia, and C. Vazquez. High performance computing of an industrial problem in tribology (best student paper award: First prize). Lecture Notes in Computer Science, 1573:652–665, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sawyer:1999:PGM

[850] W. Sawyer, L. Takacs, A. Da Silva, and P. Lyster. Parallel grid manipulations in Earth science calculations. Lecture Notes in Computer Science, 1573:666–679, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keppens:1999:SMP

[851] R. Keppens and G. Toth. Simulating magnetised plasma with the versatile advection code. Lecture Notes in Computer Science, 1573:680–690, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baraglia:1999:PGA

[852] R. Baraglia and R. Perego. Parallel genetic algorithms for hypercube machines. Lecture Notes in Computer Science, 1573:691–703, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1999:KEI

[853] W. Kim. KDD as an enterprise IT tool: Reality and agenda. Lecture Notes in Computer Science, 1574:1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Motoda:1999:CAD

[854] H. Motoda. Computer assisted discovery of first principle equations from numeric data. Lecture Notes in Computer Science, 1574:2–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pawlak:1999:DMR

[855] Z. Pawlak. Data mining: A rough set perspective. Lecture Notes in Computer Science, 1574:3–12, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aggarwal:1999:DMT

[856] C. C. Aggarwal and P. S. Yu. Data mining techniques for associations, clustering and classification. Lecture Notes in Computer Science, 1574:13–23, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1999:DMG

[857] T. Y. Lin. Data mining: Granular computing approach. Lecture Notes in Computer Science, 1574:24–33, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsukimoto:1999:REP

[858] H. Tsukimoto. Rule extraction from prediction models. Lecture Notes in Computer Science, 1574:34–43, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Du:1999:MAR

[859] X. Du, Z. Liu, and N. Ishii. Mining association rules on related numeric attributes. Lecture Notes in Computer Science, 1574:44–53, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yip:1999:LLB

[860] C. L. Yip, K. K. Loo, B. Kao, and D. W. Cheung. LGen — A latticebased candidate set generation algorithm for I/O efficient association rule mining. Lecture Notes in Computer Science, 1574:54–63, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rajamani:1999:EAA

[861] K. Rajamani, S. Sung, and A. Cox. Extending the applicability of association rules. Lecture Notes in Computer Science, 1574:64–73, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsai:1999:EAI

[862] P. S. M. Tsai, C.-C. Lee, and A. L. P. Chen. An efficient approach for incremental association rule mining. Lecture Notes in Computer Science, 1574: 74–83, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kryszkiewicz:1999:ARI

[863] M. Kryszkiewicz. Association rules in incomplete databases. Lecture Notes in Computer Science, 1574:84–93, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pramudiono:1999:PSB

[864] I. Pramudiono, T. Shintani, T. Tamura, and M. Kitsuregawa. Parallel SQL based association rule mining on large scale PC cluster: Performance comparison with directly coded C implementation. Lecture Notes in Computer Science, 1574:94–98, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1999:RMH

[865] Y. Yang and M. Singhal. H-rule mining in heterogeneous databases. Lecture Notes in Computer Science, 1574: 99–103, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhou:1999:IDM

[866] A. Zhou, S. Zhou, W. Jin, and Z. Tian. An improved definition of multidimensional, inter-transaction association rule. Lecture Notes in Computer Science, 1574:104–108, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hu:1999:IDA

[867] K. Hu, Y. Lu, and C. Shi. Incremental discovering association rules: A concept lattice approach. Lecture Notes in Computer Science, 1574:109–113, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wu:1999:IPP

[868] X. Wu. Induction as pre-processing. Lecture Notes in Computer Science, 1574:114–122, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zheng:1999:SAS

[869] Z. Zheng and G. I. Webb. Stochastic attribute selection committees with multiple boosting: Learning more accurate and more stable classifier committees. Lecture Notes in Computer Science, 1574:123–132, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yao:1999:ITM

[870] Y. Y. Yao, S. K. M. Wong, and C. J. Butz. On information-theoretic measures of attribute importance. Lecture Notes in Computer Science, 1574:133– 137, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1999:TDF

[871] K.-C. Lee. A technique of dynamic feature selection using the feature group mutual information. Lecture Notes in Computer Science, 1574:138–142, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Terabe:1999:DPP

[872] M. Terabe, O. Katai, T. Sawaragi, and T. Washio. A data preprocessing method using association rules of attributes for improving decision tree. Lecture Notes in Computer Science, 1574:143–147, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Singh:1999:ACA

[873] L. Singh, B. Chen, R. Haight, and P. Scheuermann. An algorithm for constrained association rule mining in semi-structured data. Lecture Notes in Computer Science, 1574:148–158, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhou:1999:IMS

[874] A. Zhou, W. Jin, S. Zhou, and Z. Tian. Incremental mining of schema for semistructured data. Lecture Notes in Computer Science, 1574:159–168, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jiang:1999:DSD

[875] M.-F. Jiang, S.-S. Tseng, and C.-J. Tsai. Discovering structure from document databases. Lecture Notes in Computer Science, 1574:169–173, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cho:1999:CFM

[876] V. Cho and B. Wuethrich. Combining forecasts from multiple textual data sources. Lecture Notes in Computer Science, 1574:174–178, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meng:1999:DKE

[877] X. Meng, Y. Zhou, and S. Wang. Domain knowledge extracting in a Chinese natural language interface to databases: NChiql. Lecture Notes in Computer Science, 1574:179–183, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Williams:1999:EHS

[878] G. J. Williams. Evolutionary hot spots data mining: An architecture for exploring for interesting discoveries. Lecture Notes in Computer Science, 1574:184–193, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1999:ESR

[879] H. Liu, H. Lu, L. Feng, and F. Hussain. Efficient search of reliable exceptions. Lecture Notes in Computer Science, 1574:194–203, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hilderman:1999:HRI

[880] R. J. Hilderman and H. J. Hamilton. Heuristics for ranking the interestingness of discovered knowledge. Lecture Notes in Computer Science, 1574:204– 209, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsumoto:1999:ADP

[881] S. Tsumoto. Automated discovery of plausible rules based on rough sets and rough inclusion. Lecture Notes in Computer Science, 1574:210–219, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1999:DSR

[882] Z. Liu and Z. Xie. Discernibility system in rough sets. Lecture Notes in Computer Science, 1574:220–227, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rauber:1999:ALS

[883] A. Rauber and D. Merkl. Automatic labeling of self-organizing maps: Making a treasure-map reveal its secrets. Lecture Notes in Computer Science, 1574:228–237, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1999:NNB

[884] L. Zhang and B. Zhang. Neural network based classifiers for a vast amount of data. Lecture Notes in Computer Science, 1574:238–246, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prado:1999:ATC

[885] H. A. Prado, K. F. Machado, S. R. Frigeri, and P. M. Engel. Accuracy tuning on combinatorial neural model. Lecture Notes in Computer Science, 1574:247–251, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kakemoto:1999:SIA

[886] Y. Kakemoto and S. Nakasuka. A situated information articulation neural network: VSF network. Lecture Notes in Computer Science, 1574:252–257, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deng:1999:NMD

[887] C. Deng and F. Xiong. Neural method for detection of complex patterns in databases. Lecture Notes in Computer Science, 1574:258–262, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shi:1999:PDL

[888] X. Shi, M.-C. Chan, and D. Li. Preserve discovered linguistic patterns valid in volatility data environment. Lecture Notes in Computer Science, 1574:263–267, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shibata:1999:IAB

[889] D. Shibata, N. Inuzuka, S. Kato, and T. Matsui. An induction algorithm based on fuzzy logic programming. Lecture Notes in Computer Science, 1574:268–273, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsumoto:1999:RDD

[890] S. Tsumoto. Rule discovery in databases with missing values based on rough set model. Lecture Notes in Computer Science, 1574:274–278, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1999:SKM

[891] X. Wang, R. Wang, and J. Wang. Sustainability knowledge mining from human development database. Lecture Notes in Computer Science, 1574:279– 283, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wada:1999:CDK

[892] T. Wada, T. Horiuchi, H. Motoda, and T. Washio. Characterization of default knowledge in ripple down rules method. Lecture Notes in Computer Science, 1574:284–295, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ting:1999:IPB

[893] K. M. Ting and Z. Zheng. Improving the performance of boosting for naive Bayesian classification. Lecture Notes in Computer Science, 1574:296–305, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Newlands:1999:CHC

[894] D. A. Newlands and G. I. Webb. Convex hulls in concept induction. Lecture Notes in Computer Science, 1574:306– 316, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fan:1999:MCK

[895] J. Fan and D. Li. Mining classification knowledge based on cloud models. Lecture Notes in Computer Science, 1574:317–326, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Estivill-Castro:1999:RCL

[896] V. Estivill-Castro and M. E. Houle. Robust clustering of large geo-referenced data sets. Lecture Notes in Computer Science, 1574:327–337, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhou:1999:FAD

[897] B. Zhou, D. W. Cheung, and B. Kao. A fast algorithm for density-based clustering in large database. Lecture Notes in Computer Science, 1574:338–349, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Melli:1999:LMB

[898] G. Melli. A lazy model-based algorithm for on-line classification. Lecture Notes in Computer Science, 1574:350–354, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alsabti:1999:ESP

[899] K. Alsabti, S. Ranka, and V. Singh. An efficient space-partitioning based algorithm for the K-means clustering. Lecture Notes in Computer Science, 1574:355–359, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Su:1999:FCP

[900] C.-M. Su, S.-S. Tseng, M.-F. Jiang, and J. C. S. Chen. A fast clustering process for outliers and remainder clusters. Lecture Notes in Computer Science, 1574:360–364, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

He:1999:ODM

[901] H. He and S. Hawkins. Optimising the distance metric in the nearest neighbour algorithm on a real-world patient classification problem. Lecture Notes in Computer Science, 1574:365–369, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zheng:1999:CUC

[902] Z. Zheng and B. T. Low. Classifying unseen cases with many missing values. Lecture Notes in Computer Science, 1574:370–374, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ho:1999:SMS

[903] T. B. Ho, N. B. Nguyen, and T. Morita. Study of a mixed similarity measure for classification and clustering. Lecture Notes in Computer Science, 1574:375– 379, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1999:VAE

[904] B. Liu, W. Hsu, K. Wang, and S. Chen. Visually aided exploration of interesting association rules. Lecture Notes in Computer Science, 1574:380–389, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Han:1999:DSV

[905] J. Han and N. Cercone. DVIZ: A system for visualizing data mining. Lecture Notes in Computer Science, 1574:390–399, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dai:1999:MCM

[906] H. Dai. A minimal causal model learner. Lecture Notes in Computer Science, 1574:400–408, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1999:EGB

[907] K. L. Lee, G. Lee, and A. L. P. Chen. Efficient graph-based algorithm for discovering and maintaining knowledge in large databases. Lecture Notes in Computer Science, 1574:409–419, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Inokuchi:1999:BAG

[908] A. Inokuchi, T. Washio, H. Motoda, and K. Kumasawa. Basket analysis for graph structured data. Lecture Notes in Computer Science, 1574:420–431, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neil:1999:ECM

[909] J. R. Neil and K. B. Korb. The evolution of causal models: A comparison of Bayesian metrics and structure priors. Lecture Notes in Computer Science, 1574:432–437, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miyahara:1999:KFK

[910] T. Miyahara, T. Uchida, T. Kuboyama, and T. Yamamoto. KD-FGS: A knowledge discovery system from graph data using formal graph system. Lecture Notes in Computer Science, 1574:438– 442, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guo:1999:PKD

[911] Y. Guo and J. Sutiwaraphun. Probing knowledge in distributed data mining. Lecture Notes in Computer Science, 1574:443–452, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ras:1999:DES

[912] Z. W. Ras and J. M. Zytkow. Discovery of equations and the shared operational semantics in distributed autonomous databases. Lecture Notes in Computer Science, 1574:453–463, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zigbed:1999:DMT

[913] A. Zigbed, M. Cote, and N. Troudi. The data-mining and the technology of agents to fight the illicit electronic messages. Lecture Notes in Computer Science, 1574:464–468, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zighed:1999:DMT

[914] A. Zighed, M. Cote, and N. Troudi. The data-mining and the technology of agents to fight the illicit electronic messages. Lecture Notes in Computer Science, 1574:464–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lu:1999:KDS

[915] H. Lu, L. Sterling, and A. Wyatt. Knowledge discovery in SportsFinder: An agent to extract sports results from the Web. Lecture Notes in Computer Science, 1574:469–473, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perrochon:1999:EME

[916] L. Perrochon, W. Mann, S. Kasriel, and D. C. Luckham. Event mining with event processing networks. Lecture Notes in Computer Science, 1574:474– 478, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yao:1999:AQM

[917] Y. Y. Yao and N. Zhong. An analysis of quantitative measures associated with rules. Lecture Notes in Computer Science, 1574:479–488, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheng:1999:SRL

[918] J. Cheng. A strong relevant logic model of epistemic processes in scientific discovery. Lecture Notes in Computer Science, 1574:489–493, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yoshida:1999:DCD

[919] T. Yoshida, T. Kondo, and S. Nishida. Discovering conceptual differences among different people via diverse structures. Lecture Notes in Computer Science, 1574:494–498, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lobo:1999:OEM

[920] O. O. Lobo and M. Numao. Ordered estimation of missing values. Lecture Notes in Computer Science, 1574:499– 503, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Suzuki:1999:PRD

[921] E. Suzuki and T. Ohno. Prediction rule discovery based on dynamic bias selection. Lecture Notes in Computer Science, 1574:504–508, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

An:1999:DCA

[922] A. An and N. Cercone. Discretization of continuous attributes for learning classification rules. Lecture Notes in Computer Science, 1574:509–514, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

BenYahia:1999:BBR

[923] S. Ben Yahia and A. Jaoua. BRRA: A based relevant rectangle algorithm for mining relationships in databases. Lecture Notes in Computer Science, 1574: 515–519, 1999. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic).

Yahia:1999:BBR

[924] S. B. Yahia and A. Jaoua. BRRA: A based relevant rectangle algorithm for mining relationships in databases. Lecture Notes in Computer Science, 1574: 515–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tao:1999:MFD

[925] X. Tao, N. Wang, S. Zhou, and A. Zhou. Mining functional dependency rule of relational database. Lecture Notes in Computer Science, 1574: 520–524, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jiang:1999:TSP

[926] R. Jiang, D. Li, and H. Chen. Timeseries prediction with cloud models in DMKD. Lecture Notes in Computer Science, 1574:525–530, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saraiva:1999:DSF

[927] J. Saraiva and D. Swierstra. Data structure free compilation. Lecture Notes in Computer Science, 1575:1–16, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sloane:1999:DEG

[928] A. M. Sloane. Debugging Eli-generated compilers with Noosa. Lecture Notes in Computer Science, 1575:17–31, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aycock:1999:FGL

[929] J. Aycock and N. Horspool. Faster generalized LR parsing. Lecture Notes in Computer Science, 1575:32–46, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Melski:1999:IPP

[930] D. Melski and T. Reps. Interprocedural path profiling. Lecture Notes in Computer Science, 1575:47–62, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martin:1999:ECC

[931] F. Martin. Experimental comparison of call string and functional approaches to interprocedural analysis. Lecture Notes in Computer Science, 1575:63– 75, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Debray:1999:LTI

[932] S. Debray, R. Muth, and S. Watterson. Link-time improvement of scheme programs. Lecture Notes in Computer Science, 1575:76–90, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knoop:1999:EBR

[933] J. Knoop, O. Ruething, and B. Steffen. Expansion-based removal of semantic partial redundancies. Lecture Notes in Computer Science, 1575:91–106, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1999:RPS

[934] R. Gupta and R. Bodik. Register pressure sensitive redundancy elimination. Lecture Notes in Computer Science, 1575:107–121, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaestner:1999:COI

[935] D. Kaestner and M. Langenbach. Code optimization by integer linear programming. Lecture Notes in Computer Science, 1575:122–136, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liberatore:1999:EAL

[936] V. Liberatore, M. Farach-Colton, and U. Kremer. Evaluation of algorithms for local register allocation. Lecture Notes in Computer Science, 1575:137– 152, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1999:ESD

[937] C. Zhang, R. Govindarajan, S. Ryan, and G. R. Gao. Efficient statediagram construction methods for software pipelining. Lecture Notes in Computer Science, 1575:153–167, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rivera:1999:CCT

[938] G. Rivera and C.-W. Tseng. A comparison of compiler tiling algorithms. Lecture Notes in Computer Science, 1575:168–182, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rauchwerger:1999:IIL

[939] L. Rauchwerger and D. Patel. Implementation issues of loop-level speculative run-time parallelization. Lecture Notes in Computer Science, 1575:183– 197, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VandenBrand:1999:CMM

[940] M. Van den Brand, P. Klint, and P. Olivier. Compilation and memory management for ASF+SDF. Lecture Notes in Computer Science, 1575:198– 213, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saito:1999:DPC

[941] H. Saito, N. Stavrakos, S. Carroll, and C. Polychronopoulos. The design of the PROMIS compiler. Lecture Notes in Computer Science, 1575:214–228, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cilio:1999:FPF

[942] A. Cilio and H. Corporaal. Floating point to fixed point conversion of C code. Lecture Notes in Computer Science, 1575:229–243, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tourwe:1999:OOO

[943] T. Tourwe and W. De Meuter. Optimizing object-oriented languages through architectural transformations. Lecture Notes in Computer Science, 1575:244–258, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rubin:1999:VCL

[944] S. Rubin, D. Bernstein, and M. Rodeh. Virtual cache line: A new technique to improve cache exploitation for recursive data structures. Lecture Notes in Computer Science, 1575:259–273, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeDinechin:1999:EMS

[945] B. D. De Dinechin. Extending modulo scheduling with memory reference merging. Lecture Notes in Computer Science, 1575:274–287, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ernst:1999:TMP

[946] T. Ernst. TRAPping Modelica with Python. Lecture Notes in Computer Science, 1575:288–291, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boekhold:1999:PAC

[947] M. Boekhold, I. Karkowski, H. Corporaal, and A. Cilio. A programmable ANSI C code transformation engine. Lecture Notes in Computer Science, 1575:292–295, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anlauff:1999:TSL

[948] M. Anlauff, P. W. Kutter, and A. Pierantonio. Tool support for language design and prototyping with montages. Lecture Notes in Computer Science, 1575:296–300, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hudak:1999:FRP

[949] P. Hudak. Functional reactive programming. Lecture Notes in Computer Science, 1576:1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benedikt:1999:DLD

[950] M. Benedikt, T. Reps, and M. Sagiv. A decidable logic for describing linked data structures. Lecture Notes in Computer Science, 1576:2–19, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nielson:1999:ICF

[951] F. Nielson and H. R. Nielson. Interprocedural control flow analysis. Lecture Notes in Computer Science, 1576: 20–39, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sabelfeld:1999:PMS

[952] A. Sabelfeld and D. Sands. A per model of secure information flow in sequential programs. Lecture Notes in Computer Science, 1576:40–58, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

King:1999:QSD

[953] A. King, J.-G. Smaus, and P. Hill. Quotienting share for dependency analysis. Lecture Notes in Computer Science, 1576:59–73, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gay:1999:TSC

[954] S. Gay and M. Hole. Types and subtypes for client-server interactions. Lecture Notes in Computer Science, 1576: 74–90, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Flanagan:1999:TSL

[955] C. Flanagan and M. Abadi. Types for safe locking. Lecture Notes in Computer Science, 1576:91–108, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barthe:1999:CS

[956] G. Barthe and M. J. Frade. Constructor subtyping. Lecture Notes in Computer Science, 1576:109–127, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Trifonov:1999:SPL

[957] V. Trifonov and Z. Shao. Safe and principled language interoperation. Lecture Notes in Computer Science, 1576:128– 146, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Norrish:1999:DEC

[958] M. Norrish. Deterministic expressions in C. Lecture Notes in Computer Science, 1576:147–161, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poetzsch-Heffter:1999:PLS

[959] A. Poetzsch-Heffter and P. Mueller. A programming logic for sequential Java. Lecture Notes in Computer Science, 1576:162–176, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Podelski:1999:SBF

[960] A. Podelski, W. Charatonik, and M. Mueller. Set-based failure analysis for logic programs and concurrent constraint programs. Lecture Notes in Computer Science, 1576:177–192, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moggi:1999:IMS

[961] E. Moggi, W. Taha, Z. El-Abidine Benaissa, and T. Sheard. An idealized MetaML: Simpler, and more expressive. Lecture Notes in Computer Science, 1576:193–207, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mycroft:1999:TBD

[962] A. Mycroft. Type-based decompilation. Lecture Notes in Computer Science, 1576:208–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Danvy:1999:OIC

[963] O. Danvy and Z. Yang. An operational investigation of the CPS hierarchy. Lecture Notes in Computer Science, 1576:224–242, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thiemann:1999:HOC

[964] P. Thiemann. Higher-order code splicing. Lecture Notes in Computer Science, 1576:243–257, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krishnamurthi:1999:ESP

[965] S. Krishnamurthi, Y.-D. Erlich, and M. Felleisen. Expressing structural properties as language constructs. Lecture Notes in Computer Science, 1576: 258–272, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jansson:1999:PCP

[966] P. Jansson and J. Jeuring. Polytypic compact printing and parsing. Lecture Notes in Computer Science, 1576:273– 287, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1999:DPS

[967] Y. A. Liu and S. D. Stoller. Dynamic programming via static incrementalization. Lecture Notes in Computer Science, 1576:288–305, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanDeursen:1999:RIR

[968] A. Van Deursen, P. Klint, and C. Verhoef. Research issues in the renovation of legacy systems. Lecture Notes in Computer Science, 1577:1–21, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weber:1999:CEI

[969] H. Weber. Continuous engineering of information and communication infrastructures. Lecture Notes in Computer Science, 1577:22–29, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ancona:1999:FFL

[970] D. Ancona, M. Cerioli, and E. Zucca. A formal framework with late binding. Lecture Notes in Computer Science, 1577:30–44, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Attanasio:1999:YAR

[971] C. Attanasio, F. Corradini, and P. Inverardi. Yet another real-time specification for the steam boiler: Local clocks to statically measure systems performance. Lecture Notes in Computer Science, 1577:45–59, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruni:1999:ETS

[972] R. Bruni, J. Meseguer, and U. Montanari. Executable tile specifications for process calculi. Lecture Notes in Computer Science, 1577:60–76, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coscia:1999:JJT

[973] E. Coscia and G. Reggio. JTN: A Java-targeted graphic formal notation for reactive and concurrent systems. Lecture Notes in Computer Science, 1577:77–97, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fokkink:1999:CEP

[974] W. Fokkink and C. Verhoef. Conservative extension in positive/negative conditional term rewriting with applications to software renovation factories. Lecture Notes in Computer Science, 1577:98–113, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gaudel:1999:DSI

[975] M.-C. Gaudel, C. Khoury, and A. Zamulin. Dynamic systems with implicit state. Lecture Notes in Computer Science, 1577:114–128, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lano:1999:RDU

[976] K. Lano and A. Evans. Rigorous development in UML. Lecture Notes in Computer Science, 1577:129–143, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lopes:1999:UES

[977] A. Lopes and J. L. Fiadeiro. Using explicit state to describe architectures. Lecture Notes in Computer Science, 1577:144–160, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mueller-Olm:1999:ERC

[978] M. Mueller-Olm, B. Steffen, and R. Cleaveland. On the evolution of reactive components — A processalgebraic approach. Lecture Notes in Computer Science, 1577:161–175, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nepomniaschy:1999:VDI

[979] V. A. Nepomniaschy. Verification of definite iteration over hierarchical data structures. Lecture Notes in Computer Science, 1577:176–187, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nipkow:1999:OGI

[980] T. Nipkow and L. Prensa Nieto. Owicki/Gries in Isabelle/HOL. Lecture Notes in Computer Science, 1577:188– 203, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nottegar:1999:SDP

[981] C. Nottegar, C. Priami, and P. Degano. Semantic-driven performance evaluation. Lecture Notes in Computer Science, 1577:204–218, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tapken:1999:IHG

[982] J. Tapken. Implementing hierarchical graph-structures. Lecture Notes in Computer Science, 1577:219–233, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blair:1999:TSM

[983] L. Blair, T. Jones, and G. Blair. A tool suite for multi-paradigm specification. Lecture Notes in Computer Science, 1577:234–238, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lueth:1999:TIT

[984] C. Lueth, H. Tej Kolyang, and B. Krieg-Brueckner. TAS and IsaWin: Tools for transformational program development and theorem proving. Lecture Notes in Computer Science, 1577: 239–244, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abadi:1999:SPS

[985] M. Abadi. Security protocols and specifications. Lecture Notes in Computer Science, 1578:1–13, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Esparza:1999:ATA

[986] J. Esparza and J. Knoop. An automata-theoretic approach to interprocedural data-flow analysis. Lecture Notes in Computer Science, 1578:14– 30, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sangiorgi:1999:RAC

[987] D. Sangiorgi. Reasoning about concurrent systems using types. Lecture Notes in Computer Science, 1578:31– 40, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aceto:1999:THM

[988] L. Aceto and A. Ingolfsdottir. Testing Hennessy–Milner logic with recursion. Lecture Notes in Computer Science, 1578:41–55, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arruabarrena:1999:SLP

[989] R. Arruabarrena, P. Lucio, and M. Navarro. A strong logic programming view for static embedded implications. Lecture Notes in Computer Science, 1578:56–72, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baldan:1999:UES

[990] P. Baldan, A. Corradini, and U. Montanari. Unfolding and event structure semantics for graph grammars. Lecture Notes in Computer Science, 1578:73– 89, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barthe:1999:EC

[991] G. Barthe. Expanding the cube. Lecture Notes in Computer Science, 1578: 90–103, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benke:1999:ACT

[992] M. Benke. An algebraic characterization of typability in ML with subtyping. Lecture Notes in Computer Science, 1578:104–119, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bodei:1999:SAP

[993] C. Bodei, P. Degano, F. Nielson, and H. R. Nielson. Static analysis of processes for no read-up and no writedown. Lecture Notes in Computer Science, 1578:120–134, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeBoer:1999:WCO

[994] F. S. De Boer. A WP-calculus for OO. Lecture Notes in Computer Science, 1578:135–149, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bogaert:1999:RPT

[995] B. Bogaert, F. Seynhaeve, and S. Tison. The recognizability problem for tree automata with comparisons between brothers. Lecture Notes in Computer Science, 1578:150–164, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boreale:1999:TMT

[996] M. Boreale, R. De Nicola, and R. Pugliese. A theory of “may” testing for asynchronous languages. Lecture Notes in Computer Science, 1578:165– 179, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dantsin:1999:NPT

[997] E. Dantsin and A. Voronkov. A nondeterministic polynomial-time unification algorithm for bags, sets and trees. Lecture Notes in Computer Science, 1578:180–196, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghani:1999:CME

[998] N. Ghani, V. De Paiva, and E. Ritter. Categorical models of explicit substitutions. Lecture Notes in Computer Science, 1578:197–211, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gordon:1999:EPM

[999] A. D. Gordon and L. Cardelli. Equational properties of mobile ambients. Lecture Notes in Computer Science, 1578:212–226, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huhn:1999:MCL

[1000] M. Huhn, P. Niebert, and F. Wallner. Model checking logics for communicating sequential agents. Lecture Notes in Computer Science, 1578:227–242, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lenisa:1999:CCL

[1001] M. Lenisa. A complete coinductive logical system for bisimulation equivalence on circular objects. Lecture Notes in Computer Science, 1578:243–257, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maneth:1999:SLG

[1002] S. Maneth. String languages generated by total deterministic macro tree transducers. Lecture Notes in Computer Science, 1578:258–272, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muscholl:1999:MSM

[1003] A. Muscholl. Matching specifications for message sequence charts. Lecture Notes in Computer Science, 1578:273– 287, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Narasimha:1999:PTL

[1004] M. Narasimha, R. Cleaveland, and P. Iyer. Probabilistic temporal logics via the modal mu-calculus. Lecture Notes in Computer Science, 1578:288– 305, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roeckl:1999:CPS

[1005] C. Roeckl and D. Sangiorgi. A λcalculus process semantics of concurrent idealised ALGOL. Lecture Notes in Computer Science, 1578:306–321, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kramer:1999:MMM

[1006] Jeff Kramer and Jeff Magee. Modelling for mere mortals. Lecture Notes in Computer Science, 1579:1– 18, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1579/15790001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1579/15790001.
pdf.

Friese:1999:IPO

[1007] Thomas Friese, Tiziana Margaria, and Thomas Rakow. Integrating printed and online information. Lecture Notes in Computer Science, 1579: ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1579/1579doi.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/1579/1579doi.
pdf.

Hsiung:1999:SSV

[1008] Pao-Ann Hsiung, Farn Wang, and Yue-Sun Kuo. Scheduling system verification. Lecture Notes in Computer Science, 1579:19–33, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1579/15790019.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1579/15790019.
pdf.

Ryu:1999:PAA

[1009] Minsoo Ryu and Seongsoo Hong. A period assignment algorithm for realtime system design. Lecture Notes in Computer Science, 1579:34–43, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1579/15790034.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1579/15790034.
pdf.

Gardner:1999:ASF

[1010] Mark K. Gardner and Jane W. S.

Liu. Analyzing stochastic fixedpriority real-time systems. Lecture Notes in Computer Science, 1579:44– 58, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1579/15790044.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1579/15790044.
pdf.

Tripakis:1999:TDR

[1011] Stavros Tripakis. Timed diagnostics for reachability properties. Lecture Notes in Computer Science, 1579:59– 73, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1579/15790059.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1579/15790059.
pdf.

Dong:1999:FLP

[1012] Yifei Dong, Xiaoqun Du, Y. S. Ramakrishna, C. R. Ramakrishnan, I. V. Ramakrishnan, Scott A. Smolka, Oleg Sokolsky, Eugene W. Stark, and David S. Warren. Fighting livelock in the i-protocol: A comparative study of verification tools. Lecture Notes in Computer Science, 1579:74– 88, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1579/15790074.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1579/15790074.
pdf.

Pusch:1999:PSJ

[1013] Cornelia Pusch. Proving the soundness of a Java bytecode verifier specification in Isabelle/HOL. Lecture Notes in Computer Science, 1579:89– 103, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1579/15790089.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1579/15790089.
pdf.

Kreitz:1999:AFT

[1014] Christoph Kreitz. Automated fasttrack reconfiguration of group communication systems. Lecture Notes in Computer Science, 1579:104–118, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1579/15790104.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/1579/15790104.
pdf.

Hickey:1999:SPE

[1015] Jason Hickey, Nancy Lynch, and Robbert van Renesse. Specifications and proofs for ensemble layers. Lecture Notes in Computer Science, 1579:119– 133, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1579/15790119.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1579/15790119.
pdf.

Bergeron:1999:AAP

[1016] Anne Bergeron and Jean-Christophe Manzoni. An automated analysis of ping-pong interactions in Email services. Lecture Notes in Computer Science, 1579:134–147, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1579/15790134.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1579/15790134.
pdf.

Marchignoli:1999:AVC

[1017] Davide Marchignoli and Fabio Martinelli. Automatic verification of cryptographic protocols through compositional analysis techniques. Lecture Notes in Computer Science, 1579:148– 162, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1579/15790148.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1579/15790148.
                                        pdf.

Behrmann:1999:VHS

[1018] Gerd Behrmann, Kim G. Larsen, Henrik R. Andersen, Henrik Hulgaard, and Jørn Lind-Nielsen. Verification of hierarchical state/event systems using reusability and compositionality. Lecture Notes in Computer Science, 1579:163–177, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1579/15790163.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1579/15790163.
                                        pdf.

Rusu:1999:PSP

[1019] Vlad Rusu and Eli Singerman. On proving safety properties by integrating static analysis, theorem proving and abstraction. Lecture Notes in Computer Science, 1579:178–192, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1579/15790178.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1579/15790178.
                                        pdf.

Biere:1999:SMC

[1020] Armin Biere, Alessandro Cimatti, Edmund Clarke, and Yunshan Zhu. Symbolic model checking without BDDs. Lecture Notes in Computer Science, 1579:193–207, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1579/15790193.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1579/15790193.pdf.

Abdulla:1999:SVL

[1021] Parosh Abdulla, Aurore Annichini, and Ahmed Bouajjani. Symbolic verification of lossy channel systems: Application to the bounded retransmission protocol. Lecture Notes in Computer Science, 1579:208–222, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1579/15790208.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1579/15790208.
pdf.

Delzanno:1999:MCC

[1022] Giorgio Delzanno and Andreas Podelski. Model checking in CLP. Lecture Notes in Computer Science, 1579:223– 239, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1579/15790223.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1579/15790223.
pdf.

Heljanko:1999:ULP

[1023] Keijo Heljanko. Using logic programs with stable model semantics to solve deadlock and reachability problems for 1-safe Petri nets. Lecture Notes in Computer Science, 1579:240– 254, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1579/15790240.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1579/15790240.
                                        pdf.

Montanari:1999:FSV

[1024] Ugo Montanari and Marco Pistore. Finite state verification for the asynchronous pi-calculus. Lecture Notes in Computer Science, 1579:255–269, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1579/15790255.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1579/15790255.
                                        pdf.

Basten:1999:PAP

[1025] Twan Basten and Jozef Hooman. Process algebra in PVS. Lecture Notes in Computer Science, 1579:270–284, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1579/15790270.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1579/15790270.
                                        pdf.

Hirschkoff:1999:BUT

[1026] Daniel Hirschkoff. On the benefits of using the up to techniques for bisimulation verification. Lecture Notes in Computer Science, 1579:285–299, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1579/15790285.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1579/15790285.
pdf.

Li:1999:CSW

[1027] Zhoujun Li and Huowang Chen. Computing strong/weak bisimulation equivalences and observation congruence for value-passing processes. Lecture Notes in Computer Science, 1579:300– 314, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1579/15790300.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1579/15790300.
pdf.

Kahlouche:1999:HTU

[1028] Hakim Kahlouche, César Viho, and Massimo Zendri. Hardware testing using a communication protocol conformance testing tool. Lecture Notes in Computer Science, 1579:315–329, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1579/15790315.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1579/15790315.
pdf.

Kern:1999:LWF

[1029] Christoph Kern, Tarik Ono-Tesfaye, and Mark R. Greenstreet. A light-weight framework for hardware verification. Lecture Notes in Computer Science, 1579:330–344, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1579/15790330.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1579/15790330.
                                        pdf.

Parnas:1999:EET

[1030] David Lorge Parnas and Dennis K. Peters. An easily extensible toolset for tabular mathematical expressions. Lecture Notes in Computer Science, 1579: 345–359, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1579/15790345.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1579/15790345.
                                        pdf.

Knoop:1999:DFD

[1031] Jens Knoop. From DFA-frameworks to DFA-Generators: A unifying multiparadigm approach. Lecture Notes in Computer Science, 1579:360–374, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1579/15790360.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1579/15790360.
                                        pdf.

Spelt:1999:TPB

[1032] David Spelt and Susan Even. A theorem prover-based analysis tool for object-oriented databases. Lecture Notes in Computer Science, 1579:375– 389, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1579/15790375.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1579/15790375.
pdf.

Bakhmurov:1999:DEE

[1033] A. G. Bakhmurov, A. P. Kapitonova, and R. L. Smeliansky. DYANA: An environment for embedded system design and analysis. Lecture Notes in Computer Science, 1579:390– 404, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1579/15790390.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1579/15790390.
pdf.

Gunter:1999:PET

[1034] Elsa L. Gunter and Doron Peled.

Path exploration tool. Lecture Notes in Computer Science, 1579:405–419, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1579/15790405.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1579/15790405.
pdf.

Buchholz:1999:MSL

[1035] Peter Buchholz and Peter Kemper. Modular state level analysis of distributed systems techniques and tool support. Lecture Notes in Computer Science, 1579:420–434, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1579/15790420.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1579/15790420.
                                        pdf.

Stevens:1999:SIS

[1036] Perdita Stevens. Some issues in the software engineering of verification tools. Lecture Notes in Computer Science, 1579:435–438, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1579/15790435.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1579/15790435.
                                        pdf.

Braun:1999:EOS

[1037] Volker Braun, Jürgen Kreileder, Tiziana Margaria, and Bernhard Steffen. The ETI online service in action. Lecture Notes in Computer Science, 1579:439–443, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1579/15790439.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1579/15790439.
                                        pdf.

Wiederhold:1999:MDH

[1038] G. Wiederhold. Mediation to deal with heterogeneous data sources. Lecture Notes in Computer Science, 1580:1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cuthbert:1999:OTS

[1039] A. Cuthbert. OpenGIS: Tales from a small market town. Lecture Notes in Computer Science, 1580:17–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Landgraf:1999:EEG

[1040] G. Landgraf. Evolution of EO/GIS interoperability towards an integrated application infrastructure. Lecture Notes in Computer Science, 1580:29– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sargent:1999:FID

[1041] P. Sargent. Feature identities, descriptors and handles. Lecture Notes in Computer Science, 1580:41–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bishr:1999:GUP

[1042] Y. A. Bishr. A global unique persistent object ID for geospatial information sharing. Lecture Notes in Computer Science, 1580:55–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramroop:1999:IAN

[1043] S. Ramroop and R. Pascoe. Implementation architecture for a national data center. Lecture Notes in Computer Science, 1580:65–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moeller:1999:ACA

[1044] H. L. Moeller, M. Mariucci, and B. Mitschang. Architecture considerations for advanced Earth observation application systems. Lecture Notes in Computer Science, 1580:75–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bitter:1999:SDI

[1045] P. Bitter. A spatial data infrastructure for the hindu kush — himalayan region. Lecture Notes in Computer Science, 1580:91–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kemp:1999:IGE

[1046] K. K. Kemp, D. E. Reeve, and D. I. Heywood. Interoperability for GI-Science education. Lecture Notes in Computer Science, 1580:101–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cranston:1999:AIS

[1047] C. B. Cranston, F. Brabec, G. R. Hjaltason, and D. Nebert. Adding an interoperable server interface to a spatial database: Implementation experiences with OpenMap[TM]. Lecture Notes in Computer Science, 1580:115–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Camara:1999:IPP

[1048] G. Camara, R. Thome, U. Freitas, and A. M. V. Monteiro. Interoperability in practice: Problems in semantic conversion from current technology to OpenGIS. Lecture Notes in Computer Science, 1580:129–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Doyle:1999:BPO

[1049] A. Doyle, D. Dietrick, J. Ebbinghaus, and P. Ladstaetter. Building a prototype OpenGIS[R] demonstration from interoperable GIS components. Lecture Notes in Computer Science, 1580:139–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keller:1999:MSG

[1050] S. F. Keller and H. Thalmann. Modeling and sharing graphic presentations of geospatial data. Lecture Notes in Computer Science, 1580:151–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cattenstart:1999:TOS

[1051] F. C. Cattenstart and H. J. Scholten.

Towards OpenGIS systems. Lecture Notes in Computer Science, 1580:163– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Uitermark:1999:SGA

[1052] H. Uitermark, A. Vogels, and P. Van Oosterom. Semantic and geometric aspects of integrating road networks. Lecture Notes in Computer Science, 1580:177–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rodriguez:1999:ASS

[1053] M. A. Rodriguez, M. J. Egenhofer, and R. D. Rugg. Assessing semantic similarities among geospatial feature class definitions. Lecture Notes in Computer Science, 1580:189–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bishr:1999:PRS

[1054] Y. A. Bishr, H. Pundt, and C. Ruether. Proceeding on the road of semantic interoperability — design of a semantic mapper based on a case study from transportation. Lecture Notes in Computer Science, 1580:203–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Riedemann:1999:WSG

[1055] C. Riedemann and W. Kuhn. What are sports grounds? or: Why semantics requires interoperability. Lecture Notes in Computer Science, 1580:217– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stock:1999:ISS

[1056] K. Stock and D. Pullar. Identifying semantically similar elements in heterogeneous spatial databases using predicate logic expressions. Lecture Notes in Computer Science, 1580:231–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kemppainen:1999:DMM

[1057] H. Kemppainen. Designing a mediator for managing relationships between distributed objects. Lecture Notes in Computer Science, 1580:253–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shimada:1999:GMF

[1058] S. Shimada and H. Fukui. Geospatial mediator functions and containerbased fast transfer interface in SI3CO test-bed. Lecture Notes in Computer Science, 1580:265–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Relly:1999:PPI

[1059] L. Relly and U. Roehm. Plug and play:

Interoperability in CONCERT. Lecture Notes in Computer Science, 1580:277– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Winter:1999:FER

[1060] S. Winter and A. U. Frank. Functional extensions of a raster representation for topological relations. Lecture Notes in Computer Science, 1580:293–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ladner:1999:SER

[1061] R. Ladner, K. Shaw, and M. Abdelguerfi. 3D synthetic environment representation using the “non-manifold 3D winged-edge” data structure. Lecture Notes in Computer Science, 1580: 305–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martinoni:1999:PTM

[1062] D. Martinoni and B. Schneider. Pluggable terrain module — moving digital terrain modelling to a distributed geoprocessing environment. Lecture Notes in Computer Science, 1580:315– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andreoli:1999:CLF

[1063] J.-M. Andreoli. The coordination language facility and applications. Lecture Notes in Computer Science, 1581:1–5, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eidorff:1999:APT

[1064] P. H. Eidorff, F. Henglein, C. Mossin, and H. Niss. AnnoDomini in practice: A type-theoretic approach to the year 2000 problem. Lecture Notes in Computer Science, 1581:6–13, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abrusci:1999:MNC

[1065] V. M. Abrusci. Modules in noncommutative logic. Lecture Notes in Computer Science, 1581:14–24, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baillot:1999:ECG

[1066] P. Baillot and M. Pedicini. Elementary complexity and geometry of interaction. Lecture Notes in Computer Science, 1581:25–39, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barreiro:1999:QSR

[1067] N. Barreiro and T. Ehrhard. Quantitative semantics revisited. Lecture Notes in Computer Science, 1581:40– 53, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berardi:1999:TFW

[1068] S. Berardi and U. De’Liguoro. Total functionals and well-founded strategies. Lecture Notes in Computer Science, 1581:54–68, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Broda:1999:CTP

[1069] S. Broda and L. Damas. Counting a type’s principal inhabitants. Lecture Notes in Computer Science, 1581:69– 82, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damiani:1999:UCD

[1070] F. Damiani. Useless-code detection and elimination for PCF with algebraic data types. Lecture Notes in Computer Science, 1581:83–97, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

David:1999:EUL

[1071] R. David. Every unsolvable lambda term has a decoration. Lecture Notes in Computer Science, 1581:98–113, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiGianantonio:1999:GSU

[1072] P. Di Gianantonio, G. Franco, and F. Honsell. Game semantics for untyped lambdaeta-calculus. Lecture Notes in Computer Science, 1581:114– 128, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dybjer:1999:FAI

[1073] P. Dybjer and A. Setzer. A finite axiomatization of inductive-recursive definitions. Lecture Notes in Computer Science, 1581:129–146, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fiore:1999:LDS

[1074] M. Fiore and A. Simpson. Lambda definability with sums via grothendieck logical relations. Lecture Notes in Computer Science, 1581:147–161, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fujita:1999:ETL

[1075] K. Fujita. Explicitly typed lambdacalculus for polymorphism and call-by-value. Lecture Notes in Computer Science, 1581:162–176, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goguen:1999:SLF

[1076] H. Goguen. Soundness of the logical framework for its typed operational semantics. Lecture Notes in Computer Science, 1581:177–197, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hasegawa:1999:LPI

[1077] M. Hasegawa. Logical predicates for intuitionistic linear type theories. Lecture Notes in Computer Science, 1581:198– 212, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laurent:1999:PPN

[1078] O. Laurent. Polarized proof-nets: Proof-nets for LC. Lecture Notes in Computer Science, 1581:213–227, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Levy:1999:CPV

[1079] P. B. Levy. Call-by-push-value: A subsuming paradigm. Lecture Notes in Computer Science, 1581:228–242, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mikami:1999:SAL

[1080] S. Mikami and Y. Akama. A study of Abramsky’s linear chemical abstract machine. Lecture Notes in Computer Science, 1581:243–257, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

OHearn:1999:RIB

[1081] P. W. O’Hearn. Resource interpretations, bunched implications and the lambda-calculus. Lecture Notes in Computer Science, 1581:258–279, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohori:1999:CHI

[1082] A. Ohori. A Curry–Howard isomorphism for compilation and program execution. Lecture Notes in Computer Science, 1581:280–294, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Polakow:1999:NDI

[1083] J. Polakow and F. Pfenning. Natural deduction for intuitionistic noncommutative linear logic. Lecture Notes in Computer Science, 1581:295–309, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poll:1999:LAD

[1084] E. Poll and J. Zwanenburg. A logic for abstract data types as existential types. Lecture Notes in Computer Science, 1581:310–324, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ritter:1999:CES

[1085] E. Ritter. Characterising explicit substitutions which preserve termination. Lecture Notes in Computer Science, 1581:325–339, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sato:1999:EE

[1086] M. Sato, T. Sakurai, and R. Burstall. Explicit environments. Lecture Notes in Computer Science, 1581:340–354, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Statman:1999:CJT

[1087] R. Statman. Consequences of Jacopini’s theorem: Consistent equalities and equations. Lecture Notes in Computer Science, 1581:355–364, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Urban:1999:SNC

[1088] C. Urban and G. M. Bierman. Strong normalisation of cut-elimination in classical logic. Lecture Notes in Computer Science, 1581:365–380, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zwanenburg:1999:PTS

[1089] J. Zwanenburg. Pure type systems with subtyping. Lecture Notes in Computer Science, 1581:381–396, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lambek:1999:TGR

[1090] J. Lambek. Type grammar revisited. Lecture Notes in Computer Science, 1582:1–??, 1999. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1582/15820001.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1582/15820001.
pdf.

Bouchard:1999:OP

[1091] D. Bouchard. Optimal parameters.

Lecture Notes in Computer Science, 1582:28–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1582/15820028.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1582/15820028.
pdf.

Kandulski:1999:SEG

[1092] Maciej Kandulski. Strong equivalence of generalized Ajdukiewicz and Lambek grammars. Lecture Notes in Computer Science, 1582:54–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1582/15820054.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1582/15820054.
pdf.

Shaumyan:1999:LPP

[1093] Sebastian Shaumyan and Paul Hudak. Linguistic, philosophical, and pragmatic aspects of type-directed natural language parsing. Lecture Notes in Computer Science, 1582:70– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1582/15820070.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1582/15820070.
                                        pdf.

Cornell:1999:DRV

[1094] Thomas Cornell. Derivational and representational views of minimalist transformational grammar. Lecture Notes in Computer Science, 1582:92– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1582/15820092.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1582/15820092.
                                        pdf.

Morawietz:1999:MLA

[1095] Frank Morawietz and Tom Cornell. The MSO logic-automaton connection in linguistics. Lecture Notes in Computer Science, 1582:112–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1582/15820112.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1582/15820112.
                                        pdf.

Hendriks:1999:LTP

[1096] Herman Hendriks. The logic of tune, a proof-theoretic analysis of intonation. Lecture Notes in Computer Science, 1582:132–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1582/15820132.htm;   http:

//link.springer-ny.com/link/service/
series/0558/papers/1582/15820132.
pdf.

Hodas:1999:LLT

[1097] Joshua S. Hodas. A linear logic treatment of phrase structure grammars for unbounded dependencies. Lecture Notes in Computer Science, 1582:160– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1582/15820160.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1582/15820160.
pdf.

Heylen:1999:UTL

[1098] Dirk Heylen. Underspecification in type-logical grammars. Lecture Notes in Computer Science, 1582:180–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1582/15820180.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1582/15820180.
pdf.

Francez:1999:FFL

[1099] Nissim Francez. On fibring feature logics with concatenation logics. Lecture Notes in Computer Science, 1582:200– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1582/15820200.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1582/15820200.
pdf.

Vilares:1999:OMP

[1100] Manuel Vilares, Miguel A. Alonso, and David Cabrero. An operational model for parsing definite clause grammars with infinite terms. Lecture Notes in Computer Science, 1582:212– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1582/15820212.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1582/15820212.
                                        pdf.

Luo:1999:MVC

[1101] Zhaohui Luo and Paul Callaghan. Mathematical vernacular and conceptual well-formedness in mathematical language. Lecture Notes in Computer Science, 1582:231–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1582/15820231.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1582/15820231.
                                        pdf.

Anonymous:1999:AId

[1102] Anonymous. Author index. Lecture Notes in Computer Science, 1582:251– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/1582/15820251.pdf.

Scharstein:1999:I

[1103] Daniel Scharstein. 1. introduction. Lecture Notes in Computer Science, 1583:

1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1583/15830001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1583/15830001.
pdf.

Scharstein:1999:SIB

[1104] Daniel Scharstein. 2. A survey of image-based rendering and stereo. Lecture Notes in Computer Science, 1583: 23–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1583/15830023.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1583/15830023.
pdf.

Scharstein:1999:VS

[1105] Daniel Scharstein. 3. view synthesis.

Lecture Notes in Computer Science, 1583:41–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1583/15830041.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1583/15830041.
pdf.

Scharstein:1999:RES

[1106] Daniel Scharstein. 4. re-evaluating stereo. Lecture Notes in Computer Science, 1583:63–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1583/15830063.
htm;   http://link.springer-ny.

                                        com/link/service/series/0558/papers/
                                        1583/15830063.pdf.

Scharstein:1999:GBS

[1107] Daniel Scharstein. 5. gradient-based stereo. Lecture Notes in Computer Science, 1583:87–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1583/15830087.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1583/15830087.pdf.

Scharstein:1999:SUD

[1108] Daniel Scharstein. 6. stereo using diffusion. Lecture Notes in Computer Science, 1583:113–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1583/15830113.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1583/15830113.pdf.

Scharstein:1999:C

[1109] Daniel Scharstein. 7. conclusion. Lecture Notes in Computer Science, 1583: 145–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1583/15830145.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1583/15830145.
                                        pdf.

Scharstein:1999:B

[1110] Daniel Scharstein. Bibliography. Lecture Notes in Computer Science, 1583: 149–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/1583/1583bibs.pdf.

Scharstein:1999:VSU

[1111] D. Scharstein. View synthesis using stereo vision. Lecture Notes in Computer Science, 1583:i–xiv + 1–163, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hajek:1999:TTF

[1112] P. Hajek. Trakhtenbrot theorem and fuzzy logic. Lecture Notes in Computer Science, 1584:1–8, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schwentick:1999:DCL

[1113] T. Schwentick. Descriptive complexity, lower bounds and linear time. Lecture Notes in Computer Science, 1584:9–28, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yannakakis:1999:TFS

[1114] M. Yannakakis and D. Lee. Testing of finite state systems. Lecture Notes in Computer Science, 1584:29–44, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kempe:1999:PQF

[1115] D. Kempe and A. Schoenegge. On the power of quantifiers in first-order algebraic specification. Lecture Notes in Computer Science, 1584:45–57, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giannotti:1999:ESN

[1116] F. Giannotti, G. Manco, M. Nanni, and D. Pedreschi. On the effective semantics of nondeterministic, nonmonotonic, temporal logic databases. Lecture Notes in Computer Science, 1584: 58–72, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marek:1999:RPL

[1117] V. Marek, I. Pivkina, and M. Truszczynski. Revision programming = logic programming + integrity constraints. Lecture Notes in Computer Science, 1584: 73–89, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Egly:1999:QSS

[1118] U. Egly. Quantifiers and the system KE: Some surprising results. Lecture Notes in Computer Science, 1584:90– 104, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoang:1999:CCL

[1119] H. K. Hoang. Choice construct and lindstroem logics. Lecture Notes in Computer Science, 1584:105–125, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kreidler:1999:MNG

[1120] M. Kreidler and D. Seese. Monadic NP and graph minors. Lecture Notes in Computer Science, 1584:126–141, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Makowsky:1999:IDP

[1121] J. A. Makowsky. Invariant definability and P/poly. Lecture Notes in Computer Science, 1584:142–158, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pezzoli:1999:CCE

[1122] E. Pezzoli. Computational complexity of Ehrenfeucht-Fraiesse games on finite structures. Lecture Notes in Computer Science, 1584:159–170, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buening:1999:UBM

[1123] H. Kleine Buening. An upper bound for minimal resolution refutations. Lecture Notes in Computer Science, 1584:171– 178, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sadowski:1999:ODA

[1124] Z. Sadowski. On an optimal deterministic algorithm for SAT. Lecture Notes in Computer Science, 1584:179–187, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Korovina:1999:CPM

[1125] M. V. Korovina and O. V. Kudinov.

Characteristic properties of majorantcomputability over the reals. Lecture Notes in Computer Science, 1584:188– 203, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Komara:1999:TPP

[1126] J. Komara and P. J. Voda. Theorems of Peter and Parsons in computer programming. Lecture Notes in Computer Science, 1584:204–223, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Riche:1999:KBU

[1127] J. Riche and R. K. Meyer. Kripke, Belnap, Urquhart and relevant decidability and complexity. Lecture Notes in Computer Science, 1584:224–240, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barthe:1999:EUN

[1128] G. Barthe. Existence and uniqueness of normal forms in pure type systems with eta-conversion. Lecture Notes in Computer Science, 1584:241–259, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khasidashvili:1999:NTT

[1129] Z. Khasidashvili and A. Piperno. Normalization of typable terms by superdevelopments. Lecture Notes in Computer Science, 1584:260–282, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vorobyov:1999:SFN

[1130] S. Vorobyov. Subtyping functional + nonempty record types. Lecture Notes in Computer Science, 1584:283–297, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matthes:1999:MFP

[1131] R. Matthes. Monotone fixed-point types and strong normalization. Lecture Notes in Computer Science, 1584: 298–312, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Statman:1999:MPS

[1132] R. Statman. Morphisms and partitions of V -sets. Lecture Notes in Computer Science, 1584:313–322, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Simpson:1999:CAE

[1133] A. K. Simpson. Computational adequacy in an elementary topos. Lecture Notes in Computer Science, 1584:323– 342, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Altenkirch:1999:LRI

[1134] T. Altenkirch. Logical relations and inductive/coinductive types. Lecture Notes in Computer Science, 1584:343– 354, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pichler:1999:CS

[1135] R. Pichler. On the complexity of H-subsumption. Lecture Notes in Computer Science, 1584:355–371, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonfante:1999:CCR

[1136] G. Bonfante, A. Cichon, J.-Y. Marion, and H. Touzet. Complexity classes and rewrite systems with polynomial interpretation. Lecture Notes in Computer Science, 1584:372–384, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Narendran:1999:RCS

[1137] P. Narendran, M. Rusinowitch, and R. Verma. RPO constraint solving is in NP. Lecture Notes in Computer Science, 1584:385–398, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baaz:1999:QEF

[1138] M. Baaz and H. Veith. Quantifier elimination in fuzzy logic. Lecture Notes in Computer Science, 1584:399–414, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lukasiewicz:1999:MVF

[1139] T. Lukasiewicz. Many-valued firstorder logics with probabilistic semantics. Lecture Notes in Computer Science, 1584:415–429, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schwefel:1999:NC

[1140] Hans-Paul Schwefel. Natural computation. Lecture Notes in Computer Science, 1585:1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1585/15850001.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1585/15850001.pdf.

Myung:1999:MLM

[1141] Hyun Myung and Jong-Hwan Kim. Multiple Lagrange multiplier method for constrained evolutionary optimization. Lecture Notes in Computer Science, 1585:2–9, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1585/15850002.
htm;   http://link.springer-ny.

com/link/service/series/0558/papers/
1585/15850002.pdf.

Ohkura:1999:RES

[1142] Kazuhiro Ohkura, Yoshiyuki Matsumura, and Kanji Ueda. Robust evolution strategies. Lecture Notes in Computer Science, 1585:10–17, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1585/15850010.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1585/15850010.
pdf.

Estivill-Castro:1999:HGA

[1143] Vladimir Estivill-Castro and Rodolfo Torres-Velázquez. Hybrid genetic algorithm for solving the p-median problem. Lecture Notes in Computer Science, 1585:18–25, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1585/15850018.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1585/15850018.pdf.

Pham:1999:CRL

[1144] Binh Pham and Zhongwei Zhang.

Correction of reflection lines using genetic algorithms. Lecture Notes in Computer Science, 1585:26–33, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1585/15850026.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/1585/15850026.
pdf.

Sasaki:1999:AUC

[1145] Takahiro Sasaki and Mario Tokoro. Adaptation under changing environments with various rates of inheritance of acquired characters: Comparison between darwinian and lamarckian evolution. Lecture Notes in Computer Science, 1585:34–41, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1585/15850034.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1585/15850034.
pdf.

Liang:1999:DCA

[1146] Ko-Hsin Liang, Xin Yao, and Charles Newton. Dynamic control of adaptive parameters in evolutionary programming. Lecture Notes in Computer Science, 1585:42–49, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1585/15850042.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1585/15850042.pdf.

Yamada:1999:IOS

[1147] Takeshi Yamada, Kazuyuki Yoshimura, and Ryohei Nakano. Information operator scheduling by genetic algorithms. Lecture Notes in Computer Science, 1585:50–57, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1585/15850050.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1585/15850050.
pdf.

Carvalho:1999:SRT

[1148] P. M. S. Carvalho, L. A. F. M. Ferreira, and L. M. F. Barruncho. Solving radial topology constrained problems with evolutionary algorithms. Lecture Notes in Computer Science, 1585:58– 65, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1585/15850058.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1585/15850058.
pdf.

Burke:1999:ASA

[1149] E. K. Burke and D. B. Varley. Automating space allocation in higher education. Lecture Notes in Computer Science, 1585:66–73, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1585/15850066.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1585/15850066.
pdf.

He:1999:AGA

[1150] Hongxing He, Warwick Graco, and Xin Yao. Application of genetic algorithm and k-nearest neighbour method in medical fraud detection. Lecture Notes in Computer Science, 1585:74– 81, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1585/15850074.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1585/15850074.
                                        pdf.

Ishibuchi:1999:ERS

[1151] Hisao Ishibuchi and Tomoharu Nakashima. Evolution of reference sets in nearest neighbor classification. Lecture Notes in Computer Science, 1585:82– 89, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1585/15850082.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1585/15850082.
                                        pdf.

Kirley:1999:ICG

[1152] Michael Kirley, Xiaodong Li, and David G. Green. Investigation of a cellular genetic algorithm that mimics landscape ecology. Lecture Notes in Computer Science, 1585:90–97, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1585/15850090.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1585/15850090.
                                        pdf.

Garionis:1999:QNP

[1153] Ralf Garionis. Quantifying neighborhood preservation: Joint properties of evolutionary and unsupervised neural learning. Lecture Notes in Computer Science, 1585:98–105, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1585/15850098.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1585/15850098.
pdf.

Mandischer:1999:NNE

[1154] Martin Mandischer, Hannes Geyer, and Peter Ulbig. Neural networks and evolutionary algorithms for the prediction of thermodynamic properties for chemical engineering. Lecture Notes in Computer Science, 1585:106– 113, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1585/15850106.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1585/15850106.
pdf.

Porter:1999:EFB

[1155] Reid Porter and Neil Bergmann.

Evolving FPGA based cellular automata. Lecture Notes in Computer Science, 1585:114–121, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1585/15850114.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1585/15850114.pdf.

Horii:1999:AIP

[1156] Hirosuke Horii, Susumu Kunifuji, and Teruo Matsuzawa. Asynchronous island parallel GA using multiform subpopulations. Lecture Notes in Computer Science, 1585:122–129, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1585/15850122.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1585/15850122.
                                        pdf.

Anbarasu:1999:MSA

[1157] L. A. Anbarasu, P. Narayanasamy, and V. Sundararajan. Multiple sequence alignment using parallel genetic algorithms. Lecture Notes in Computer Science, 1585:130–137, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1585/15850130.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1585/15850130.
                                        pdf.

Reiser:1999:ELP

[1158] Philip G. K. Reiser and Patricia J. Riddle. Evolving logic programs to classify chess-endgame positions. Lecture Notes in Computer Science, 1585:138– 145, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1585/15850138.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1585/15850138.
                                        pdf.

Zhang:1999:GPA

[1159] Byoung-Tak Zhang and Dong-Yeon Cho. Genetic programming with active data selection. Lecture Notes in Computer Science, 1585:146–153, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1585/15850146.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1585/15850146.
pdf.

Kim:1999:EPB

[1160] Yong-Jae Kim, Dong-Han Kim, and Jong-Hwan Kim. Evolutionary programming-based uni-vector field method for fast mobile robot navigation. Lecture Notes in Computer Science, 1585:154–161, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1585/15850154.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1585/15850154.pdf.

Ishinishi:1999:ELA

[1161] Masayuki Ishinishi and Akira Namatame. Evolution with learning adaptive functions. Lecture Notes in Computer Science, 1585:162–170, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1585/15850162.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1585/15850162.
pdf.

Podlich:1999:MPB

[1162] D. W. Podlich and M. Cooper.

Modelling plant breeding programs as search strategies on a complex response surface. Lecture Notes in Computer Science, 1585:171–178, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1585/15850171.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1585/15850171.
                                        pdf.

Shimooka:1999:GEG

[1163] Hiroaki Shimooka and Yoshiji Fujimoto. Generating equations with genetic programming for control of a movable inverted pendulum. Lecture Notes in Computer Science, 1585:179– 186, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1585/15850179.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1585/15850179.
                                        pdf.

Shiraooka:1999:GEG

[1164] H. Shiraooka and Y. Fujimoto. Generating equations with genetic programming for control of a movable inverted pendulum. Lecture Notes in Computer Science, 1585:179–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burke:1999:HTS

[1165] Edmund Burke, Patrick De Causmaecker, and Greet Vanden Berghe. A hybrid tabu search algorithm for the nurse rostering problem. Lecture Notes in Computer Science, 1585:187– 194, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1585/15850187.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1585/15850187.
pdf.

Sutton:1999:RLP

[1166] Richard S. Sutton. Reinforcement learning: Past, present and future. Lecture Notes in Computer Science, 1585: 195–197, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1585/15850195.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1585/15850195.
pdf.

Kawakami:1999:RLC

[1167] Hiroshi Kawakami, Osamu Katai, and Tadataka Konishi. A reinforcement learning with condition reduced fuzz rules. Lecture Notes in Computer Science, 1585:198–205, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1585/15850198.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1585/15850198.
pdf.

Tachibana:1999:GCS

[1168] Kanta Tachibana and Takeshi Furuhashi. Generality and conciseness of submodels in hierarchical fuzzy modeling. Lecture Notes in Computer Science, 1585:206–214, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1585/15850206.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1585/15850206.
                                        pdf.

Porto:1999:UEP

[1169] V. William Porto. Using evolutionary programming to optimize the allocation of surveillance assets. Lecture Notes in Computer Science, 1585:215– 222, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1585/15850215.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1585/15850215.
                                        pdf.

Kaise:1999:AEN

[1170] Naoki Kaise and Yoshiji Fujimoto. Applying the evolutionary neural networks with genetic algorithms to control a rolling inverted pendulum. Lecture Notes in Computer Science, 1585: 223–230, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1585/15850223.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1585/15850223.
                                        pdf.

Fuginaga:1999:ECA

[1171] T. Fuginaga, K. Moriwaki, N. Inuzuka, and H. Itoh. Evolving cooperative actions among heterogeneous agents by an evolutionary programming method. Lecture Notes in Computer Science, 1585:231–239, 1999.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fujinaga:1999:ECA

[1172] Takayuki Fujinaga, Kousuke Moriwaki, Nobuhiro Inuzuka, and Hidenori Itoh. Evolving cooperative actions among heterogeneous agents by an evolutionary programming method. Lecture Notes in Computer Science, 1585:231– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1585/15850231.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1585/15850231.
pdf.

Yamaguchi:1999:CWW

[1173] Toru Yamaguchi, Makoto Sato, Tomohiro Takagi, and Hideki Hashimoto. Cooperative works for welfare agent robot and human using chaotic evolutionary computation. Lecture Notes in Computer Science, 1585:240–250, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1585/15850240.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1585/15850240.
pdf.

Kohata:1999:ECI

[1174] Naoki Kohata, Makoto Sato, Toru Yamaguchi, Takanobu Baba, and Hideki Hashimoto. Evolutionary computation for intelligent agents based on chaotic retrieval and soft DNA. Lecture Notes in Computer Science, 1585:251– 259, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1585/15850251.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1585/15850251.
                                        pdf.

Aoki:1999:SBC

[1175] Keiko Aoki, Kazunori Matsumoto, Keiichiro Hoashi, and Kazuo Hashimoto. A study of Bayesian clustering of a document set based on GA. Lecture Notes in Computer Science, 1585:260– 267, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1585/15850260.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1585/15850260.
                                        pdf.

Husbands:1999:EAQ

[1176] Phil Husbands and Pedro P. B. de Oliveira. An evolutionary approach in quantitative spectroscopy. Lecture Notes in Computer Science, 1585:268– 275, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1585/15850268.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1585/15850268.
                                        pdf.

Tsujimura:1999:ERF

[1177] Yasuhiro Tsujimura and Mitsuo Gen. Evolutionary recognition of features from CAD data. Lecture Notes in Computer Science, 1585:276–284, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1585/15850276.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1585/15850276.
pdf.

Johansson:1999:MSG

[1178] Stefan Johansson, Bengt Carlsson, and Magnus Boman. Modeling strategies as generous and greedy in Prisoner’s Dilemma like games. Lecture Notes in Computer Science, 1585:285– 292, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1585/15850285.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1585/15850285.
pdf.

Chen:1999:UGA

[1179] Shu-Heng Chen and Chih-Chi Ni. Using genetic algorithms to simulate the evolution of an oligopoly game. Lecture Notes in Computer Science, 1585:293– 300, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1585/15850293.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1585/15850293.
pdf.

Chert:1999:UGA

[1180] S.-H. Chert and C.-C. Ni. Using genetic algorithms to simulate the evolution of an oligopoly game. Lecture Notes in Computer Science, 1585:293–??, 1999.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seo:1999:ESC

[1181] Yeon-Gyu Seo and Sung-Bae Cho. An evolutionary study on cooperation in N-person iterated Prisoner’s Dilemma game. Lecture Notes in Computer Science, 1585:301–308, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1585/15850301.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1585/15850301.
pdf.

Iwasaki:1999:SPM

[1182] Atsushi Iwasaki, Sobei H. Oda, and Kanji Ueda. Simulating a Nperson multi-stage game for making a state. Lecture Notes in Computer Science, 1585:309–316, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1585/15850309.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1585/15850309.
pdf.

Tanaka:1999:LLR

[1183] Kimiko Tanaka, Manabu Nii, and Hisao Ishibuchi. Learning from linguistic rules and rule extraction for function approximation by neural networks. Lecture Notes in Computer Science, 1585:317–324, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1585/15850317.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1585/15850317.
pdf.

Imada:1999:CNM

[1184] Akira Imada and Keijiro Araki. Can a niching method locate multiple attractors embedded in the Hopfield network? Lecture Notes in Computer Science, 1585:325–332, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1585/15850325.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1585/15850325.
pdf.

Liu:1999:TSP

[1185] Yong Liu and Xin Yao. Time series prediction by using negatively correlated neural networks. Lecture Notes in Computer Science, 1585:333– 340, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1585/15850333.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1585/15850333.
pdf.

Li:1999:AEP

[1186] An Li and Kit Po Wong. Animating the evolution process of genetic algorithms. Lecture Notes in Computer Science, 1585:341–348, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1585/15850341.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1585/15850341.
                                        pdf.

Niwa:1999:AIM

[1187] Tatsuya Niwa and Masaru Tanaka. Analysis on the island model parallel genetic algorithms for the genetic drifts. Lecture Notes in Computer Science, 1585:349–356, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1585/15850349.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1585/15850349.
                                        pdf.

Tonkes:1999:PNE

[1188] Bradley Tonkes, Alan Blair, and Janet Wiles. A paradox of neural encoders and decoders or why don’t we talk backwards? Lecture Notes in Computer Science, 1585:357–364, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1585/15850357.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1585/15850357.
                                        pdf.

Djurisic:1999:COU

[1189] Aleksandra B. Djurisic, Aleksandar D. Rakic, E. Herbert Li, Marijan L. Majewski, Nenad Bundaleski, and Bozidar V. Stanic. Continuous optimization using elite genetic algorithms with adaptive mutations. Lecture Notes in Computer Science, 1585:365–372, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1585/15850365.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1585/15850365.
pdf.

Zebulum:1999:ESA

[1190] Ricardo S. Zebulum, Marco Aurélio Pacheco, and Marley Vellasco. Evolutionary systems applied to the synthesis of a CPU controller. Lecture Notes in Computer Science, 1585:373– 380, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1585/15850373.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1585/15850373.
pdf.

Gero:1999:NME

[1191] John S. Gero. Novel models in evolutionary designing. Lecture Notes in Computer Science, 1585:381–388, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1585/15850381.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1585/15850381.
pdf.

Blair:1999:CED

[1192] Alan D. Blair, Elizabeth Sklar, and Pablo Funes. Co-evolution, determinism and robustness. Lecture Notes in Computer Science, 1585:389–396, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1585/15850389.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1585/15850389.
                                        pdf.

Hallinan:1999:COE

[1193] Jennifer Hallinan and Paul Jackway. Co-operative evolution of a neural classifier and feature subset. Lecture Notes in Computer Science, 1585:397– 404, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1585/15850397.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1585/15850397.
                                        pdf.

Wong:1999:OPF

[1194] Kit Po Wong and Jason Yuryevich. Optimal power flow method using evolutionary programming. Lecture Notes in Computer Science, 1585:405– 412, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1585/15850405.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1585/15850405.
                                        pdf.

Cho:1999:GDE

[1195] Sung-Bae Cho and Katsunori Shimohara. Grammatical development of evolutionary modular neural networks. Lecture Notes in Computer Science, 1585:413–420, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1585/15850413.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1585/15850413.
pdf.

Gen:1999:HNN

[1196] Mitsuo Gen, Kenichi Ida, and Chang-Yun Lee. Hybridized neural network and genetic algorithms for solving nonlinear integer programming problem. Lecture Notes in Computer Science, 1585:421–430, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1585/15850421.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1585/15850421.
pdf.

Runarsson:1999:EGC

[1197] Thomas Philip Runarsson and Magnus Thor Jonsson. Evolution of gene coordination networks. Lecture Notes in Computer Science, 1585:430–437, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1585/15850430.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1585/15850430.
pdf.

Hall:1999:ASI

[1198] Richard Hall, Binh Pham, and John Yearwood. Adaptive simulation: An implementation framework. Lecture Notes in Computer Science, 1585:438– 445, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1585/15850438.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1585/15850438.
                                        pdf.

Akira:1999:MMA

[1199] Yoshida Akira. A model of mutual associative memory for simulations of evolution and learning. Lecture Notes in Computer Science, 1585:446– 453, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1585/15850446.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1585/15850446.
                                        pdf.

Oda:1999:ACA

[1200] Sobei H. Oda, Kouhei Iyori, Miura Ken, and Kanji Ueda. The application of cellular automata to the consumer’s theory: Simulating a duopolistic market. Lecture Notes in Computer Science, 1585:454–461, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1585/15850454.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1585/15850454.
                                        pdf.

Lai:1999:OOG

[1201] L. L. Lai, H. Subasinghe, N. Rajkumar, E. Vaseekar, B. J. Gwyn, and V. K. Sood. Object-oriented genetic algorithm based artificial neural network for load forecasting. Lecture Notes in Computer Science, 1585:462– 469, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1585/15850462.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1585/15850462.
pdf.

Mueller:1999:WHL

[1202] F. Mueller. Workshop on high-level parallel programming models and supportive environments. Lecture Notes in Computer Science, 1586:1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Subramanian:1999:EPP

[1203] R. Subramanian and S. Pande. Efficient program partitioning based on compiler controlled communication. Lecture Notes in Computer Science, 1586:4–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schulz:1999:SVF

[1204] M. Schulz. SCI-VM: A flexible base for transparent shared memory programming models on clusters of PCs. Lecture Notes in Computer Science, 1586: 19–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nolte:1999:FCO

[1205] J. Nolte. Flexible collective operations for distributed object groups. Lecture Notes in Computer Science, 1586:34– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sun:1999:SFP

[1206] X.-H. Sun, M. Pantano, T. Fahringer, and Z. Zhan. SCALA: A framework for performance evaluation of scalable computing. Lecture Notes in Computer Science, 1586:49–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mou:1999:RID

[1207] Z. G. Mou. Recursive individually distributed objects. Lecture Notes in Computer Science, 1586:63–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leberecht:1999:MSF

[1208] M. Leberecht. The MuSE system: A flexible combination of on-stack execution and work-stealing. Lecture Notes in Computer Science, 1586:79– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spiegel:1999:PAD

[1209] A. Spiegel. Pangaea: An automatic distribution front-end for Java. Lecture Notes in Computer Science, 1586: 93–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fodor:1999:CLS

[1210] E. F. Fodor and R. A. Olsson. Concurrent language support for interoperable applications. Lecture Notes in Computer Science, 1586:100–??, 1999.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keller:1999:DIA

[1211] G. Keller and M. M. T. Chakravarty.

On the distributed implementation of aggregate data by program transformation structures. Lecture Notes in Computer Science, 1586:108–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gorlatch:1999:TFS

[1212] S. Gorlatch and S. Pelagatti. A transformational framework for skeletal programs: Overview and case study. Lecture Notes in Computer Science, 1586: 123–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kusakabe:1999:INS

[1213] S. Kusakabe, K. Inenaga, M. Amamiya, and X. Tang. Implementing a nonstrict functional programming language on a threaded architecture. Lecture Notes in Computer Science, 1586: 138–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zomaya:1999:WBI

[1214] A. Y. Zomaya, F. Ercal, and S. Olariu.

Workshop on biologically inspired solutions to parallel processing problems. Lecture Notes in Computer Science, 1586:153–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

King:1999:BBI

[1215] R. L. King, A. B. Lambert, S. H. Russ, and D. S. Reese. The biological basis of the immune system as a model for intelligent agents. Lecture Notes in Computer Science, 1586:156–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Szuba:1999:FDP

[1216] T. Szuba. A formal definition of the phenomenon of collective intelligence and its IQ measure. Lecture Notes in Computer Science, 1586:165–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wasiewicz:1999:IDF

[1217] P. Wasiewicz, P. Borsuk, J. J. Mulawka, and P. Weglenski. Implementation of data flow logical operations via self-assembly of DNA. Lecture Notes in Computer Science, 1586:174–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vianna:1999:PHE

[1218] D. S. Vianna, L. S. Ochi, and L. M. A. Drummond. A parallel hybrid evolutionary metaheuristic for the period vehicle routing problem. Lecture Notes in Computer Science, 1586:183–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seredynski:1999:DSD

[1219] F. Seredynski, J. Koronacki, and C. Z. Janikow. Distributed scheduling with decomposed optimization criterion: Genetic programming approach. Lecture Notes in Computer Science, 1586:192–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alaoui:1999:PGA

[1220] S. M. Alaoui, O. Frieder, and T. El-Ghazawi. A parallel genetic algorithm for task mapping on parallel machines. Lecture Notes in Computer Science, 1586:201–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jedrzejowicz:1999:EBS

[1221] P. Jedrzejowicz, I. Czarnowski, H. Szreder, and A. Skakowski. Evolution-based scheduling of fault-tolerant programs on multiple processors. Lecture Notes in Computer Science, 1586:210–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lob:1999:GBF

[1222] P. K. K. Lob and V. Shaw. A geneticbased fault-tolerant routing strategy for multiprocessor networks. Lecture Notes in Computer Science, 1586:220– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leopold:1999:RCI

[1223] C. Leopold. Regularity considerations in instance-based locality optimization. Lecture Notes in Computer Science, 1586:230–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Talbi:1999:PAC

[1224] E.-G. Talbi, O. Roux, C. Fonlupt, and D. Robillard. Parallel ant colonies for combinatorial optimization problems. Lecture Notes in Computer Science, 1586:239–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alba:1999:ASA

[1225] E. Alba and J. M. Troya. An analysis of synchronous and asynchronous parallel distributed genetic algorithms with structured and panmictic islands. Lecture Notes in Computer Science, 1586: 248–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chalermwat:1999:GBP

[1226] P. Chalermwat, T. El-Chazawi, and J. LeMoigne. GA-based parallel image registration on parallel clusters. Lecture Notes in Computer Science, 1586: 257–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sena:1999:IPG

[1227] C. Sena, G. Isem, and D. Megherbi. Implementation of a parallel genetic algorithm on a cluster of workstations: The traveling salesman problem, A case study. Lecture Notes in Computer Science, 1586:266–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boeleoni:1999:SBM

[1228] L. Boeleoni, R. Hao, K. Jun, and D. C. Marinescu. Structural biology metaphors applied to the design of a distributed object system. Lecture Notes in Computer Science, 1586:275– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ravindran:1999:WPD

[1229] B. Ravindran, J. Gustafsson, and H. Takada. Workshop on parallel and distributed real-time systems. Lecture Notes in Computer Science, 1586:284– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1999:BAM

[1230] L. Chen, S. Khan, K. F. Li, and E. G. Manning. Building an adaptive multimedia system using the utility mode. Lecture Notes in Computer Science, 1586:289–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rajagopat:1999:ERT

[1231] P. Rajagopat and A. W. Apon. Evaluation of real-time fiber communications for parallel collective operations. Lecture Notes in Computer Science, 1586:299–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dinda:1999:CPB

[1232] P. A. Dinda, L. F. Kallivokas, B. Lowekamp, and D. R. O’Hallaron. The case for prediction-based besteffort real-time systems. Lecture Notes in Computer Science, 1586:309–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mittal:1999:DRT

[1233] A. Mittal, G. Manimaran, and C. S. R.

Murthy. Dynamic real-time channel establishment in multiple access bus networks. Lecture Notes in Computer Science, 1586:319–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lam:1999:SBP

[1234] K.-Y. Lam, T.-W. Kuo, G. C. K. Law, and W.-H. Tsang. A similarity-based protocol for concurrency control in mobile distributed real-time database systems. Lecture Notes in Computer Science, 1586:329–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tovar:1999:TSS

[1235] E. Tovar and F. Vasques. From task scheduling in single processor environments to message scheduling in a PROFIBUS fieldbus network. Lecture Notes in Computer Science, 1586:339– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clark:1999:ADA

[1236] R. Clark, E. D. Jensen, A. Kanevsky, and J. Maurer. An adaptive distributed airborne tracking system. Lecture Notes in Computer Science, 1586: 353–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jonsson:1999:NPS

[1237] J. Jonsson, H. Loenn, and K. G. Shin. Non-preemptive scheduling of real-time threads on multi-level-context architectures. Lecture Notes in Computer Science, 1586:363–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nat-Abdesselam:1999:QCA

[1238] F. Nat-Abdesselam and N. Agoulmine. QoS control and adaptation in distributed multimedia systems. Lecture Notes in Computer Science, 1586:375– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Campelo:1999:DEF

[1239] J. C. Campelo, P. Yuste, F. Rodriguez, and P. J. Gil. Dependability evaluation of fault tolerant distributed industrial control systems. Lecture Notes in Computer Science, 1586:384–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chappell:1999:AMI

[1240] B. L. Chappell, D. T. Marlow, P. M.

Ivey, and K. O’Donoghue. An approach for measuring IP security performance in a distributed environment. Lecture Notes in Computer Science, 1586:389–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fagundes:1999:EGA

[1241] L. G. Fagundes, R. F. Mello, and C. E. Mdron. An environment for generating applications involving remote manipulation of parallel machines. Lecture Notes in Computer Science, 1586:395–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gentile:1999:RTI

[1242] A. Gentile, Y. L. Cruz-Rivera, D. S.

Wills, and L. Bustelo. Real-time image processing on a focal plane SIMD array. Lecture Notes in Computer Science, 1586:400–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ivey:1999:MEM

[1243] P. M. Ivey and D. T. Marlow. Metrics for the evaluation of multicast communications. Lecture Notes in Computer Science, 1586:406–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jaeyong:1999:DPW

[1244] K. Jaeyong, K. Kihan, and S. Heonshik. Distributing periodic workload uniformly across time to achieve better service quality. Lecture Notes in Computer Science, 1586:413–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:1999:DFT

[1245] J.-M. Huang and T. C. Yang. A dynamic fault-tolerant mesh architecture. Lecture Notes in Computer Science, 1586:418–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Livani:1999:EHR

[1246] Mohammad Ali Livani and J. Kaiser. Evaluation of a hybrid real-time bus scheduling mechanism for CAN. Lecture Notes in Computer Science, 1586: 425–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakajima:1999:SSM

[1247] T. Nakajima, M. T. Kone, and H. Aizu. System support for migratory continuous media applications in distributed real-time environments. Lecture Notes in Computer Science, 1586:430–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chatterjee:1999:DAS

[1248] S. Chatterjee. Dynamic application structuring on heterogeneous, distributed systems. Lecture Notes in Computer Science, 1586:442–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Niehaus:1999:ISM

[1249] D. Niehaus. Improving support for multimedia system experimentation and deployment. Lecture Notes in Computer Science, 1586:454–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Olsson:1999:WRT

[1250] R. Olsson, L. V. Kale, P. Beckman, and M. Haines. Workshop on runtime systems for parallel programming. Lecture Notes in Computer Science, 1586:466–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouge:1999:ECM

[1251] L. Bouge, J.-F. Mehaut, and R. Namyst.

Efficient communications in multithreaded runtime systems. Lecture Notes in Computer Science, 1586:468– 482, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kale:1999:APL

[1252] L. Kale, R. Brunner, J. Phillips, and K. Varadarajan. Application performance of a Linux cluster using Converse. Lecture Notes in Computer Science, 1586:483–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antoniu:1999:ETT

[1253] G. Antoniu, L. Bouge, and R. Namyst. An efficient and transparent thread migration scheme in the PM2 runtime system. Lecture Notes in Computer Science, 1586:496–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thitikamol:1999:CIP

[1254] K. Thitikamol and P. Keleher. Communication-intensive parallel applications and non-dedicated clusters of workstations. Lecture Notes in Computer Science, 1586:511–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Simitci:1999:FAS

[1255] H. Simitci, D. A. Reed, R. Fox, and M. Medina. A framework for adaptive storage input/output on computational grids. Lecture Notes in Computer Science, 1586:519–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nieplocha:1999:APR

[1256] J. Nieplocha and B. Carpenter. ARMCI: A portable remote memory copy library for distributed array libraries and compiler run-time systems. Lecture Notes in Computer Science, 1586:533–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seidmann:1999:MBR

[1257] T. Seidmann. Multicast-based runtime system for highly efficient causally consistent software-only DSM. Lecture Notes in Computer Science, 1586:547– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mueller:1999:ADR

[1258] F. Mueller. Adaptive DSM-Runtime behavior via speculative data distribution. Lecture Notes in Computer Science, 1586:553–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Elgindy:1999:RAW

[1259] H. Elgindy. Reconfigurable architectures workshop. Lecture Notes in Computer Science, 1586:568–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bondalapati:1999:DDE

[1260] K. Bondalapati, P. Diniz, P. Duncan, and J. Granacki. DEFACTO: A design environment for adaptive computing technology. Lecture Notes in Computer Science, 1586:570–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diessel:1999:WBM

[1261] O. Diessel, D. Kearney, and G. Wigley.

A Web-based multiuser operating system for reconfigurable computing. Lecture Notes in Computer Science, 1586: 579–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Srinivasan:1999:ISR

[1262] V. Srinivasan, S. Radhakrishnan, R. Vemuri, and J. Walrath. Interconnect synthesis for reconfigurable multi-FPGA architectures. Lecture Notes in Computer Science, 1586:588–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khalid:1999:HCP

[1263] M. A. S. Khalid and J. Rose. Hardwired-clusters partial-crossbar: A hierarchical routing architecture for multi-FPGA systems. Lecture Notes in Computer Science, 1586:597–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaul:1999:IBP

[1264] M. Kaul and R. Vemuri. Integrated block-processing and design-space exploration in temporal partitioning for RTR architectures. Lecture Notes in Computer Science, 1586:606–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernandez-Zepeda:1999:ISS

[1265] J. A. Fernandez-Zepeda, R. Vaidyanathan, and J. L. Trahan. Improved scaling simulation of the general reconfigurable mesh. Lecture Notes in Computer Science, 1586:616–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Middendorf:1999:BSR

[1266] M. Middendorf. Bit summation on the reconfigurable mesh. Lecture Notes in Computer Science, 1586:625–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1999:SHA

[1267] R. Lin, K. Nakano, S. Olariu, and M. C. Pinotti. Scalable hardware-algorithms for binary prefix sums. Lecture Notes in Computer Science, 1586:634–??, 1999.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wojko:1999:CSS

[1268] M. Wojko and H. ElGindy. Configuration sequencing with self configurable binary multipliers. Lecture Notes in Computer Science, 1586:643–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dandalis:1999:DSM

[1269] A. Dandalis, A. Mei, and V. K.

Prasanna. Domain specific mapping for solving graph problems on reconfigurable devices. Lecture Notes in Computer Science, 1586:652–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lu:1999:MRP

[1270] G. Lu, M.-H. Lee, H. Singh, and N. Bagherzadeh. MorphoSys: A reconfigurable processor targeted to high performance image application. Lecture Notes in Computer Science, 1586: 661–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nagano:1999:EIM

[1271] H. Nagano, A. Matsuura, and A. Nagoya. An efficient implementation method of fractal image compression on dynamically reconfigurable architecture. Lecture Notes in Computer Science, 1586:670–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakada:1999:PCA

[1272] H. Nakada, K. Oguri, N. Imlig, and M. Inamori. Plastic cell architecture: A dynamically reconfigurable hardwarebased computer. Lecture Notes in Computer Science, 1586:679–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ritter:1999:LDS

[1273] G. Ritter, J.-M. Puiatti, and E. Sanchez. Leonardo and Discipulus Simplex: An autonomous, evolvable six-legged walking robot. Lecture Notes in Computer Science, 1586:688–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agun:1999:RIH

[1274] K.-A. Agun and M. Chang. Reusable internal hardware templates. Lecture Notes in Computer Science, 1586:697– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beuchat:1999:LAB

[1275] J.-L. Beuchat and E. Sanchez. An on-line arithmetic-based reconfigurable neuroprocessor. Lecture Notes in Computer Science, 1586:700–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hardt:1999:RCD

[1276] W. Hardt, A. Rettberg, and B. Kleinjohann. The re-configurable delayintensive FLYSIG architecture. Lecture Notes in Computer Science, 1586:703– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koehler:1999:DSP

[1277] S. Koehler, S. Sawitzki, A. Gratz, and R. G. Spallek. Digital signal processing with general purpose microprocessors, DSP and reconfigurable logic. Lecture Notes in Computer Science, 1586:706–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Suyama:1999:SSP

[1278] T. Suyama, M. Yokoo, and A. Nagoya.

Solving satisfiability problems on FP-GAs using experimental unit propagation heuristic. Lecture Notes in Computer Science, 1586:709–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tiountchik:1999:FIM

[1279] A. Tiountchik and E. Trichina. FPGA implementation of modular exponentiation. Lecture Notes in Computer Science, 1586:712–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caromel:1999:WJP

[1280] D. Caromel, S. Chaumette, and G. Fox.

Workshop on Java for parallel and distributed computing. Lecture Notes in Computer Science, 1586:716–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Philippsen:1999:MEO

[1281] M. Philippsen and B. Haumacher.

More efficient object serialization. Lecture Notes in Computer Science, 1586: 718–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breg:1999:CIR

[1282] F. Breg and D. Cannon. A customizable implementation of RMI for high performance computing. Lecture Notes in Computer Science, 1586:733– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baker:1999:MOO

[1283] M. Baker, B. Carpenter, G. Fox, and Sung Hoon Koo. mpiJava: An objectoriented Java interface to MPI. Lecture Notes in Computer Science, 1586: 748–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sarmenta:1999:AFT

[1284] L. F. G. Sarmenta. An adaptive, faulttolerant implementation of BSP for Java-based volunteer computing systems. Lecture Notes in Computer Science, 1586:763–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clement:1999:HPC

[1285] M. Clement, Q. Snell, and G. Judd. High performance computing for the masses. Lecture Notes in Computer Science, 1586:781–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Webb:1999:PNH

[1286] D. Webb, A. Wendelborn, and K. Maciunas. Process networks as a high-level notation for metacomputing. Lecture Notes in Computer Science, 1586:797– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Galatopoullos:1999:DPA

[1287] D. G. Galatopoullos and E. S.

Manolakos. Developing parallel applications using the JavaPorts environment. Lecture Notes in Computer Science, 1586:813–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yi:1999:WOC

[1288] P. Yi and L. Keqin. Workshop on optics and computer science. Lecture Notes in Computer Science, 1586:829–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weifa:1999:PRA

[1289] L. Weifa and S. Xiaojun. Permutation routing in all-optical product networks. Lecture Notes in Computer Science, 1586:831–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carrera:1999:NOD

[1290] E. V. Carrera and R. Bianchini.

NWCache: Optimizing disk accesses via an optical network/write cache hybrid. Lecture Notes in Computer Science, 1586:845–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carrera:1999:NNC

[1291] E. V. Carrera and R. Bianchini. Net-Cache: A network/cache hybrid for multiprocessors. Lecture Notes in Computer Science, 1586:859–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Choo:1999:MWO

[1292] Peng Yin Choo, A. Detofsky, and A. Louri. A multi-wavelength optical content-addressable parallel processor (MW-OCAPP) for high-speed parallel relational database processing: Architectural concepts and preliminary experimental system. Lecture Notes in Computer Science, 1586:873–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hongjin:1999:OSA

[1293] Y. Hongjin, W. Kyubum, and H. Manpyo. Optimal scheduling algorithms in WDM optical passive star networks. Lecture Notes in Computer Science, 1586:887–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coudert:1999:OBM

[1294] D. Coudert, A. Ferreira, and X. Muñoz. OTIS-based multi-hop multi-OPS lightwave networks. Lecture Notes in Computer Science, 1586:897–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1999:SGT

[1295] K. Li, Y. Pan, and M. Hamdi. Solving graph theory problems using reconfigurable pipelined optical buses. Lecture Notes in Computer Science, 1586: 911–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feldman:1999:HSH

[1296] M. Feldman, R. Vaidyanathan, and A. El-Amawy. High speed, high capacity bused interconnects using optical slab waveguides. Lecture Notes in Computer Science, 1586:924–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jaekel:1999:NAM

[1297] A. Jaekel, S. Bandyopadhyay, and A. Sengupta. A new architecture for multihop optical networks. Lecture Notes in Computer Science, 1586:938– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xijun:1999:PVN

[1298] Z. Xijun and Q. Chunming. Pipelined versus non-pipelined traffic scheduling in unidirectional WDM rings. Lecture Notes in Computer Science, 1586:950– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tao:1999:WSI

[1299] Y. Tao. Workshop on solving irregularly structured problems in parallel. Lecture Notes in Computer Science, 1586:964–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heber:1999:SAW

[1300] C. Heber, R. Biswas, and C. R. Gao.

Self-avoiding walks over adaptive unstructured grids. Lecture Notes in Computer Science, 1586:968–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holzrichter:1999:GBM

[1301] M. Holzrichter and S. Oliveira. A graph based method for generating the Fiedler vector of irregular problems. Lecture Notes in Computer Science, 1586:978–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pellegrini:1999:HND

[1302] F. Pellegrini, J. Roman, and P. Amestoy. Hybridizing nested dissection and halo approximate minimum degree for efficient sparse matrix ordering. Lecture Notes in Computer Science, 1586:986– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1999:PPM

[1303] J. Chen and V. E. Taylor. ParaPART: Parallel mesh partitioning tool for distributed systems. Lecture Notes in Computer Science, 1586:996–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voisin:1999:SCP

[1304] F. Voisin and G.-R. Perrin. Sparse computations with PEI. Lecture Notes in Computer Science, 1586:1006–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benkner:1999:OIH

[1305] S. Benkner. Optimizing irregular HPF applications using halos. Lecture Notes in Computer Science, 1586:1015–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gao:1999:EHE

[1306] G. R. Gao. From EARTH to HTMT: An evolution of a multitheaded architecture model. Lecture Notes in Computer Science, 1586:1025–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blount:1999:IPA

[1307] B. Blount, S. Chatterjee, and M. Philippsen. Irregular parallel algorithms in Java. Lecture Notes in Computer Science, 1586:1026–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1999:SFC

[1308] Y. Lin and D. Padua. A simple framework to calculate the reaching definition of array references and its use in subscript array analysis. Lecture Notes in Computer Science, 1586:1036–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Everaars:1999:DPC

[1309] C. T. H. Everaars, B. Koren, and F. Arbab. Dynamic process composition and communication patterns in irregularly structured applications. Lecture Notes in Computer Science, 1586: 1046–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rhodes:1999:SPH

[1310] D. L. Rhodes and A. Gerasoulis. Scalable parallelization of harmonic balance simulation. Lecture Notes in Computer Science, 1586:1055–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boeres:1999:TET

[1311] C. Boeres, A. Nascimento, and V. E. F.

Rebello. Towards an effective task clustering heuristic for LogP machines. Lecture Notes in Computer Science, 1586:1065–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mongelli:1999:RMP

[1312] H. Mongelli and W. Song. A range minima parallel algorithm for coarse grained multicomputers. Lecture Notes in Computer Science, 1586:1075–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herley:1999:DBB

[1313] K. T. Herley, A. Pietracaprina, and G. Pucci. Deterministic branch-and-bound on distributed memory machines. Lecture Notes in Computer Science, 1586:1085–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chiola:1999:WPC

[1314] G. Chiola, G. Conte, and L. V. Mancini. Workshop on personal computer based networks of workstations. Lecture Notes in Computer Science, 1586:1095–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Donaldson:1999:PRR

[1315] S. R. Donaldson, J. M. D. Hill, and D. B. Skillicorn. Performance results for a reliable low-latency cluster communication protocol. Lecture Notes in Computer Science, 1586:1097–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karl:1999:CTS

[1316] H. Karl. Coscheduling through synchronized scheduling servers A prototype and experiments. Lecture Notes in Computer Science, 1586:1115–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anglano:1999:HPK

[1317] C. Anglano, A. Giordana, and G. Lo Bello. High-performance knowledge extraction from data on PC-based networks of workstations. Lecture Notes in Computer Science, 1586:1130–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rajasekaran:1999:ACL

[1318] U. K. V. Rajasekaran, M. Chetlur, G. D. Sharma, and R. Radhakrishnan. Addressing communication latency issues on clusters for fine grained asynchronous applications — A case study. Lecture Notes in Computer Science, 1586:1145–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takahashi:1999:IEM

[1319] T. Takahashi, F. O’Carroll, H. Tezuka, and A. Hori. Implementation and evaluation of MPI on an SMP cluster. Lecture Notes in Computer Science, 1586:1178–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mery:1999:WFM

[1320] D. Mery and B. Sanders. Workshop on formal methods for parallel programming: Theory and applications. Lecture Notes in Computer Science, 1586:1193–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Attali:1999:SEP

[1321] I. Attali, D. Caromel, and S. Lippi.

From a specification to an equivalence proof in object-oriented parallelism. Lecture Notes in Computer Science, 1586:1197–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Charpentier:1999:EPC

[1322] M. Charpentier and K. M. Chandy. Examples of program composition illustrating the use of universal properties. Lecture Notes in Computer Science, 1586:1215–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chernyakhovsky:1999:FFS

[1323] V. Chernyakhovsky, P. Frey, R. Radhakrishnan, and P. A. Wilsey. A formal framework for specifying and verifying time warp optimizations. Lecture Notes in Computer Science, 1586:1228–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Creese:1999:VEE

[1324] S. J. Creese and J. Reed. Verifying end-to-end protocols using induction with CSP/FDR. Lecture Notes in Computer Science, 1586:1243–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Havelund:1999:MVG

[1325] K. Havelund. Mechanical verification of a garbage collector. Lecture Notes in Computer Science, 1586:1258–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Massingill:1999:SAP

[1326] B. L. Massingill. A structured approach to parallel programming: Methodology and models. Lecture Notes in Computer Science, 1586:1284–??, 1999.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Simpson:1999:BCE

[1327] A. C. Simpson, J. M. D. Hill, and S. R. Donaldson. BSP in CSP: Easy as ABC. Lecture Notes in Computer Science, 1586:1299–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhatt:1999:WEH

[1328] D. Bhatt and V. Prasanna. Workshop on embedded HPC systems and applications. Lecture Notes in Computer Science, 1586:1314–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Welch:1999:DSR

[1329] L. R. Welch, M. W. Masters, L. A.

Madden, and D. T. Marlow. A distributed system reference architecture for adaptive QoS and resource management. Lecture Notes in Computer Science, 1586:1316–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Millan:1999:BFD

[1330] William Millan, Andrew Clark, and Ed Dawson. Boolean function design using hill climbing methods. Lecture Notes in Computer Science, 1587:1– 11, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1587/15870001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1587/15870001.
pdf.

Maitra:1999:ECI

[1331] Subhamoy Maitra and Palash Sarkar. Enumeration of correlation immune Boolean functions. Lecture Notes in Computer Science, 1587:12–25, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1587/15870012.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1587/15870012.
pdf.

Qu:1999:SPH

[1332] Chengxin Qu, Jennifer Seberry, and Josef Pieprzyk. On the symmetric property of homogeneous Boolean functions. Lecture Notes in Computer Science, 1587:26–35, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1587/15870026.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1587/15870026.
pdf.

Viswanathan:1999:PVK

[1333] Kapali Viswanathan, Colin Boyd, and Ed Dawson. Publicly verifiable key escrow with limited time span. Lecture Notes in Computer Science, 1587:36– 50, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1587/15870036.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1587/15870036.
pdf.

Lee:1999:AKE

[1334] Seungwon Lee, Seong-Min Hong, Hyunsoo Yoon, and Yookun Cho. Accelerating key establishment protocols for mobile communication. Lecture Notes in Computer Science, 1587:51– 63, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1587/15870051.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1587/15870051.
pdf.

Li:1999:CKA

[1335] Chih-Hung Li and Josef Pieprzyk.

Conference key agreement from secret sharing. Lecture Notes in Computer Science, 1587:64–76, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1587/15870064.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1587/15870064.pdf.

Fung:1999:PPS

[1336] W. W. Fung and J. W. Gray III.

On m-permutation protection scheme against modification attack. Lecture Notes in Computer Science, 1587:77– 87, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1587/15870077.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1587/15870077.
pdf.

Golic:1999:IAB

[1337] Jovan Dj. Golić, Andrew Clark, and Ed Dawson. Inversion attack and branching. Lecture Notes in Computer Science, 1587:88–102, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1587/15870088.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1587/15870088.
pdf.

Susilo:1999:FST

[1338] Willy Susilo, Rei Safavi-Naini, and Josef Pieprzyk. Fail-stop threshold signature schemes based on elliptic curves. Lecture Notes in Computer Science, 1587:103–116, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1587/15870103.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1587/15870103.
pdf.

Nguyen:1999:DZK

[1339] Khanh Quoc Nguyen, Yi Mu, and Vijay Varadharajan. Divertible zeroknowledge proof of polynomial relations and blind group signature. Lecture Notes in Computer Science, 1587: 117–128, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1587/15870117.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1587/15870117.
pdf.

Ghodosi:1999:RCN

[1340] Hossein Ghodosi and Josef Pieprzyk.

Repudiation of cheating and nonrepudiation of Zhang’s proxy signature schemes. Lecture Notes in Computer Science, 1587:129–134, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1587/15870129.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1587/15870129.
pdf.

Mueller:1999:SRB

[1341] S. Mueller. On the security of an RSA based encryption scheme. Lecture Notes in Computer Science, 1587:135– 148, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muller:1999:SRB

[1342] Siguna Müller. On the security of an RSA based encryption scheme. Lecture Notes in Computer Science, 1587: 135–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1587/15870135.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1587/15870135.
pdf.

Gysin:1999:GCA

[1343] Marc Gysin and Jennifer Seberry. Generalised cycling attacks on RSA and strong RSA primes. Lecture Notes in Computer Science, 1587:149–163, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1587/15870149.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1587/15870149.
pdf.

Tiountchik:1999:RAF

[1344] Alexander Tiountchik and Elena Trichina. RSA acceleration with field programmable gate arrays. Lecture Notes in Computer Science, 1587:164– 176, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1587/15870164.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1587/15870164.
pdf.

Martin:1999:CTA

[1345] Keith M. Martin, Josef Pieprzyk, Rei Safavi-Naini, and Huaxiong Wang. Changing thresholds in the absence of secure channels. Lecture Notes in Computer Science, 1587:177–191, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1587/15870177.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1587/15870177.
pdf.

Saeednia:1999:SCG

[1346] Shahrokh Saeednia and Hossein Ghodosi. A self-certified group-oriented cryptosystem without a combiner. Lecture Notes in Computer Science, 1587: 192–201, 1999. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1587/15870192.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1587/15870192.
pdf.

Horton:1999:CVM

[1347] Jeffrey Horton and Jennifer Seberry.

Companion viruses and the macintosh: Threats and countermeasures. Lecture Notes in Computer Science, 1587:202– 212, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1587/15870202.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1587/15870202.
pdf.

Ashley:1999:ISV

[1348] Paul Ashley, Bradley Broom, and Mark Vandenwauver. An implementation of a secure version of NFS including RBAC. Lecture Notes in Computer Science, 1587:213–227, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1587/15870213.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1587/15870213.
pdf.

Traore:1999:GST

[1349] Jacques Traoré. Group signatures and their relevance to privacy-protecting off-line electronic cash systems. Lecture Notes in Computer Science, 1587:228– 243, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1587/15870228.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1587/15870228.
                                        pdf.

Boyd:1999:EEC

[1350] Colin Boyd, Ernest Foo, and Chris Pavlovski. Efficient electronic cash using batch signatures. Lecture Notes in Computer Science, 1587:244–257, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1587/15870244.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1587/15870244.
                                        pdf.

Zhou:1999:EFN

[1351] Jianying Zhou, Robert Deng, and Feng Bao. Evolution of fair nonrepudiation with TTP. Lecture Notes in Computer Science, 1587:258–243, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1587/15870258.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1587/15870258.
                                        pdf.

Bai:1999:AOO

[1352] Yun Bai and Vijay Varadharajan. Authorization in object oriented databases. Lecture Notes in Computer Science, 1587:270–280, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1587/15870270.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1587/15870270.
pdf.

Saunders:1999:AAC

[1353] Gregory Saunders, Michael Hitchens, and Vijay Varadharajan. An analysis of access control models. Lecture Notes in Computer Science, 1587:281– 293, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1587/15870281.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1587/15870281.
pdf.

Lenstra:1999:EIB

[1354] Arjen K. Lenstra. Efficient identity based parameter selection for elliptic curve cryptosystems. Lecture Notes in Computer Science, 1587:294– 302, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1587/15870294.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1587/15870294.
pdf.

Pei:1999:COA

[1355] Dingyi Pei, Yuqiang Li, Yejing Wang, and Rei Safavi-Naini. Characterization of optimal authentication codes with arbitration. Lecture Notes in Computer Science, 1587:303–313, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1587/15870303.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1587/15870303.
                                        pdf.

Yamamura:1999:FCU

[1356] Akihiro Yamamura. A functional cryptosystem using a group action. Lecture Notes in Computer Science, 1587:314– 326, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1587/15870314.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1587/15870314.
                                        pdf.

Anonymous:1999:AIe

[1357] Anonymous. Author index. Lecture Notes in Computer Science, 1587:327– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/1587/1587auth.pdf.

Ancona:1999:AFS

[1358] D. Ancona. An algebraic framework for separate type-checking. Lecture Notes in Computer Science, 1589:1–15, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borzyszkowski:1999:MSS

[1359] T. Borzyszkowski. Moving specification structures between logical systems. Lecture Notes in Computer Science, 1589:16–30, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruni:1999:NFP

[1360] R. Bruni, F. Gadducci, and U. Montanari. Normal forms for partitions and relations. Lecture Notes in Computer Science, 1589:31–47, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caleiro:1999:PL

[1361] C. Caleiro, C. Sernadas, and A. Sernadas. Parameterisation of logics. Lecture Notes in Computer Science, 1589: 48–62, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cirstea:1999:SCH

[1362] C. Cirstea. Semantic constructions for hidden algebra. Lecture Notes in Computer Science, 1589:63–78, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corradini:1999:FSM

[1363] A. Corradini and F. Gadducci. Functorial semantics for multi-algebras. Lecture Notes in Computer Science, 1589: 79–91, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corradini:1999:AGD

[1364] A. Corradini, M. Grosse-Rhode, and R. Heckel. An algebra of graph derivations using finite (co-) limit double theories. Lecture Notes in Computer Science, 1589:92–106, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coudert:1999:HHS

[1365] S. Coudert, G. Bernot, and P. Le Gall. Hierarchical heterogeneous specifications. Lecture Notes in Computer Science, 1589:107–121, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Echahed:1999:PAG

[1366] R. Echahed and J.-C. Janodet. Parallel admissible graph rewriting. Lecture Notes in Computer Science, 1589:122– 137, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grosse-Rhode:1999:RMT

[1367] M. Grosse-Rhode, F. P. Presicce, and M. Simeoni. Refinements and modules for typed graph transformation systems. Lecture Notes in Computer Science, 1589:138–151, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habel:1999:CST

[1368] A. Habel and D. Plump. Complete strategies for term graph narrowing. Lecture Notes in Computer Science, 1589:152–167, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kirchner:1999:NDC

[1369] H. Kirchner and P.-E. Moreau. Nondeterministic computations in ELAN. Lecture Notes in Computer Science, 1589:168–182, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Konikowska:1999:RSD

[1370] B. Konikowska. Rasiowa-Sikorski deduction systems: A handy tool for computer science logics. Lecture Notes in Computer Science, 1589:183–197, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mossakowski:1999:TOC

[1371] T. Mossakowski. Translating OBJ3 into CASL: The institution level. Lecture Notes in Computer Science, 1589: 198–215, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mosses:1999:CGT

[1372] P. D. Mosses. CASL: A guided tour of its design. Lecture Notes in Computer Science, 1589:216–240, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Padberg:1999:APN

[1373] J. Padberg. Abstract Petri nets as a uniform approach to high-level Petri nets. Lecture Notes in Computer Science, 1589:241–260, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pita:1999:URS

[1374] I. Pita and N. Marti-Oliet. Using reflection to specify transaction sequences in rewriting logic. Lecture Notes in Computer Science, 1589:261–276, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poizat:1999:CDT

[1375] P. Poizat, C. Choppy, and J.-C. Royer.

Concurrency and data types: A specification method an example with LO-TOS. Lecture Notes in Computer Science, 1589:277–292, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramos:1999:SSC

[1376] J. Ramos and A. Sernadas. The situation and state calculus versus branching temporal logic. Lecture Notes in Computer Science, 1589:293–309, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Resende:1999:MSC

[1377] P. Resende. Modular specification of concurrent systems with observational logic. Lecture Notes in Computer Science, 1589:310–325, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wirsing:1999:PNS

[1378] M. Wirsing, J. N. Crossley, and H. Peterreins. Proof normalization of structured algebraic specifications is convergent. Lecture Notes in Computer Science, 1589:326–340, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shim:1999:UAC

[1379] J. Shim, P. Scheuermann, and R. Vingralek. A unified algorithm for cache replacement and consistency in Web proxy servers. Lecture Notes in Computer Science, 1590:1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Billard:1999:TSI

[1380] D. Billard. Transactional services for the Internet. Lecture Notes in Computer Science, 1590:14–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Connor:1999:UPP

[1381] R. Connor, K. Sibson, and P. Manghi.

On the unification of persistent programming and the World Wide Web. Lecture Notes in Computer Science, 1590:34–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goldman:1999:IQS

[1382] R. Goldman and J. Widom. Interactive query and search in semistructured databases. Lecture Notes in Computer Science, 1590:52–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Konopnicki:1999:BDF

[1383] D. Konopnicki and O. Shmueli. Bringing database functionality to the WWW. Lecture Notes in Computer Science, 1590:63–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bidoit:1999:FCQ

[1384] N. Bidoit and M. Ykhlef. Fixpoint calculus for querying semistructured data. Lecture Notes in Computer Science, 1590:78–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sindoni:1999:IMH

[1385] G. Sindoni. Incremental maintenance of hypertext views. Lecture Notes in Computer Science, 1590:98–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Simeon:1999:UYB

[1386] J. Simeon and S. Gluer. Using YAT to build a Web server. Lecture Notes in Computer Science, 1590:118–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Falquet:1999:LTS

[1387] G. Falquet, J. Guyot, and L. Nerima. Languages and tools to specify hypertext views on databases. Lecture Notes in Computer Science, 1590:136– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beeri:1999:WTS

[1388] C. Beeri, G. Elber, T. Milo, and Y. Sagiv. WebSuite — A tool suite for harnessing Web data. Lecture Notes in Computer Science, 1590:152–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brin:1999:EPR

[1389] S. Brin. Extracting patterns and relations from the World Wide Web. Lecture Notes in Computer Science, 1590: 172–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spiliopoulou:1999:WTW

[1390] M. Spiliopoulou and L. C. Faulstich. WUM: A tool for Web utilization analysis. Lecture Notes in Computer Science, 1590:184–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shivakumar:1999:FNR

[1391] N. Shivakumar and H. Garcia-Molina.

Finding near-replicas of documents on the Web. Lecture Notes in Computer Science, 1590:204–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duke:1999:CPS

[1392] David J. Duke, Ivan Herman, and M. Scott Marshall. Chapter 1. PREMO: A standard for distributed multimedia. Lecture Notes in Computer Science, 1591:1–??, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.com/link/
service/series/0558/bibs/1591/15910001.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1591/15910001.pdf.

Duke:1999:COP

[1393] David J. Duke, Ivan Herman, and M. Scott Marshall. Chapter 2. an overview of PREMO. Lecture Notes in Computer Science, 1591:9– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1591/15910009.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1591/15910009.
pdf.

Duke:1999:CFP

[1394] David J. Duke, Ivan Herman, and M. Scott Marshall. Chapter 3. the fundamentals of PREMO. Lecture Notes in Computer Science, 1591:35– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1591/15910035.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1591/15910035.
pdf.

Duke:1999:CGI

[1395] David J. Duke, Ivan Herman, and M. Scott Marshall. Chapter 4. general implementation issues. Lecture Notes in Computer Science, 1591:49– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1591/15910049.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1591/15910049.
pdf.

Duke:1999:CFC

[1396] David J. Duke, Ivan Herman, and M. Scott Marshall. Chapter 5. the foundation component. Lecture Notes in Computer Science, 1591:59– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1591/15910059.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1591/15910059.
pdf.

Duke:1999:CMS

[1397] David J. Duke, Ivan Herman, and M. Scott Marshall. Chapter 6. multimedia systems services component. Lecture Notes in Computer Science, 1591:125–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1591/15910125.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1591/15910125.
pdf.

Duke:1999:CMR

[1398] David J. Duke, Ivan Herman, and M. Scott Marshall. Chapter 7. the modelling, rendering, and interaction component. Lecture Notes in Computer Science, 1591:165–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1591/15910165.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1591/15910165.
pdf.

Duke:1999:CDJ

[1399] David J. Duke, Ivan Herman, and M. Scott Marshall. Chapter 8. detailed Java specifications of the PREMO objects. Lecture Notes in Computer Science, 1591:205–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1591/15910205.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1591/15910205.
pdf.

Duke:1999:ASI

[1400] David J. Duke, Ivan Herman, and M. Scott Marshall. Appendix. selected implementation issues. Lecture Notes in Computer Science, 1591:231– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1591/15910231.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1591/15910231.
                                        pdf.

Anonymous:1999:R

[1401] Anonymous. References. Lecture Notes in Computer Science, 1591:243– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/1591/1591refe.pdf.

Anonymous:1999:I

[1402] Anonymous. Index. Lecture Notes in Computer Science, 1591:251–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/1591/1591auth.pdf.

Boneh:1999:CRP

[1403] Dan Boneh and Glenn Durfee. Cryptanalysis of RSA with private key d less than N^{0.292}. Lecture Notes in Computer Science, 1592:1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1592/15920001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1592/15920001.
                                        pdf.

Biham:1999:CSR

[1404] Eli Biham, Alex Biryukov, and Adi Shamir. Cryptanalysis of Skipjack reduced to 31 rounds using impossible differentials. Lecture Notes in Computer Science, 1592:12–23, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1592/15920012.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1592/15920012.
pdf.

Nevelsteen:1999:SPU

[1405] Wim Nevelsteen and Bart Preneel. Software performance of universal hash functions. Lecture Notes in Computer Science, 1592:24–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1592/15920024.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1592/15920024.
pdf.

Dodis:1999:LBO

[1406] Yevgeniy Dodis and Silvio Micali. Lower bounds for oblivious transfer reductions. Lecture Notes in Computer Science, 1592:42–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1592/15920042.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1592/15920042.
pdf.

Damgaard:1999:IPB

[1407] Ivan Damgård, Joe Kilian, and Louis Salvail. On the (im)possibility of basing oblivious transfer and bit commitment on weakened security assumptions. Lecture Notes in Computer Science, 1592:56–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1592/15920056.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1592/15920056.
                                        pdf.

Crescenzo:1999:COT

[1408] Giovanni Di Crescenzo, Rafail Ostrovsky, and Sivaramakrishnan Rajagopalan. Conditional oblivious transfer and timed-release encryption. Lecture Notes in Computer Science, 1592: 74–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1592/15920074.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1592/15920074.
                                        pdf.

DiCrescenzo:1999:COT

[1409] G. Di Crescenzo, R. Ostrovsky, and S. Rajagopalan. Conditional oblivious transfer and timed-release encryption. Lecture Notes in Computer Science, 1592:74–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Canetti:1999:ETP

[1410] Ran Canetti and Shafi Goldwasser. An efficient Threshold public key cryptosystem secure against adaptive chosen ciphertext attack. Lecture Notes in Computer Science, 1592:90– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1592/15920090.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1592/15920090.
pdf.

Camenisch:1999:PZK

[1411] Jan Camenisch and Markus Michels.

Proving in zero-knowledge that a number is the product of two safe primes. Lecture Notes in Computer Science, 1592:107–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1592/15920107.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1592/15920107.
pdf.

Gennaro:1999:SHS

[1412] Rosario Gennaro, Shai Halevi, and Tal Rabin. Secure hash-and-sign signatures without the random oracle. Lecture Notes in Computer Science, 1592:123– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1592/15920123.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1592/15920123.
pdf.

Ergun:1999:NLC

[1413] Funda Ergun, Joe Kilian, and Ravi Kumar. A note on the limits of collusion-resistant watermarks. Lecture Notes in Computer Science, 1592:140– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1592/15920140.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1592/15920140.
                                        pdf.

Pfitzmann:1999:CBA

[1414] Birgit Pfitzmann and Ahmad-Reza Sadeghi. Coin-based anonymous fingerprinting. Lecture Notes in Computer Science, 1592:150–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1592/15920150.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1592/15920150.pdf.

Smart:1999:PHC

[1415] Nigel P. Smart. On the performance of hyperelliptic cryptosystems. Lecture Notes in Computer Science, 1592:165– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1592/15920165.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1592/15920165.
                                        pdf.

Kobayashi:1999:FEC

[1416] Tetsutaro Kobayashi, Hikaru Morita, Kunio Kobayashi, and Fumitaka Hoshino. Fast elliptic curve algorithm combining Frobenius map and table reference to adapt to higher characteristic. Lecture Notes in Computer Science, 1592:176–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1592/15920176.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1592/15920176.
pdf.

Harasawa:1999:CMF

[1417] Ryuichi Harasawa, Junji Shikata, Joe Suzuki, and Hideki Imai. Comparing the MOV and FR reductions in elliptic curve cryptography. Lecture Notes in Computer Science, 1592:190– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1592/15920190.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1592/15920190.
pdf.

Kipnis:1999:UOV

[1418] Aviad Kipnis, Jacques Patarin, and Louis Goubin. Unbalanced oil and vinegar signature schemes. Lecture Notes in Computer Science, 1592:206– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1592/15920206.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1592/15920206.
pdf.

Paillier:1999:PKC

[1419] Pascal Paillier. Public-key cryptosystems based on composite degree residuosity classes. Lecture Notes in Computer Science, 1592:223–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1592/15920223.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1592/15920223.
                                        pdf.

Pointcheval:1999:NPK

[1420] David Pointcheval. New public key cryptosystems based on the dependent — RSA problems. Lecture Notes in Computer Science, 1592:239– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1592/15920239.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1592/15920239.
                                        pdf.

Vaudenay:1999:RAG

[1421] Serge Vaudenay. Resistance against general iterated attacks. Lecture Notes in Computer Science, 1592:255– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1592/15920255.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1592/15920255.
                                        pdf.

Hawkes:1999:XNX

[1422] Philip Hawkes and Luke O’Connor. XOR and non-XOR differential probabilities. Lecture Notes in Computer Science, 1592:272–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/

series/0558/bibs/1592/15920272.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1592/15920272.pdf.

Cheon:1999:BCN

[1423] Jung Hee Cheon, Seongtaek Chee, and Choonsik Park. S-boxes with controllable nonlinearity. Lecture Notes in Computer Science, 1592:286–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1592/15920286.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1592/15920286.
pdf.

Gennaro:1999:SDK

[1424] Rosario Gennaro, Stanisław Jarecki, Hugo Krawczyk, and Tal Rabin. Secure distributed key generation for discrete-log based cryptosystems. Lecture Notes in Computer Science, 1592: 295–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1592/15920295.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1592/15920295.
pdf.

Cramer:1999:EMC

[1425] Ronald Cramer, Ivan Damgård, Stefan Dziembowski, Martin Hirt, and Tal Rabin. Efficient multiparty computations secure against an adaptive adversary. Lecture Notes in Computer Science, 1592:311–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1592/15920311.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1592/15920311.
                                        pdf.

Naor:1999:DPR

[1426] Moni Naor, Benny Pinkas, and Omer Reingold. Distributed pseudo-random functions and KDCs. Lecture Notes in Computer Science, 1592:327–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1592/15920327.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1592/15920327.
                                        pdf.

Johansson:1999:IFC

[1427] Thomas Johansson and Fredrik Jönsson. Improved fast correlation attacks on stream ciphers via convolutional codes. Lecture Notes in Computer Science, 1592:347–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1592/15920347.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1592/15920347.
                                        pdf.

Knudsen:1999:CIS

[1428] Lars R. Knudsen and Willi Meier. Cryptanalysis of an identification scheme based on the permuted perceptron problem. Lecture Notes in Computer Science, 1592:363–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1592/15920363.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1592/15920363.
pdf.

OConnor:1999:AEB

[1429] Luke O’Connor. An analysis of exponentiation based on formal languages. Lecture Notes in Computer Science, 1592:375–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1592/15920375.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1592/15920375.
pdf.

Mizuki:1999:DNS

[1430] Takaaki Mizuki, Hiroki Shizuya, and Takao Nishizeki. Dealing necessary and sufficient numbers of cards for sharing a one-bit secret key. Lecture Notes in Computer Science, 1592:389– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1592/15920389.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1592/15920389.
pdf.

Cachin:1999:CPI

[1431] Christian Cachin, Silvio Micali, and Markus Stadler. Computationally private information retrieval with polylogarithmic communication. Lecture Notes in Computer Science, 1592:402– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1592/15920402.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1592/15920402.
                                        pdf.

Richardson:1999:CCZ

[1432] Ransom Richardson and Joe Kilian. On the concurrent composition of zero-knowledge proofs. Lecture Notes in Computer Science, 1592:415– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1592/15920415.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1592/15920415.
                                        pdf.

Fischlin:1999:PFT

[1433] Marc Fischlin. Pseudorandom function tribe ensembles based on oneway permutations: Improvements and applications. Lecture Notes in Computer Science, 1592:432–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1592/15920432.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1592/15920432.
                                        pdf.

Wang:1999:SCB

[1434] Yongge Wang and Yvo Desmedt. Secure communication in broadcast channels: The answer to Franklin and Wright’s question. Lecture Notes in Computer Science, 1592:446–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1592/15920446.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1592/15920446.
pdf.

Canetti:1999:ECS

[1435] Ran Canetti, Tal Malkin, and Kobbi Nissim. Efficient communicationstorage tradeoffs for multicast encryption. Lecture Notes in Computer Science, 1592:459–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1592/15920459.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1592/15920459.pdf.

Sanders:1999:HPI

[1436] P. Sanders, T. Takkula, and D. Wedelin.

High performance integer optimization for crew scheduling. Lecture Notes in Computer Science, 1593:3–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jung:1999:SSP

[1437] B. Jung, H.-P. Lenhof, P. Mueller, and C. Rueb. Simulating synthetic polymer chains in parallel. Lecture Notes in Computer Science, 1593:13–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reijns:1999:RTS

[1438] G. L. Reijns, A. J. C. Van Gemund, J. Schier, and P. J. F. Swart. Real-time signal processing in a collision avoidance radar system using parallel computing. Lecture Notes in Computer Science, 1593:23–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Riedl:1999:IWS

[1439] R. Riedl. The impact of workload on simulation results for distributed transaction processing. Lecture Notes in Computer Science, 1593:33–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maksymowicz:1999:CSA

[1440] A. Z. Maksymowicz, M. Bubak, K. Zajac, and M. Magdon. Computer simulation of ageing with an extended penna model. Lecture Notes in Computer Science, 1593:43–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanSterkenburg:1999:SMT

[1441] R. P. Van Sterkenburg and A. A. Ten Dam. The scenario management tool SMARTFED for real-time interactive high performance networked simulations. Lecture Notes in Computer Science, 1593:50–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lebas:1999:OHA

[1442] F.-X. Lebas and T. Uslaender. OPERA: An HPCN architecture for distributed component-based real-time simulations. Lecture Notes in Computer Science, 1593:60–??, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rackl:1999:ASU

[1443] G. Rackl, F. De Stefani, F. Heran, and A. Pasquarelli. Airport simulation using CORBA and DIS. Lecture Notes in Computer Science, 1593:70–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chang:1999:IRG

[1444] C.-H. Chang, H.-K. Wu, M.-H. Jin, and Y.-O. Tseng. Intelligent routing for global broadband satellite Internet. Lecture Notes in Computer Science, 1593:80–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kellar:1999:ICC

[1445] W. P. Kellar, A. M. Savill, and W. N.

Dawes. Integrated CAD/CFD visualisation of a generic formula I car front wheel flowfield. Lecture Notes in Computer Science, 1593:90–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fritzson:1999:ASS

[1446] D. Fritzson and P. Nordling. Adaptive scheduling strategy optimizer for parallel rolling bearing simulation. Lecture Notes in Computer Science, 1593: 99–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Radhakrishna:1999:MBP

[1447] H. Radhakrishna, S. Divakar, N. Magotra, and S. R. J. Brueck. MPIbased parallel implementation of a lithography pattern simulation algorithm. Lecture Notes in Computer Science, 1593:109–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schuele:1999:PHR

[1448] J. Schuele and T. Wilhelmsson. Parallelizing an high resolution operational ocean model. Lecture Notes in Computer Science, 1593:120–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roads:1999:WCF

[1449] J. Roads, S. Chen, C. McCord, and W. Smith. Weather and climate forecasts and analyses at MHPCC. Lecture Notes in Computer Science, 1593: 130–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garatani:1999:GHP

[1450] K. Garatani, H. Nakamura, H. Okuda, and G. Yagawa. GeoFEM: High performance parallel FEM for solid earth. Lecture Notes in Computer Science, 1593:133–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Modersitzki:1999:EMV

[1451] J. Modersitzki, W. Obeloeer, O. Schmitt, and G. Lustig. Elastic matching of very large digital images on high performance clusters. Lecture Notes in Computer Science, 1593:141–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1999:DID

[1452] J. Lee, B. Tierney, and W. Johnston. Data intensive distributed computing: A medical application example. Lecture Notes in Computer Science, 1593:150–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papazis:1999:UHT

[1453] N. Papazis and D. Dimitrelos. Utilizing HPC technology in 3D cardiac modeling. Lecture Notes in Computer Science, 1593:159–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rivas:1999:PAR

[1454] R. Rivas, M. B. Ibañez, Y. Cardinale, and P. Windyga. A parallel algorithm for 3D reconstruction of angiographic images. Lecture Notes in Computer Science, 1593:168–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maniatis:1999:DTM

[1455] T. A. Maniatis, K. S. Nikita, and K. Voliotis. A diffraction tomography method for medical imaging implemented on high performance computing environment. Lecture Notes in Computer Science, 1593:178–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kalinov:1999:HDC

[1456] A. Kalinov and A. Lastovetsky. Heterogeneous distribution of computations while solving linear algebra problems on networks of heterogeneous computers. Lecture Notes in Computer Science, 1593:191–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heras:1999:MIL

[1457] D. B. Heras, V. B. Perez, J. C. C. Dominguez, and F. F. Rivera. Modeling and improving locality for irregular problems: Sparse matrix-vector product on cache memories as a case study. Lecture Notes in Computer Science, 1593:201–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Satoh:1999:PSC

[1458] S. Satoh, K. Kusano, Y. Tanaka, and M. Matsuda. Parallelization of sparse Cholesky factorization on an SMP cluster. Lecture Notes in Computer Science, 1593:211–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schenk:1999:SPS

[1459] O. Schenk, K. Gaertner, and W. Fichtner. Scalable parallel sparse factorization with left-right looking strategy on shared memory multiprocessors. Lecture Notes in Computer Science, 1593:221–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bogdanov:1999:ATL

[1460] A. V. Bogdanov, I. A. Grishin, G. O. Khanlarov, and G. A. Lukianov. Algorithm of two-level parallelization for direct simulation Monte Carlo of unsteady flows in molecular gasdynamics. Lecture Notes in Computer Science, 1593:231–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Briguglio:1999:PGF

[1461] S. Briguglio, G. Vlad, G. Fogaccia, and B. Di Martino. Parallelization of gridless finite-size-particle plasma simulation codes. Lecture Notes in Computer Science, 1593:241–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sun:1999:DOO

[1462] Y. Sun, Z. Liang, and C.-L. Wang. A distributed object-oriented method for particle simulations on clusters. Lecture Notes in Computer Science, 1593: 251–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sato:1999:SDL

[1463] N. Sato and J.-M Jezequel. A simple dynamic load-balancing scheme for parallel molecular dynamics simulation on distributed memory machines. Lecture Notes in Computer Science, 1593: 260–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keller:1999:RMH

[1464] A. Keller, M. Brune, and A. Reinefeld. Resource management for highperformance PC clusters. Lecture Notes in Computer Science, 1593:270–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mahmoud:1999:WGC

[1465] Q. H. Mahmoud. The Web as a global computing platform. Lecture Notes in Computer Science, 1593:281–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haupt:1999:WFW

[1466] T. Haupt, E. Akarsu, and G. Fox. WebFlow: A framework for Web based metacomputing. Lecture Notes in Computer Science, 1593:291–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanAlbada:1999:DBO

[1467] G. D. Van Albada, J. Clinckemaillie, A. H. L. Emmen, and J. Gehring. Dynamite-blasting obstacles to parallel cluster computing. Lecture Notes in Computer Science, 1593:300–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kandhai:1999:IMR

[1468] D. Kandhai, A. Koponen, A. Hoekstra, and P. M. A. Sloot. Iterative momentum relaxation for fast lattice-Boltzmann simulations. Lecture Notes in Computer Science, 1593:311–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dupuis:1999:LGE

[1469] A. Dupuis and B. Chopard. Lattice gas: An efficient and reusable parallel library based on a graph partitioning technique. Lecture Notes in Computer Science, 1593:319–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kraeva:1999:APR

[1470] M. A. Kraeva and V. E. Malyshkin. Algorithms of parallel realisation of the PIC method with assembly technology. Lecture Notes in Computer Science, 1593:329–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dubbeldam:1999:CAM

[1471] D. Dubbeldam, A. G. Hoekstra, and P. M. A. Sloot. Computational aspects of multi-species lattice-gas automata. Lecture Notes in Computer Science, 1593:339–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Patten:1999:TSM

[1472] C. J. Patten, K. A. Hawick, and J. F.

Hercus. Towards a scalable metacomputing storage service. Lecture Notes in Computer Science, 1593:350–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1999:DVC

[1473] J. Li and E. De Doncker. Dynamic visualization of computations on the Internet. Lecture Notes in Computer Science, 1593:360–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferrari:1999:FSS

[1474] A. Ferrari, F. Knabe, M. Humphrey, and S. Chapin. A flexible security system for metacomputing environments. Lecture Notes in Computer Science, 1593:370–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Malony:1999:CEU

[1475] A. D. Malony, J. L. Skidmore, and M. J. Sottile. Computational experiments using distributed tools in a Web-based electronic notebook environment. Lecture Notes in Computer Science, 1593:381–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Casavant:1999:PDA

[1476] T. L. Casavant, T. E. Scheetz, T. A. Braun, and K. J. Munn. A parallel/ distributed architecture for hierarchically heterogeneous Web-based cooperative applications. Lecture Notes in Computer Science, 1593:391–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1999:EDL

[1477] D. A. Smith. Effective dynamic load balancing of the UKMO tracer advection routines. Lecture Notes in Computer Science, 1593:402–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schoneveld:1999:DLB

[1478] A. Schoneveld, M. Lees, E. Karyadi, and P. M. A. Sloot. Dynamic load balancing in parallel finite element simulations. Lecture Notes in Computer Science, 1593:409–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ford:1999:LBR

[1479] R. W. Ford and M. O’Brien. A load balancing routine for the NAG parallel library. Lecture Notes in Computer Science, 1593:420–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Munz:1999:PAP

[1480] F. Munz, T. Ludwig, S. Ziegler, and P. Bartenstein. Performance assessment of parallel spectral analysis: Towards a practical performance model for parallel medical applications. Lecture Notes in Computer Science, 1593:

430–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yu:1999:PAP

[1481] S. Yu, M. Clement, Q. Snell, and B. Morse. Parallel algorithm and processor selection based on fuzzy logic. Lecture Notes in Computer Science, 1593:440–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kechadi:1999:RNN

[1482] M.-T. Kechadi. Recurrent neural network approach for partitioning irregular graphs. Lecture Notes in Computer Science, 1593:450–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hu:1999:JSD

[1483] W. Hu, W. Shi, and Z. Tang. JIA-JIA: A software DSM system based on a new cache coherence protocol. Lecture Notes in Computer Science, 1593: 463–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harper:1999:EAM

[1484] J. S. Harper, D. J. Kerbyson, and G. R. Nudd. Efficient analytical modelling of multi-level set-associative caches. Lecture Notes in Computer Science, 1593:473–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kotapati:1999:BMW

[1485] K. Kotapati and S. P. Dandamudi.

Buffer management in wormholerouted torus multicomputer networks.

Lecture Notes in Computer Science, 1593:483–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pan:1999:PAB

[1486] K.-H. Pan, H.-K. Wu, R.-J. Shang, and F. Lai. Performance analysis of broadcast in synchronized multihop wireless networks. Lecture Notes in Computer Science, 1593:493–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolf:1999:EPE

[1487] F. Wolf and B. Mohr. EARL — A programmable and extensible toolkit for analyzing event traces of message passing programs. Lecture Notes in Computer Science, 1593:503–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blackburn:1999:XES

[1488] K. Blackburn, A. Lazzarini, T. Prince, and R. Williams. XSIL: Extensible scientific interchange language. Lecture Notes in Computer Science, 1593:513– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kessler:1999:FCS

[1489] C. W. Kessler and H. Seidl. ForkLight: A control-synchronous parallel programming language. Lecture Notes in Computer Science, 1593:525–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiMartino:1999:HPM

[1490] B. Di Martino, M. Celino, and V. Rosato. HPF parallelization of a molecular dynamics code: Strategies and performances. Lecture Notes in Computer Science, 1593:535–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bubak:1999:DHP

[1491] M. Bubak, J. T. Moscicki, and J. Shiers. Design of high-performance C++ package for handling of multidimensional histograms. Lecture Notes in Computer Science, 1593:543–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oguchi:1999:DRM

[1492] M. Oguchi and M. Kitsuregawa. Dynamic remote memory acquiring for parallel data mining on PC cluster: Preliminary performance results. Lecture Notes in Computer Science, 1593: 553–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aloisio:1999:DPP

[1493] G. Aloisio, M. Cafaro, and R. Williams.

The digital puglia project: An active digital library of remote sensing data. Lecture Notes in Computer Science, 1593:563–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chattratichat:1999:ADE

[1494] J. Chattratichat, J. Darlington, Y. Guo, and S. Hedvall. An architecture for distributed enterprise data mining. Lecture Notes in Computer Science, 1593:573–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1999:RSM

[1495] S. Li and A. L. Ananda. Representatives selection in multicast group. Lecture Notes in Computer Science, 1593: 583–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zambonelli:1999:DPI

[1496] F. Zambonelli. Deadlock prevention in incremental replay of message-passing programs. Lecture Notes in Computer Science, 1593:593–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hobbs:1999:RCP

[1497] M. Hobbs and A. Goscinski. Remote and concurrent process duplication for SPMD based parallel processing on COWs. Lecture Notes in Computer Science, 1593:603–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zavenella:1999:UBO

[1498] A. Zavenella and S. Pelagatti. Using BSP to optimize data distribution in skeleton programs. Lecture Notes in Computer Science, 1593:613–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brauss:1999:STC

[1499] S. Brauss, M. Frey, A. Gunzinger, and M. Lienhard. Swiss-Tx communication libraries. Lecture Notes in Computer Science, 1593:623–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zingirian:1999:FOU

[1500] N. Zingirian and M. Maresca. Finding the optimal unroll-and-jam. Lecture Notes in Computer Science, 1593: 633–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cilio:1999:LEW

[1501] A. G. M. Cilio and H. Corporaal. A linker for effective whole-program optimizations. Lecture Notes in Computer Science, 1593:643–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Silber:1999:NLT

[1502] G.-A. Silber and A. Darte. The Nestor library: A tool for implementing Fortran source to source transformations. Lecture Notes in Computer Science, 1593:653–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takabatake:1999:PMS

[1503] M. Takabatake, H. Honda, and T. Yuba. Performance measurements on sandglass-type parallelization of doacross loops. Lecture Notes in Computer Science, 1593:663–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boekhold:1999:TPA

[1504] M. Boekhold, I. Karkowski, and H. Corporaal. Transforming and parallelizing ANSI C programs using pattern recognition. Lecture Notes in Computer Science, 1593:673–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zeng:1999:CAP

[1505] S. Zeng and S. P. Dandamudi. Centralized architecture for parallel query processing on networks of workstations. Lecture Notes in Computer Science, 1593:683–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kitowski:1999:OOD

[1506] J. Kitowski, D. Wajs, and P. Trzeciak. Object-oriented database system for large-scale molecular dynamics simulations. Lecture Notes in Computer Science, 1593:693–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deshpande:1999:VEM

[1507] V. Deshpande, L. Fornasier, E. A. Gerteisen, and N. Hilbrink. Virtual engineering of multi-disciplinary applications and the significance of seamless accessibility of geometry data. Lecture Notes in Computer Science, 1593:702– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tomov:1999:SRN

[1508] N. Tomov, E. Dempster, M. H. Williams, and A. Burger. Some results from a new technique for response time estimation in parallel DBMS. Lecture Notes in Computer Science, 1593:713– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vinter:1999:PDS

[1509] B. Vinter, O. J. Anshus, and T. Larsen. PastSet — A distributed structured shared memory system. Lecture Notes in Computer Science, 1593:722–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Piriyakurnar:1999:OSI

[1510] D. A. L. Piriyakurnar, P. Levi, and C. S. R. Murthy. Optimal scheduling of iterative data-flow programs onto multiprocessors with non-negligible interprocessor communication. Lecture Notes in Computer Science, 1593:732– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baude:1999:OCC

[1511] F. Baude, D. Caromel, N. Furmento, and D. Sagnol. Overlapping communication with computation in distributed object systems. Lecture Notes in Computer Science, 1593:744–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marcuello:1999:EST

[1512] P. Marcuello and A. Gonzalez. Exploiting speculative thread-level parallelism on a SMT processor. Lecture Notes in Computer Science, 1593:754–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matsuda:1999:NIA

[1513] M. Matsuda, Y. Tanaka, K. Kubota, and M. Sato. Network interface active messages for low overhead communication on SMP PC clusters. Lecture Notes in Computer Science, 1593:764–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bernaschi:1999:ERA

[1514] M. Bernaschi, G. Iannello, and M. Lauria. Experimental results about MPI collective communication operations. Lecture Notes in Computer Science, 1593:774–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dimitrelos:1999:MSD

[1515] D. Dimitrelos and C. Halatsis. MaD-CoWS: A scalable distributed shared memory environment for massively parallel multiprocessors. Lecture Notes in Computer Science, 1593:784–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cleary:1999:VGV

[1516] D. Cleary and D. O’Donoghue. Visual-Expresso: Generating a virtual reality Internet. Lecture Notes in Computer Science, 1593:797–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boyd:1999:VUC

[1517] D. R. S. Boyd, J. R. Gallop, K. E. V. Pahnen, and R. T. Platon. VIVRE: User-centred visualization. Lecture Notes in Computer Science, 1593:807– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Belleman:1999:GGP

[1518] R. G. Belleman, J. A. Kaandorp, D. Dijkman, and P. M. A. Sloot. GEO-PROVE: Geometric probes for virtual environments. Lecture Notes in Computer Science, 1593:817–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moreira:1999:GSS

[1519] J. E. Moreira, H. Franke, W. Chan, and L. L. Fong. A gang-scheduling system for ASCI Blue-Pacific. Lecture Notes in Computer Science, 1593:831–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karl:1999:TQS

[1520] H. Karl. Towards quality of service for parallel computing: An overview of the MILAN project. Lecture Notes in Computer Science, 1593:841–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schwiegelshohn:1999:RAS

[1521] U. Schwiegelshohn and R. Yahyapour.

Resource allocation and scheduling in metasystems. Lecture Notes in Computer Science, 1593:851–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Walker:1999:UJH

[1522] D. Walker and O. Rana. The use of Java in high performance computing: A data mining example. Lecture Notes in Computer Science, 1593:863– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coddington:1999:IIR

[1523] P. D. Coddington, J. A. Mathew, and K. A. Hawick. Interfaces and implementations of random number generators for Java Grande applications. Lecture Notes in Computer Science, 1593:873–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gimbel:1999:JBP

[1524] M. Gimbel, M. Philippsen, B. Haumacher, and P. C. Lockemann. Java as a basis for parallel data mining in workstation clusters. Lecture Notes in Computer Science, 1593:884–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krall:1999:GCL

[1525] A. Krall and P. Tomsich. Garbage collection for large memory Java applications. Lecture Notes in Computer Science, 1593:895–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marsh:1999:EVM

[1526] A. Marsh, T. Kauranne, G. Zahlmann, and A. Emmen. The emergence of virtual medical worlds. Lecture Notes in Computer Science, 1593:909–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moore:1999:CUM

[1527] M. Moore. Characteristics of users of medical innovations. Lecture Notes in Computer Science, 1593:912–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blobel:1999:SAD

[1528] B. Blobel, P. Pharow, and F. Roger-France. Security analysis and design based on a general conceptual security model and UML. Lecture Notes in Computer Science, 1593:919–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1999:IDA

[1529] Anonymous. 3DHeartView: Introducing 3-dimensional angiographical modelling. Lecture Notes in Computer Science, 1593:931–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohlsson:1999:WBS

[1530] M. Ohlsson, A. Jaerund, and L. Edenbrandt. WWW based service for automated interpretation of diagnostic images: The AIDI-Heart project. Lecture Notes in Computer Science, 1593:941– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kokol:1999:DTC

[1531] P. Kokol, M. Zorman, K. Podgorelec, and A. Habjanic. Decision trees — A CIM tool in nursing education. Lecture Notes in Computer Science, 1593:951–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Samiotakis:1999:HII

[1532] Y. Samiotakis, S. Anagnostopoulou, and A. Alexakis. HealthLine: Integrated information provision to telemedicine networks. Lecture Notes in Computer Science, 1593:959–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jovahov:1999:MMP

[1533] E. Jovahov, D. Starcevic, A. Marsh, and Z. Obrenovic. Multi modal presentation in virtual telemedical environments. Lecture Notes in Computer Science, 1593:964–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stamatakos:1999:UWT

[1534] G. S. Stamatakos, E. I. Zacharaki, N. A. Mouravliansky, and K. K. Delibasis. Using Web technologies and metacomputing to visualise a simplified simulation model of tumor growth in vitro. Lecture Notes in Computer Science, 1593:973–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Polemi:1999:ECC

[1535] D. Polemi and A. Marsh. The electronic commerce component in telemedicine. Lecture Notes in Computer Science, 1593:983–??, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Delibasis:1999:EIM

[1536] K. K. Delibasis, G. K. Matsopoulos, N. A. Mouravliansky, and K. S. Nikita. Efficient implementation of the marching cubes algorithm for rendering medical data. Lecture Notes in Computer Science, 1593:989–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chan:1999:MAE

[1537] T. F. Chan and P. Vanek. Multilevel algebraic elliptic solvers. Lecture Notes in Computer Science, 1593:1001–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rokicki:1999:PPC

[1538] J. Rokicki, D. Drikakis, J. Majewski, and J. Zoltak. Parallel performance of chimera overlapping mesh technique. Lecture Notes in Computer Science, 1593:1015–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacques:1999:ESB

[1539] T. Jacques, L. Nicolas, and C. Voilaire.

Electromagnetic scattering with the boundary integral method on MIMD systems. Lecture Notes in Computer Science, 1593:1025–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roest:1999:DCN

[1540] M. R. T. Roest and E. A. H. Vollebregt. Decomposition of complex numerical software into cooperating components. Lecture Notes in Computer Science, 1593:1032–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baird:1999:MBP

[1541] S. Baird and J. J. McGuirk. Multiblock parallel simulation of fluid flow in a fuel cell. Lecture Notes in Computer Science, 1593:1042–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vuik:1999:PIB

[1542] C. Vuik and J. Frank. A parallel implementation of the block preconditioned GCR method. Lecture Notes in Computer Science, 1593:1052–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brok:1999:CTP

[1543] S. W. Brok and L. Dekker. Comparison of two parallel analytic simulation models of inhomogeneous distributed parameter systems. Lecture Notes in Computer Science, 1593:1061–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cooper:1999:CSF

[1544] T. Cooper. Case studies of four industrial meta-applications. Lecture Notes in Computer Science, 1593:1077–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aretzaz:1999:PAS

[1545] M. Aretzaz, R. Doallo, J. Touriño, and C. Vazquez. A parallel approach for solving a lubrication problem in industrial devices. Lecture Notes in Computer Science, 1593:1087–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kandemir:1999:RIC

[1546] M. Kandemir, A. Choudhary, and J. Ramanujam. Restructuring I/Ointensive computations for locality. Lecture Notes in Computer Science, 1593:1097–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caron:1999:VMM

[1547] E. Caron, O. Cozette, D. Lazure, and G. Utard. Virtual memory management in data parallel applications. Lecture Notes in Computer Science, 1593: 1107–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

No:1999:HPP

[1548] J. No, J. Carretero, and A. Choudhary. High performance parallel I/O schemes for irregular applications on clusters of workstations. Lecture Notes in Computer Science, 1593:1117–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brezany:1999:ADR

[1549] P. Brezany and M. Winslett. Advanced data repository support for Java scientific programming. Lecture Notes in Computer Science, 1593:1127–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agrawal:1999:ACO

[1550] G. Agrawal. Advanced communication optimizations for data-parallel programs. Lecture Notes in Computer Science, 1593:1139–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Telford:1999:CAS

[1551] S. O. Telford. A cellular automata simulation environment for modelling soil bioremediation. Lecture Notes in Computer Science, 1593:1143–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shi:1999:HEP

[1552] W. Shi, J. Ma, and Z. Tang. High efficient parallel computation of resonant frequencies of waveguide loaded cavities on JIAJIA software DSMs. Lecture Notes in Computer Science, 1593:1147– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iikura:1999:SPI

[1553] M. Iikura, K. Kobayashi, T. Yoshioka, and S. Ihara. A study of parallel image processing in a distributed processing environment. Lecture Notes in Computer Science, 1593:1151–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manoharan:1999:DPD

[1554] S. Manoharan. Data prefetching for digital alpha. Lecture Notes in Computer Science, 1593:1155–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1999:CSB

[1555] S.-Y. Wang, Y.-C. Tseng, S.-Y. Ni, and J.-P Sheu. Circuit-switched broadcast in multi-port 2D tori. Lecture Notes in Computer Science, 1593:1159–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koranne:1999:DAE

[1556] S. Koranne. A distributed algorithm for the estimation of average switching activity in combinational circuits. Lecture Notes in Computer Science, 1593:1163–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ranawake:1999:BTM

[1557] U. A. Ranawake and J. E. Dorband. BVIEW: A tool for monitoring distributed systems. Lecture Notes in Computer Science, 1593:1167–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Girona:1999:QSW

[1558] S. Girona, S. Bello, J. Labarta, and P. Ribes. The queue system within PHASE. Lecture Notes in Computer Science, 1593:1171–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kasyanov:1999:STS

[1559] V. N. Kasyanov, V. A. Evstigneev, J. V. Malinina, and J. V. Birjukova. Support tools for supercomputing and networking. Lecture Notes in Computer Science, 1593:1175–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeNicola:1999:CAC

[1560] R. De Nicola. Coordination and access control of mobile agents. Lecture Notes in Computer Science, 1594:1–2, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lange:1999:CAS

[1561] D. B. Lange. Characteristics of an agent scripting language and its execution environment. Lecture Notes in Computer Science, 1594:3–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bryce:1999:CMA

[1562] C. Bryce, M. Oriol, and J. Vitek. A coordination model for agents based on secure spaces. Lecture Notes in Computer Science, 1594:4–20, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wood:1999:CA

[1563] A. M. Wood. Coordination with attributes. Lecture Notes in Computer Science, 1594:21–36, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mascolo:1999:MSL

[1564] C. Mascolo. MobiS: A specification language for mobile systems. Lecture Notes in Computer Science, 1594:37– 52, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Murillo:1999:CRP

[1565] J. M. Murillo, J. Hernandez, F. Sanchez, and L. A. Alvarez. Coordinated roles: Promoting re-usability of coordinated active objects using event notification protocols. Lecture Notes in Computer Science, 1594:53–68, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weichert:1999:PMS

[1566] M. Weichert. Pipelining the molecule soup: A plumber’s approach to gamma. Lecture Notes in Computer Science, 1594:69–84, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moran:1999:EFS

[1567] A. Moran, D. Sands, and M. Carlsson. Erratic fudgets: A semantic theory for an embedded coordination language. Lecture Notes in Computer Science, 1594:85–102, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Budde:1999:CSP

[1568] R. Budde, G. M. Pinna, and A. Poigne. Coordination of synchronous programs. Lecture Notes in Computer Science, 1594:103–117, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Montangero:1999:CSC

[1569] C. Montangero and L. Semini. Composing specifications for coordination. Lecture Notes in Computer Science, 1594:118–133, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brogi:1999:ECM

[1570] A. Brogi and J.-M. Jacquet. On the expressiveness of coordination models. Lecture Notes in Computer Science, 1594:134–149, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonsangue:1999:CSA

[1571] M. M. Bonsangue, J. N. Kok, and G. Zavattaro. Comparing software architectures for coordination languages. Lecture Notes in Computer Science, 1594:150–165, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Varela:1999:HMC

[1572] C. Varela and G. Agha. A hierarchical model of coordination of concurrent activities. Lecture Notes in Computer Science, 1594:166–182, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tschudin:1999:SDE

[1573] C. F. Tschudin. A self-deploying election service for active networks. Lecture Notes in Computer Science, 1594:183– 195, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rowstron:1999:MCO

[1574] A. Rowstron. Mobile co-ordination: Providing fault tolerance in tuple space based co-ordination languages. Lecture Notes in Computer Science, 1594:196– 210, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sekiguchi:1999:SEJ

[1575] T. Sekiguchi, H. Masuhara, and A. Yonezawa. A simple extension of Java language for controllable transparent migration and its portable implementation. Lecture Notes in Computer Science, 1594:211–226, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petre:1999:CAM

[1576] L. Petre and K. Sere. Coordination among mobile objects. Lecture Notes in Computer Science, 1594:227–242, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scutella:1999:SCM

[1577] A. Scutella. Simulation of conference management using an event-driven coordination language. Lecture Notes in Computer Science, 1594:243–258, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rossi:1999:IBC

[1578] D. Rossi and F. Vitali. Internetbased coordination environments and document-based applications: A case study. Lecture Notes in Computer Science, 1594:259–274, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Everaars:1999:CPP

[1579] C. T. H. Everaars and B. Lisser. Coordination of a parallel proposition solver. Lecture Notes in Computer Science, 1594:275–290, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sample:1999:CCL

[1580] N. Sample, D. Beringer, L. Melloul, and G. Wiederhold. CLAM: Composition language for autonomous megamodules. Lecture Notes in Computer Science, 1594:291–306, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Podorozhny:1999:MRA

[1581] R. M. Podorozhny, B. Staudt Lerner, and L. J. Osterweil. Modeling resources for activity coordination and scheduling. Lecture Notes in Computer Science, 1594:307–322, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Attanasio:1999:SAR

[1582] C. Attanasio, F. Corradini, and P. Inverardi. Static analysis of real-time component-based systems configurations. Lecture Notes in Computer Science, 1594:323–339, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garlan:1999:ABS

[1583] D. Garlan and Z. Wang. Acme-based software architecture interchange. Lecture Notes in Computer Science, 1594: 340–354, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cruz:1999:GBA

[1584] J. C. Cruz and S. Ducasse. A group based approach for coordinating active objects. Lecture Notes in Computer Science, 1594:355–370, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Assmann:1999:ICC

[1585] U. Assmann and A. Ludwig. Introducing connections into classes with static meta-programming. Lecture Notes in Computer Science, 1594:371–383, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jamison:1999:TAC

[1586] W. C. Jamison and D. Lea. TRUCE: Agent coordination through concurrent interpretation of role-based protocols. Lecture Notes in Computer Science, 1594:384–398, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schumacher:1999:SCL

[1587] M. Schumacher, F. Chantemargue, and B. Hirsbrunner. The STL++ coordination language: A base for implementing distributed multi-agent applications. Lecture Notes in Computer Science, 1594:399–414, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buffo:1999:DSI

[1588] M. Buffo and D. Buchs. A distributed semantics for a IWIM-based coordination language. Lecture Notes in Computer Science, 1594:415–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cremonini:1999:CCA

[1589] M. Cremonini, A. Omicini, and F. Zambonelli. Coordination in context: Authentication, authorisation and topology in mobile agent applications. Lecture Notes in Computer Science, 1594:416–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Day:1999:PIM

[1590] M. Day. Presence and instant messaging via HTTP/1.1: A coordination perspective. Lecture Notes in Computer Science, 1594:417–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hirsch:1999:TPT

[1591] D. Hirsch, S. Uchitel, and D. Yankelevich. Towards a periodic table of connectors. Lecture Notes in Computer Science, 1594:418, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Assmann:1999:PRI

[1592] C. Assmann. Performance results for an implementation of the process coordination language K2. Lecture Notes in Computer Science, 1595:1–19, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Charles:1999:IAP

[1593] N. Charles and C. Runciman. An interactive approach to profiling parallel functional programs. Lecture Notes in Computer Science, 1595:20–37, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grelck:1999:SMM

[1594] C. Grelck. Shared memory multiprocessor support for SAC. Lecture Notes in Computer Science, 1595:38– 53, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hall:1999:TOS

[1595] J. G. Hall, C. Baker-Finch, P. Trinder, and D. J. King. Towards an operational semantics for a parallel non-strict functional language. Lecture Notes in Computer Science, 1595:54–71, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holyer:1999:CMI

[1596] I. Holyer and E. Spiliopoulou. Concurrent monadic interfacing. Lecture Notes in Computer Science, 1595:72– 87, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

King:1999:SPG

[1597] D. J. King, J. G. Hall, and P. Trinder. A strategic profiler for Glasgow Parallel Haskell. Lecture Notes in Computer Science, 1595:88–102, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klusik:1999:IED

[1598] U. Klusik, Y. Ortega, and R. Peña. Implementing Eden — or: Dreams become reality. Lecture Notes in Computer Science, 1595:103–119, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koopman:1999:ECP

[1599] P. Koopman and R. Plasmeijer. Efficient combinator parsers. Lecture Notes in Computer Science, 1595:120–136, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McAdam:1999:UST

[1600] B. J. McAdam. On the unification of substitutions in type inference. Lecture Notes in Computer Science, 1595:137–152, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pape:1999:HOD

[1601] D. Pape. Higher order demand propagation. Lecture Notes in Computer Science, 1595:153–168, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pil:1999:DTT

[1602] M. Pil. Dynamic types and type dependent functions. Lecture Notes in Computer Science, 1595:169–185, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reid:1999:PSB

[1603] A. Reid. Putting the spine back in the spineless tagless G-machine: An implementation of resumable blackholes. Lecture Notes in Computer Science, 1595:186–199, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reinke:1999:THJ

[1604] C. Reinke. Towards a Haskell/Java connection. Lecture Notes in Computer Science, 1595:200–215, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scholz:1999:CSE

[1605] S.-B. Scholz. A case study: Effects of WITH-Loop-Folding on the NAS benchmark MG in SAC. Lecture Notes in Computer Science, 1595:216–228, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Serrarens:1999:EMP

[1606] P. R. Serrarens and R. Plasmeijer. Explicit message passing for concurrent clean. Lecture Notes in Computer Science, 1595:229–246, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wade:1999:TFT

[1607] Vincent P. Wade. Towards future telecommunications management architectures. Lecture Notes in Computer Science, 1597:1–2, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1597/15970001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1597/15970001.
pdf.

Davison:1999:PPT

[1608] Rob Davison and Terry Turner. A practical perspective on TMN evolution. Lecture Notes in Computer Science, 1597:3–12, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1597/15970003.htm;   http:
//link.springer-ny.com/link/service/

series/0558/papers/1597/15970003.
pdf.

Hellemans:1999:AMT

[1609] Patrick Hellemans, Cliff Redmond, Koen Daenen, and Dave Lewis. Accounting management in a TINA-based service and network environment. Lecture Notes in Computer Science, 1597: 13–24, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1597/15970013.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1597/15970013.
pdf.

Davison:1999:NAO

[1610] Rob Davison and Jim Hardwicke. A new architecture for open and distributed network management. Lecture Notes in Computer Science, 1597:25– 38, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1597/15970025.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1597/15970025.
pdf.

Athanassiou:1999:IBR

[1611] Eleutherios Athanassiou. Intelligence in brokerage and retrieval. Lecture Notes in Computer Science, 1597:39– 40, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1597/15970039.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/1597/15970039.
pdf.

Gleizes:1999:SOL

[1612] Marie Pierre Gleizes, Alain Léger, Eleutherios Athanassiou, and Pierre Glize. Self-organization and learning in MultiAgent based brokerage services. Lecture Notes in Computer Science, 1597:41–54, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1597/15970041.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1597/15970041.
pdf.

Bessonov:1999:OAD

[1613] Mikhail Bessonov, Udo Heuser, Igor Nekrestyanov, and Ahmed Patel. Open architecture for distributed search systems. Lecture Notes in Computer Science, 1597:55–69, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1597/15970055.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1597/15970055.
pdf.

Nekrestyanov:1999:BTS

[1614] Igor Nekrestyanov, Tadhg O’Meara, Ahmed Patel, and Ekaterina Romanova. Building topic-specific collections with intelligent agents. Lecture Notes in Computer Science, 1597:70– 82, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/1597/15970070.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1597/15970070.
pdf.

Start:1999:I

[1615] Keith Start. Interworking. Lecture Notes in Computer Science, 1597:83– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1597/15970083.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1597/15970083.
pdf.

Mitra:1999:DCT

[1616] Nilo Mitra and Rob Brennan. Design of the CORBA/TC interworking gateway. Lecture Notes in Computer Science, 1597:84–100, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1597/15970084.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1597/15970084.
pdf.

Fischbeck:1999:CIS

[1617] Nils Fischbeck and Olaf Kath. CORBA interworking over SS.7. Lecture Notes in Computer Science, 1597:101–113, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1597/15970101.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1597/15970101.
pdf.

Brennan:1999:SSN

[1618] Rob Brennan, Brendan Jennings, and Thomas Curran. Signalling system no. 7 as an agent platform message transport protocol. Lecture Notes in Computer Science, 1597:114–128, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1597/15970114.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1597/15970114.
                                        pdf.

Quantz:1999:IAT

[1619] Joachim Quantz. Issues in agent technology. Lecture Notes in Computer Science, 1597:129–130, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1597/15970129.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1597/15970129.pdf.

Kaasinen:1999:UCA

[1620] Eija Kaasinen. Usability challenges in agent based services. Lecture Notes in Computer Science, 1597:131–142, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1597/15970131.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1597/15970131.
                                        pdf.

Breugst:1999:MMA

[1621] Markus Breugst and Sang Choy. Management of mobile agent based services. Lecture Notes in Computer Science, 1597:143–154, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1597/15970143.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1597/15970143.pdf.

Schelderup:1999:MAS

[1622] Kristian Schelderup and Jon Ølnes.

Mobile agent security — issues and directions. Lecture Notes in Computer Science, 1597:155–167, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1597/15970155.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1597/15970155.
pdf.

Choy:1999:CES

[1623] Sang Choy, Markus Breugst, and Thomas Magedanz. A CORBA environment supporting mobile objects. Lecture Notes in Computer Science, 1597:168–180, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1597/15970168.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1597/15970168.pdf.

Donnelly:1999:CR

[1624] Willie Donnelly. Component reuse.

Lecture Notes in Computer Science, 1597:181–182, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1597/15970181.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1597/15970181.pdf.

Hubert:1999:WBS

[1625] Christian Hubert. Web-based service customization in an IN-TINA environment. Lecture Notes in Computer Science, 1597:183–193, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1597/15970183.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1597/15970183.
                                        pdf.

Maknavicius:1999:CTS

[1626] Linas Maknavicius, Gautier Koscielny, and Simon Znaty. Customizing telecommunication services: Patterns, issues, and models. Lecture Notes in Computer Science, 1597:194–209, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1597/15970194.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1597/15970194.
                                        pdf.

Lewis:1999:MMC

[1627] Dave Lewis, Chris Malbon, and Alina DaCruz. Modelling management components for reuse using UML. Lecture Notes in Computer Science, 1597:210– 222, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1597/15970210.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1597/15970210.
pdf.

Vandermeulen:1999:ADM

[1628] Filip Vandermeulen, Piet Demeester, Piet De Ceuleners, and Jean-Marc Reynders. Automated design of modular SNMP-CORBA gateways and their application for the development of an ADSL access network manager. Lecture Notes in Computer Science, 1597:223– 238, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1597/15970223.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1597/15970223.
pdf.

Zizza:1999:MM

[1629] Fabrizio Zizza. Mobility management.

Lecture Notes in Computer Science, 1597:239–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1597/15970239.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1597/15970239.
pdf.

Francis:1999:RAU

[1630] John C. Francis and Bernhard Diem.

Roaming aspects of ubiquitous mobile service support. Lecture Notes in Computer Science, 1597:240–249, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1597/15970240.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1597/15970240.
                                        pdf.

Macaire:1999:PND

[1631] Alain Macaire and David Carlier. A personal naming and directory service for UMTS users. Lecture Notes in Computer Science, 1597:250–262, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1597/15970250.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1597/15970250.
                                        pdf.

Donnelly:1999:AMA

[1632] Mark Donnelly, Ivan Griffin, Conor Morris, and John Nelson. The application of mobile agent technology to mobility management in MARINE. Lecture Notes in Computer Science, 1597:263–274, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1597/15970263.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1597/15970263.
                                        pdf.

Efremidis:1999:SC

[1633] Sofoklis Efremidis. Service creation. Lecture Notes in Computer Science, 1597:275–276, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1597/15970275.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1597/15970275.pdf.

Lodge:1999:ATS

[1634] Fiona Lodge, Kristofer Kimbler, and Manuel Hubert. Alignment of the TOSCA and SCREEN approaches to service creation. Lecture Notes in Computer Science, 1597:277–290, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1597/15970277.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1597/15970277.
pdf.

Kaltabani:1999:SLD

[1635] Anastasia Kaltabani, Elsa Cardoso, Evangelia Tzifa, Panagiotis Demestichas, António Silva, and Miltiades Anagnostou. Service logic deployment and distribution configuration. Lecture Notes in Computer Science, 1597:291– 304, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1597/15970291.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1597/15970291.
pdf.

Ing:1999:SRT

[1636] Sarom Ing and Steve Rudkin. Simplifying real-time multimedia application development using session descriptions. Lecture Notes in Computer Science, 1597:305–314, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1597/15970305.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1597/15970305.
                                        pdf.

Kimbler:1999:ESP

[1637] Kristofer Kimbler, Carl-Henrik Hagenfeldt, Jan Ellsberger, and Gustav Bergman. An environment for IN service prototyping and validation. Lecture Notes in Computer Science, 1597: 315–327, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1597/15970315.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1597/15970315.
                                        pdf.

Kind:1999:TSI

[1638] Andreas Kind, Miltiades E. Anagnostou, Malamati D. Louta, Jan Nicklisch, and Juan Tous. Towards the seamless integration of mobile agents into service creation practice. Lecture Notes in Computer Science, 1597:328– 341, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1597/15970328.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1597/15970328.
                                        pdf.

Olsen:1999:PCU

[1639] Anders Olsen, Didier Demany, Elsa Cardoso, Fiona Lodge, Mario Kolberg, Morgan Björkander, and Richard Sinnott. The pros and cons of using SDL for creation of distributed services. Lecture Notes in Computer Science, 1597:342–354, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1597/15970342.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1597/15970342.
pdf.

Delgado:1999:MSS

[1640] Jaime Delgado. Multimedia services & security. Lecture Notes in Computer Science, 1597:355–356, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1597/15970355.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1597/15970355.
pdf.

Sellin:1999:AVS

[1641] Eric Sellin, Peter Loosemore, Sohail Rana, Jürgen Dittrich, and Maarten Wegdam. Audio/video stream binding in a pan-European service management platform. Lecture Notes in Computer Science, 1597:357–372, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1597/15970357.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1597/15970357.
pdf.

Becchini:1999:JEM

[1642] Roberto Becchini and Gianluca De Petris. Java enabled MPEG-4 services: The MPEG-J framework. Lecture Notes in Computer Science, 1597:373– 384, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1597/15970373.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1597/15970373.
                                        pdf.

Hardjono:1999:KMF

[1643] Thomas Hardjono. Key management framework for IP multicast infrastructure security. Lecture Notes in Computer Science, 1597:385–398, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1597/15970385.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1597/15970385.
                                        pdf.

Trigila:1999:MT

[1644] Sebastiano Trigila. Mobility in TINA. Lecture Notes in Computer Science, 1597:399–400, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1597/15970399.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1597/15970399.pdf.

Lombardo:1999:PEA

[1645] Alfio Lombardo, Pietro Nicosia, Sergio Palazzo, Marina Samarotto, and Giovanni Schembra. Performance evaluation of an allocation strategy for TINA compliant mobile agents supporting PCS in a multiretailer environment. Lecture Notes in Computer Science, 1597:401–415, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1597/15970401.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1597/15970401.
pdf.

Kuepper:1999:LTU

[1646] A. Kuepper. Locating TINA user agents: Strategies for a broker federation and their comparison. Lecture Notes in Computer Science, 1597:416– 428, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kupper:1999:LTU

[1647] Axel Küpper. Locating TINA user agents: Strategies for a broker federation and their comparison. Lecture Notes in Computer Science, 1597:416– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1597/15970416.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1597/15970416.
pdf.

Stamoulis:1999:ABS

[1648] George D. Stamoulis, Dimitrios Kalopsikakis, Anna Kirikoglou, Vasilios A. Siris, Didoe Prevedourou, Sofoklis Efremidis, and Henryka Jormakka. Agent-based service and retailer selection in a personal mobility context. Lecture Notes in Computer Science, 1597:429–442, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1597/15970429.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1597/15970429.
pdf.

Galis:1999:DIS

[1649] Alex Galis. Distributed intelligence in services and networks. Lecture Notes in Computer Science, 1597:443– 445, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1597/15970443.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1597/15970443.
pdf.

Schoo:1999:MTR

[1650] Peter Schoo, Christian Egelhaaf, Tim Eckardt, Nazim Agoulmine, and Michael Tschichholz. Modularization of TINA reference points for information networking. Lecture Notes in Computer Science, 1597:446–458, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1597/15970446.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1597/15970446.
pdf.

Solomonides:1999:REI

[1651] Christos Solomonides and Mark Searle. Relevance of existing intelligent network infrastructure to the Internet. Lecture Notes in Computer Science, 1597:459–468, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1597/15970459.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1597/15970459.
pdf.

Faglia:1999:IDM

[1652] Lorenzo Faglia, Thomas Magedanz, and Andreas Papadakis. Introduction of DOT/MAT into a broadband IN architecture for flexible service provision. Lecture Notes in Computer Science, 1597:469–481, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1597/15970469.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1597/15970469.
pdf.

Cabell:1999:UTP

[1653] Quincy Cabell, Peter Kirstein, Theodore Pagtzis, Thanassis Tiropanis, and Lichun Wang. The use of TINA principles in the management of Internet multimedia conferences. Lecture Notes in Computer Science, 1597:482– 498, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1597/15970482.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1597/15970482.
                                        pdf.

Ndumu:1999:ASR

[1654] Divine Ndumu. Agent support for resource, network, and enterprise management. Lecture Notes in Computer Science, 1597:499–500, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1597/15970499.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1597/15970499.
                                        pdf.

Molina:1999:KMM

[1655] Martín Molina and Sascha Ossowski. Knowledge modelling in multiagent systems: The case of the management of a national network. Lecture Notes in Computer Science, 1597:501– 513, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1597/15970501.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1597/15970501.
                                        pdf.

Bigham:1999:MAS

[1656] John Bigham, Laurie Cuthbert, Alex Hayzelden, and Zhiyuan Luo. Multi-agent system for network resource management. Lecture Notes in Computer Science, 1597:514–526, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/1597/15970514.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1597/15970514.
pdf.

Sugauchi:1999:EEM

[1657] Kiminori Sugauchi, Satoshi Miyazaki, Stefan Covaci, and Tianning Zhang. Efficiency evaluation of a mobile agent based network management system. Lecture Notes in Computer Science, 1597:527–535, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1597/15970527.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1597/15970527.pdf.

Covaci:1999:MIA

[1658] Stefan Covaci, Markus Zell, and Rudi Broos. A mobile intelligent agent based virtual enterprise support environment. Lecture Notes in Computer Science, 1597:536–549, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1597/15970536.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1597/15970536.
pdf.

Yuan:1999:MBE

[1659] Soe-Tsyr Yuan. MAS building environments with product-line-architecture awareness. Lecture Notes in Computer Science, 1599:1–17, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/

                                        series/0558/bibs/1599/15990001.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1599/15990001.pdf.

Kinoshita:1999:AFF

[1660] Tetsuo Kinoshita and Kenji Sugawara. ADIPS framework for flexible distributed systems. Lecture Notes in Computer Science, 1599:18– 32, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1599/15990018.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1599/15990018.
                                        pdf.

Vally:1999:CRC

[1661] Jean-Dany Vally and Rémy Courdier. A conceptual “role-centered” model for design of multi-agent systems. Lecture Notes in Computer Science, 1599:33– 46, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1599/15990033.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1599/15990033.
                                        pdf.

Lee:1999:DAA

[1662] Jaeho Lee. Desiderata in agent architectures for coordinating multiagent systems. Lecture Notes in Computer Science, 1599:47–60, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1599/15990047.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1599/15990047.
pdf.

Lee:1999:TBN

[1663] Kyoung Jun Lee and Yong Sik Chang.

Time-bounded negotiation framework for multi-agent coordination. Lecture Notes in Computer Science, 1599:61– 75, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1599/15990061.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1599/15990061.
pdf.

Wu:1999:FGT

[1664] Shih-Hung Wu and Von-Wun Soo.

A fuzzy game theoretic approach to multi-agent coordination. Lecture Notes in Computer Science, 1599:76– 87, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1599/15990076.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1599/15990076.
pdf.

Lee:1999:PPR

[1665] Jung-Jin Lee and Robert McCartney. Partial plan recognition using predictive agents. Lecture Notes in Computer Science, 1599:88–102, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1599/15990088.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1599/15990088.
                                        pdf.

Ramamonjisoa:1999:MAC

[1666] David Ramamonjisoa. Multi-agent coordination with OASIS. Lecture Notes in Computer Science, 1599:103– 117, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1599/15990103.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1599/15990103.
                                        pdf.

Hsiang:1999:PWR

[1667] Jieh Hsiang and Hsieh-Chang Tu. Personalized Web retrieval: Three agents for retrieving Web information. Lecture Notes in Computer Science, 1599:118– 132, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1599/15990118.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1599/15990118.
                                        pdf.

Jieh:1999:PWR

[1668] H. Jieh and H.-C. Tu. Personalized Web retrieval: Three agents for retrieving Web information. Lecture Notes in Computer Science, 1599:118–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kitamura:1999:SAM

[1669] Yasuhiko Kitamura, Tomoya Noda, and Shoji Tatsumi. Single-agent and multi-agent approaches to WWW information integration. Lecture Notes in Computer Science, 1599:133–147, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1599/15990133.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1599/15990133.
pdf.

Hsu:1999:MAF

[1670] Jane Yung jen Hsu. A multiagent framework for intranet service integration. Lecture Notes in Computer Science, 1599:148–161, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1599/15990148.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1599/15990148.
pdf.

Wang:1999:SAM

[1671] Xiao Feng Wang, Kwok Yan Lam, and X. Yi. Secure agent-mediated mobile payment. Lecture Notes in Computer Science, 1599:162–173, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1599/15990162.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1599/15990162.
pdf.

Wang:1999:SCS

[1672] Hongxue Wang, Vijay Varadharajan, and Yan Zhang. A secure communication scheme for multiagent systems. Lecture Notes in Computer Science, 1599:174–186, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1599/15990174.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1599/15990174.
                                        pdf.

Anonymous:1999:AIf

[1673] Anonymous. Author index. Lecture Notes in Computer Science, 1599:187– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/1599/1599indx.pdf.